### v1

# Automatic doubly robust inference for linear functionals via calibrated debiased machine learning

Abstract
In causal inference, many estimands of interest can be expressed as a linear functional of the outcome regression function; this includes, for example, average causal effects of static, dynamic and stochastic interventions. For learning such estimands, in this work, we propose novel debiased machine learning estimators that are doubly robust asymptotically linear, thus providing not only doubly robust consistency but also facilitating doubly robust inference (e.g., confidence intervals and hypothesis tests). To do so, we first establish a key link between calibration, a machine learning technique typically used in prediction and classification tasks, and the conditions needed to achieve doubly robust asymptotic linearity. We then introduce calibrated debiased machine learning (C-DML), a unified framework for doubly robust inference, and propose a specific C-DML estimator that integrates cross-fitting, isotonic calibration, and debiased machine learning estimation. A C-DML estimator maintains asymptotic linearity when either the outcome regression or the Riesz representer of the linear functional is estimated sufficiently well, allowing the other to be estimated at arbitrarily slow rates or even inconsistently. We propose a simple bootstrap-assisted approach for constructing doubly robust confidence intervals. Our theoretical and empirical results support the use of C-DML to mitigate bias arising from the inconsistent or slow estimation of nuisance functions.
Keywords: Debiasing, isotonic calibration, double robustness, targeted learning, causal inference.
1 Introduction
1.1 Motivation
Nonparametric inference on the causal effects of dynamic and stochastic treatment interventions using flexible statistical learning techniques is a fundamental problem of causal inference. These inferential targets fall within a broader class of target parameters that can be expressed as linear functionals of the outcome regression, that is, the conditional mean of the outcome given treatment assignment and baseline covariates (Chernozhukov et al., 2018b; Rotnitzky et al., 2021; Hirshberg and Wager, 2021). To achieve inference for such parameters, various debiased machine learning frameworks are available, including one-step estimation (Pfanzagl and Wefelmeyer, 1985; Bickel et al., 1993), estimating equations and double machine learning (Robins et al., 1995, 1994; van der Laan and Robins, 2003; Chernozhukov et al., 2018a), and targeted minimum loss estimation (TMLE) (van der Laan and Rubin, 2006; van der Laan and Rose, 2011). When performing inference on linear functionals, debiased machine learning methods frequently produce estimators that demonstrate robustness against inconsistent or slow estimation of certain nuisance functions, a property known as double robustness (Bang and Robins, 2005; Smucler et al., 2019). For example, when estimating the average treatment effect (ATE), doubly robust estimators remain consistent as long as either the outcome regression or the propensity score is consistently estimated, even if the other is inconsistently estimated. This robustness property extends more generally to linear functionals of the outcome regression, necessitating consistent estimation of only the outcome regression itself or the Riesz representer of the linear functional (Chernozhukov et al., 2018b; Rotnitzky et al., 2021).
Extending the double robustness property of debiased machine learning estimators to statistical inference, that is, to uncertainty quantification, is a challenging problem. Achieving asymptotic linearity for such estimators typically requires consistent estimation of both the outcome regression and the Riesz representer at sufficiently fast rates. When one of the nuisance estimators is inconsistent or estimated too slowly, desirable properties such as asymptotic linearity, asymptotic normality, and $n^{\frac{1}{2}}$–consistency can be lost, compromising the validity of inferences (Benkeser et al., 2017). In particular, in such cases, typical confidence intervals and p-values can be expected to be poorly calibrated. One approach to achieving doubly robust inference is through the use of parametric methods, such as maximum likelihood estimation, for estimating nuisance parameters. These methods exhibit model double robustness, maintaining asymptotic linearity when at least one of the nuisance parametric models is correctly specified (Vermeulen and Vansteelandt, 2015, 2016). However, the applicability of these methods for nonparametric inference is limited due to their reliance on strong assumptions, as both models are then likely to be misspecified in practice (Kang and Schafer, 2007).
To overcome the limitations of parametric methods, several estimators exhibiting sparsity double robustness, a weaker form of model double robustness, have been proposed (Avagyan and Vansteelandt, 2017; Athey et al., 2018; Bradic et al., 2019; Smucler et al., 2019; Tan, 2020). For instance, in the context of high-dimensional generalized linear regression models, Avagyan and Vansteelandt (2017) and Bradic et al. (2019) introduced doubly robust $\ell_{1}$–regularized regression estimators of the ATE. These estimators achieve asymptotic linearity when either the outcome regression or the propensity score exhibits sufficient sparsity in the covariates, allowing for misspecification of one of the nuisance models. Additionally, Athey et al. (2018) and Tan (2020) proposed singly robust estimators that are robust to misspecification of the outcome regression model, assuming sufficient sparsity in the propensity score. However, a significant drawback of these estimators is their reliance on strong sparsity assumptions for the nuisance functions, which may fail to hold in practice. Smucler et al. (2019) developed a unifying framework for constructing sparsity doubly robust estimators using $\ell_{1}$–regularization techniques. They relax the strong sparsity assumptions by requiring that nuisance estimators based on possibly misspecified models converge to approximately sparse functions sufficiently quickly. However, as the authors note, there is limited justification to expect an inconsistent nuisance estimator derived from a misspecified model to tend to a sparse or approximately sparse limiting function, even if the true nuisance function depends on only a small subset of covariates. Furthermore, these methods are constrained by their reliance on $\ell_{1}$–regularized generalized linear models for the nuisance functions, limiting their ability to fully leverage the advantages offered by modern learning techniques.
The seminal work of van der Laan (2014) introduced a debiasing technique, based on the TMLE framework (van der Laan and Rose, 2011), for constructing estimators with which it is possible to perform doubly robust inference using generic machine learning techniques. The resulting estimators are doubly robust asymptotically linear (DRAL), in the sense that they remain asymptotically linear even when one of the nuisance functions is estimated too slowly or even inconsistently. Their approach has since been applied to nonparametric inference of the ATE (Benkeser et al., 2017), mean outcomes under informative missingness (Díaz and van der Laan, 2017), counterfactual survival curves (Díaz, 2019), and semiparametric inference on the ATE in partially linear regression models (Dukes et al., 2021). For inference on the ATE, these estimators involve iteratively debiasing initial estimators of the propensity score and outcome regression to linearize the doubly robust remainder term of the resulting parameter estimator. However, the TMLE-based framework for doubly robust inference has certain limitations. First, the iterative debiasing procedure used to obtain doubly robust inference is challenging to justify theoretically and can be computationally demanding to implement, posing challenges for its application to large-scale data and complex data structures. Second, when considering new parameters of interest, a new debiasing algorithm must be derived on a case-by-case basis, requiring careful analysis of the relevant remainder terms and novel targeting algorithms. Notably, to date, there is no unified approach for constructing DRAL estimators or obtaining doubly robust inference for the entire class of linear functionals of the outcome regression.
1.2 Contributions of this work
In this work, we aim to derive doubly robust inferential procedures for linear functionals of the outcome regression. Specifically, these procedures should facilitate calibrated inference even when one nuisance estimator converges to its intended target arbitrarily slowly or even inconsistently. We introduce a novel estimation framework, which we refer to as calibrated debiased machine learning (C-DML), that allows for the construction of DRAL estimators and confidence intervals that provide doubly robust inferences. To accomplish this, we rely on cross-fitted estimators of the outcome regression and Riesz representer obtained through any generic supervised learning algorithms. Within the C-DML procedure, calibrated nuisance estimators may be constructed, for example, using a generalization of isotonic calibration — a technique commonly used in prediction and classification tasks (Zadrozny and Elkan, 2001; Niculescu-Mizil and Caruana, 2005; van der Laan et al., 2023b) — based on loss functions tailored for the task at hand. These calibrated estimators are then used to create a calibrated one-step debiased estimator, which we refer to as a C-DML estimator. For inference on counterfactual means, an augmented inverse probability weighted estimator in which the cross-fitted propensity score and outcome regression estimators are calibrated using isotonic calibration provides an example of a C-DML estimator. We derive the closed-form influence function of our proposed estimators, allowing us to construct confidence intervals without needing to know which, if any, of the nuisance functions is possibly estimated too slowly or inconsistently. Moreover, to avoid the estimation of additional nuisance functions that arise in van der Laan (2014) and related works, we propose an automated bootstrap-assisted procedure for doubly robust inference, which is both computationally efficient and valid as long as at least one of the nuisance estimators is consistent at a fast enough rate.
Our approach is motivated by the DRAL TMLE estimators of the ATE proposed in van der Laan (2014) and Benkeser et al. (2017). While we share the objective of constructing nuisance estimators to achieve doubly robust asymptotic linearity by solving certain score equations, our strategy and implementation for doing so differs significantly, even when focusing on estimating the ATE. Specifically, our approach is a noniterative and computationally efficient two-step procedure that can be easily implemented using commonly available software for isotonic regression. Moreover, our method is not dependent on the specific form of the linear functional but rather on the availability of initial nuisance estimators and the ability to evaluate an empirical estimator of the linear functional as a black box. Furthermore, even in the special case of the ATE, our theoretical results refine those presented by van der Laan (2014) and Benkeser et al. (2017), establishing that our estimator achieves asymptotic linearity not only under consistent estimation of one nuisance function alone but also when both nuisance functions are estimated consistently, with one estimated too slowly.
Prior work has empirically investigated the finite-sample benefits of calibrating propensity score estimates in IPW and AIPW estimators of the ATE (Gutman et al., 2022; Deshpande and Kuleshov, 2023; Ballinari and Bearth, 2024). However, these studies have focused exclusively on the calibration of a single nuisance function — the propensity score — and a single parameter, the ATE. Additionally, none of these works has provided theoretical justification for the improved performance achieved through propensity score calibration beyond the observation that calibration can reduce the estimation error of nuisance functions. Our contribution extends this line of research by demonstrating that, for the general class of linear functionals, calibration of both the outcome regression and the Riesz representer is essential for further debiasing, both in finite samples and asymptotically, and for constructing DRAL estimators. In the special case of the ATE, we find that the outcome regression should also be calibrated and that it is the inverse propensity score, rather than the propensity score itself, that must be calibrated.
This paper is organized as follows. In Section 2, we introduce the class of linear functional parameters we consider and provide background on the standard efficient one-step debiased estimator and its rate double robustness properties. In Section 3, we formulate the problem of doubly robust inference, introduce a novel model-agnostic linear expansion for the bias of one-step debiased estimators, and demonstrate how further debiasing of nuisance functions can be employed to construct DRAL estimators. In Section 4, we present a novel link between nuisance estimator calibration and the doubly robust inference properties of debiased machine learning estimators, use this link to build our general calibrated debiased machine learning framework, and propose our novel isotonic-calibrated DML estimator. In Section 5, we establish the theoretical properties of our proposed C-DML estimators and bootstrap-assisted confidence intervals. Finally, in Section 6, we investigate the empirical performance of C-DML in synthetic and semi-synthetic experiments. We provide concluding remarks in Section 7.
2 Problem setup
2.1 Data structure and target parameter
[…]
Our objective is to obtain doubly robust nonparametric inference for a feature $\tau_{0}:=\psi_{0}(\mu_{0})$ of the outcome regression $\mu_{0}$ arising from a continuous linear functional $\psi_{0}:\mathcal{H}\rightarrow\mathbb{R}$. Here, $\psi_{0}$ represents the evaluation at $P_{0}$ of a smooth mapping $P\mapsto\psi_{P}(\cdot)$ defined on the statistical model $\mathcal{M}$. In other words, we explicitly allow the functional of interest to itself depend on the unknown data-generating distribution $P_{0}$. We require $\psi_{0}$ to satisfy the following conditions:
  1. A1)
    (linearity and boundedness) it holds that $\psi_{0}(c\mu_{1}+\mu_{2})=c\psi_{0}(\mu_{1})+\psi_{0}(\mu_{2})$ for each $\mu_{1},\mu_{2}\in\mathcal{H}$ and $c\in\mathbb{R}$ and that $\sup_{\mu\in\mathcal{H}}|\psi_{0}(\mu)|/\|\mu\|<\infty$;
  2. A2)
    (Hellinger continuity): $P\mapsto\psi_{P}$ is a continuous functional at $P_{0}$ with domain $\mathcal{M}$ equipped with the Hellinger distance and range $\{\psi_{P}:P\in\mathcal{M}\}$ equipped with the operator norm;
  3. A3)
    (pathwise differentiability) for each $\mu\in\mathcal{H}$, $P\mapsto\psi_{P}(\mu)$ is pathwise differentiable at $P_{0}$ relative to $\mathcal{M}$ with efficient influence function $\widetilde{\phi}_{0,\mu}:=\phi_{0,\mu}-P_{0}\phi_{0,\mu}$, where $\phi_{0,\mu}\in\mathcal{H}$ is such that, for some $L\in(0,\infty)$, the inequality $|\phi_{0,\mu_{1}}(A,W)-\phi_{0,\mu_{2}}(A,W)|\leq L|\mu_{1}(A,W)-\mu_{2}(A,W)|$ holds $P_{0,A,W}$–almost surely for each $\mu_{1},\mu_{2}\in\mathcal{H}$;
We also assume that we can construct a well-behaved estimator $\psi_{n}$ of $\psi_{0}$ based on the available data, such that $\psi_{n}(\mu)=\psi_{0}(\mu)+(P_{n}-P_{0})\widetilde{\phi}_{0,\mu}+o_{p}(n^{-\frac{1}{2}})$ locally uniformly around each $\mu\in\mathcal{H}$ — this requirement is stated precisely as Condition C7 later.
The target parameter mapping $P\mapsto\Psi(P):=\psi_{P}(\mu_{P})$ we consider encompasses various interesting parameters that exhibit linear dependence on the outcome regression. Notably, many estimands of interest in causal inference, such as counterfactual means and the ATE, correspond to functionals of the form
$\mu\mapsto\psi_{0}(\mu)=E_{0}\left\{m(A,W,\mu)\right\}$ (1)
for a known map $m:\mathcal{A}\times\mathcal{W}\times\mathcal{H}\rightarrow\mathbb{R}$. For example, the counterfactual mean outcome under an intervention that sets treatment to level $a_{0}\in\mathcal{A}$ is obtained by taking $m:(a,w,\mu)\mapsto\mu(a_{0},w)$, whereas the ATE is obtained by taking $m:(a,w,\mu)\mapsto\mu(1,w)-\mu(0,w)$. For such functionals, $\psi_{0}$ satisfies A1 if $m$ is linear in its last argument and if, for example, there exists some $M\in(0,\infty)$ such that, for each $\mu\in\mathcal{H}$, $|m(A,W,\mu)|\leq M|\mu(A,W)|$ $P_{0,A,W}$–almost surely. Furthermore, A3 is satisfied with $\phi_{0,\mu}:=m(\cdot,\mu)$, and an asymptotically linear estimator of $\psi_{0}(\mu)$ is given by $\psi_{n}(\mu):=\frac{1}{n}\sum_{i=1}^{n}m(A_{i},W_{i},\mu)$. Specifically, for the ATE functional, this yields $\mu\mapsto\psi_{n}(\mu)=\frac{1}{n}\sum_{i=1}^{n}\{\mu(1,W_{i})-\mu(0,W_{i})\}$. Additional examples of such functionals can be found in Chernozhukov et al. (2018b) and Rotnitzky et al. (2021).
By the Riesz representation theorem, there exists a unique nonparametric Riesz representer $\alpha_{0}\in\mathcal{H}$ such that $\psi_{0}$ can be expressed as an inner product $\mu\mapsto\psi_{0}(\mu)=\langle\alpha_{0},\mu\rangle$. The Riesz representer can be written as a minimizer of the objective function $\alpha\mapsto\|\alpha\|^{2}-2\psi_{0}(\alpha)$ over $\mathcal{H}$ (Chernozhukov et al., 2018b, 2021). It plays a crucial role in the construction of asymptotically linear and nonparametric efficient estimators of $\tau_{0}$. In particular, $\alpha_{0}$ appears in the efficient influence function of the parameter $P\mapsto\psi_{P}(\mu_{P})$, viewed as a functional on the nonparametric statistical model $\mathcal{M}$, and this plays a critical role in the debiasing of plug-in estimators based on the functional $\psi_{n}$. Below, for any $\mu,\alpha\in\mathcal{H}$, we denote by $D_{\mu,\alpha}$ the mapping $z\mapsto\alpha(a,w)\left\{y-\mu(a,w)\right\}$.
Theorem 1.
Under conditions A1A3, $P\mapsto\psi_{P}(\mu_{P})$ is pathwise differentiable at $P_{0}$ under the model $\mathcal{M}$ and has nonparametric efficient influence function $D_{0}:=\widetilde{\phi}_{0,\mu_{0}}+D_{\mu_{0},\alpha_{0}}$ with $\widetilde{\phi}_{0,\mu_{0}}$ defined in A3. Furthermore, for any $\mu,\alpha\in\mathcal{H}$, $\psi_{0}$ admits the representation
$\psi_{0}(\mu)-\psi_{0}(\mu_{0})=-P_{0}D_{\mu,\alpha}+\langle\mu-\mu_{0},\alpha_{0}-\alpha\rangle\,.$
As indicated above, our framework covers various pathwise differentiable parameters in semiparametric and nonparametric statistics of particular interest in causal inference. Below, we provide various examples.
Example 1 (counterfactual means under general interventions).
Let $Y^{a_{0}}$ denote the potential outcome that would be observed if treatment $a_{0}\in\mathcal{A}$ were administered. Under certain causal assumptions (Rubin, 1974), the mean of $Y^{a_{0}}$ is identified by $\tau_{0}:=E_{0}\{\mu_{0}(a_{0},W)\}$, which corresponds to the linear functional $\psi_{0}:\mu\mapsto E_{0}\{\mu(a_{0},W)\}$. Doubly robust inference for this parameter was considered in van der Laan (2014) and Benkeser et al. (2017). More generally, the mean of the counterfactual outcome under a stochastic intervention that, given $W=w$, draws a treatment value $A$ at random from the distribution with probability mass function $a\mapsto\omega_{0}(a|w)$ depending only on $P_{0,A,W}$ can be identified by $\tau_{0}:=\int\mu_{0}(a,w)\omega_{0}(a|w)P_{0,A,W}(da,dw)$. If $(a,w)\mapsto\alpha_{0}(a,w):=\omega_{0}(a|w)/\pi_{0}(a|w)$ has finite $\mathcal{H}$–norm, then conditions A1A3 hold and $\alpha_{0}$ is the nonparametric Riesz representer of $\tau_{0}$.
Example 2 (average treatment effect under covariate-dependent outcome missingness).
Suppose that the observed data unit consists of $(W,S,\Delta,\Delta Y)$, where $S\in\{0,1\}$ is a binary treatment, $\Delta\in\{0,1\}$ is a missingness indicator taking value $1$ if $Y$ is observed, and $\Delta Y$ is the observed outcome that equals $Y$ if $\Delta=1$ and 00 otherwise. Under certain causal assumptions, the ATE under covariate-dependent outcome missingness is identified by the estimand $\tau_{0}:=E_{0}\{E_{0}(\Delta Y\,|\,S=1,\Delta=1,W)-E_{0}(\Delta Y\,|\,S=0,\Delta=1,W)\}$. Doubly robust inference for this estimand was considered in Díaz and van der Laan (2017). Using the notation of this section, we can write $Z:=(W,A,Y^{\prime})$, where $A:=(S,\Delta)$ is a multi-valued interventional variable and $Y^{\prime}:=\Delta Y$ is an outcome with missingness. The linear functional corresponding with $\tau_{0}$ is $\psi_{0}:\mu\mapsto E_{0}\{\mu(1,1,W)-\mu(0,1,W)\}$. The nonparametric Riesz representer of $\psi_{0}$ is given by $\alpha_{0}:(s,\delta,w)\mapsto(2s-1)\delta/P_{0}(\Delta=1,S=s\,|\,W=w)$.
Example 3 (projection onto semiparametric working model).
Suppose now that the parameter of interest is $\tau_{0}=E_{0}\{m(A,W,\mu_{0})\}$ for a known function $m:\mathcal{A}\times\mathcal{W}\times\mathcal{H}\rightarrow\mathbb{R}$. If we believe that $\mu_{0}$ can be approximated well by elements in a closed linear subspace $\Theta\subset\mathcal{H}$, we could consider instead of $\tau_{0}$ the working estimand $E_{0}\{m(A,W,\Pi_{0}\mu_{0})\}$ with $\Pi_{0}:\mu\mapsto\operatorname*{argmin}_{\theta\in\Theta}\|\mu-\theta\|$ defined on $\mathcal{H}$. For example, we may take $\Theta$ to be a partially linear regression model that assumes a known parametric form for the conditional average treatment effect (CATE). Doubly robust inference for the ATE in partially linear regression models was considered in Dukes et al. (2021). The nonparametric Riesz representer of the working functional $\mu\mapsto E_{0}\{m(A,W,\Pi_{0}\mu)\}$ is given by $\Pi_{0}\alpha_{0}$, where $\alpha_{0}$ is the nonparametric Riesz representer of the original functional $\mu\mapsto E_{0}\{m(A,W,\mu)\}$. For a finite dimensional sieve $\Theta_{n}\subset\Theta$ approximating $\Theta$, a possible estimator of this functional is $\mu\mapsto\psi_{n}(\mu):=\frac{1}{n}\sum_{i=1}^{n}m(A_{i},W_{i},\Pi_{n}\mu)$, where $\Pi_{n}\mu:=\operatorname*{argmin}_{\mu\in\Theta_{n}}\sum_{i=1}^{n}\{Y_{i}-\mu(A_{i},W_{i})\}^{2}$.
2.2 One-step debiased estimation and rate double robustness
Suppose we are given arbitrary estimators $\mu_{n}$ and $\alpha_{n}$ of the outcome regression $\mu_{0}$ and Riesz representer $\alpha_{0}$. For example, as in Chernozhukov et al. (2018b, 2021), the estimators $\mu_{n}$ and $\alpha_{n}$ can be learned directly using any supervised machine learning algorithm based on the least-squares loss $(z,\mu)\mapsto\{y-\mu(a,w)\}^{2}$ and the Riesz loss $(z,\alpha)\mapsto\alpha^{2}(a,w)-2\psi_{n}(\alpha)$, respectively. Given these nuisance estimators, the plug-in estimator $\psi_{n}(\mu_{n})$ could be used to estimate $\tau_{0}$. However, when $\mu_{n}$ is obtained using flexible statistical learning tools, this estimator typically lacks both $n^{\frac{1}{2}}$–consistency and asymptotic normality due to excessive bias that arises from the first-order dependence of $\psi_{n}(\mu_{n})-\psi_{0}(\mu_{0})$ on the nuisance estimation error $\mu_{n}-\mu_{0}$ (van der Laan and Rose, 2011; Chernozhukov et al., 2018b). To overcome this sensitivity, debiasing methods must typically be used to eliminate the first-order bias of the plug-in estimator $\psi_{n}(\mu_{n})$. For example, the one-step debiased estimator (Chernozhukov et al., 2018b, 2021)
$\tau_{n}:=\psi_{n}(\mu_{n})+\frac{1}{n}\sum_{i=1}^{n}\alpha_{n}(A_{i},W_{i})\{Y_{i}-\mu_{n}(A_{i},W_{i})\}$
could be used. The incorporation of this debiasing step renders the resulting one-step debiased estimator $\tau_{n}$ insensitive not only to the nuisance estimation error $\mu_{n}-\mu_{0}$ but also to $\alpha_{n}-\alpha_{0}$, thereby allowing for fast estimation rates and inference for $\tau_{0}$, even when $\alpha_{0}$ and $\mu_{0}$ are estimated at comparatively slow rates. The above estimator is a generalization of the well-known augmented inverse probability weighted estimator (Robins et al., 1994, 1995; Bang and Robins, 2005) and has also been referred to as an augmented balancing weighted (ABW) estimator (Bruns-Smith et al., 2023).
Suppose that both $\alpha_{n}$ and $\mu_{n}$ tend to possibly incorrect limits $\overline{\alpha}_{0}$ and $\overline{\mu}_{0}$, in the sense that $\|\alpha_{n}-\overline{\alpha}_{0}\|=o_{p}(1)$ and $\|\mu_{n}-\overline{\mu}_{0}\|=o_{p}(1)$, with either $\overline{\mu}_{0}=\mu_{0}$ or $\overline{\alpha}_{0}=\alpha_{0}$. To understand the asymptotic properties of the one-step debiased estimator $\tau_{n}$, it is useful to scrutinize the expansion in Theorem 1, from which the doubly robust decomposition of the estimation error
$\tau_{n}-\tau_{0}=(P_{n}-P_{0})(\widetilde{\phi}_{0,\mu_{n}}+D_{\mu_{n},\alpha_{n}})+\text{Rem}_{\mu_{n}}(\psi_{n},\psi_{0})+\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ (2)
can be derived, where $D_{\mu_{n},\alpha_{n}}$ is an estimator of the efficient influence function component $D_{\mu_{0},\alpha_{0}}$ in Theorem 1, $\text{Rem}_{\mu_{n}}(\psi_{n},\psi_{0}):=\psi_{n}(\mu_{n})-\psi_{0}(\mu_{n})-(P_{n}-P_{0})\widetilde{\phi}_{0,\mu_{n}}$ is a linearization remainder, and $\widetilde{\phi}_{0,\mu_{n}}$ is as defined in A3. Under weak conditions, the first term in the decomposition is asymptotically linear, satisfying $(P_{n}-P_{0})(\widetilde{\phi}_{0,\mu_{n}}+D_{\mu_{n},\alpha_{n}})=(P_{n}-P_{0})\overline{D}_{0}+o_{p}(n^{-\frac{1}{2}})$, where the influence function $\overline{D}_{0}:=\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}$ reduces to the efficient influence function in Theorem 1 when both $\overline{\mu}_{0}=\mu_{0}$ and $\overline{\alpha}_{0}=\alpha_{0}$. The asymptotic linearity of $\psi_{n}$ in a locally uniform sense implies that the second term is asymptotically negligible, in that $\text{Rem}_{\mu_{n}}(\psi_{n},\psi_{0})=o_{p}(n^{-\frac{1}{2}})$, and is, in fact, identically zero for functionals of the form given in (1). Consequently, even in the case of inconsistent estimation of one nuisance function, it can typically be established, under regularity conditions, that
$\tau_{n}-\tau_{0}=(P_{n}-P_{0})\overline{D}_{0}+o_{p}(n^{-\frac{1}{2}})+\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle\,.$ (3)
As a result, the bias of the one-step debiased estimator is primarily driven by the cross-product (or mixed) bias term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$, whose absolute value is bounded above by $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|$ in view of the Cauchy-Schwarz inequality. Most of the existing literature aims to show that $\tau_{n}$ is asymptotically linear with influence function $\overline{D}_{0}$. When $\tau_{n}$ satisfies the above, this is only possible if $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle=o_{p}(n^{-\frac{1}{2}})$. To ensure that this rate condition holds, the product rate condition $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{2}})$ is often assumed, which typically requires consistent estimation of both nuisance functions at sufficiently fast rates (Robins et al., 1994, 1995; van der Laan and Robins, 2003; Bang and Robins, 2005; Chernozhukov et al., 2018a). In particular, this condition holds if both $\|\alpha_{n}-\alpha_{0}\|=o_{p}(n^{-\frac{1}{4}})$ and $\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{4}})$.
As a consequence of (2), the one-step debiased estimator $\tau_{n}$ is said to be rate doubly robust (Rotnitzky et al., 2021) in the nuisance estimation errors in that asymptotic linearity and efficiency can be achieved if $\|\alpha_{n}-\alpha_{0}\|=o_{p}(1)$, $\|\mu_{n}-\mu_{0}\|=o_{p}(1)$ and $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{2}})$. Moreover, $\tau_{n}$ is doubly robust in terms of consistency: it converges in probability to $\tau_{0}$ as long as $\|\alpha_{n}-\alpha_{0}\|=o_{p}(1)$ or $\|\mu_{n}-\mu_{0}\|=o_{p}(1)$. While it maintains consistency under inconsistent estimation of either the outcome regression $\mu_{0}$ or Riesz representer $\alpha_{0}$, the one-step debiased estimator typically lacks desirable properties necessary for inference, such as $n^{\frac{1}{2}}$–consistency and asymptotic normality, in such scenarios. This is also the case when both nuisance estimators are estimated consistently but at insufficiently fast rates so that the condition $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{2}})$ does not hold. Consequently, despite rate double robustness, $\tau_{n}$ is typically only asymptotically linear when $\mu_{0}$ and $\alpha_{0}$ are sufficiently structured (e.g., smooth or sparse) to be estimated at fast enough rates, and this represents a stringent condition that may fail to hold in practice.
3 Preliminaries on doubly robust inference
3.1 Statistical objective: doubly robust asymptotic linearity
In this work, we seek to construct estimators of $\tau_{0}$ that are doubly robust asymptotically linear (DRAL), in the sense that they are asymptotically linear if any of the following hold: $\|\alpha_{n}-\alpha_{0}\|^{2}=o_{p}(n^{-\frac{1}{2}})$, $\|\mu_{n}-\mu_{0}\|^{2}=o_{p}(n^{-\frac{1}{2}})$ or $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{2}})$. The DRAL property substantially improves upon rate double robustness as it only requires that $\|\alpha_{n}-\alpha_{0}\|\wedge\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{4}})$ and therefore allows one of the nuisance estimators to converge arbitrarily slowly to its target or even be inconsistent. To illustrate this improvement, suppose that $\mu_{0}$ and $\alpha_{0}$ are known only to be Hölder smooth with respective exponents $\beta>0$ and $\gamma>0$. In such cases, achieving asymptotic linearity under rate double robustness generally requires that $\beta+\gamma>d$, where $d$ is the dimension of the covariate vector $W$ (Robins et al., 2008). Thus, under rate double robustness, if one of the nuisance functions is insufficiently smooth so that the minimum nuisance smoothness satisfies $\beta\wedge\gamma\ll d/2$, achieving asymptotic linearity typically necessitates the other nuisance function to exhibit a high degree of smoothness, which may not be satisfied by the data-generating mechanism in many applications. In contrast, achieving asymptotic linearity under the DRAL property only requires that the maximum nuisance smoothness satisfy $\beta\vee\gamma>d/2$. As a result, DRAL estimators can maintain asymptotic linearity even when one of the nuisance functions is arbitrarily non-smooth, as long as the other nuisance function has smoothness exponent larger than $d/2$.
As illustrated in the previous section, the lack of normality of the one-step debiased estimator under inconsistent or insufficiently fast estimation of one of the nuisance functions occurs because in such case the cross-product remainder term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ becomes sensitive to the estimation rate of the consistent nuisance estimator. This sensitivity leads to the cross-product remainder term contributing to the asymptotic bias and variance of $n^{\frac{1}{2}}(\tau_{n}-\tau_{0})$. Therefore, to construct a DRAL estimator, it is critical to ensure that the bias from the cross-product term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ in (2) is corrected for, making it either asymptotically linear or negligible, even if one of the nuisance estimators converges to its target too slowly or inconsistently, in the sense that $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|\neq o_{p}(n^{-\frac{1}{2}})$. In the remainder of this section, we examine this bias term in greater detail, illustrating how additional debiasing efforts can render it asymptotically linear in such cases.
3.2 General strategy: debiasing to linearize the cross-product remainder term
In this section, we illustrate how additional debiasing of the nuisance estimators $\alpha_{n}$ and $\mu_{n}$ linearizes the cross-product remainder term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$. Our key idea, rooted in van der Laan (2014), is to express the cross-product remainder term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ solely in terms of the most favorable nuisance estimator and the corresponding nuisance function. Here, we use most favorable to refer to the nuisance estimator that estimates its intended target at the fastest rate. We can then show that if the most favorable nuisance estimator satisfies certain empirical orthogonality (or score) equations, the cross-product remainder term is debiased and asymptotically linear under relatively weak conditions. This strategy was originally used in van der Laan (2014) and Benkeser et al. (2017) to construct DRAL estimators of the ATE for the special case where one of the nuisance functions is estimated inconsistently. However, the important regimes in which both nuisance functions are consistently estimated but one is estimated too slowly were not studied. In this section, in addition to studying a broad class of statistical parameters, we establish the asymptotic linearity of the remainder in general, and provide a refined analysis to establish the asymptotic negligibility of the remainder in such regimes.
To provide intuition and elucidate key ideas, we first present an informal analysis of the cross-product remainder term for the case where $\mu_{n}$ is most favorable, which will be useful when the outcome regression is consistently estimated but the Riesz representer is estimated inconsistently or at a slower rate. We first introduce the following notation. For any vector-valued function $v:\mathcal{A}\times\mathcal{W}\rightarrow\mathbb{R}^{k}$ with $k\in\mathbb{N}$, we define the conditional expectation projection operator $\Pi_{v}:\mathcal{H}\rightarrow\mathcal{H}$ as the mapping $f\mapsto\Pi_{v}f:=\operatorname*{argmin}_{\theta\in\Theta_{v}}\|f-\theta\|$, where $\Theta_{v}$ is the set of functions of the form $h\circ v$ for any $h:\mathbb{R}\rightarrow\mathbb{R}$. Whenever $v$ and $f$ are nonrandom functions, we have that $\Pi_{v}f:(a,w)\mapsto E_{0}[f(A,W)\,|\,v(A,W)=v(a,w)]$. For brevity, we denote the orthogonal projections of the error terms $\alpha_{0}-\alpha_{n}$ and $\mu_{0}-\mu_{n}$ onto $\mu_{n}$ and $\alpha_{n}$, respectively, as $s_{n,0}:=\Pi_{\mu_{n}}(\alpha_{0}-\alpha_{n})$ and $r_{n,0}:=\Pi_{\alpha_{n}}(\mu_{0}-\mu_{n})$. Following the strategy of van der Laan (2014), when $\mu_{n}$ is consistent for $\mu_{0}$, and $\alpha_{n}$ is inconsistent or converges to $\alpha_{0}$ slower than $\mu_{n}$ tends to $\mu_{0}$, we aim to express the cross-product remainder $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ solely in terms of $\mu_{n}$ and $\mu_{0}$ by applying the law of total expectation to condition on the random vector $(\mu_{n}(A,W),\mu_{0}(A,W))$ within the $P_{0}$–expectation. Doing so, we find that the cross-product remainder term can be written as
$\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ $=\ \langle\Pi_{(\mu_{n},\mu_{0})}(\alpha_{0}-\alpha_{n}),\mu_{n}-\mu_{0}\rangle$
[…]
where, informally, the higher-order terms arise because $\Pi_{(\mu_{n},\mu_{0})}\approx\Pi_{\mu_{n}}$ whenever $\mu_{n}\approx\mu_{0}$ under regularity conditions. The key bias term precluding asymptotic linearity of the one-step debiased estimator under inconsistent or insufficiently fast estimation of $\alpha_{0}$ is driven by the population moment expression $\langle s_{n,0},\mu_{n}-\mu_{0}\rangle$ for the outcome nuisance estimator $\mu_{n}$. To analyze this term further, we note that
[…]
$=\ (P_{n}-P_{0})A_{n,0}-P_{n}A_{n,0}\,,$
where $A_{n,0}$ is defined as the map $z\mapsto s_{n,0}(a,w)\{y-\mu_{n}(a,w)\}$. The first term is a so-called drift term that is asymptotically linear under conditions, satisfying $(P_{n}-P_{0})A_{n,0}=P_{n}A_{0}+o_{p}(n^{-\frac{1}{2}})$ with $A_{0}:z\mapsto s_{0}(a,w)\{y-\mu_{0}(a,w)\}$ and $s_{0}:=\Pi_{\mu_{0}}(\alpha_{0}-\overline{\alpha}_{0})$. The second term is typically a first-order bias term that precludes the asymptotic linearity of the cross-product remainder in this case.
As a consequence, to debias the cross-product remainder with respect to the estimation error of $\alpha_{n}$, it suffices that the residuals $Y_{1}-\mu_{n}(A_{1},W_{1}),\ldots,Y_{n}-\mu_{n}(A_{n},W_{n})$ for $\mu_{0}$ be orthogonalized in $L^{2}(P_{n})$ to the projected residuals $s_{n,0}(A_{1},W_{1}),\ldots,s_{n,0}(A_{n},W_{n})$ for $\alpha_{0}$, in the sense that
$P_{n}A_{n,0}=\frac{1}{n}\sum_{i=1}^{n}s_{n,0}(A_{i},W_{i})\{Y_{i}-\mu_{n}(A_{i},W_{i})\}=0\,.$ (4)
A similar analysis can be applied to the cross-product remainder term for the case where the Riesz representer estimator $\alpha_{n}$ is most favorable by instead conditioning on $(\alpha_{n},\alpha_{0})$. In particular, when $\alpha_{n}$ is consistent for $\alpha_{0}$ but $\mu_{n}$ is either inconsistent or converges to $\mu_{0}$ slower than $\alpha_{n}$ tends to $\alpha_{0}$, the estimator $\alpha_{n}$ must satisfy the empirical orthogonality condition
$\psi_{n}(r_{n,0})-\frac{1}{n}\sum_{i=1}^{n}r_{n,0}(A_{i},W_{i})\alpha_{n}(A_{i},W_{i})=0$ (5)
for the cross-product remainder term to be debiased with respect to the estimation error of $\mu_{n}$. To understand why this also represents an orthogonality condition, we note that (5) can be rewritten as $\frac{1}{n}\sum_{i=1}^{n}r_{n,0}(A_{i},W_{i})\{V_{n,i}-\alpha_{n}(A_{i},W_{i})\}=0,$ where $V_{n,1},\ldots,V_{n,n}$ together form the empirical Riesz representer of $\psi_{n}$, defined implicitly by the equation $\psi_{n}(h)=\frac{1}{n}\sum_{i=1}^{n}h(A_{i},W_{i})V_{n,i}$ for each $h\in\mathcal{H}$.
We now formalize the heuristic arguments above and establish that properties (4) and (5) indeed suffice to debias the cross-product remainder term with respect to the error in the Riesz representer and outcome regression estimators, respectively. In the next section, we provide a simple method for post-hoc construction of nuisance estimators that satisfy properties (4) and (5) exactly. For the moment, we take for granted that nuisance estimators $\mu_{n}$ and $\alpha_{n}$ satisfying these properties can be constructed. We define $B_{0}:=\psi_{0}(r_{0})+\widetilde{\phi}_{r_{0}}-r_{0}\alpha_{0}$ with $r_{0}:=\Pi_{\alpha_{0}}(\mu_{0}-\overline{\mu}_{0})$, and $B_{n,0}:=\psi_{0}(r_{n})+\widetilde{\phi}_{r_{n}}-r_{n}\alpha_{n}$, and recall the definitions $A_{0}:z\mapsto s_{0}(a,w)\{y-\mu_{0}(a,w)\}$ and $A_{n,0}:z\mapsto s_{n,0}(a,w)\{y-\mu_{n}(a,w)\}$. We also denote by $\Delta_{\alpha,n}:=\alpha_{n}-\alpha_{0}$ and $\Delta_{\mu,n}:=\mu_{n}-\mu_{0}$ the nuisance estimation errors. The theorem we now state relies on the following condition to relate the estimation error in the projection operators to the nuisance estimation errors $\|\mu_{n}-\mu_{0}\|$ and $\|\alpha_{n}-\alpha_{0}\|$:
  1. B1)
    Coupling between errors:
    1. i)
      $\|(\Pi_{\mu_{n}}-\Pi_{\mu_{0}})\Delta_{\alpha,n}\|+\|(\Pi_{\mu_{n},\mu_{0}}-\Pi_{\mu_{n}})\Delta_{\alpha,n}\|=O_{p}(\|\mu_{n}-\mu_{0}\|)$ ;
    2. ii)
      $\|(\Pi_{\alpha_{n}}-\Pi_{\alpha_{0}})\Delta_{\mu,n}\|+\|(\Pi_{\alpha_{n},\alpha_{0}}-\Pi_{\alpha_{n}})\Delta_{\mu,n}\|=O_{p}(\|\alpha_{n}-\alpha_{0}\|)$
Theorem 2.
If $\mu_{n}$ satisfies (4) and condition B1B1i holds, the outcome regression-favorable decomposition
$\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ $=P_{n}A_{0}+(P_{n}-P_{0})(A_{n,0}-A_{0})+O_{p}\left(\|\mu_{n}-\mu_{0}\|^{2}\wedge\|\mu_{n}-\mu_{0}\|\|\alpha_{n}-\alpha_{0}\|\right)$
holds. If $\alpha_{n}$ satisfies (5) and condition B1B1ii holds, the Riesz representer-favorable decomposition
$\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ $=P_{n}B_{0}+(P_{n}-P_{0})(B_{n,0}-B_{0})+O_{p}\left(\|\alpha_{n}-\alpha_{0}\|^{2}\wedge\|\mu_{n}-\mu_{0}\|\|\alpha_{n}-\alpha_{0}\|\right)+Rem_{\psi_{n}}(r_{n})$
holds with $Rem_{\psi_{n}}(r_{n}):=\psi_{n}(r_{n})-\psi_{0}(r_{n})-(P_{n}-P_{0})\widetilde{\phi}_{r_{n}}$.
Importantly, the second-order terms in both decompositions above are second-order in the estimation error of their respective favorable nuisance function. As a consequence, when all conditions of the theorem hold, we find that the bias term is second-order in the estimation rate of whichever nuisance function estimator converges fastest to its intended target. When $\|\mu_{n}-\mu_{0}\|=o_{p}(1)$, the outcome-favorable decomposition establishes that asymptotic linearization of the cross-product remainder requires that either $\|\mu_{n}-\mu_{0}\|^{2}=o_{p}(n^{-\frac{1}{2}})$ or $\|\mu_{n}-\mu_{0}\|\|\alpha_{n}-\alpha_{0}\|=o_{p}(n^{-\frac{1}{2}})$. Similarly, when $\|\alpha_{n}-\alpha_{0}\|=o_{p}(1)$, the Riesz representer-favorable decomposition establishes that asymptotic linearization of the cross-product remainder requires that either $\|\alpha_{n}-\alpha_{0}\|^{2}=o_{p}(n^{-\frac{1}{2}})$ or $\|\mu_{n}-\mu_{0}\|\|\alpha_{n}-\alpha_{0}\|=o_{p}(n^{-\frac{1}{2}})$. Notably, if both nuisance functions are consistently estimated, that is, $\overline{\alpha}_{0}=\alpha_{0}$ and $\overline{\mu}_{0}=\mu_{0}$, then $A_{0}=B_{0}=0$ and the cross-product remainder can be expressed purely in terms of second-order remainders that are $o_{p}(n^{-\frac{1}{2}})$ under usual rate conditions. As a consequence, provided that $\mu_{n}$ and $\alpha_{n}$ simultaneously satisfy (4) and (5) and that the remainders are $o_{p}(n^{-\frac{1}{2}})$, Theorem 2 establishes that the cross-product remainder $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ is a DRAL estimator of 0 with a possibly degenerate influence function $1(\overline{\alpha}_{0}\neq\alpha_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})B_{0}$. Furthermore, if the estimation error for $\tau_{n}$ satisfies the asymptotic expansion in (3), then the one-step debiased estimator $\tau_{n}$ of $\tau_{0}$ with suitably orthogonalized nuisance estimators $\mu_{n}$ and $\alpha_{n}$ will be DRAL with influence function $\overline{D}_{0}+1(\overline{\mu}_{0}\neq\mu_{0})A_{0}+1(\overline{\alpha}_{0}\neq\alpha_{0})B_{0}$.
To formally establish that the cross-product remainder is DRAL, we need to show that the remainder terms in the outcome regression- and Riesz representer-favorable decompositions are asymptotically negligible, that is, of rate $o_{p}(n^{-\frac{1}{2}})$, even under inconsistent or insufficiently fast estimation of $\mu_{0}$ or $\alpha_{0}$. Asymptotic negligibility holds for the empirical process remainders $(P_{n}-P_{0})(A_{n,0}-A_{0})$ and $(P_{n}-P_{0})(B_{n,0}-B_{0})$, as long as $\|A_{n,0}-A_{0}\|=o_{p}(1)$, $\|B_{n,0}-B_{0}\|=o_{p}(1)$, and the functions $A_{n,0}-A_{0}$ and $B_{n,0}-B_{0}$ fall, for example, in a Donsker function class with probability tending to one. Consistency of $A_{n,0}$ and $B_{n,0}$ only requires consistent estimation of the associated favorable nuisance functions, as under B1 and boundedness of the nuisance estimators, we have that $\|A_{n,0}-A_{0}\|=O_{p}(\|\mu_{n}-\mu_{0}\|)$ and $\|B_{n,0}-B_{0}\|=O_{p}(\|\alpha_{n}-\alpha_{0}\|)$.
High-level conditions similar to B1 have appeared in prior work on doubly robust inference, such as van der Laan (2014), Benkeser et al. (2017), and Dukes et al. (2021). To illustrate this condition heuristically, we may expect the projection $\Pi_{\mu_{n},\mu_{0}}\Delta_{\alpha,n}$ to be approximated well by $\Pi_{\mu_{n}}\Delta_{\alpha,n}$ as long as $\mu_{n}$ converges to $\mu_{0}$ in an appropriate sense. This is because conditioning on the ‘information’ contained in $\{\mu_{n}(A,W),\mu_{0}(A,W),Z_{1},\ldots,Z_{n}\}$ should be similar to conditioning on the ‘information’ contained in $\{\mu_{n}(A,W),Z_{1},\ldots,Z_{n}\}$. The following lemma establishes that Lipschitz continuity of two bivariate functions is sufficient for B1 to hold, a novel result that is relevant also to prior works requiring a similar condition:
  1. B1enumi
    Lipschitz continuity: Denoting $\mathcal{D}_{n}:=\{Z_{1},\ldots,Z_{n}\}$, it holds that, for all $n$ large enough:
    1. i)
      $(\widehat{m},m)\mapsto E_{0}[\Delta_{\alpha,n}(A,W)\,|\,\mu_{n}(A,W)=\widehat{m},\mu_{0}(A,W)=m,\mathcal{D}_{n}]$ is $P_{0}^{n}$–almost surely Lipschitz continuous with fixed constant $L<\infty$;
    2. ii)
      $(\widehat{a},a)\mapsto E_{0}[\Delta_{\mu,n}(A,W)\,|\,\alpha_{n}(A,W)=\widehat{a},\alpha_{0}(A,W)=a,\mathcal{D}_{n}]$ is $P_{0}^{n}$–almost surely Lipschitz continuous with fixed constant $L<\infty$.
Lemma 1.
Conditions B1B1i and B1B1ii are implied, respectively, by conditions B1enumiB1enumii and B1enumiB1enumiii.
Conditions analogous to B1enumiB1enumii and B1enumiB1enumiii were also imposed in Bonvini et al. (2024) to construct a variant of the DRAL estimator of van der Laan (2014) for the ATE parameter, and in Wang et al. (2023) to analyze regressions involving estimated features. Condition B1enumi implies that a small change in the value of one nuisance function at a point should, on average, lead to a small change in the estimation error for the other nuisance function at that point. Bonvini et al. (2024) verified this condition in a toy example in which $\mu_{n}(A,W)$, $\mu_{0}(A,W)$, $\alpha_{n}(A,W)$ and $\alpha_{0}(A,W)$ jointly follow a multivariate normal distribution conditional on $\mathcal{D}_{n}$. Lipschitz continuity is a fairly mild smoothness condition, and so, we hypothesize that these conditions or the higher-level conditions of B1 are plausibly satisfied by real data-generating processes. Our experimental results support this hypothesis empirically.
3.3 Debiasing via targeted minimum loss-based estimation
For the special case of doubly robust inference on the average treatment effect, van der Laan (2014) and Benkeser et al. (2017) propose a debiasing procedure based on targeted minimum loss estimation (TMLE) to refine given nuisance estimators to asymptotically satisfy (4) and (5). The idea behind the TMLE-based approach is to iteratively debias the initial nuisance estimators $\mu_{n}$ and $\alpha_{n}$ using a reweighted generalized linear regression algorithm with offset to obtain targeted estimators $\mu_{n}^{*}$ and $\alpha_{n}^{*}$ that satisfy the empirical score equations
$\frac{1}{n}\sum_{i=1}^{n}s_{n}^{*}(A_{i},W_{i})\{Y_{i}-\mu_{n}^{*}(A_{i},W_{i})\}\,=\,o(n^{-\frac{1}{2}})\,;$ (6)
$\frac{1}{n}\sum_{i=1}^{n}r_{n}^{*}(A_{i},W_{i})\alpha_{n}^{*}(A_{i},W_{i})-\psi_{n}(r_{n}^{*})\,=\,o(n^{-\frac{1}{2}})\,.$ (7)
Here, $s_{n}^{*}$ and $r_{n}^{*}$ are estimators of the (data-dependent) nuisance functions $s_{n,0}^{*}:=\Pi_{\mu_{n}^{*}}(\alpha_{0}-\alpha_{n}^{*})$ and $r_{n,0}^{*}:=\Pi_{\alpha_{n}^{*}}(\mu_{0}-\mu_{n}^{*})$, which can be obtained, for example, using univariate empirical risk minimization with features $\mu_{n}^{*}(A_{1},W_{1}),\ldots,\mu_{n}^{*}(A_{n},W_{n})$ and $\alpha_{n}^{*}(A_{1},W_{1}),\ldots,\alpha_{n}^{*}(A_{n},W_{n})$, respectively — we provide details in Appendix B.1. Under certain regularity conditions, it can be shown that (6) and (7) imply that (4) and (5) are satisfied asymptotically, so that the one-step debiased estimator based on $\mu_{n}^{*}$ and $\alpha_{n}^{*}$ is indeed DRAL.
As discussed in the Introduction, the TMLE-based algorithm for doubly robust inference involves estimating two univariate nuisance functions at each iteration, with the final iteration involving estimation of nuisances $r_{n,0}^{*}$ and $s_{n,0}^{*}$; the process can be computationally quite prohibitive. In addition, the final estimation of $r_{n,0}^{*}$ and $s_{n,0}^{*}$ possibly induces additional (unnecessary) finite-sample biases on the one-step debiased estimator $\tau_{n}$ by only approximately solving the empirical moment equations (4) and (5). We also note that neither van der Laan (2014) nor Benkeser et al. (2017) formally demonstrate that their iterative refinement procedures maintain the convergence properties of the original nuisance estimators in pursuit of satisfying the empirical orthogonality equations. Such a formal analysis is important to ensure that, under consistent estimation of both nuisance functions, their refinement procedure does not harm the quality of estimators and inferences based on the one-step debiased estimator. In the following section, we propose a novel procedure for constructing one-step debiased estimators that is free from these issues and, unlike prior works, simultaneously tackles a whole class of target parameters.
4 Calibrated debiased machine learning
4.1 Link between calibration and doubly robust inference
[…]
To formalize this claim, let $(z,\nu)\mapsto\ell(z,\nu)$ be a valid loss function for some nuisance function $\nu_{0}\in\mathcal{H}$. We say that an estimator $\nu_{n}$ is $\ell$–empirically calibrated for $\nu_{0}$ if it satisfies the self-consistency property
$\frac{1}{n}\sum_{i=1}^{n}\ell(Z_{i},\nu_{n})=\min_{\theta}\frac{1}{n}\sum_{i=1}^{n}\ell(Z_{i},\theta\circ\nu_{n})\,,$ (8)
where the minimization is over the class of all real-valued functions on $\mathbb{R}$. In other words, an estimator $\nu_{n}$ is $\ell$–empirically calibrated for $\nu_{0}$ if its empirical risk cannot be improved through transformation by any one-dimensional mapping $\theta:\mathbb{R}\to\mathbb{R}$. This notion of $\ell$–empirical calibration generalizes existing notions of calibration, including calibration for regression and quantile functions (Roth, 2022), which respectively correspond to the least-squares and hinge loss functions (Deng et al., 2023). A related notion of calibration was also considered in Whitehouse et al. (2024) in the context of calibration of orthogonal machine learning estimators. When the loss $\ell$ is sufficiently smooth, the first-order equations characterizing the empirical risk minimization problem imply that $\ell$–empirical calibration is equivalent to requiring that, for each map $\theta:\mathbb{R}\rightarrow\mathbb{R}$,
[…]
The following theorem establishes that the empirical calibration of nuisance estimators is sufficient for the orthogonality conditions (4) and (5) to be exactly met, and is a direct consequence of (9) for suitable loss functions.
  1. C1)
    (linearity and boundedness of estimated functional) it holds $P_{0}$–almost surely that $\psi_{n}(c\mu_{1}+\mu_{2})=c\psi_{n}(\mu_{1})+\psi_{n}(\mu_{2})$ for each $\mu_{1},\mu_{2}\in\mathcal{H}$ and $c\in\mathbb{R}$ and that $\sup_{\mu\in\mathcal{H}}|\psi_{n}(\mu)|/\|\mu\|<\infty$;
Theorem 3.
Suppose that Condition C1 holds. Suppose also that $\mu_{n}$ is empirically calibrated with respect to the least-squares loss $(z,\mu)\mapsto\{y-\mu(a,w)\}^{2}$ and that $\alpha_{n}$ is empirically calibrated with respect to the Riesz loss $(z,\alpha)\mapsto\alpha^{2}(a,w)-2\psi_{n}(\alpha)$, so that the calibration conditions
[…]
hold for any $\theta:\mathbb{R}\rightarrow\mathbb{R}$. Then, $\mu_{n}$ and $\alpha_{n}$ satisfy the empirical orthogonality conditions (4) and (5).
Theorem 3 demonstrates that empirical calibration suffices to simultaneously debias the cross-product remainder of the debiased machine learning estimator $\tau_{n}$ with respect to the nuisance estimation errors for $\mu_{0}$ and $\alpha_{0}$, so that the error decompositions in Theorem 2 can then be applied to study the behavior of $\tau_{n}$.
The DRAL properties that arise from the calibration conditions in Theorem 3 relate to the fact that adjustment for either $\mu_{0}(A,W)$ or $\alpha_{0}(A,W)$ is sufficient for confounding bias adjustment. Notably, when $\psi_{0}$ is a counterfactual mean, adjusting for $\alpha_{0}(A,W)$ is equivalent to adjusting for the propensity score $\pi_{0}(A,W)$, a well-known balancing score (Imai and Ratkovic, 2014). When $\mu_{n}$ is consistent at a sufficiently fast rate, the first calibration condition implies that the plug-in estimator $\psi_{n}(\mu_{n})$ is ‘debiased’ for $\psi_{0}$ under the model that utilizes $\mu_{0}(A,W)$ as a dimension reduction of $(A,W)$ (Benkeser et al., 2020). Similarly, when $\alpha_{n}$ is consistent, the second calibration condition implies that the balancing weighted outcome estimator $\frac{1}{n}\sum_{i=1}^{n}\alpha_{n}(A_{i},W_{i})Y_{i}$ is ‘debiased’ for $\psi_{0}$ under the model that employs $\alpha_{0}(A,W)$ as a dimension reduction of $(A,W)$. The calibrated debiased machine learning estimator inherits the debiasedness of both the calibrated plug-in and weighted outcome estimators. This property arises because it can leverage either $\alpha_{0}(A,W)$ or $\mu_{0}(A,W)$ to adjust for confounding bias, as long as at least one of the associated nuisance estimators is consistent. We formalize this intuition later in Section 5.3.
4.2 Isotonic-calibrated debiased machine learning
We refer to the implementation of debiased machine learning with calibrated nuisance functions as C-DML. Theorem 3 establishes that calibration of nuisance functions bestows DML estimators with the DRAL property. We now describe a particular implementation of our general framework for C-DML, which outlines how to construct DRAL estimators using isotonic calibration and one-step debiased estimation.
Our post-hoc procedure employs isotonic regression (Barlow and Brunk, 1972) to empirically calibrate any given set of nuisance estimators, $\mu_{n}$ and $\alpha_{n}$, such that they satisfy the empirical orthogonality equations (4) and (5). In view of Theorem 3, this calibration ensures that the cross-product remainder term is asymptotically linear even when one nuisance function is poorly estimated, leading the resulting debiased machine learning estimator to be DRAL. Our procedure is automatic, requiring no knowledge of the specific form of the linear functional $\psi_{0}$; it merely necessitates the ability to evaluate as a black box an asymptotically linear estimator $\psi_{n}(\mu)$ of $\psi_{0}(\mu)$ for any given fixed $\mu$. Furthermore, our method is non-iterative and can be efficiently implemented as a two-step procedure using widely available software for isotonic regression.
To this end, let $\mathcal{F}_{\text{iso}}$ denote the space of real-valued monotone non-decreasing functions defined over $\mathbb{R}$. Given initial estimators $\mu_{n}$ and $\alpha_{n}$ for $\mu_{0}$ and $\alpha_{0}$, respectively, we define the empirically calibrated versions of the outcome regression estimator as $\mu_{n}^{*}:=f_{n}\circ\mu_{n}$ and of the Riesz representer estimator as $\alpha_{n}^{*}:=g_{n}\circ\alpha_{n}$, where $f_{n}$ and $g_{n}$ are elements of $\mathcal{F}_{\text{iso}}$ obtained through a generalization of isotonic regression as
$f_{n}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}Y_{i}-f(\mu_{n}(A_{i},W_{i}))\big{\}}^{2};$ (10)
$g_{n}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}g(\alpha_{n}(A_{i},W_{i}))\big{\}}^{2}-2\psi_{n}(g\circ\alpha_{n})\,.$
Our proposed isotonic-calibrated debiased machine learning (IC-DML) estimator is then simply
$\tau_{n}^{*}:=\psi_{n}(\mu_{n}^{*})+\frac{1}{n}\sum_{i=1}^{n}\alpha_{n}^{*}(A_{i},W_{i})\{Y_{i}-\mu_{n}^{*}(A_{i},W_{i})\}\,,$ (11)
the debiased machine learning estimator based on these calibrated nuisance function estimators. Notably, the properties of the isotonic regression problem ensure that the isotonic regression solutions $f_{n}$ and $g_{n}$ are monotone non-decreasing and act as histogram estimators over a data-dependent binning of their input space. This property can be used to show that the isotonic calibrated nuisance estimators $\mu_{n}^{*}$ and $\alpha_{n}^{*}$ are $\ell$–empirically calibrated, as defined in (8), with respect to their respective loss functions and thereby satisfy the conditions of Lemma 3. In a prediction setting, this technique was used in van der Laan et al. (2023b) and van der Laan and Alaa (2024) to derive distribution-free calibration guarantees for isotonic calibration. Importantly, the monotonicity constraint in isotonic calibration provides intrinsic regularization, ensuring that the quality of the initial nuisance estimators is not compromised in the pursuit of empirical calibration. In light of the infinite-dimensional score equations they solve, the calibrated nuisance estimators $\alpha_{n}^{*}$ and $\mu_{n}^{*}$ can be viewed as infinite-dimensional targeted minimum loss-based estimators of $\alpha_{0}$ and $\mu_{0}$ — see, e.g., van der Laan and Gruber (2016), Luedtke et al. (2017), Qiu et al. (2021) and van der Laan et al. (2024a).
In the context of causal inference using debiased machine learning, cross-fitting techniques for nuisance estimation are often used to mitigate overfitting and relax constraints on nuisance estimator complexity (van der Laan et al., 2011; Chernozhukov et al., 2018a). To obtain strong theoretical guarantees under weak conditions, we also recommend cross-fitting the initial nuisance estimators. We note, however, that the isotonic calibration step of (10) should not be cross-fitted. Algorithm 1 outlines the procedure for the cross-fitted variant of the calibrated nuisance constructions. Specifically, it involves exhaustively partitioning the available data into $J$ disjoint folds, where for each fold $j$ initial estimators $\alpha_{n,j}$ and $\mu_{n,j}$ are computed using only the data in the complementary folds. The resulting estimators are then combined to obtain out-of-fold nuisance function estimators for each observation. The pooled out-of-fold estimators are then used to learn the calibrators $f_{n},g_{n}\in\mathcal{F}_{\text{iso}}$ in a similar manner as in (10). Finally, these calibrators are composed with the initial fold-specific nuisance estimators to construct a calibrated one-step debiased estimator of $\tau_{0}$.
0:  dataset $\mathcal{D}_{n}=\{O_{i}:i=1,\ldots,n\}$, number $J$ of cross-fitting splits
1:  partition $\mathcal{D}_{n}$ into datasets $\mathcal{C}^{(1)},\mathcal{C}^{(2)},\ldots,\mathcal{C}^{(J)}$;
2:  for $s=1,\ldots,J$ do
3:      get initial estimators $\alpha_{n,s}$ of $\alpha_{0}$ and $\mu_{n,s}$ of $\mu_{0}$ from $\mathcal{E}^{(s)}:=\mathcal{D}_{n}\backslash\mathcal{C}^{(s)}$;
4:      get initial estimators $\mu\mapsto\psi_{n,s}(\mu)$ of $\mu\mapsto\psi_{0}(\mu)$ from $\mathcal{D}_{n}$, where sample-splitting is performed as needed to satisfy Condition C7;
5:  end for
6:  for $s=1,\ldots,J$, set $j(i):=s$ for each $i\in\mathcal{C}^{(s)}$;
7:  compute calibrators $f_{n}$ and $g_{n}$ using pooled out-of-fold estimates as
$f_{n}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}Y_{i}-f(\mu_{n,j(i)}(A_{i},W_{i}))\big{\}}^{2}$
$g_{n}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}g(\alpha_{n,j(i)}(A_{i},W_{i}))^{2}-2\psi_{n,j(i)}(g\circ\alpha_{n,j(i)})\big{\}};$
8:  for $s=1,\ldots,J$, set $\mu_{n,s}^{*}:=f_{n}\circ\mu_{n,s}$ and $\alpha_{n,s}^{*}:=g_{n}\circ\alpha_{n,s}$;
9:  compute calibrated one-step debiased estimator
$\tau_{n}^{*}:=\frac{1}{J}\sum_{j=1}^{J}\left[\psi_{n,j}(\mu_{n,j}^{*})+\frac{1}{n}\sum_{i=1}^{n}\alpha_{n,j(i)}^{*}(A_{i},W_{i})\left\{Y_{i}-\mu_{n,j(i)}^{*}(A_{i},W_{i})\right\}\right].$ (12)
10:  return  $\tau_{n}^{*}$
Algorithm 1 IC-DML: calibrated debiased machine learning using isotonic calibration
The solution to an isotonic regression problem is typically nonunique. As in Groeneboom and Lopuhaa (1993), we take $f_{n}$ and $g_{n}$ to be the unique càdlàg piecewise constant solution of the isotonic regression problem with jumps occurring only at observed values of $\mu_{n}$ and $\alpha_{n}$. The computation of $g_{n}$ in (10) may correspond to solving an isotonic regression problem for a nonstandard loss function. However, the calibrator $g_{n}$ can equivalently be represented, for the same loss function, as the empirical risk minimizer over all univariate regression trees of unbounded depth subject to the constraint that the trees are monotone nondecreasing. Hence, the isotonic solutions of (10) can be computed using any software that supports generalized regression trees with monotonicity constraints and custom losses, such as the R and Python implementations of xgboost (Chen and Guestrin, 2016). Even so, the following example illustrates that for a wide range of causal parameters of interest, the estimator can be computed using readily available standard implementations of isotonic regression.
Example 1 (continued).
Consider the $G$-computation identification $\tau_{0}(a_{0}):=E_{0}\{\mu_{0}(a_{0},W)\}$ of the mean of the counterfactual outcome $Y(a_{0})$ for $a_{0}\in\mathcal{A}$. In this case, the Riesz representer $\alpha_{0}$, given by $(a,w)\mapsto 1(a=a_{0})/\pi_{0}(a_{0}|w)$, is determined by the propensity score $\pi_{0}$. Given estimators $\mu_{n}$ of $\mu_{0}$ and $\pi_{n}$ of $\pi_{0}$, an IC-DML estimator of $\tau_{0}(a_{0})$ (without cross-fitting) is given by
$\tau_{n}^{*}(a_{0}):=\frac{1}{n}\sum_{i=1}^{n}\left[\mu_{n}^{*}(a_{0},W_{i})+\frac{1(A_{i}=a_{0})}{\pi_{n}^{*}(a_{0}|W_{i})}\{Y_{i}-\mu_{n}^{*}(a_{0},W_{i})\}\right],$
where, for each $w\in\mathcal{W}$, we set $\mu_{n}^{*}(a_{0},w):=f_{n,a_{0}}\circ\mu_{n}(a_{0},w)$ and $\pi_{n}^{*}(a_{0}|w):=g_{n,a_{0}}\circ\pi_{n}(a_{0}|w)$ with
$f_{n,a_{0}}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\textnormal{iso}}}\sum_{i=1}^{n}1(A_{i}=a_{0})\left\{Y_{i}-f(\mu_{n}(a_{0},W_{i}))\right\}^{2};$
$g_{n,a_{0}}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\textnormal{iso}}}\sum_{i=1}^{n}\left[\left\{\frac{1(A_{i}=a_{0})}{g(\pi_{n}(A_{i}|W_{i}))}\right\}^{2}-\frac{2}{g(\pi_{n}(a_{0}|W_{i}))}\right].$
The first-order conditions characterizing the isotonic regression solution $g_{n,a_{0}}$ show that $g_{n,a_{0}}$ can be equivalently obtained directly from propensity score calibration (van der Laan et al., 2024b) as
$g_{n,a_{0}}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\textnormal{iso}}}\sum_{i=1}^{n}\left\{1(A_{i}=a_{0})-g(\pi_{n}(a_{0}\mid W_{i}))\right\}^{2}.$
For parameters that can be expressed as smooth transformations of counterfactual means, we can use the above to construct DRAL estimators of $\{\tau_{0}(a_{0}):a_{0}\in\mathcal{A}\}$, say $\{\tau_{n}^{*}(a_{0}):a_{0}\in\mathcal{A}\}$, and then compute the corresponding plug-in estimator of the estimand of interest, which will itself be DRAL. Inference can be obtained using either the delta method or the bootstrap, as detailed in the following sections. For example, DRAL estimators of the ATE $\tau_{0}(1)-\tau_{0}(0)$ and relative ATE $\tau_{0}(1)/\tau_{0}(0)$ are given by $\tau_{n}^{*}(1)-\tau_{n}^{*}(0)$ and $\tau_{n}^{*}(1)/\tau_{n}^{*}(0)$, respectively. Python and R implementations of the calibration procedure are provided in Appendix C.
Our approach shares some similarities with the automatic debiased machine learning estimators of Chernozhukov et al. (2018b, 2021, 2022), where similar tailored loss functions were used to construct cross-fitted estimators of the Riesz representer and outcome regression. However, our framework differs in its objective and use of these loss functions. Specifically, we make use of these loss functions to calibrate user-supplied cross-fitted nuisance estimators to ensure that the resulting one-step debiased estimator is DRAL. In contrast, Chernozhukov et al. (2018b) uses these loss functions to obtain nuisance estimators for the standard one-step debiased estimator. Notably, C-DML can be seamlessly integrated within the automatic DML framework by calibrating the cross-fitted nuisance estimators using isotonic regression.
4.3 Automated bootstrap-assisted confidence intervals
In previous work, doubly robust confidence intervals were obtained through a standard Wald construction based on the asymptotic normality of the corrected estimator using an adjusted estimator of the asymptotic variance. This estimator is complex because it involves the estimation of additional, complicated nuisance functions. Here, we propose an automated approach for the construction of confidence intervals associated with the estimator odescribed in Algorithm 1 using a variant of the bootstrap (Efron and Tibshirani, 1994). Our proposed bootstrap-assisted inference approach, which we refer to as boot-IC-DML, is straightforward to describe, computationally efficient, and valid if either of the nuisance estimators is consistent at a sufficiently fast rate. Importantly, boot-IC-DML can be fully automated since, unlike existing approaches for doubly robust inference, it does not require the estimation of additional nuisance functions beyond those already used to compute the IC-DML estimator. Additionally, boot-IC-DML uses the same initial cross-fitted nuisance estimators, namely $(\alpha_{n,1},\mu_{n,1}),\ldots,(\alpha_{n,J},\mu_{n,J})$, across all bootstrap samples. As a result, the initial nuisance estimators only need to be computed once using the original data, greatly saving computational time. Boot-IC-DML only requires recomputation of the isotonic regression problems outlined in (10) for each bootstrap sample, which can be efficiently solved in linear time using standard software.
For simplicity, we focus the description of our automated algorithm on the case where the linear functional $\psi_{0}$ is of the form $\mu\mapsto E_{0}\{m(A,W,\mu)\}$ for a known map $m:\mathcal{A}\times\mathcal{W}\times\mathcal{H}\rightarrow\mathbb{R}$ although a similar algorithm can be devised for the general case by appropriately bootstrapping the estimation of the functional $\psi_{n}$. In such cases, we can consider the empirical plug-in estimator $\psi_{n}$ of $\psi_{0}$ defined as $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(A_{i},W_{i},\mu)$. Because this setting is simpler, we are able to prove the validity of the proposed bootstrap scheme under weak conditions. As in the previous subsection, we first describe boot-IC-DML in the simpler case where the initial estimators are not cross-fitted. The general procedure allowing for cross-fitting is provided in Algorithm 2. To this end, let $\alpha_{n}$ and $\mu_{n}$ be initial estimators $\alpha_{0}$ and $\mu_{0}$ computed once using the original sample. Our approach involves bootstrapping the meta isotonic fitting step of (10) used originally to obtain the calibrators $f_{n}$ and $g_{n}$. Specifically, we generate a bootstrap sample $\{Z_{i}^{\#}:i=1,\ldots,n\}$ from the original sample $\{Z_{i}:i=1,\ldots,n\}$ and compute the bootstrapped calibrators
$f_{n}^{\#}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}Y_{i}^{\#}-f(\mu_{n}(A_{i}^{\#},W_{i}^{\#}))\big{\}}^{2};$
$g_{n}^{\#}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}g(\alpha_{n}(A_{i}^{\#},W_{i}^{\#}))^{2}-2m(A_{i}^{\#},W_{i}^{\#},g\circ\alpha_{n})\big{\}}\,.$
Next, we obtain the bootstrap-calibrated nuisance estimators $\alpha_{n}^{*\#}:=f_{n}^{\#}\circ\alpha_{n}$ and $\mu_{n}^{*\#}:=f_{n}^{\#}\circ\mu_{n}$, and compute the bootstrapped IC-DML estimator $\tau_{n}^{*\#}$ of $\tau_{0}$ from $\mu_{n}^{*\#}$ and $\alpha_{n}^{*\#}$ following the same steps as in (11). This procedure is repeated $K$ times to obtain $K$ bootstrap estimators $\tau_{n,1}^{*\#},\ldots,\tau_{n,K}^{*\#}$ of $\tau_{0}$. In practice, a good rule of thumb is to set $K\approx 10,000$ (Efron and Tibshirani, 1994).
To construct a bootstrap-assisted confidence interval, we first compute the bootstrap average $\overline{\tau}_{n}^{*\#}:=\frac{1}{K}\sum_{k=1}^{K}\tau_{n,k}^{*\#}$. For given confidence level $\rho\in(0,1)$, we may consider the interval
$\text{CI}_{n}(\rho):=\left(\tau_{n}^{*}-q(\rho)\sigma_{n}^{\#},\,\tau_{n}^{*}+q(\rho)\sigma_{n}^{\#}\right),$
where $\sigma_{n}^{\#}$ is the empirical standard deviation of $\tau_{n,1}^{*\#},\ldots,\tau_{n,K}^{*\#}$ and, for each $\gamma\in(0,1)$, $q(\gamma)$ denotes the $\gamma/2$–quantile of the standard normal distribution. Alternatively, we could consider a percentile bootstrap interval $\left(\tau_{n}^{*}-q_{n}(\rho),\,\tau_{n}^{*}-q_{n}(2-\rho)\right)$, where for each $\gamma\in(0,1)$ $q_{n}(\gamma)$ denotes the empirical $\gamma/2$–quantile of the centered bootstrapped estimates $\tau_{n,1}^{*\#}-\overline{\tau}_{n}^{*\#},\ldots,\tau_{n,K}^{*\#}-\overline{\tau}_{n}^{*\#}$; this interval may be more robust to finite sample deviations from normality. We refer to Algorithm 2 for the pseudo-code implementing the general cross-fitted form of boot-C-DML.
Algorithm 2 Bootstrap IC-DML: bootstrap-assisted doubly robust CIs with cross-fitting
0:  cross-fitted estimators $(\alpha_{n,j},\mu_{n,j})$ excluding fold $\mathcal{C}^{(j)}$ for $j=1,\ldots,J$, IC-DML estimator $\tau_{n}^{*}$; confidence level $1-\rho$; bootstrap replication number $K$;
1:  for $k=1,\ldots,K$ do
2:      for $s=1,\ldots,J$, draw bootstrap sample $\mathcal{C}_{k}^{(s)\#}$ from $\mathcal{C}_{k}^{(s)}$;
3:      set $\mathcal{D}_{n,k}^{\#}:=\{Z_{1}^{\#},\dots,Z_{n}^{\#}\}=\{\mathcal{C}_{k}^{(s)\#}:s=1,\ldots,J\}$;
4:      for $s=1,\ldots,J$, set $j(i)=s$ for each $i\in\{1,\ldots,n\}$ such that $Z_{i}^{\#}\in\mathcal{C}_{k}^{(s)\#}$;
5:      compute calibrators $f^{\#(k)}_{n}$ and $g^{\#(k)}_{n}$ using pooled out-of-fold estimates as
$f_{n}^{\#(k)}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}Y_{i}^{\#}-f(\mu_{n,j(i)}(A_{i}^{\#},W_{i}^{\#}))\big{\}}^{2};$
$g_{n}^{\#(k)}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}g(\alpha_{n,j(i)}(A_{i}^{\#},W_{i}^{\#}))^{2}-2m(A_{i}^{\#},W_{i}^{\#},g\circ\alpha_{n,j(i)})\big{\}}\,;$
6:      for $s=1,\ldots,J$, set $\mu_{n,s}^{(k)\#}:=f_{n}^{(k)\#}\circ\mu_{n,s}$ and $\alpha_{n,s}^{(k)\#}:=g_{n}^{(k)\#}\circ\alpha_{n,s}$;
7:      compute fold-specific calibrated one-step debiased estimator
$\tau_{n,k}^{*\#}:=\frac{1}{n}\sum_{i=1}^{n}\left[m(A_{i}^{\#},W_{i}^{\#},\mu_{n,j(i)}^{(k)\#})+\alpha_{n,j(i)}^{(k)\#}(A_{i}^{\#},W_{i}^{\#})\left\{Y_{i}^{\#}-\mu_{n,j(i)}^{(k)\#}(A_{i}^{\#},W_{i}^{\#})\right\}\right];$
8:  end for
9:  set $\sigma_{n}^{\#2}:=\frac{1}{K}\sum_{k=1}^{K}(\tau_{n,k}^{*\#}-\overline{\tau}_{n}^{*\#})^{2}$ with $\overline{\tau}_{n}^{*\#}:=\frac{1}{K}\sum_{k=1}^{K}\tau_{n,k}^{*\#};$
10:  construct interval $\text{CI}_{n}(\rho):=(\tau_{n}^{*}-q(\rho)\sigma_{n}^{\#},\tau_{n}^{*}+q(\rho)\sigma_{n}^{\#});$
11:  return  bootstrap-assisted confidence interval $\text{CI}_{n}(\rho)$ for $\tau_{0}$.
5 Theoretical properties
5.1 Doubly robust asymptotic linearity and inference
We now present our main theorem on the doubly robust asymptotic linearity of the IC-DML estimator. Before doing so, we provide an informal lemma that establishes that the isotonic calibrated nuisance estimators perform as well as the best monotone non-decreasing transformation of the uncalibrated, cross-fitted nuisance estimators, up to a typically negligible error of order $O_{p}(n^{-\frac{1}{3}})$. The formal results and regularity conditions are provided in Section D.5.
Lemma 2.
Suppose that $\psi_{0}$ has the form (1) with mapping $m:\mathcal{A}\times\mathcal{W}\times\mathcal{H}\rightarrow\mathbb{R}$, and that $\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}$ is taken to be the plug-in estimator $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(A_{i},W_{i},\mu)$. Under regularity conditions, it holds that
$\sum_{j=1}^{J}\|\mu_{n,j}^{*}-\mu_{0}\|^{2}\,$ $=\,\min_{f\in\mathcal{F}_{\textnormal{iso}}}\sum_{j=1}^{J}\|f\circ\mu_{n,j}-\mu_{0}\|^{2}+O_{p}(n^{-\frac{2}{3}})\,;$
$\sum_{j=1}^{J}\|\alpha_{n,j}^{*}-\alpha_{0}\|^{2}\,$ $=\,\min_{g\in\mathcal{F}_{\textnormal{iso}}}\sum_{j=1}^{J}\|g\circ\alpha_{n,j}-\alpha_{0}\|^{2}+O_{p}(n^{-\frac{2}{3}})\,.$
Since the identity transform is itself monotone non-decreasing, the above lemma implies that the isotonic calibrated nuisance estimators typically perform no worse than the uncalibrated ones, with potentially improved performance when the uncalibrated estimators are poorly calibrated or inconsistent up to a monotone transformation. As such, in addition to providing doubly robust asymptotic linearity, the use of isotonic calibration weakens nuisance estimator consistency requirements by only requiring that the initial cross-fitted nuisance estimators be consistent up to a monotone transformation.
We now present our main result. The limiting distribution of the scaled and centered version of the cross-fitted IC-DML estimator involves the additional nuisance functions $s_{0}$ and $r_{0}$, define pointwise as
$s_{0}(a,w)$ $:=E_{0}\left\{\alpha_{0}(A,W)-\overline{\alpha}_{0}(A,W)\,|\,\mu_{0}(A,W)=\mu_{0}(a,w)\right\};$
$r_{0}(a,w)$ $:=E_{0}\left\{\mu_{0}(A,W)-\overline{\mu}_{0}(A,W)\,|\,\alpha_{0}(A,W)=\alpha_{0}(a,w)\right\}.$
Denoting by $\Delta_{\alpha,n,j}:=\alpha_{n,j}^{*}-\alpha_{0}$ and $\Delta_{\mu,n,j}:=\mu_{n,j}^{*}-\mu_{0}$ for $j=1,\ldots,J$ the nuisance estimation errors, and defining $r_{n,j}^{*}:=\Pi_{\alpha_{n,j}^{*}}(\mu_{0}-\mu_{n,j}^{*})$, we make use of the following conditions:
  1. C2)
    Limits of nuisance estimators: there exist $\overline{\alpha}_{0},\overline{\mu}_{0}\in\mathcal{H}$ such that $\|\alpha_{n,j}^{*}-\overline{\alpha}_{0}\|+\|\mu_{n,j}^{*}-\overline{\mu}_{0}\|=o_{p}(1)$ with either $\overline{\mu}_{0}=\mu_{0}$ or $\overline{\alpha}_{0}=\alpha_{0}$ for $j=1,\ldots,J$;
  2. C3)
    Boundedness of nuisance functions: for some $M\in(0,\infty)$, the functions $\mu_{n,j}^{*}$, $\alpha_{n,j}^{*}$, $\mu_{0}$, $\alpha_{0}$, $\overline{\mu}_{0}$ and $\overline{\alpha}_{0}$ are uniformly bounded by $M$ with $P_{0}$–probability tending to one for each $j=1,\ldots,J$;
  3. C4)
    Sparsity of calibrators: with $P_{0}$–probability tending to one, the isotonic calibrators $f_{n}$ and $g_{n}$ take at most $k(n)>0$ distinct values for some deterministic sequence $k(n)=o(n^{\frac{1}{2}})$;
  4. C5)
    Lipschitz continuity: the random functions $(\widehat{m},m)\mapsto E_{0}\{\Delta_{\alpha,n,j}\,|\,\mu_{n,j}^{*}(X)=\widehat{m},\mu_{0}(X)=m,\mathcal{D}_{n}\}$ and $(\widehat{a},a)\mapsto E_{0}\{\Delta_{\mu,n,j}\,|\,\alpha_{n,j}^{*}(X)=\widehat{a},\alpha_{0}(X)=a,\mathcal{D}_{n}\}$ are $P_{0}$–almost surely Lipschitz continuous with some fixed constant $L\in(0,\infty)$ for each $j=1,\ldots,J$ with $P_{0}$–probability tending to one;
  5. C6)
    Rate of nuisance estimation: $\|\mu_{n,j}^{*}-\mu_{0}\|\wedge\|\alpha_{n,j}^{*}-\alpha_{0}\|=o_{p}(n^{-\frac{1}{4}})$ for each $j=1,\ldots,J$.
  6. C7)
    Negligibility of remainders: it holds that $\frac{1}{J}\sum_{j=1}^{J}\{\psi_{n,j}(\mu_{n,j}^{*})-\psi_{0}(\mu_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{0,\mu_{n,j}^{*}}\}=o_{p}(n^{-\frac{1}{2}})$ and $\frac{1}{J}\sum_{j=1}^{J}\{\psi_{n,j}(r_{n,j}^{*})-\psi_{0}(r_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{0,r_{n,j}^{*}}\}=o_{p}(n^{-\frac{1}{2}})$.
Theorem 4 (doubly robust asymptotic linearity).
Assume that conditions A1-A3 and C1-C7 hold. Then, the cross-fitted IC-DML estimator $\tau_{n}^{*}$ of $\tau_{0}$ is DRAL with influence function
$\chi_{0}={\phi}_{0,\overline{\mu}_{0}}-\psi_{0}(\overline{\mu}_{0})+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}+1(\overline{\alpha}_{0}\neq\alpha_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})B_{0}$
with $A_{0}:=(w,a,y)\mapsto s_{0}(a,w)\{y-\mu_{0}(a,w)\}$ and $B_{0}:=\widetilde{\phi}_{0,r_{0}}-r_{0}\alpha_{0}+\psi_{0}(r_{0})$, in the sense that $\tau_{n}^{*}-\tau_{0}=(P_{n}-P_{0})\chi_{0}+o_{p}(n^{-\frac{1}{2}})$. If also $\overline{\alpha}_{0}=\alpha_{0}$ and $\overline{\mu}_{0}=\mu_{0}$, then $\tau_{n}^{*}-\tau_{0}=(P_{n}-P_{0})D_{0}+o_{p}(n^{-\frac{1}{2}})$, so that $\tau_{n}^{*}$ is a regular and efficient estimator of $\tau_{0}$ under the nonparametric statistical model $\mathcal{M}$.
Theorem 4 establishes that the IC-DML estimator is asymptotically linear with influence function $\chi_{0}$, so that $n^{\frac{1}{2}}(\tau_{n}^{*}-\tau_{0})$ tends in distribution to a mean-zero normal random variable with variance $\sigma_{0}^{2}:=var_{0}\{\chi_{0}(Z)\}$; this paves the way for confidence interval construction provided $\sigma_{0}^{2}$ can be estimated consistently. The form of $\chi_{0}$ differs from the efficient influence function $\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}$ by either $A_{0}$ or $B_{0}$ under the inconsistent limits. In the special case where $\psi_{0}$ equals $\mu\mapsto E_{0}\{m(A,W,\mu)\}$ for a known map $m:\mathcal{A}\times\mathcal{W}\times\mathcal{H}\rightarrow\mathbb{R}$, the influence function component $B_{0}$ simplifies to $m(\cdot,r_{0})-r_{0}\alpha_{0}$.
Notably, when both nuisance estimators are consistent, Theorem 4 indicates that the IC-DML estimator is a regular, asymptotically linear, and efficient estimator of $\tau_{0}$ under weaker conditions than standard doubly robust estimators, allowing one nuisance function to be estimated arbitrarily slowly. Specifically, condition C6 requires that either $\|\mu_{n,j}^{*}-\mu_{0}\|=o_{p}(n^{-\frac{1}{4}})$ or $\|\alpha_{n,j}^{*}-\alpha_{0}\|=o_{p}(n^{-\frac{1}{4}})$, which substantially relaxes the standard condition $\|\mu_{n,j}^{*}-\mu_{0}\|\|\alpha_{n,j}^{*}-\alpha_{0}\|=o_{p}(n^{-\frac{1}{2}})$ for rate double robustness. This relaxation can also be shown to hold for the DRAL estimators proposed in van der Laan (2014), Benkeser et al. (2017), Dukes et al. (2021), and Bonvini et al. (2024). Higher-order influence function-based estimators, such as those proposed in Robins et al. (2008), have also been shown to achieve asymptotic efficiency under weaker conditions than standard doubly robust estimators. However, these estimators are usually not model-agnostic, relying, for example, on strong Hölder smoothness assumptions on the nuisance functions, and require the selection of tuning parameters for which there is limited practical guidance.
As detailed in the next section, the regularity and efficiency of the IC-DML estimator under the consistency of both nuisance estimators are important, as they address a potential criticism of DRAL estimators — namely, that they are irregular (Dukes et al., 2021). In particular, provided one nuisance function is estimated consistently at a sufficiently fast rate, Theorem 4 finds that irregularity only ever occurs when the other nuisance function is inconsistently estimated. In other words, if the other nuisance function is consistently estimated but at a slow rate, the IC-DML estimator remains regular and asymptotically linear, in contrast to standard doubly robust estimators, which are then neither regular nor asymptotically linear. Since many machine learning methods — including random forests, neural networks, and gradient-boosted trees — are known to be universal approximators (Hornik et al., 1989; Biau et al., 2008), while slow nuisance estimation may well occur, inconsistent nuisance estimation is less likely to arise. Even in the special case of the ATE parameter, Theorem 4 assumes substantially weaker conditions than prior works on doubly robust inference (van der Laan, 2014; Benkeser et al., 2017; Dukes et al., 2021). Specifically, these previous works imposed that the nuisance estimators converge to their (possibly incorrect) limits quickly, requiring that $\|\mu_{n,j}^{*}-\overline{\mu}_{0}\|\|\alpha_{n,j}^{*}-\overline{\alpha}_{0}\|=o_{P}(n^{-\frac{1}{2}})$, which offers no improvement over rate double robustness when both nuisance estimators are consistent.
It can be shown that Theorem 4 remains true even when only the consistent nuisance estimators are calibrated. In particular, if $\overline{\mu}_{0}=\mu_{0}$ and we take $\alpha_{n,j}^{*}=\overline{\alpha}_{0}:=0$ for each $j=1,\ldots,J$, we can establish that the isotonic calibrated plug-in estimator $\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(\mu_{n,j}^{*})$ is asymptotically linear under similar conditions. In the case of the ATE, this debiased plug-in estimator is asymptotically equivalent to the superefficient TML estimator that leverages the outcome regression for confounding bias adjustment, as proposed in Benkeser et al. (2020). Similarly, if $\overline{\alpha}_{0}=\alpha_{0}$ and we take $\mu_{n,j}^{*}=\overline{\mu}_{0}:=0$, we can show that the balancing weighted outcome estimator $\frac{1}{n}\sum_{i=1}^{n}\alpha_{n,j(i)}^{*}(A_{i},W_{i})Y_{i}$ is asymptotically linear, leveraging the Riesz representer for confounding bias adjustment. As a result, isotonic calibration offers a simple and automated debiasing approach for constructing asymptotically linear (although irregular) inverse-probability-of-treatment weighting and G-computation plug-in estimators, eliminating the requirement to derive the efficient influence function of the target parameter.
We now discuss the conditions of the above theorem. Condition C2 ensures that the (possibly inconsistent) nuisance estimators have in-probability limits in $\mathcal{H}$ and is commonly required for doubly robust inference (Benkeser et al., 2017; Smucler et al., 2019; Dukes et al., 2021). Condition C3 requires bounded nuisance functions and estimators. The boundedness of the Riesz representer $\alpha_{0}$ in C3 is a general form of the strong positivity assumption (Chernozhukov et al., 2022). If the $P_{0}$–support of the outcome $Y$ is bounded, then it is typically not difficult to ensure boundedness of $\mu_{0}$ and its estimator. Condition C4 imposes a weak upper bound on the number of plateaus in the (piecewise constant) isotonic regression solution; this is expected to hold in any setting with noisy outcomes where isotonic regression cannot perfectly fit the data. In view of Deng et al. (2021), continuous differentiability of $t\mapsto E_{0}\{Y\,|\,\mu_{n,j}(A,W)=t,\mathcal{D}_{n}\}$ implies that the number of observations in a given plateau of an isotonic regression solution grows at a rate $n^{\frac{2}{3}}$ with probability tending to one; heuristically, this suggests that the number of plateaus grows at rate $n^{\frac{1}{3}}=o(n^{\frac{1}{2}})$. Regardless, this condition can be ensured by constraining the maximum number of plateaus of the isotonic regression solution via the maximum tree depth argument in xgboost. Condition C5 is a smoothness condition on the data-generating distribution that allows us to deploy Lemma 1 to bound the projection estimation error terms in B1 by the cross-fitted nuisance estimation errors. This condition can be potentially relaxed by directly verifying that B1 holds for each cross-fitted nuisance estimator. Condition C6 ensures that the IC-DML estimator is DRAL, allowing for valid inference even with inconsistent or insufficiently fast estimation of one of the nuisance functions, as long as the other nuisance function is consistently estimated at a sufficiently fast rate. Notably, Bonvini et al. (2024) recently showed that, under a hybrid smoothness and structure-agnostic model on the nuisance functions, which assumes only C5, the estimation rate achieved by the IC-DML estimator is minimax optimal. As a consequence, the rate conditions for asymptotic normality imposed by C6 cannot be meaningfully improved without making additional structural assumptions. Condition C7 requires that the linearization remainder term of $\psi_{n,j}(r_{n,j}^{*})$ as an estimator of $\psi_{0}(r_{n,j}^{*})$ be asymptotically negligible. This condition is trivially satisfied for functionals of the form given in (1), that is, with $\psi_{0}:=\mu\mapsto E_{0}\{m(A,W,\mu)\}$ for some map $(a,w,\mu)\mapsto m(a,w,\mu)$, when $\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}$ is taken to be the empirical plug-in estimator $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(A_{i},W_{i},\mu)$.
5.2 Doubly robust validity of bootstrap-assisted confidence intervals
We now study the properties of the automated bootstrap-assisted confidence intervals obtained from Algorithm 2, showing, in particular, that they provide doubly robust asymptotically valid inference for $\tau_{0}$ under weak conditions. Importantly, Algorithm 2 does not require prior knowledge of whether any of the nuisance estimators is inconsistent and circumvents entirely estimation of the nuisance functions $r_{0}$ and $s_{0}$ appearing in the limiting distribution of Theorem 4. Hence, provided the IC-DML estimator $\tau_{n}^{*}$ can be computed within each bootstrap sample, this bootstrap procedure is fully automated. Nevertheless, while the bootstrap-assisted inference procedure does not require estimation of $r_{0}$ and $s_{0}$ to construct confidence intervals and conduct hypothesis testing, if desired, these nuisance functions can be estimated using univariate empirical risk minimization — see Appendix B.1 for details.
We focus our results on the case where $\psi_{0}$ has the form $\mu\mapsto E_{0}\{m(A,W,\mu)\}$ and $\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}$ is taken to be the empirical plug-in estimator $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(A_{i},W_{i},\mu)$. Below, we define $P_{n,j}^{\#}$ as the empirical distribution of an arbitrary bootstrap sample $\mathcal{C}^{\#(j)}$ drawn from the empirical distribution $P_{n,j}$ of $\mathcal{C}^{(j)}$. Moreover, we denote the bootstrapped calibrated nuisance estimators corresponding to $P_{n,j}^{\#}$ as $\mu_{n,j}^{\#}$ and $\alpha_{n,j}^{\#}$, which are obtained as in Algorithm 2. Finally, we define $\tau_{n}^{*\#}$ as the corresponding bootstrapped IC-DML estimator obtained from $\mu_{n,1}^{\#},\ldots,\mu_{n,J}^{\#}$ and $\alpha_{n,1}^{\#},\ldots,\alpha_{n,J}^{\#}$. The following conditions are used in the upcoming theorem:
  1. D1)
    Boundedness of bootstrap nuisance estimators: for some $M\in(0,\infty)$, the functions $\mu_{n,j}^{*\#}$ and $\alpha_{n,j}^{*\#}$ are uniformly bounded by $M$ with $P_{0}$–probability tending to one for each $j=1,\ldots,J$;
  2. D2)
    Sparsity of bootstrap calibrators: with $P_{0}$–probability tending to one, the bootstrap isotonic calibrators $f_{n}^{\#}$ and $g_{n}^{\#}$ take at most $k(n)>0$ distinct values for some deterministic sequence $k(n)=o(n^{\frac{1}{2}})$;
  3. D3)
    Bootstrap Lipschitz continuity: the random functions $(\widehat{m},m)\mapsto E_{0}\{\Delta_{\alpha,n,j}^{*\#}\,|\,\mu_{n,j}^{*\#}(X)=\widehat{m},\mu_{0}(X)=m,\mathcal{D}_{n}\}$ and $(\widehat{a},a)\mapsto E_{0}\{\Delta_{\mu,n,j}^{*\#}\,|\,\alpha_{n,j}^{*\#}(X)=\widehat{a},\alpha_{0}(X)=a,\mathcal{D}_{n}\}$ is $P_{0}$–almost surely Lipschitz continuous with some constant $L\in(0,\infty)$ for each $j=1,\ldots,J$ and all $n$ large enough.
Theorem 5.
Under the conditions of Theorem 4 and D1D3, it holds that
$\tau_{n}^{*\#}-\tau_{n}^{*}=\frac{1}{J}\sum_{j=1}^{J}(P_{n,j}^{\#}-P_{n,j})\chi_{0}+o_{p}(n^{-\frac{1}{2}})\ .$
Moreover, the conditional law of $n^{\frac{1}{2}}(\tau_{n}^{*\#}-\tau_{n}^{*})$ given $\{P_{n,j}:j=1,\ldots,J\}$ converges marginally in $P_{0}$–probability to the law of a mean-zero normal random variable with variance $\sigma_{0}^{2}:=var_{0}\{\chi_{0}(Z)\}$.
Under regularity conditions, as the number of bootstrap samples $K\rightarrow\infty$, the scaled version $n^{\frac{1}{2}}\sigma_{n}^{\#}$ of the bootstrap standard error of Algorithm 2 is a consistent estimator of the limiting standard deviation $\sigma_{0}$ of $n^{\frac{1}{2}}(\tau_{n}^{*}-\tau_{0})$. Hence, by Theorem 5, the bootstrap-assisted confidence interval $\text{CI}_{n}(\rho)$ from Algorithm 2 provides doubly robust asymptotically exact coverage, in the sense that $P_{0}\{\tau_{0}\in\text{CI}_{n}(\rho)\}\rightarrow 1-\rho$ as $n\rightarrow\infty$, provided that $K\rightarrow\infty$ sufficiently quickly.
5.3 On estimator irregularity under inconsistent nuisance estimation
As emphasized before, Theorem 4 establishes that the isotonic calibrated one-step debiased estimator is DRAL, and that it is regular and efficient when both nuisance functions are consistently estimated, even if one nuisance function is estimated insufficiently fast. However, under inconsistent estimation of one of the nuisance functions, the resulting estimator is typically irregular under a nonparametric model, as it is asymptotically linear with an influence function differing from the nonparametric efficient influence function of the target parameter (van der Vaart, 2000; Dukes et al., 2021). This irregularity implies that inference for $\tau_{0}$ is not uniformly valid when considering sampling from local perturbations of the data-generating distribution. Irregularity is typically viewed as undesirable due to its potential to cause poor performance in finite samples, since it is possible to find a distribution such that an arbitrarily large number of samples may be required for the asymptotic limiting distribution to accurately approximate finite-sample behavior (Leeb and Pötscher, 2005). This criticism is arguably less relevant in our case since standard one-step debiased estimators are, under inconsistent estimation of one nuisance function, typically neither regular nor $n^{\frac{1}{2}}$–consistent, let alone asymptotically linear (Dukes et al., 2021). We stress here that the set of laws under which the IC-DML estimator is regular and asymptotically efficient is a strict superset of the corresponding set of laws for standard one-step debiased estimators, and furthermore, that the IC-DML estimator is asymptotically linear and permits valid pointwise (as opposed to uniform) inference at $n^{\frac{1}{2}}$–rate at laws at which the standard estimators neither converge at $n^{\frac{1}{2}}$–rate nor readily allow even pointwise inference.
The following theorem characterizes the local asymptotic behavior of the IC-DML estimator in the irregular case where one of the nuisance functions is inconsistently estimated. It establishes that the estimator is a locally regular and efficient estimator of a projection-based oracle parameter $\Psi_{0}$, which agrees with the original target parameter $P\mapsto\Psi(P):=\psi_{P}(\mu_{P})$ at $P_{0}$. As a result, the IC-DML procedure offers regular and locally uniformly valid inference for the corresponding projection estimand, even when sampling from least favorable local perturbations of $P_{0}$.
To define this oracle parameter, we first introduce additional notation. We define the affine spaces $\overline{\Theta}_{0,P}:=\{\overline{\mu}_{0}+f\circ\alpha_{0}:f\}\cap L^{2}(P)$ and $\overline{\Theta}_{0,P}^{*}:=\{\overline{\alpha}_{0}+g\circ\mu_{0}:g\}\cap L^{2}(P)$, where the indexing functions $f$ and $g$ vary over all one-dimensional, real-valued transformations. We denote the orthogonal projection of $\mu_{P}$ onto $\overline{\Theta}_{0,P}$ by $\overline{\Pi}_{0,P}\mu_{P}:=\operatorname*{argmin}_{\mu\in\overline{\Theta}_{0,P}}\|\mu_{P}-\mu\|_{P}$, and the orthogonal projection of $\alpha_{P}$ onto $\overline{\Theta}_{0,P}^{*}$ by $\overline{\Pi}_{0,P}^{*}\alpha_{P}:=\operatorname*{argmin}_{\alpha\in\overline{\Theta}_{0,P}^{*}}\|\alpha_{P}-\alpha\|_{P}$. We may then define the projection-based oracle parameter $\Psi_{0}$, which varies depending on which nuisance function is estimated inconsistently, as
$P\mapsto\Psi_{0}(P)$ $:=1(\overline{\mu}_{0}=\mu_{0})\overline{\psi}_{0,P}(\mu_{P})+1(\overline{\alpha}_{0}=\alpha_{0})\psi_{P}(\overline{\Pi}_{0,P}\mu_{P})\ ,$
where $\overline{\psi}_{0,P}:h\mapsto\langle\overline{\Pi}_{0,P}^{*}\alpha_{P},h\rangle_{P}$ is an oracle approximation of the linear functional $\psi_{P}:h\mapsto\langle\alpha_{P},h\rangle_{P}$. The orthogonality conditions characterizing the projections $\overline{\Pi}_{0}^{*}\alpha_{0}$ and $\overline{\Pi}_{0,P}\mu_{0}$ imply that $\Psi_{0}(P_{0})=\Psi(P_{0})$, so that the oracle parameter agrees with the target parameter at $P_{0}$. In the case where $\mu_{0}$ is consistently estimated, $\Psi_{0}$ reduces to the parameter $P\mapsto\overline{\psi}_{0,P}(\mu_{P})$, which leverages the outcome regression $\mu_{0}(A,W)$ as a dimension reduction of $(A,W)$. Alternatively, in the case where $\alpha_{0}$ is consistently estimated, $\Psi_{0}$ reduces to the parameter $P\mapsto\psi_{P}(\overline{\Pi}_{0,P}\mu_{P})$, which leverages the Riesz representer $\alpha_{0}(A,W)$ as a dimension reduction of $(A,W)$.
Theorem 6 (Behavior under local perturbations).
Suppose that the conditions of Theorem 4 hold and that either $\overline{\mu}_{0}=\mu_{0}$ or $\overline{\alpha}_{0}=\alpha_{0}$ but not both. Then, $\tau_{n}^{*}$ is a locally regular and asymptotically linear at $P_{0}$ for the oracle projection parameter $\Psi_{0}$ with influence function $\chi_{0}$. Consequently, under sampling from any local perturbation $P_{0,hn^{-\frac{1}{2}}}$ of $P_{0}$, it holds that $n^{\frac{1}{2}}\{\tau_{n}^{*}-\Psi_{0}(P_{0,hn^{-\frac{1}{2}}})\}$ tends to a mean-zero normal random variable with variance $\sigma_{0}^{2}$ as defined in Theorem 4.
Theorem 6 demonstrates that even under sampling from least favorable local perturbations, the IC-DML estimator still provides regular and locally uniformly valid inference for the projection-based oracle parameter $\Psi_{0}$. Notably, the magnitude of the oracle bias $\Psi_{0}(P)-\Psi(P)$ for $\Psi$ is driven by the degree of nuisance estimator inconsistency via projection distances $\|\overline{\Pi}_{\mu_{0}}\mu_{P}-\mu_{P}\|_{P}$ and $\|\overline{\Pi}_{\alpha_{0}}\alpha_{P}-\alpha_{P}\|_{P}$. Consequently, under sampling from a local perturbation, when the nuisance functions are only mildly inconsistently estimated, the local asymptotic bias of the IC-DML estimator for the target parameter $\Psi$ may still be relatively small. A similar result was derived in van der Laan et al. (2023a) for superefficient and irregular, adaptive debiased machine learning estimators obtained using data-driven model selection techniques. As a special case, Theorem 6 implies that $\tau_{n}^{*}$ is regular and asymptotically linear for the actual target parameter $\Psi$ over all local perturbations that do not perturb the consistently estimated nuisance function, which agrees with Theorem 2 of Dukes et al. (2021).
6 Numerical experiments
6.1 Experiment 1: performance under inconsistent nuisance estimation
We now evaluate the numerical performance of the IC-DML estimator through simulation studies. First, we conduct an experiment based on synthetic data to empirically verify the theoretical properties of the IC-DML estimator introduced in Algorithm 1 for inference on the ATE $\tau_{0}=E_{0}\{\mu_{0}(1,W)-\mu_{0}(0,W)\}$. The estimator we consider is defined as $\tau_{n}^{*}:=\tau_{n}^{*}(1)-\tau_{n}^{*}(0)$, where $\tau_{n}^{*}(a)$ is an IC-DML estimator of the counterfactual mean $E_{0}\{\mu_{0}(a,W)\}$ obtained, as in Example 1, using Algorithm 1. As a benchmark, we compare the performance of the proposed estimator with the standard cross-fitted AIPW estimator, outlined in Section 2.1, as well as the DR-TMLE proposed in Benkeser et al. (2017), described in Section 3.3 and implemented in the R package drtmle (Benkeser and Hejazi, 2023). Notably, the IC-DML estimator differs from the AIPW estimator solely through the additional step of isotonic calibration of the cross-fitted propensity score and outcome regression estimators. To quantify uncertainty, we utilize the bootstrap-assisted confidence intervals outlined in Algorithm 2 for the IC-DML estimator, and Wald-type confidence intervals derived from the influence function-based variance estimator for both the AIPW estimator and DR-TMLE.
We consider the same data-generating process used in the experiments of Benkeser et al. (2017). The covariate vector $W=(W_{1},W_{2})$ is first drawn as independent $W_{1}\sim\text{Uniform}(-2,2)$ and $W_{2}\sim\text{Bernoulli}(0.5)$. Then, given $(W_{1},W_{2})=(w_{1},w_{2})$, the treatment assignment variable $A$ is drawn from a Bernoulli distribution with success probability $\pi_{0}(w_{1},w_{2}):=\text{expit}(-w_{1}+2w_{1}w_{2})$. Given $A=a$ and $(W_{1},W_{2})=(w_{1},w_{2})$, the outcome variable $Y$ is drawn from a Bernoulli distribution with success probability $\mu(a,w_{1},w_{2}):=\text{expit}(0.2a-w_{1}+2w_{1}w_{2})$. To evaluate the doubly robust inference property, we examine three scenarios of inconsistent nuisance estimation for both the IC-DML, AIPW and DR-TMLE estimators:
  • (a)
    when both the outcome regression and treatment mechanism are estimated consistently;
  • (b)
    when only the treatment mechanism is estimated consistently;
  • (c)
    when only the outcome regression is estimated consistently.
In settings (b) and (c), to induce inconsistent estimation of the outcome regression and treatment mechanisms, we estimate the propensity score $\pi_{0}$ and outcome regressions $\mu_{0}$ using (misspecified) main-term logistic regression models. In all settings, any consistently estimated nuisance function is estimated using a Nadaraya–Watson kernel smoother with bandwidth selected via cross-validation and with stratification by both $A$ and $W_{2}$; the implementation available in R package fksum was used (Hofmeyr, 2020). For the DR-TMLE estimator, we used this same kernel estimator to estimate the nuisance functions $r_{0}$ and $s_{0}$. For both the IC-DML, DR-TMLE, and AIPW estimators, we employ 5-fold cross-fitting when estimating the initial nuisance functions. To avoid positivity issues, the cross-fitted propensity score estimators used to construct the AIPW estimator were truncated to lie in $[c_{n},1-c_{n}]$, where $c_{n}:=25/(n^{\frac{1}{2}}\log n)$, as suggested in Gruber et al. (2022). Details for reproducing these analyses can be found in the associated GitHub repository, accessible through a link provided in the Appendix.
For each scenario and sample sizes ranging from 250 to 10,000, a total of 5,000 datasets were randomly generated and analyzed using all estimation procedures considered. Empirical estimates of the standard error of the IC-DML estimator against the AIPW estimator as well as the empirical coverage of nominal 95% confidence interval are displayed in Figure 1. Both the IC-DML and DR-TMLE estimators appeared to exhibit DRAL, showing less bias and better coverage than the AIPW estimator in settings with inconsistent estimation. Despite being easier and faster to implement, across all inconsistent settings, the IC-DML estimator showed less bias and resulted in confidence intervals with better coverage than DR-TMLE, while also having smaller or equal variance. In the setting where both nuisance functions were consistently estimated, all estimators performed similarly.
Figure 1: Relative bias, standard error, and 95% confidence interval coverage for IC-DML, DR-TMLE and AIPW estimators under both doubly and singly consistent estimation of the outcome regression and treatment mechanism.
(a) Both treatment and outcome mechanisms are correct.
(b) Only outcome mechanism is correct.
(c) Only treatment mechanism is correct.
6.2 Experiment 2: benchmarking on semi-synthetic datasets
In this experiment, we evaluate the performance of our approach on various semi-synthetic datasets that are commonly used for benchmarking causal inference methods. The objective of this experiment is to evaluate the performance of the IC-DML estimator in challenging and complex real-world scenarios. These scenarios are characterized by moderately high-dimensional covariates, limited overlap, realistic confounding, and class imbalance between treatment arms. Among these semi-synthetic datasets are the Twins, Lalonde PSID, and Lalonde CPS datasets introduced in Neal et al. (2020), which are based on generative models trained on the original data sources (Louizos et al., 2017; LaLonde, 1986) using the realCause framework. Several semi-synthetic datasets are derived from the 2017 Atlanta Causal Inference Challenge (ACIC-2017, settings 17-24) (Hahn et al., 2019), which uses covariate information from the Infant Health and Development Program (IHDP) (Brooks-Gunn et al., 1992). We also examine a total of 680 datasets from the scaling component of ACIC-2018 (Shimoni et al., 2018), which include sample sizes of 1,000, 2,500, 5,000, and 10,000. Notably, ACIC-2018 is based on a diverse set of randomly generated data-generating mechanisms and has also been used as a benchmark in Shi et al. (2019). The final dataset (IHDP) is a semi-synthetic dataset based on data from the IHDP, which was introduced in Hill (2011) (Setting B) and also analyzed in Shalit et al. (2017). Detailed information about sample sizes, covariate dimensions, and the sources of these datasets is provided in Table 1. For all datasets, our inferential target is the sample average treatment effect (SATE) $\tau_{n,0}:=\frac{1}{n}\sum_{i=1}^{n}(Y_{1,i}-Y_{0,i})$, that is, the empirical mean of the individual treatment effects among individuals in the observed sample. We compare the IC-DML estimator against the AIPW estimator, where both estimators are defined identically as in the previous section. The cross-fitted propensity score and outcome regression estimates are obtained using gradient-boosted regression trees as implemented by xgboost (Chen et al., 2015) with maximum tree depth tuned using 10-fold cross-validation. None of the nuisance estimators are deliberately inconsistently estimated, unlike in Experiment 1.
Dataset $n_{total}$ $n_{treated}$ $d$ source
ACIC-2017 4302 $2233$ 58 Hahn et al. (2019)
ACIC-2018 varied varied varied Shimoni et al. (2018)
IHDP 672 123 25 Shalit et al. (2017)
Lalonde CPS 16177 $167$ 9 Neal et al. (2020)
Lalonde PSID 2675 $175$ 9 Neal et al. (2020)
Twins 11984 8935 76 Neal et al. (2020)
Table 1: Sample size of total study ($n_{total}$), expected size of treatment arm $(n_{treated})$, and covariate dimension $(d)$ for the semi-synthetic datasets of second experiment. For the ACIC-2017 synthetic dataset, we employed simulation setting 24, characterized by additive errors, a large signal, large noise, and strong confounding.
For each collection of $k$ dataset realizations, Table 2 reports empirical absolute bias, scaled root mean square error $\frac{1}{k}\sum_{i=1}^{k}\{\widehat{\tau}_{n}^{(i)}/\tau_{n,0}^{(i)}-1\}^{2}$, and nominal 95% confidence interval coverage, where $\widehat{\tau}_{n}^{(i)}$ and $\tau_{n,0}^{(i)}$ are the estimated and true SATEs for the $i$th dataset realization. The IC-DML estimator generally exhibited lower absolute bias compared to the AIPW estimator across several datasets. Notably, in the ACIC-2017 settings 20 and 24, and the Lalonde CPS dataset, the IC-DML estimator significantly outperformed the AIPW estimator, achieving bias reductions by factors ranging from approximately 1.5 to 8. In other instances, such as the ACIC-2017 setting 22 and the Lalonde PSID dataset, the biases of both estimators were comparable. The IC-DML estimator consistently demonstrated superior performance in terms of scaled RMSE, achieving lower values across most datasets. For example, in the ACIC-2017 setting 24, the IC-DML estimator had a scaled RMSE of 0.252 compared to 0.345 for the AIPW estimator. Similarly, in the ACIC-2017 setting 20, the IC-DML estimator had a scaled RMSE of 1.39 compared to 1.95 for the AIPW estimator, indicating greater accuracy in estimating the treatment effect. The IC-DML estimator generally maintained better or comparable coverage rates for the 95% confidence intervals. For instance, in the ACIC-2017 settings 18, 20, and 24, and the Lalonde CPS and Lalonde PSID datasets, the IC-DML estimator achieved empirical coverage rates of 64.4%, 90%, 90%, 75%, and 80%, respectively, compared to significantly lower rates for the AIPW estimator, ranging from 10% to 50%.
Overall, the IC-DML estimator outperformed the AIPW estimator in terms of absolute bias and scaled RMSE across a range of semi-synthetic datasets. The isotonic calibration step in the IC-DML estimator improved the accuracy of point estimates, while the bootstrap-assisted confidence intervals provided better coverage properties.
Table 2: Evaluation of Absolute Bias, Scaled Root Mean Square Error (RMSE), and 95% Confidence Interval coverage on semi-synthetic benchmark datasets. Both outcome regression and propensity scores are estimated using an ensemble of gradient-boosted trees. For comparability across rows, the absolute bias is scaled by a constant equal to the average effect size across dataset realizations for that benchmark. For clarity, we report only the sample-size aggregated results for ACIC-2018 and the settings with strong confounding for ACIC-2017. The full results, which also include misspecification of the outcome and treatment mechanisms, can be found in Appendix. A.2.
Dataset Bias RMSE Coverage
IC-DML AIPW IC-DML AIPW IC-DML AIPW
ACIC-2017 (18) 0.28 0.21 0.58 0.70 0.64 0.27
ACIC-2017 (20) 0.20 1.6 1.4 2.0 0.90 0.32
ACIC-2017 (22) 0.035 0.004 0.10 0.11 0.81 0.56
ACIC-2017 (24) 0.04 0.30 0.25 0.35 0.90 0.32
ACIC-2018 (Aggr) 7.1 9.0 110 97 0.69 0.58
IHDP 0.13 0.13 0.46 0.46 0.57 0.57
Lalonde CPS 0.084 0.14 0.34 0.22 0.75 0.16
Lalonde PSID 0.039 0.038 0.44 0.19 0.84 0.46
Twins 0.21 0.22 0.23 0.24 0.54 0.51
7 Conclusion
In this work, we introduced a unified machine learning framework for doubly robust inference on linear functionals of the outcome regression. These functionals belong to a subset of parameters characterized by the mixed bias property (Rotnitzky et al., 2021). For the case of linear functionals, our framework provides an affirmative answer to the open question raised by Smucler et al. (2019), regarding the feasibility of achieving doubly robust inference for mixed-bias parameters using black-box nuisance function estimators. Another significant contribution of this work lies in establishing a close relationship between calibration and the doubly robust inference properties of debiased machine learning estimators. Importantly, our proof techniques rely solely on the nuisance estimators satisfying the calibration score equations stated in Section 5.1. While our IC-DML framework uses isotonic calibration, these score equations can also be satisfied when nuisance functions are calibrated using other histogram-type estimators, such as (honest) regression trees (Wager and Athey, 2018), the highly adaptive lasso (van der Laan, 2017; Benkeser and van der Laan, 2016; Fang et al., 2021), and Venn-Abers calibration (Vovk and Petej, 2012; van der Laan and Alaa, 2024). In general, both our theoretical and empirical results support the use of calibrated nuisance estimators in causal inference to mitigate bias arising from inconsistent nuisance estimation and to ensure calibrated inferences.
There remain several open questions along this line of research. First, whether our framework can be extended to handle general parameters characterized by the mixed bias property (Rotnitzky et al., 2021) remains to be established. Preliminary research we have done on this question suggests that this may indeed be possible, although further work is required. Second, it may be of interest to develop an isotonic calibrated targeted minimum loss estimator that inherits both the plug-in property and doubly robust asymptotic linearity, as considered in van der Laan (2014) and Benkeser et al. (2017). Such an estimator could outperform non-plug-in alternatives such as one-step debiased estimators (Porter et al., 2011). A potential approach for this could be to iteratively apply the isotonic calibration and plug-in debiasing steps until the relevant score equations for both DRAL and the plug-in property are solved. Lastly, extending our framework to accommodate multiple and sequentially robust estimators in longitudinal data structures presents further exciting research opportunities (Rotnitzky et al., 2017; Luedtke et al., 2017).
Acknowledgements. Research reported in this publication was supported by NIH grants DP2-LM013340 and R01-HL137808, and NSF grant DMS-2210216. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding agencies.
Appendix A Additional details on numerical experiments
A.1 Simulation design for experiment 1
We generate $n$ iid draws of $Z=(W,A,Y)$ with $n\in\{500,1000,2000,3000,4000,5000\}$ to construct each simulated dataset as follows. First, we generate $W=(W_{1},W_{2},W_{3})$ as a vector of three independent draws from the the uniform distribution over $(-1,1)$. Next, given $W=w=(w_{1},w_{2},w_{3})$, we generate $A$ from a Bernoulli distribution with mean $\pi_{0}(w):=\operatorname*{expit}[1.5\{\text{sign}(w_{1})\sqrt{|w_{1}|}+\sin(3.14w_{2})+w_{3}\sin(w_{3})-0.5\}]$. Finally, given $A=a$ and $W=w$, we generate $Y$ from a Bernoulli distribution with mean $\mu_{0}(a,w)=\operatorname*{expit}[-1+\cos(3.14W_{1})+W_{2}\sin(W_{2})+\text{sign}(W_{3})\sqrt{|W_{3}|}+a\{1+\text{sign}(W_{1})\sqrt{|W_{1}|}+\sin(3.14W_{2})+W_{3}\sin(W_{3})\}]$.
A.2 Complete results for experiment 2
Table 3: Bias, root mean square error (RMSE), and 95% confidence interval coverages on semi-synthetic datasets without deliberate inconsistent estimation of the outcome regression or treatment mechanism.
Dataset Bias RMSE Coverage
IC-DML AIPW IC-DML AIPW IC-DML AIPW
ACIC-2017 (17) 0.028 0.031 0.097 0.096 0.95 0.95
ACIC-2017 (18) 0.28 0.21 0.58 0.70 0.64 0.27
ACIC-2017 (19) 0.057 0.082 0.43 0.43 0.96 0.96
ACIC-2017 (20) 0.2 1.6 1.4 2.0 0.90 0.32
ACIC-2017 (21) 0.0042 0.0057 0.017 0.018 1.00 1.00
ACIC-2017 (22) 0.035 0.004 0.10 0.11 0.81 0.56
ACIC-2017 (23) 0.010 0.014 0.080 0.080 0.96 0.97
ACIC-2017 (24) 0.04 0.30 0.25 0.35 0.90 0.32
ACIC-2018 (1000) 1.5 4.9 190 120 0.67 0.58
ACIC-2018 (10000) 0.50 0.11 13 12 0.71 0.58
ACIC-2018 (2500) 11 13 240 240 0.67 0.60
ACIC-2018 (5000) 16 20 78 56 0.66 0.56
ACIC-2018 (Aggr) 7.1 9.0 110 97 0.69 0.58
IHDP 0.13 0.13 0.46 0.46 0.57 0.57
Lalonde CPS 0.084 0.14 0.34 0.22 0.75 0.16
Lalonde PSID 0.039 0.038 0.44 0.19 0.84 0.46
Twins 0.21 0.22 0.23 0.24 0.54 0.51
Table 4: Bias, root mean square error (RMSE), and 95% confidence interval coverages on semi-synthetic datasets with deliberate misspecification of outcome regression (empirical mean estimator).
Dataset Bias RMSE Coverage
C-DML AIPW C-DML AIPW C-DML AIPW
ACIC-2017 (17) 0.30 0.58 0.33 0.69 0.86 0.73
ACIC-2017 (18) 3.1 6.9 3.3 6.9 0.06 0.00
ACIC-2017 (19) 0.32 0.59 0.53 0.81 0.94 0.83
ACIC-2017 (20) 3.1 6.9 3.5 7.0 0.29 0.00
ACIC-2017 (21) 0.05 0.10 0.07 0.13 0.91 0.77
ACIC-2017 (22) 0.6 1.2 0.6 1.2 0.14 0.00
ACIC-2017 (23) 0.06 0.10 0.10 0.15 0.92 0.86
ACIC-2017 (24) 0.5 1.2 0.6 1.2 0.35 0.00
ACIC-2018 (1000) 1.6 3.8 120 120 0.65 0.58
ACIC-2018 (10000) 7.8 10 30 29 0.63 0.54
ACIC-2018 (2500) 51 93 150 120 0.64 0.55
ACIC-2018 (5000) 6.5 24 160 81 0.57 0.53
ACIC-2018 (AGGR) 15 30 110 70 0.62 0.54
IHDP 0.14 0.14 0.49 0.50 0.58 0.53
Lalonde CPS 0.14 0.37 0.36 0.38 0.49 0.00
Lalonde PSID 0.041 0.013 0.44 0.06 0.87 0.71
Twins 0.31 0.33 0.33 0.35 0.25 0.19
Table 5: Bias, root mean square error (RMSE), and 95% confidence interval coverages on semi-synthetic datasets with deliberate misspecification of propensity score (empirical mean estimator).
Dataset Bias RMSE Coverage
C-DML AIPW C-DML AIPW C-DML AIPW
ACIC-2017 (17) 0.08 0.13 0.12 0.16 0.89 0.67
ACIC-2017 (18) 0.78 0.87 2.0 2.0 0.47 0.32
ACIC-2017 (19) 0.25 0.45 0.49 0.62 0.93 0.78
ACIC-2017 (20) 5.4 5.8 5.7 6.1 0.044 0.024
ACIC-2017 (21) 0.010 0.025 0.019 0.031 1.00 0.98
ACIC-2017 (22) 0.18 0.21 0.36 0.37 0.55 0.44
ACIC-2017 (23) 0.041 0.081 0.09 0.11 0.95 0.83
ACIC-2017 (24) 0.9 1.0 1.0 1.1 0.028 0.008
ACIC-2018 (1000) 3.0 3.9 190 110 0.64 0.63
ACIC-2018 (10000) 5.4 9.3 15 19 0.66 0.61
ACIC-2018 (2500) 15 17 210 210 0.64 0.60
ACIC-2018 (5000) 24 22 65 58 0.61 0.57
ACIC-2018 (Aggr) 12 14 95 87 0.64 0.60
IHDP 0.14 0.14 0.46 0.47 0.54 0.57
Lalonde CPS 0.21 0.14 0.26 0.22 0.56 0.25
Lalonde PSID 0.025 0.052 0.11 0.20 0.92 0.77
Twins 0.23 0.25 0.25 0.27 0.43 0.40
Appendix B Additional details on IC-DML estimator
B.1 Estimation of nuisance functions in limiting distribution
Using the bootstrap-assisted inference method of Algorithm 2, it is not necessary to obtain estimates of the univariate nuisance functions $f_{0,\overline{\mu}_{0}}$ and $g_{0,\overline{\alpha}_{0}}$, which appear in the limited distribution of Theorem 4. Nonetheless, if desired, we note that these nuisance functions can be estimated using univariate empirical risk minimization. To see this, observe that $f_{0,\overline{\mu}_{0}}$ is the Riesz-representer of the linear functional $\alpha\mapsto\langle\alpha_{0}-\overline{\alpha}_{0},\alpha\rangle_{P_{0}}=\psi_{0}(\alpha)-\langle\overline{\alpha}_{0},\alpha\rangle_{P_{0}}$ with respect to the linear space $\left\{f\circ\overline{\mu}_{0}:f:\mathbb{R}\rightarrow\mathbb{R},\|f\circ\mu\|<\infty\right\}$. Hence, as a consequence of the Riesz-defined above Theorem 1, the function $f_{0,\overline{\mu}_{0}}$ can be obtained via univariate empirical risk minimization, after substituting $\overline{\mu}_{0}$ and $\overline{\alpha}_{0}$ with their respective estimators. Specifically, the function $f_{0,\overline{\mu}_{0}}$ can be directly estimated via offset univariate empirical risk minimization over $\{f\circ\mu_{n,\diamond}^{*}\in L^{2}(\overline{P}_{0,A,W})\}$ based on an empirical estimator of the Riesz-loss risk:
$f\mapsto\frac{1}{n}\sum_{i=1}^{n}\left[\left\{\alpha_{n,j(i)}^{*}-(f\circ\mu_{n,j(i)}^{*})\right\}^{2}(A_{i},W_{i})-2\left\{(f\circ\mu_{n,j(i)}^{*}\cdot\alpha_{n,j(i)}^{*})(A_{i},W_{i})-\psi_{n,j(i)}\left(f\circ\mu_{n,j(i)}^{*}\right)\right\}\right].$
Similarly, noting that $\overline{\mu}_{0,\alpha}(a,w)=E_{0}\left[Y-\overline{\mu}_{0}(A,W)|\alpha(A,W)=\alpha(a,w)\right]$, the function $g_{0,\overline{\alpha}_{0}}$ can be estimated by regressing the residual $\{Y_{i}-\mu_{n,j(i)}^{*}(A_{i},W_{i})\}_{i\in[n]}$ on $\{\alpha_{n,j(i)}^{*}(A_{i},W_{i})\}_{i\in[n]}$ using, for instance, locally adaptive regression splines (Mammen and Van De Geer, 1997) or random forests (Breiman, 2001).
B.2 Extension to parameters with the mixed bias property
In this section, we present a brief overview of how our doubly robust inference framework can be extended to a general class of parameters with the mixed bias property, as introduced in Rotnitzky et al. (2021). While our primary focus in this paper remains on the class of linear functionals discussed in Section 2.1, it is important to highlight that our results can be easily generalized to parameters exhibiting the mixed bias property.
To this end, let $Z\sim P_{0}$ be an arbitrary data-structure drawn from a distribution $P_{0}$ belonging to a nonparametric statistical model $\mathcal{M}$. We consider doubly robust inference for a feature $\tau_{0}:=\Psi(P_{0})$ for a parameter $\Psi:\mathcal{M}\rightarrow\mathbb{R}$. We assume that $\Psi$ is pathwise differentiable with efficient influence function $\chi_{0}\in L^{2}_{0}(P_{0})$ at $P_{0}$. We further assume that $\Psi$ has the mixed bias property at $P_{0}$ meaning that, for all $\overline{P}\in\mathcal{M}$, the following condition holds:
$\Psi(\overline{P})-\Psi(P_{0})+E_{0}[\chi_{\overline{P}}(Z)]=E_{0}[S(Z)\{\overline{\nu}(Z)-\nu_{0}(Z)\}\{\overline{\eta}(Z)-\eta_{0}(Z)\}],$ (13)
where $z\mapsto S(z)$ is a fixed function, $\overline{\nu}:=\nu_{\overline{P}}$, $\nu_{0}:=\nu_{P_{0}}$, $\overline{\eta}:=\eta_{\overline{P}}$, and $\eta_{0}:=\eta_{P_{0}}$ are nuisance functions. Under the regularity conditions of Theorem 2 in Rotnitzky et al. (2021), it can be shown that
$(\nu_{0},\eta_{0})=\operatorname*{argmin}_{\nu,\eta\in L^{2}(P_{0})}E_{0}\left[S(Z)\left\{\nu^{2}(Z)+\eta^{2}(Z)\right\}-2\{m_{1}(Z,\nu)+m_{2}(Z,\eta)\}\right],$
for fixed functions $(z,\nu)\mapsto m_{1}(z,\nu)$ and $(z,\eta)\mapsto m_{2}(z,\eta)$. In the special case, where $S=1$, the optimization problem can be decoupled as follows:
$\nu_{0}=\operatorname*{argmin}_{\nu\in L^{2}(P_{0})}E_{0}\left[\nu^{2}(Z)-2m_{1}(Z,\nu)\right];\;\eta_{0}=\operatorname*{argmin}_{\eta\in L^{2}(P_{0})}E_{0}\left[\eta^{2}(Z)-2m_{2}(Z,\eta)\right].$
We note that the setup of Section 2.1 is recovered by taking $Z=(W,A,Y)$, $P\mapsto\Psi(P)=\psi_{P}(\mu_{P}):=E_{P}\{m(A,W,\mu_{P})\}$, $\mu\mapsto\psi_{n}(\mu):=\frac{1}{n}\sum_{i=1}^{m}m(W_{i},A_{i},\mu)$, $(\nu,\eta,z)\mapsto S:=1$, $P\mapsto\nu(P):=\mu_{P}$, $P\mapsto\eta(P):=\alpha_{P}$, $(z,\nu)\mapsto m_{1}(z,\nu):=y\nu(a,w)$, and $(z,\eta)\mapsto m_{2}(z,\eta):=m(a,w,\eta)$.
We now introduce the generalization of the calibrated one-step debiased estimator proposed in Section 4. Given initial estimators $\nu_{n}$ and $\eta_{n}$ of $\nu_{0}$ and $\eta_{0}$, we define the calibrated estimators $\nu_{n}^{*}=f_{n}\circ\nu_{n}$ and $\eta_{n}^{*}=g_{n}\circ\eta_{n}$ where $(f_{n},g_{n})$ are obtained as:
$(f_{n},g_{n})$ $\in\operatorname*{argmin}_{f,g\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\left[S(Z_{i})\left\{(f\circ\nu_{n})^{2}(Z_{i})+(g\circ\eta_{n})^{2}(Z_{i})\right\}-2\{m_{1}(Z_{i},f\circ\nu_{n})+m_{2}(Z_{i},g\circ\eta_{n})\}\right].$ (14)
Notably, assuming suitable smoothness in the loss function, the first order conditions of the minimizing solutions guarantees that
$\frac{1}{n}\sum_{i=1}^{n}\left[S(Z_{i})\left\{(f\circ\nu_{n})(Z_{i})\nu_{n}^{*}(Z_{i})-\frac{m_{1}(Z_{i},f\circ\nu_{n})}{S(Z_{i})}+(g\circ\eta_{n})(Z_{i})\eta_{n}^{*}(Z_{i})-\frac{m_{2}(Z_{i},g\circ\eta_{n})}{S(Z_{i})}\right\}\right]=0,$
for all transformations $f:\mathbb{R}\rightarrow\mathbb{R}$ and $g:\mathbb{R}\rightarrow\mathbb{R}$.
Given the calibrated nuisance estimators $\nu_{n}^{*}$ and $\eta_{n}^{*}$, a calibrated one-step debiased estimator of $\Psi(P_{0})$ can be constructed as in Rotnitzky et al. (2021) (see also the companion paper Smucler et al. (2019)). As in Algorithms 1 and 2, the initial nuisance estimators can be cross-fitted and confidence intervals can be constructed using the bootstrap. While our focus in this work is on linear functionals, our proof techniques can be readily extended to establish rate and inference double robustness for the proposed calibrated one-step debiased estimator of $\Psi(P_{0})$, subject to conditions.
Appendix C Implementation of IC-DML
An R package CDML as well as Python code implementing CDML can be found on GitHub at the following link: https://github.com/Larsvanderlaan/CDML.
The following subsections provide R and Python code for calibrating inverse propensity weights and outcome regressions. These calibrated models can be directly used within DML frameworks to construct IC-DML estimators of the average treatment effect. Here, we implement isotonic regression using xgboost, allowing control over the maximum tree depth and the minimum number of observations in each constant segment of the isotonic regression fit.
C.1 R Code
# Function: isoreg_with_xgboost
# Purpose: Fits isotonic regression using XGBoost.
# Inputs:
#   - x: A vector or matrix of predictor variables.
#   - y: A vector of response variables.
#   - max_depth: Maximum depth of the trees in XGBoost (default = 15).
#   - min_child_weight: Minimum sum of instance weights (Hessian)
#        needed in a child node (default = 20).
# Returns:
#   - A function that takes a new predictor variable x
#     and returns the model’s predicted values.
isoreg_with_xgboost <- function(x, y, max_depth = 15, min_child_weight = 20) {
  # Create an XGBoost DMatrix object from the data
  data <- xgboost::xgb.DMatrix(data = as.matrix(x), label = as.vector(y))
  # Set parameters for the monotonic XGBoost model
  params <- list(max_depth = max_depth,
                 min_child_weight = min_child_weight,
                 monotone_constraints = 1,  # Enforce monotonic increase
                 eta = 1, gamma = 0,
                 lambda = 0)
  # Train the model with one boosting round
  iso_fit <- xgboost::xgb.train(params = params,
                                data = data,
                                nrounds = 1)
  # Prediction function for new data
  fun <- function(x) {
    data_pred <- xgboost::xgb.DMatrix(data = as.matrix(x))
    pred <- predict(iso_fit, data_pred)
    return(pred)
  return(fun)
# Function: calibrate_inverse_weights
# Purpose: Calibrates inverse weights using isotonic regression
#          with XGBoost for two propensity scores.
# Inputs:
#   - A: Binary indicator variable.
#   - pi1: Estimated propensity score for treatment group A = 1.
#   - pi0: Estimated propensity score for control group A = 0.
# Returns:
#   - A list containing calibrated inverse weights for each group:
#       - alpha1_star: Calibrated inverse weights for A = 1.
#       - alpha0_star: Calibrated inverse weights for A = 0.
calibrate_inverse_weights <- function(A, pi1, pi0) {
  # Calibrate pi1 using monotonic XGBoost
  calibrator_pi1 <- isoreg_with_xgboost(pi1, A)
  pi1_star <- calibrator_pi1(pi1)
  # Set minimum truncation level for treated group
  c1 <- min(pi1_star[A == 1])
  pi1_star = pmax(pi1_star, c0)
  alpha1_star <- 1 / pi1_star
  # Calibrate pi0 using monotonic XGBoost
  calibrator_pi0 <- isoreg_with_xgboost(pi0, 1 - A)
  pi0_star <- calibrator_pi0(pi0)
  # Set minimum truncation level for control group
  c0 <- min(pi0_star[A == 0])
  pi0_star = pmax(pi0_star, c0)
  alpha0_star <- 1 / pi0_star
  # Return calibrated inverse weights for both groups
  return(list(alpha1_star = alpha1_star, alpha0_star = alpha0_star))
# Function: calibrate_outcome_regression
# Purpose: Calibrates outcome regression predictions using isotonic
#          regression with XGBoost.
# Inputs:
#   - Y: Observed outcomes.
#   - mu1: Predicted outcome for treated group.
#   - mu0: Predicted outcome for control group.
#   - A: Binary indicator variable.
# Returns:
#   - A list containing calibrated predictions for each group:
#       - mu1_star: Calibrated predictions for A = 1.
#       - mu0_star: Calibrated predictions for A = 0.
calibrate_outcome_regression <- function(Y, mu1, mu0, A) {
  # Calibrate mu1 using monotonic XGBoost for treated group
  calibrator_mu1 <- isoreg_with_xgboost(mu1[A==1], Y[A==1])
  mu1_star <- calibrator_mu1(mu1)
  # Calibrate mu0 using monotonic XGBoost for control group
  calibrator_mu0 <- isoreg_with_xgboost(mu0[A==0], Y[A==0])
  mu0_star <- calibrator_mu0(mu0)
  # Return calibrated values for both groups
  return(list(mu1_star = mu1_star, mu0_star = mu0_star))
C.2 Python code
import xgboost as xgb
import numpy as np
def isoreg_with_xgboost(x, y, max_depth=15, min_child_weight=20):
    Fits isotonic regression using XGBoost with monotonic constraints to ensure
    non-decreasing predictions as the predictor variable increases.
    Args:
        x (np.array): A vector or matrix of predictor variables.
        y (np.array): A vector of response variables.
        max_depth (int, optional): Maximum depth of the trees in XGBoost.
                                   Default is 15.
        min_child_weight (float, optional): Minimum sum of instance weights
                                            needed in a child node. Default is 20.
    Returns:
        function: A prediction function that takes a new predictor variable x
                  and returns the model’s predicted values.
    Example:
        $>$ x = np.array([[1], [2], [3]])
        $>$ y = np.array([1, 2, 3])
        $>$ model = isoreg_with_xgboost(x, y)
        $>$ model(np.array([[1.5], [2.5]]))
    # Create an XGBoost DMatrix object from the data
    data = xgb.DMatrix(data=np.asarray(x), label=np.asarray(y))
    # Set parameters for the monotonic XGBoost model
    params = {
        ’max_depth’: max_depth,
        ’min_child_weight’: min_child_weight,
        ’monotone_constraints’: "(1)",  # Enforce monotonic increase
        ’eta’: 1,
        ’gamma’: 0,
        ’lambda’: 0
    # Train the model with one boosting round
    iso_fit = xgb.train(params=params, dtrain=data, num_boost_round=1)
    # Prediction function for new data
    def predict_fn(x):
        Predicts output for new input data using the trained isotonic regression model.
        Args:
            x (np.array): New predictor variables as a vector or matrix.
        Returns:
            np.array: Predicted values.
        data_pred = xgb.DMatrix(data=np.asarray(x))
        pred = iso_fit.predict(data_pred)
        return pred
    return predict_fn
def calibrate_inverse_weights(A, pi1, pi0):
    Calibrates inverse weights using isotonic regression with XGBoost for two propensity scores.
    Args:
        A (np.array): Binary indicator variable.
        pi1 (np.array): Propensity score for treatment group (A = 1).
        pi0 (np.array): Propensity score for control group (A = 0).
    Returns:
        dict: Contains calibrated inverse weights:
              - alpha1_star: Inverse weights for A = 1.
              - alpha0_star: Inverse weights for A = 0.
    calibrator_pi1 = isoreg_with_xgboost(pi1, A)
    pi1_star = calibrator_pi1(pi1)
    c1 = np.min(pi1_star[A == 1])
    pi1_star = np.maximum(pi1_star, c1)
    alpha1_star = 1 / pi1_star
    calibrator_pi0 = isoreg_with_xgboost(pi0, 1 - A)
    pi0_star = calibrator_pi0(pi0)
    c0 = np.min(pi0_star[A == 0])
    pi0_star = np.maximum(pi0_star, c0)
    alpha0_star = 1 / pi0_star
    return {’alpha1_star’: alpha1_star, ’alpha0_star’: alpha0_star}
def calibrate_outcome_regression(Y, mu1, mu0, A):
    Calibrates outcome regression predictions using isotonic regression with XGBoost.
    Args:
        Y (np.array): Observed outcomes.
        mu1 (np.array): Predicted outcome for the treated group (A = 1).
        mu0 (np.array): Predicted outcome for the control group (A = 0).
        A (np.array): Binary treatment indicator (1 for treatment, 0 for control).
    Returns:
        dict: Calibrated predictions for each group:
              - ’mu1_star’: Calibrated predictions for A = 1.
              - ’mu0_star’: Calibrated predictions for A = 0.
    # Calibrate mu1 using treated group data (A = 1)
    calibrator_mu1 = isoreg_with_xgboost(mu1[A == 1], Y[A == 1])
    mu1_star = calibrator_mu1(mu1)  # Apply calibrator to all mu1 values
    # Calibrate mu0 using control group data (A = 0)
    calibrator_mu0 = isoreg_with_xgboost(mu0[A == 0], Y[A == 0])
    mu0_star = calibrator_mu0(mu0)  # Apply calibrator to all mu0 values
    return {’mu1_star’: mu1_star, ’mu0_star’: mu0_star}
Appendix D Proofs
D.1 Notation
Let $J\in\mathbb{N}$ denote a fixed number of cross-fitting splits. Let $\mathcal{D}_{n}^{1},\mathcal{D}_{n}^{2},\ldots,\mathcal{D}_{n}^{J}$ be a partition of the available data $\mathcal{D}_{n}$ into $J$ datasets of approximately equal size, corresponding to index sets $\mathcal{I}^{1}_{n},\mathcal{I}^{2}_{n},\ldots,\mathcal{I}^{J}_{n}$. For each $i\in[n]$, let $j(i)\in[J]$ be the index of the data fold containing observation $i$, so that $i\in\mathcal{I}_{n}^{j(i)}$. For ease of presentation, we will use the following cross-fitting notation. For a collection of fold-specific functions $\nu_{n,\diamond}:=\{\nu_{n,j}:j\in[J]\}$ that may depend on $n$, we use the fold-averaged empirical process notation $\overline{P}_{0}\nu_{n,\diamond}:=\frac{1}{J}\sum_{j=1}^{J}P_{0}\nu_{n,j}$ and $\overline{P}_{n}:=\frac{1}{J}\sum_{j=1}^{J}P_{n,j}\nu_{n,j}$. This notation allows us to view the collections of calibrated cross-fitted estimators $\alpha_{n,\diamond}^{*}:=\{\alpha_{n,j}^{*}:j\in[J]\}$ and $\mu_{n,\diamond}^{*}:=\{\mu_{n,j}^{*}:j\in[J]\}$ as extended functions defined on $\mathcal{A}\times\mathcal{W}\times[J]$.
For a uniformly bounded function class $\mathcal{F}$, let $N(\epsilon,\mathcal{F},L_{2}(P))$ denote the $\epsilon-$covering number (van der Vaart and Wellner, 1996) of $\mathcal{F}$ relative to the $L^{2}(P)$ metric and define the uniform entropy integral of $\mathcal{F}$ by
$\mathcal{J}(\delta,\mathcal{F}):=\int_{0}^{\delta}\sup_{Q}\sqrt{\log N(\epsilon,\mathcal{F},L_{2}(Q))}\,d\epsilon\ ,$
where the supremum is taken over all discrete probability distributions $Q$. In contrast to the definition provided in van der Vaart and Wellner (1996), we do not define the uniform entropy integral relative to an envelope function for the function class $\mathcal{F}$. We can do this since all function classes we consider are uniformly bounded. Thus, any uniformly bounded envelope function will only change the uniform entropy integral as defined in van der Vaart and Wellner (1996) by a constant.
Let $k(n)=o(\sqrt{n})$ be the upper bound on the number of jump points of $f_{n}$ and $g_{n}$ in C4. Since C4 states that $k(n)$ is an upper bound with probability tending to one, we can, without loss of generality, assume that $k(n)$ is an upper bound. Let $M/2<\infty$ be an upper bound on the nuisance functions and estimators in C3. Let $\mathcal{F}_{\text{iso}}\subset\{f:[-M,M]\rightarrow\mathbb{R}|\text{$f$ is monotone nondecreasing}\}$ consist of all univariate, isotonic functions with supremum norm smaller than $M<\infty$. Let $\mathcal{F}_{n}\subset\{f:[-M,M]\rightarrow\mathbb{R}|\text{$f$ is piecewise constant with at most $k(n)$ jumps}\}$ consist of all univariate functions with supremum norm smaller than $M<\infty$, that are left-continuous with right-hand limits and have at most $k(n)<\infty$ jump points (i.e., discontinuities). For $j\in[J]$, we define $\mathcal{S}_{n,j}:=\left\{f\circ\mu_{n,j}|f\in\mathcal{F}_{n}\cup\mathcal{F}_{\text{iso}}\right\}$ and $\mathcal{R}_{n,j}:=\left\{g\circ\alpha_{n,j}|g\in\mathcal{F}_{n}\cup\mathcal{F}_{\text{iso}}\right\}$. By C4 and C3, we know $\mu_{n,j}^{*}=f_{n}\circ\mu_{n,j}$ and $s_{n,j}^{*}$ fall in $\mathcal{S}_{n,j}$ (with probability tending to one). Similarly, we know $\alpha_{n,j}^{*}=g_{n}\circ\alpha_{n,j}$ and $r_{n,j}^{*}$ fall in $\mathcal{R}_{n,j}$. Note that both $\mathcal{S}_{n,j}$ and $\mathcal{R}_{n,j}$ are fixed function classes conditional on the $j$-th training set $\mathcal{D}_{n}\backslash\mathcal{C}_{j}$.
For two quantities $x$ and $y$, we use the expression $x\lesssim y$ to mean that $x$ is upper bounded by $y$ times a universal constant that may only depend on global constants that appear in C3.
D.2 Derivation of efficient influence function
Proof of Theorem 1.
We will first show that $(P,\mu)\mapsto\psi_{P}(\mu)$ is total pathwise differentiable in the sense of Appendix A.4. in Luedtke (2024). To do so, we apply Lemma S6 in Luedtke (2024). To do so, we need to show that the total operator $(P,\mu)\mapsto\psi_{P}(\mu)$ is partially differentiable in its first argument and continuously partially differentiable in its second argument at $(P_{0},\mu_{0})$, where these notions of differentiability are defined in Appendix A.4 of Luedtke (2024). The former differentiability holds since, by assumption A3, $P\mapsto\psi_{P}(\mu_{0})$ is pathwise differentiable at $P_{0}$. To establish continuous partial differentiability in the second argument, note that $\mu\mapsto\psi_{P}(\mu)$ is a linear functional and, thus, the partial derivative of $(P,\mu)\mapsto\psi_{P}(\mu)$ in its second argument is simply $\mu\mapsto\psi_{P}(\mu)$. Hence, to establish continuous partial differentiability, it remains to show that $P\mapsto\psi_{P}$ is a continuous functional at $P_{0}$ with the domain equipped with the Hellinger distance and the range equipped with the operator norm. This is true by assumption A2. By Lemma S6 in Luedtke (2024), the total pathwise derivative at $(P_{0},\mu_{0})$ in the direction of $(s,h)\in T_{\mathcal{M}}(P_{0})\times\mathcal{H}$ is then given by $\langle\widetilde{\phi}_{0,\mu_{0}},s\rangle+\langle\alpha_{0},h\rangle$.
Under our assumptions, $P\mapsto\mu_{P}$ is pathwise differentiable at $P_{0}$ in a Hilbert-valued sense by Example 5 of Luedtke and Chung (2024), with pathwise derivative $d\mu_{0}(s)$ in the direction $s\in T_{\mathcal{M}}(P_{0})$ given by $h\mapsto\langle h\{\mathcal{I}_{y}-\mu_{0}\},s\rangle$, where $\mathcal{I}_{y}:o\mapsto y$. As a consequence, by the chain rule, $P\mapsto\psi_{P}(\mu_{P})$ is pathwise differentiable at $P_{0}$ with pathwise derivative given by $\langle\widetilde{\phi}_{0,\mu_{0}},s\rangle+\langle\alpha_{0}\{\mathcal{I}_{y}-\mu_{0}\},s\rangle.$ It follows that the efficient influence function is $\widetilde{\phi}_{0,\mu_{0}}+D_{\mu_{0},\alpha_{0}}$.
D.3 Decomposing the estimation error and cross-product remainder
In this section, we let $\alpha_{n,\diamond}=\{\alpha_{n,j}:j\in[J]\}$ and $\mu_{n,\diamond}=\{\mu_{n,j}:j\in[J]\}$ be arbitrary fold-dependent estimators of $\alpha_{0}$ and $\mu_{0}$ that are potentially uncalibrated. We denote the associated one-step debiased estimator by
$\tau_{n}:=\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(\mu_{n,j})+\frac{1}{n}\sum_{i=1}^{n}\alpha_{n,j(i)}(A_{i},W_{i})\{Y_{i}-\mu_{n,j(i)}(A_{i},W_{i})\}.$
The following expansion of the estimation error $\tau_{n}-\tau_{0}$ is a standard starting point in the analysis of doubly robust estimators.
Lemma 3 (Standard doubly robust bias expansion).
It holds that
$\tau_{n}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}}+D_{\mu_{n,\diamond},\alpha_{n,\diamond}}\right\}+\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}+Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}).$
where $Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}):=\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(\mu_{n,j})-\psi_{0}(\mu_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{0,\mu_{n,j}}\right\}$.
Proof.
We can write
$\frac{1}{J}\sum_{j=1}^{J}\{\psi_{n,j}(\mu_{n,j})-\psi_{0}(\mu_{n,j})\}=(\overline{P}_{n}-\overline{P}_{0})\widetilde{\phi}_{0,\mu_{n,\diamond}}+Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}),$
where $Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}):=\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(\mu_{n,j})-\psi_{0}(\mu_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{0,\mu_{n,j}}\right\}$. The estimation error of $\tau_{n}$ can be decomposed as:
$\tau_{n}-\tau_{0}$ $=\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(\mu_{n,j})+\overline{P}_{n}D_{\mu_{n,\diamond},\alpha_{n,\diamond}}-\psi_{0}(\mu_{0})$
$=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}}+D_{\mu_{n,\diamond},\alpha_{n,\diamond}}\right\}+\overline{P}_{0}D_{\mu_{n,\diamond},\alpha_{n,\diamond}}+\frac{1}{J}\sum_{j=1}^{J}\{\psi_{0}(\mu_{n,j})-\psi_{0}(\mu_{0})\}+Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}).$
By the Reisz representation $\psi_{0}(\cdot)=\langle\alpha_{0},\cdot\rangle_{P_{0}}$, it holds that $\psi_{0}(\mu_{n,j})-\psi_{0}(\mu_{0})=\langle\alpha_{0},\mu_{n,j}-\mu_{0}\rangle_{P_{0}}$. Moreover, by the law of iterated expectations, we have
[…]
$=\langle\alpha_{n,\diamond},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}.$
Hence, $\overline{P}_{0}D_{\mu_{n,\diamond},\alpha_{n,\diamond}}+\frac{1}{J}\sum_{j=1}^{J}\{\psi_{0}(\mu_{n,j})-\psi_{0}(\mu_{0})\}=\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}.$ Therefore,
$\tau_{n}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}}+D_{\mu_{n,\diamond},\alpha_{n,\diamond}}\right\}+\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}+Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}).$
The next lemma establishes the leading order behavior of the cross-product remainder term in both the outcome- and Riesz-favorable settings. We introduce the following notation. For a fixed vector-valued function $v:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}^{k}$ with $k\in\{1,2\}$, we define the conditional expectation projection operator $\Pi_{v}:L^{2}(P_{0,A,X})\rightarrow L^{2}(P_{0,A,X})$ through the mapping $f\mapsto\Pi_{v}f:=(a,w)\mapsto\mathbb{E}_{0}[f(A,W)|v(A,W)=v(a,w)]$. For brevity, we denote the orthogonal projections of the error terms $\alpha_{0}-\alpha_{n,j}$ and $\mu_{0}-\mu_{n,j}$ onto $\mu_{n,j}$ and $\alpha_{n,j}$, respectively, as $s_{n,j}:=\Pi_{(\mu_{n,j})}\{\alpha_{0}-\alpha_{n}\}$ and $r_{n,j}:=\Pi_{(\alpha_{n,j})}\{\mu_{0}-\mu_{n,j}\}$. Denote $A_{n,j}:z\mapsto s_{n,j}(a,w)\{y-\mu_{n,j}(a,w)\}$ and define $B_{n,j}:=\psi_{0}(r_{0})+\widetilde{\phi}_{r_{n,j}}-r_{n,j}\alpha_{n,j}$.
Lemma 4 (Decompositions of cross-product remainder).
We have the outcome-favorable decomposition:
$\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})A_{n,\diamond}-\overline{P}_{n}A_{n,\diamond}+O_{p}\left(\|(\Pi_{(\mu_{n,\diamond},\mu_{0})}-\Pi_{\mu_{n,\diamond}})(\alpha_{n,\diamond}-\alpha_{0})\|_{\overline{P}_{0}}\|\mu_{0}-\mu_{n,\diamond}\|_{\overline{P}_{0}}\right),$
Similarly, we have the Riesz-favorable decomposition:
$\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})B_{n,\diamond}-\left\{\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(r_{n,j})-\overline{P}_{n}r_{n,\diamond}\alpha_{n,\diamond}\right\}$
$\quad+\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}}\right\}$
$\quad+O_{p}(\|\alpha_{n,\diamond}-\alpha_{0}\|_{\overline{P}_{0}}\|(\Pi_{(\alpha_{n,\diamond},\alpha_{0})}-\Pi_{\alpha_{n,\diamond}})(\mu_{0}-\mu_{n,\diamond})\|_{\overline{P}_{0}},$
Proof.
By definition, $\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}=\frac{1}{J}\sum_{j=1}^{J}\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$. Next, by the law of total expectation, we have, for each $j\in[J]$, that
$\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$ $=\langle\Pi_{(\mu_{n,j},\mu_{0})}\{\alpha_{n,j}-\alpha_{0}\},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$
$=\langle\Pi_{\mu_{n,j}}\{\alpha_{n,j}-\alpha_{0}\},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}+\langle(\Pi_{(\mu_{n,j},\mu_{0})}-\Pi_{\mu_{n,j}})(\alpha_{n,j}-\alpha_{0}),\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$
$=\langle s_{n,j},\mu_{n,j}-\mu_{0}\rangle_{P_{0}}+\langle(\Pi_{(\mu_{n,j},\mu_{0})}-\Pi_{\mu_{n,j}})(\alpha_{n,j}-\alpha_{0}),\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$
$=\langle s_{n,j},\mu_{n,j}-\mu_{0}\rangle_{P_{0}}+O_{p}\left(\|(\Pi_{(\mu_{n,j},\mu_{0})}-\Pi_{\mu_{n,j}})(\alpha_{n,j}-\alpha_{0})\|\|\mu_{0}-\mu_{n,j}\|_{P_{0}}\right),$
where the final equality follows from the Cauchy-Schwarz inequality. The leading term on the right-hand side of the above display can be expressed as
[…]
$=(P_{n,j}-P_{0})A_{n,j}-P_{n,j}A_{n,j}.$
Hence,
$\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$ $=(P_{n,j}-P_{0})A_{n,j}-P_{n,j}A_{n,j}+O_{p}\left(\|(\Pi_{(\mu_{n,j},\mu_{0})}-\Pi_{\mu_{n,j}})(\alpha_{n,j}-\alpha_{0})\|\|\mu_{0}-\mu_{n,j}\|_{P_{0}}\right),$
Averaging the above display over $j\in[J]$, the first equality then follows.
We now derive the second equality. Similarly, by the law of total expectation, we have, for each $j\in[J]$, that
$\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$ $=\langle\alpha_{n,j}-\alpha_{0},\Pi_{(\alpha_{n,j},\alpha_{0})}(\mu_{0}-\mu_{n,j})\rangle_{P_{0}}$
$=\langle\alpha_{n,j}-\alpha_{0},\Pi_{\alpha_{n,j}}\{\mu_{0}-\mu_{n,j}\}\rangle_{P_{0}}+\langle\alpha_{n,j}-\alpha_{0},(\Pi_{(\alpha_{n,j},\alpha_{0})}-\Pi_{\alpha_{n,j})})(\mu_{0}-\mu_{n,j})\rangle_{P_{0}}$
$=\langle\alpha_{n,j}-\alpha_{0},r_{n,j}\rangle_{P_{0}}+O_{p}(\|\alpha_{n,j}-\alpha_{0}\|\|(\Pi_{(\alpha_{n,j},\alpha_{0})}-\Pi_{\alpha_{n,j})})(\mu_{0}-\mu_{n,j})\|,$
where $\langle\alpha_{n,j}-\alpha_{0},r_{n,j}\rangle_{P_{0}}=\langle r_{n,j},\alpha_{n,j}\rangle_{P_{0}}-\psi_{0}(r_{n,j})$ by the Cauchy-Schwarz inequality. The leading term on the right-hand side of the above display can be expressed as
$\langle r_{n,j},\alpha_{n,j}\rangle_{P_{0}}-\psi_{0}(r_{n,j})$ $=P_{0}r_{n,j}\alpha_{n,j}-\psi_{0}(r_{n,j})$
$=P_{n,j}r_{n,j}\alpha_{n,j}-\psi_{n,j}(r_{n,j})+\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})+(P_{0}-P_{n,j})\left\{r_{n,j}\alpha_{n,j}\right\}$
$=P_{n,j}r_{n,j}\alpha_{n,j}-\psi_{n,j}(r_{n,j})+(P_{0}-P_{n,j})\left\{r_{n,j}\alpha_{n,j}-\widetilde{\phi}_{r_{n,j}}\right\}$
$\quad+\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}}$
$=P_{n,j}r_{n,j}\alpha_{n,j}-\psi_{n,j}(r_{n,j})+(P_{n,j}-P_{0})\left\{\widetilde{\phi}_{r_{n,j}}-r_{n,j}\alpha_{n,j}\right\}$
$\quad+\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}}.$
We note that $(P_{n,j}-P_{0})B_{n,j}=(P_{n,j}-P_{0})\left\{\widetilde{\phi}_{r_{n,j}}-r_{n,j}\alpha_{n,j}\right\}$ since $\psi_{0}(r_{0})$ is a constant and, therefore, satisfies $(P_{n,j}-P_{0})(\psi_{0}(r_{0}))=0$. Hence,
$\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$ $=P_{n,j}r_{n,j}\alpha_{n,j}-\psi_{n,j}(r_{n,j})+(P_{n,j}-P_{0})\left\{\widetilde{\phi}_{r_{n,j}}-r_{n,j}\alpha_{n,j}\right\}$
$\quad+\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}}$
$\quad+O_{p}(\|\alpha_{n,j}-\alpha_{0}\|\|(\Pi_{(\alpha_{n,j},\alpha_{0})}-\Pi_{\alpha_{n,j})})(\mu_{0}-\mu_{n,j})\|,$
Averaging the above displays over $j\in[J]$, the second equality then follows.
Proof of Lemma 1.
Denote $X:=(W,A)$. Let $g:\mathbb{R}^{2}\rightarrow\mathbb{R}$ be a Lipschitz continuous function with constant $L>0$. By Lipschitz continuity, we have that
$\left|g(\mu_{n}(x),\mu_{0}(x))-E[g(\mu_{n}(X),\mu_{0}(X))|\mu_{n}(X)=\mu_{n}(x)]\right|$ $=\left|E[g(\mu_{n}(x),\mu_{0}(x))-g(\mu_{n}(x),\mu_{0}(X))|\mu_{n}(X)=\mu_{n}(x)]\right|$
$\leq E[\left|g(\mu_{n}(x),\mu_{0}(x))-g(\mu_{n}(x),\mu_{0}(X))\right||\mu_{n}(X)=\mu_{n}(x)]$
$\leq LE[\left|\mu_{0}(x)-\mu_{0}(X)\right||\mu_{n}(X)=\mu_{n}(x)].$
On the event $\{\mu_{n}(X)=\mu_{n}(x)\}$, we know
$|\mu_{0}(x)-\mu_{0}(X)|$ $\leq|\mu_{0}(x)-\mu_{n}(x)|+|\mu_{n}(x)-\mu_{n}(X)|+|\mu_{0}(X)-\mu_{n}(X)|$
$\leq|\mu_{0}(x)-\mu_{n}(x)|+|\mu_{0}(X)-\mu_{n}(X)|.$
Therefore,
$\left|g(\mu_{n}(x),\mu_{0}(x))-E[g(\mu_{n}(X),\mu_{0}(X))|\mu_{n}(X)=\mu_{n}(x)]\right|$ $\lesssim E[\left|\mu_{0}(x)-\mu_{0}(X)\right|\mid\mu_{n}(X)=\mu_{n}(x)]$
$\lesssim E|\mu_{0}(x)-\mu_{n}(x)|+E[|\mu_{0}(X)-\mu_{n}(X)|\mid\mu_{n}(X)=\mu_{n}(x)].$
Now, for some function $f\in L^{2}(P_{X})$, suppose that $(\mu_{n}(x),\mu_{0}(x))\mapsto(\Pi_{\mu_{n},\mu_{0}}f)(x)$ is Lipschitz continuous. Then, defining $g:(\widehat{m},m)\mapsto E_{0}[f(X)|\mu_{n}(X)=\widehat{m},\mu_{0}(X)=m,\mathcal{D}_{n}]$ and noting by the law of iterated expectation that $\Pi_{\mu_{n}}\Pi_{\mu_{n},\mu_{0}}f=\Pi_{\mu_{n}}f$, we obtain the following pointwise error bound:
$\left|\Pi_{\mu_{n},\mu_{0}}f-\Pi_{\mu_{n}}f\right|\lesssim\left|\mu_{n}-\mu_{0}\right|+\Pi_{\mu_{n}}(|\mu_{n}-\mu_{0}|).$
Since $\|\Pi_{\mu_{n}}(|\mu_{n}-\mu_{0}|)\|_{L^{2}(P_{0})}\leq\|\mu_{n}-\mu_{0}\|_{L^{2}(P_{0})}$ by the properties of projections, it follows that
$\|\Pi_{\mu_{n},\mu_{0}}f-\Pi_{\mu_{n}}f\|_{L^{2}(P)}\lesssim\|\mu_{n}-\mu_{0}\|_{L^{2}(P)}.$
Taking $f$ to be $\Delta_{\alpha_{n}}$, we conclude that Condition B1B1i implies Condition B1enumiB1enumii. By an identical argument, Condition B1B1ii implies Condition B1enumiB1enumiii. The result then follows.
D.4 Maximal inequalities for empirical processes
Lemma 5.
For $j\in[J]$, it holds that $\mathcal{J}(\delta,\mathcal{R}_{n,j})+\mathcal{J}(\delta,\mathcal{S}_{n,j})\lesssim\sqrt{\delta}\left(1+\sqrt{k(n)\delta\log(1/\delta)}\right)$.
Proof.
We establish this explicitly for $\mathcal{S}_{n,j}$ below, as the result for $\mathcal{R}_{n,j}$ follows from an identical argument. To this end, we have
$\mathcal{J}(\delta,\mathcal{S}_{n,j})$ $=\int_{0}^{\delta}\sup_{Q}\sqrt{N(\varepsilon,\mathcal{S}_{n,j},\norm{\,\cdot\,}_{Q})}\,d\varepsilon=\int_{0}^{\delta}\sup_{Q}\sqrt{N(\varepsilon,\mathcal{F}_{n}\cup\mathcal{F}_{\text{iso}},\norm{\,\cdot\,}_{Q\circ\mu_{n,j}^{-1}})}\,d\varepsilon$
$\leq\int_{0}^{\delta}\sup_{Q}\sqrt{N(\varepsilon,\mathcal{F}_{n},\norm{\,\cdot\,}_{Q\circ\mu_{n,j}^{-1}})}\,d\varepsilon+\int_{0}^{\delta}\sup_{Q}\sqrt{N(\varepsilon,\mathcal{F}_{\text{iso}},\norm{\,\cdot\,}_{Q\circ\mu_{n,j}^{-1}})}\,d\varepsilon$
$\leq\mathcal{J}(\delta,\mathcal{F}_{n})+\mathcal{J}(\delta,\mathcal{F}_{iso}),$
where $Q\circ\mu_{n,j}^{-1}$ is the push-forward probability measure for the random variable $\mu_{n,j}(W)$ conditional on $\mathcal{E}_{n}^{(j)}=\mathcal{D}_{n}\backslash\mathcal{C}^{(j)}$. Since $\mathcal{J}(\delta,\mathcal{F}_{n})$ consists of univariate functions with at most $k(n)$ constant segments, it has VC-subgraph dimension $O(k(n))$. As a consequence, $\mathcal{J}(\delta,\mathcal{F}_{n})\lesssim\delta\sqrt{k(n)\log(1/\delta)}$, as desired. Moreover, by ???, we have $\mathcal{J}(\delta,\mathcal{F}_{iso})\lesssim\sqrt{\delta}$. ∎
Lemma 6.
For each $j\in[J]$, let $\mathcal{H}_{n,j}=\mathcal{F}_{n,j}\times\mathcal{F}_{n,j}\times\mathcal{F}_{n,j}\times\mathcal{F}_{n,j}$ for a function class $\mathcal{F}_{n,j}$ that is deterministic conditional on the $j$-th training set $\mathcal{D}_{n}\backslash\mathcal{C}^{(j)}$. Assume that $\mathcal{J}(\delta,\mathcal{F}_{n,j})\leq t_{n}(\delta)$ for some map $\delta\mapsto t_{n}(\delta)$. For each $j\in[J]$, let $T_{n,j}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ be a bounded Lipchitz function that is deterministic conditional on $\mathcal{D}_{n}\backslash\mathcal{C}^{(j)}$. Then, for all $\delta>0$,
$E_{0}^{n}\left[\sup_{h_{\diamond}\in\mathcal{H}_{n,\diamond}:\|h_{\diamond}\|_{\overline{P}_{0}}\leq\delta}\left(\overline{P}_{n}-\overline{P}_{0}\right)T_{n,\diamond}(h_{\diamond})\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right).$
Furthermore, for any $\sqrt{n}\delta^{2}>t_{n}(\delta)$,
$E_{0}^{n}\left[\sup_{h_{\diamond}\in\mathcal{H}_{n,\diamond}:\|h_{\diamond}\|_{\overline{P}_{0}}\leq\delta}\left(\overline{P}_{n}^{\#}-\overline{P}_{n}\right)T_{n,\diamond}(h_{\diamond})\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right).$
Proof.
To establish the first bound, by the triangle inequality, it suffices to show, for each $j\in[J]$, that
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|h\|\leq\delta}\left(P_{n,j}-P_{0}\right)T_{n,j}(h)\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right).$
By Theorem 2.10.20 of van der Vaart and Wellner (1996) on preservation of entropy integrals under Lipschitz transformation and Lemma 5, we have that $\mathcal{H}_{n,j}$ is a uniformly bounded function class with uniform entropy integral satisfying $\mathcal{J}(\delta,\mathcal{H}_{n,j})\lesssim 4\mathcal{J}(\delta,\mathcal{F}_{n,j})\lesssim t_{n}(\delta)$. Hence, by Theorem 2.1 of van der Vaart and Wellner (2011),
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|h\|\leq\delta}\left(P_{n,j}-P_{0}\right)T_{n,j}(h)\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right).$
Next, noting that $P_{n,j}^{\#}$ is the empirical distribution of $O(n/J)$ i.i.d. samples from $P_{n,j}$, we can apply the same argument conditional on $P_{n,j}$ with $\delta_{n}:=\sup_{h\in\mathcal{H}_{n,j}:\|h\|_{P_{0}}\leq\delta}\|h\|_{P_{n,j}}$ to find:
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|h\|_{P_{0}}\leq\delta}\left(P_{n,j}^{\#}-P_{n,j}\right)T_{n,j}(h)|P_{n,j}\right]$ $\lesssim n^{-\frac{1}{2}}\{\mathcal{J}(\delta_{n},\mathcal{S}_{n,j})+\mathcal{J}(\delta_{n},\mathcal{R}_{n,j})\}\left(1+\frac{\mathcal{J}(\delta_{n},\mathcal{S}_{n,j})+\mathcal{J}(\delta_{n},\mathcal{R}_{n,j})}{\sqrt{n}\delta_{n}^{2}}\right)$
$\lesssim n^{-\frac{1}{2}}t_{n}(\delta_{n})\left(1+\frac{t_{n}(\delta_{n})}{\sqrt{n}\delta_{n}^{2}}\right)$
We can show that $\varepsilon\mapsto t_{n}(\sqrt{\varepsilon})\left(1+\frac{t_{n}(\sqrt{\varepsilon})}{\sqrt{n}\varepsilon}\right)$ is a concave function. Thus, taking the expectation over both sides of the previous display and applying Jenson’s inequality, we find for $\widetilde{\delta}_{n}:=\sqrt{E\delta_{n}^{2}}$ that
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|h\|_{P_{0}}\leq\delta}\left(P_{n,j}^{\#}-P_{n,j}\right)T_{n,j}(h)\right]\lesssim n^{-\frac{1}{2}}t_{n}(\widetilde{\delta}_{n})\left(1+\frac{t_{n}(\widetilde{\delta}_{n})}{\sqrt{n}\widetilde{\delta}_{n}^{2}}\right)$
Moreover, by Theorem 2.2 of van de Geer (2014), we have $\widetilde{\delta}_{n}^{2}=E\delta_{n}^{2}\lesssim\delta^{2}$ for any $\delta>0$ with $\sqrt{n}\delta^{2}>t_{n}(\delta)$. Combining this with the previous display, we obtain
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|h\|_{P_{0}}\leq\delta}\left(P_{n,j}^{\#}-P_{n,j}\right)T_{n,j}(h)\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right),$
for any $\delta^{2}>n^{-\frac{1}{2}}t_{n}(\delta)$. ∎
Lemma 7.
For each $j\in[J]$, let $T_{n,j}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ be a Lipschitz continuous function that is deterministic conditional on $\mathcal{D}_{n}\backslash\mathcal{C}^{(j)}$. Moreover, let $\rho_{n,j}:=T_{n,j}(\mu_{n,j}^{*},s_{n,j}^{*},\alpha_{n,j}^{*},r_{n,j}^{*})$ be a random function by applying the Lipschitz transformation $T_{n,j}$ to the estimated nuisance functions $(\mu_{n,j}^{*},s_{n,j}^{*},\alpha_{n,j}^{*},r_{n,j}^{*})$. Suppose $\|\rho_{n,\diamond}\|_{\overline{P}_{0}}=O_{P}(\varepsilon_{n})$ for a deterministic sequence $\varepsilon_{n}\rightarrow 0$. Then, $\left(\overline{P}_{n}-\overline{P}_{0}\right)\rho_{n,\diamond}=o_{p}(n^{-\frac{1}{2}})$. Moreover, when $\overline{P}_{n},\alpha_{n,j}^{*}$, $\mu_{n,j}^{*}$, $\rho_{n,j}$ are replaced with the bootstrap estimators $\overline{P}_{n}^{\#},\alpha_{n,j}^{*\#}$, $\mu_{n,j}^{*\#}$, and $\rho_{n,j}^{\#}$, we also have $\left(\overline{P}_{n}^{\#}-\overline{P}_{n}\right)\rho_{n,\diamond}^{\#}=o_{p}(n^{-\frac{1}{2}})$ whenever $\|\rho_{n,\diamond}^{\#}\|_{\overline{P}_{0}}=O_{P}(\varepsilon_{n})$.
Proof.
Since $\|\rho_{n,\diamond}\|_{\overline{P}_{0}}=O_{P}(\varepsilon_{n})$ for some $\varepsilon_{n}\rightarrow 0$, there exists a deterministic sequence $\delta_{n}\rightarrow 0$ such that $\|\rho_{n,\diamond}\|_{\overline{P}_{0}}\lesssim\delta_{n}$ with probability tending to one. Let $A_{n}$ denote the event that $\|\rho_{n,\diamond}\|_{\overline{P}_{0}}\lesssim\delta_{n}$ holds. On this event, we have that
$\left(\overline{P}_{n}-\overline{P}_{0}\right)\rho_{n,\diamond}\leq\sup_{h_{\diamond}\in\mathcal{H}_{n,\diamond}:\|h_{\diamond}\|_{\overline{P}_{0}}\leq\delta_{n}}\left(\overline{P}_{n}-\overline{P}_{0}\right)T_{n,\diamond}(h_{\diamond}).$
Moreover, Lemma 6 applied on $A_{n}$ implies
$E_{0}^{n}\left[1(A_{n})\sup_{h_{\diamond}\in\mathcal{H}_{n,\diamond}:\|h_{\diamond}\|_{\overline{P}_{0}}\leq\delta_{n}}\left(\overline{P}_{n}-\overline{P}_{0}\right)T_{n,\diamond}(h_{\diamond})\right]=o(n^{-\frac{1}{2}}).$
Hence, by Markov’s inequality, $1(A_{n})\left(\overline{P}_{n}-\overline{P}_{0}\right)\rho_{n,\diamond}=o(n^{-\frac{1}{2}})$. Since $A_{n}$ occurs with probability tending to one, we also have $\left(\overline{P}_{n}-\overline{P}_{0}\right)\rho_{n,\diamond}=o_{p}(n^{-\frac{1}{2}})$, as desired. The statement for the bootstrap empirical process follows from an identical argument, using the second part of Lemma 6, noting that $P_{n,j}^{\#}$ is the empirical distribution of $\approx n/J$ i.i.d. samples from $P_{n,j}$. ∎
D.5 Convergence Rates for Isotonic Calibrated Nuisance Estimators
Let $\alpha_{n,\diamond}^{*}:=\{\alpha_{n,j}^{*}:j\in[J]\}$ and $\mu_{n,\diamond}^{*}:=\{\mu_{n,j}^{*}:j\in[J]\}$ be the isotonic calibrated nuisance estimators obtained via Algorithm 1. Let $\overline{\alpha}_{0}$ and $\overline{\mu}_{0}$ be their respective potentially misspecified limits, as in C2. In this section, we determine the mean square error rates for the calibrated nuisance estimators $\alpha_{n,j}^{*}$ and $\mu_{n,j}^{*}$ with respect to $\alpha_{0}$ and $\mu_{0}$. We focus our convergence rate results on the case where the linear functional $\psi_{0}$ takes the form $\mu\mapsto E_{0}\{m(A,W,\mu)\}$ for a known map $m:\mathcal{A}\times\mathcal{W}\times\mathcal{H}\rightarrow\mathbb{R}$. Additionally, we assume that the linear functional estimator $\psi_{n}$ is taken as the empirical plug-in estimator $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(A_{i},W_{i},\mu)$. Our convergence rate results can be extended to general estimators $\psi_{n}$, provided that certain regularity conditions hold for the second-order remainder.
We begin by providing a result for a more general setting that holds for general loss functions. Consider a loss function $\ell(x,y,\nu):=\nu^{2}(x)-m(x,y,\nu)$ for some mapping $m$. We assume $|m(x,y,\nu_{1})-m(x,y,\nu_{2})|\leq L|\nu_{1}(x)-\nu_{2}(x)|$ for some $L>0$, and that, for any bounded, convex function class $\mathcal{F}$, the risk is locally quadratic around its minimizer in that $\|\nu-\nu_{0}(\mathcal{F})\|^{2}_{P_{0}}\lesssim P_{0}\{\ell(\cdot,\nu)-\ell(\cdot,\nu_{0}(\mathcal{F})\}$, where $\nu_{0}(\mathcal{F}):=\operatorname*{argmin}_{\nu\in\mathcal{F}}\|\ell(X,Y,\nu)\|_{P_{0}}$. Let $\{\nu_{n,j}:j\in[J]\}$ be cross-fitted estimators of the population risk minimizer $\nu_{0}:=\operatorname*{argmin}_{\nu\in L^{2}(P_{0})}P_{0}\ell(\cdot,\nu)$. Define $\nu_{n,\diamond}^{*}:=h_{n}\circ\nu_{n,\diamond}$ and $\nu_{n,\diamond}^{*\#}:=h_{n}^{\#}\circ\nu_{n,\diamond}$ where we define the isotonic regression solutions:
$h_{n}$ $:=\operatorname*{argmin}_{h\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\ell(X_{i},Y_{i},h\circ\nu_{n,j(i)});$
$h_{n}^{\#}$ $:=\operatorname*{argmin}_{h\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\ell(X_{i}^{\#},Y_{i}^{\#},h\circ\nu_{n,j(i)});$
$h_{n,0}$ $:=\operatorname*{argmin}_{h\in\mathcal{F}_{iso}}\|h\circ\nu_{n,\diamond}-\nu_{0}\|_{\overline{P}_{0}}.$
Theorem 7.
Suppose that $h_{n}$, $h_{n}^{\#}$, and $h_{n,0}$ fall in a uniformly bounded subset of $\mathcal{F}_{iso}$ with probability tending to one. Then, $\|h_{n}\circ\nu_{n,\diamond}-h_{n,0}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}=O_{p}(n^{-1/3})$. Hence,
$\|h_{n}\circ\nu_{n,\diamond}-\nu_{0}\|_{\overline{P}_{0}}$ $\leq\inf_{h\in\mathcal{F}_{iso}}\|h\circ\nu_{n,\diamond}-\nu_{0}\|+O_{p}(n^{-1/3})$
Moreover, the bootstrap calibrated nuisance estimators satisfy, for any function $\overline{\nu}_{0}\in L^{2}(P_{0})$ and as $n\rightarrow\infty$, the following root mean square error bounds:
$\|h_{n}^{\#}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}$ $\leq\|h_{n}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3}).$
Proof.
Let ~$\widetilde{\mathcal{F}}_{iso}$ be a uniformly bounded subset of $\mathcal{F}_{iso}$ that contains $h_{n}$, $h_{n}^{\#}$, and $h_{n,0}$ with probability tending to one. Define $\mathcal{F}_{n,j}:=\{f\circ\nu_{n,j}:f\in\widetilde{\mathcal{F}}_{iso}\}$. Since $h_{n}$ and $h_{n,0}$ fall in $\mathcal{F}_{iso}$ with probability tending to one, it holds that $\nu_{n,j}^{*}$ and $\nu_{n,0}$ fall in $\mathcal{F}_{n,j}$ with probability tending to one. The property $|m(x,y,\nu_{1})-m(x,y,\nu_{2})|\leq L|\nu_{1}(x)-\nu_{2}(x)|$ for some $L>0$ implies that $|\ell(x,y,\nu_{1})-\ell(x,y,\nu_{2})|\leq|\nu_{1}(x,y)-\nu_{2}(x,y)|\left\{\|\nu_{1}+\nu_{2}\|_{\infty}+L\right\}$.
We prove the convergence rates for the bootstrap estimators, as the rates for the non-bootstrap estimators follow along similar lines. To this end, for $j\in[J]$, we define the random rate $\delta_{n,j}^{\#}:=\|h_{n}^{\#}\circ\nu_{n,j}-h_{n,0}\circ\nu_{n,j}\|_{P_{0}}$ and let $(\delta_{n}^{\#})^{2}:=\frac{1}{J}\sum_{j=1}^{J}(\delta_{n,j}^{\#})^{2}$. By convexity of $\mathcal{F}_{iso}$, the locally quadratic assumption on $\ell$, and $h_{n,0}$ being a population risk minimizer, we have that
$(\delta_{n}^{\#})^{2}\lesssim\overline{P}_{0}\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\overline{P}_{0}\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond}).$
Furthermore, since $h_{n}^{\#}\circ\nu_{n}$ is the bootstrap empirical risk minimizer, it holds that
$\overline{P}_{n}^{\#}\left\{\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})\right\}\leq 0.$
Combining the above two displays, we obtain that:
$\overline{P}_{0}\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\overline{P}_{0}\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})$ $=(\overline{P}_{0}-\overline{P}_{n}^{\#})\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})+\overline{P}_{n}^{\#}\left\{\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})\right\}$
$\quad-(\overline{P}_{0}-\overline{P}_{n}^{\#})\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})$
$\leq(\overline{P}_{0}-\overline{P}_{n}^{\#})\left\{\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})\right\},$
and, therefore,
$(\delta_{n}^{\#})^{2}$ $\lesssim(\overline{P}_{0}-\overline{P}_{n}^{\#})\left\{\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})\right\}$
$\leq\frac{1}{J}\sum_{j=1}^{J}\left[\gamma_{n,j}(\delta_{n,j}^{\#})+\gamma_{n,j}^{\#}(\delta_{n,j}^{\#})\right].$
where we define:
$\gamma_{n,j}(\delta)$ $:=\sup_{\nu_{1},\nu_{2}\in\mathcal{F}_{n,j}:\|\nu_{1}-\nu_{2}\|\leq\delta}\left|({P}_{0}-{P}_{n,j})\left\{\ell(\cdot,\nu_{1})-\ell(\cdot,\nu_{2})\right\}\right|;$
$\gamma_{n,j}^{\#}(\delta)$ $:=\sup_{\nu_{1},\nu_{2}\in\mathcal{F}_{n,j}:\|\nu_{1}-\nu_{2}\|\leq\delta}\left|(P_{n,j}-P_{n,j}^{\#})\left\{\ell(\cdot,\nu_{1})-\ell(\cdot,\nu_{2})\right\}\right|.$
By our assumptions on the loss $\ell$, we have, for each $j\in[J]$ and $\nu_{1},\nu_{2}\in\mathcal{F}_{n,j}$, that $\ell(\cdot,\nu_{1})-\ell(\cdot,\nu_{2})$ falls in a fixed bounded Lipschitz transformation of $\mathcal{F}_{n,j}\times\mathcal{F}_{n,j}$. Hence, from the proof of Lemma 6 with $t_{n}(\delta):=\sqrt{\delta}$, we have for any fixed $\delta>0$ and conditional on $\mathcal{E}_{n,j}:=\mathcal{D}_{n}\backslash\mathcal{C}_{n,j}$ that:
$E_{0}^{n}[\gamma_{n,j}(\delta)|\mathcal{E}_{n,j}]$ $\lesssim n^{-\frac{1}{2}}\sqrt{\delta}\left(1+\frac{\sqrt{\delta}}{\sqrt{n}\delta^{2}}\right);$
$E_{0}^{n}[\gamma_{n,j}^{\#}(\delta)|\mathcal{E}_{n,j}]$ $\lesssim n^{-\frac{1}{2}}\sqrt{\delta}\left(1+\frac{\sqrt{\delta}}{\sqrt{n}\delta^{2}}\right),$
where the bound on the right-hand side is deterministic.
We are now ready to obtain a convergence rate bound for $\delta_{n}^{\#}=\|h_{n}^{\#}\circ\nu_{n,\diamond}-h_{n,0}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}$. To do so, we use a peeling argument similar to the proof of Theorem 3.2.5 in van der Vaart and Wellner (1996). We will also use the following fact: if, for given vectors $x,y,u,v\in\mathbb{R}^{J}$, both $\sum_{j=1}^{J}x_{j}\leq\sum_{j=1}^{J}y_{j}$ and $\sum_{j=1}^{J}u_{j}\leq\sum_{j=1}^{J}v_{j}$ hold then there exists some $j^{*}\in[J]$ such that both $x_{j^{*}}\leq y_{j^{*}}$ and $u_{j^{*}}\leq v_{j^{*}}$. To verify this claim, suppose the contrapositive that $x_{j}>y_{j}$ and $u_{j}>v_{j}$ for all $j\in[J]$. Then, we must have $\sum_{j=1}^{J}x_{j}>\sum_{j=1}^{J}y_{j}$ and $\sum_{j=1}^{J}u_{j}<\sum_{j=1}^{J}v_{j}$, which is a contradiction. The claim follows. Proceeding with the proof and letting $\varepsilon_{n}\in\mathbb{R}$ be a deterministic rate to be set later, note
$P(2^{k+1}\varepsilon_{n}\geq\delta_{n}^{\#}\geq 2^{k}\varepsilon_{n})$ $=P\left(2^{k+1}\varepsilon_{n}\geq\delta_{n}^{\#}\geq 2^{k}\varepsilon_{n},\sum_{j\in[J]}(\delta_{n,j}^{\#})^{2}\leq\sum_{j\in[J]}\left[\gamma_{n,j}(\delta_{n,j}^{\#})+\gamma_{n,j}^{\#}(\delta_{n,j}^{\#})\right]\right)$
$\leq P\left(2^{k+1}\varepsilon_{n}\geq\delta_{n,j_{n}}\geq 2^{k}\varepsilon_{n},\delta_{n,j_{n}}^{2}\leq\left[\theta_{n,j_{n}}(\delta_{n,j_{n}})+\theta_{n,j_{n}}^{\#}(\delta_{n,j_{n}})\right]\text{ for some }j_{n}\in[J]\right)$
$\leq\sum_{j\in[J]}P\left(2^{k+1}\varepsilon_{n}\geq\delta_{n,j}^{\#}\geq 2^{k}\varepsilon_{n},(\delta_{n,j}^{\#})^{2}\leq\left[\gamma_{n,j}(\delta_{n,j}^{\#})+\gamma_{n,j}^{\#}(\delta_{n,j}^{\#})\right]\right)$
$\leq\sum_{j\in[J]}P\left(2^{k+1}\varepsilon_{n}\geq\delta_{n,j}^{\#}\geq 2^{k}\varepsilon_{n},2^{2k}\varepsilon_{n}^{2}\leq\left[\gamma_{n,j}(2^{k+1}\varepsilon_{n})+\gamma_{n,j}^{\#}(2^{k+1}\varepsilon_{n})\right]\right)$
$\leq\sum_{j\in[J]}E_{0}^{n}\left[P\left(2^{2k}\varepsilon_{n}^{2}\leq\left[\gamma_{n,j}(2^{k+1}\varepsilon_{n})+\gamma_{n,j}^{\#}(2^{k+1}\varepsilon_{n})\right]\right)|\mathcal{E}_{n,j}\right]$
$\lesssim E_{0}^{n}\left[\frac{E_{0}^{n}[\widetilde{\gamma}_{n,j}^{\#}(2^{k+1}\varepsilon_{n})|\mathcal{E}_{n,j}]+E_{0}^{n}[\widetilde{\gamma}_{n,j}^{\#}(2^{k+1}\varepsilon_{n})|\mathcal{E}_{n,j}]}{2^{2k}\varepsilon_{n}^{2}}\right]$
$\leq C\frac{n^{-\frac{1}{2}}\sqrt{2^{k+1}\varepsilon_{n}}(1+\frac{\sqrt{2^{k+1}\varepsilon_{n}}}{\sqrt{n}\varepsilon_{n}^{2}})}{2^{2k}\varepsilon_{n}^{2}},$
for some fixed constant $C>0$. Taking $\varepsilon_{n}:=n^{-1/3}$ in the above, it is straightforward to show, for some other fixed constant $C^{\prime}>0$, that
$P(2^{k+1}n^{-1/3}\geq\delta_{n}^{\#}\geq 2^{k}n^{-1/3})$ $\lesssim\frac{n^{-\frac{1}{2}}\sqrt{2^{k+1}\varepsilon_{n}}(1+\frac{\sqrt{2^{k+1}\varepsilon_{n}}}{\sqrt{n}\varepsilon_{n}^{2}})}{2^{2k}\varepsilon_{n}^{2}}\leq C^{\prime}\frac{1}{2^{k}}.$
Therefore, for any constant $K>0$, we have
$P(\delta_{n}^{\#}\geq 2^{K}n^{-1/3})$ $=\sum_{k=K}^{\infty}P(2^{k+1}n^{-1/3}\geq\delta_{n}^{\#}\geq 2^{k}n^{-1/3})\leq C^{\prime}\sum_{k=K}^{\infty}\frac{1}{2^{k}},$
where $\sum_{k=K}^{\infty}\frac{1}{2^{k}}$ tends to zero as $K\rightarrow\infty$. Hence, for all $\varepsilon>0$, we can find a constant $K>0$ such that $P(\delta_{n}^{\#}\geq 2^{K}n^{-1/3})\leq\varepsilon$. By definition, it follows that $\delta_{n}^{\#}=O_{p}(n^{-1/3})$ and, therefore, $\|h_{n}^{\#}\circ\nu_{n,\diamond}-h_{n,0}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}^{2}=o_{p}(n^{-2/3})$. An analogous argument shows $\|h_{n}\circ\nu_{n,\diamond}-h_{n,0}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}^{2}=O_{p}(n^{-2/3})$. Next, the triangle inequality implies that $\|h_{n}^{\#}\circ\nu_{n,\diamond}-h_{n}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}^{2}=O_{p}(n^{-2/3})$. Therefore, for any function $\overline{\nu}_{0}$, it holds that
$\|h_{n}^{\#}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}$ $\leq\|h_{n}^{\#}\circ\nu_{n,\diamond}-h_{n}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}+\|h_{n}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}$
$\leq O_{p}(n^{-1/3})+\|h_{n}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}.$
The following theorem is a direct application of Theorem 7. In the following, we define the population minimizers corresponding to the calibrators $f_{n}$ and $g_{n}$ as:
$f_{n,0}$ $:=\operatorname*{argmin}_{f\in\mathcal{F}_{iso}}\sum_{j=1}^{J}\|\mu_{0}-f\circ\mu_{n,j}\|_{P_{0}}^{2};$
$g_{n,0}$ $:=\operatorname*{argmin}_{g\in\mathcal{F}_{iso}}\sum_{j=1}^{J}\left\{\|g\circ\alpha_{n,j}\|_{P_{0}}^{2}-2\psi_{0}(g\circ\alpha_{n,j})\right\}.$
Theorem 8.
Suppose that Condition C3 holds. Then, $\|\alpha_{n,\diamond}^{*}-g_{n,0}\circ\alpha_{n}\|_{\overline{P}_{0}}=O_{p}(n^{-1/3})$ and $\|\mu_{n,\diamond}^{*}-f_{n,0}\circ\mu_{n}\|_{\overline{P}_{0}}=O_{p}(n^{-1/3})$. Hence,
$\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}$ $\leq\min_{g\in\mathcal{F}_{iso}}\|g\circ\alpha_{n,\diamond}-\alpha_{0}\|+O_{p}(n^{-1/3});$
$\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}$ $\leq\min_{f\in\mathcal{F}_{iso}}\|f\circ\mu_{n,\diamond}-\mu_{0}\|+O_{p}(n^{-1/3}).$
Additionally, assume D1 holds. Then, the bootstrap calibrated nuisance estimators satisfy, for any functions $\overline{\alpha}_{0},\overline{\mu}_{0}\in L^{2}(P_{0})$ and as $n\rightarrow\infty$, the following root mean square error bounds:
$\|\alpha_{n}^{*\#}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}$ $\leq\|\alpha_{n,\diamond}^{*}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3});$
$\|\mu_{n}^{*\#}-\overline{\mu}_{0}\|_{\overline{P}_{0}}$ $\leq\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3}).$
Proof.
The proof follows from two direct applications of Theorem 7. To obtain the claimed rates for $\alpha_{n,\diamond}^{*}$, we take $\ell(x,y,\nu):=\nu^{2}(a,w)-2m(a,w,\nu)$, $h_{n}:=g_{n}$, $h_{n,0}:=g_{n,0}$, $h_{n}^{\#}:=g_{n}^{\#}$, $\nu_{0}:=\alpha_{0}$, $\nu_{n,j}:=\alpha_{n,j}$. Similarly, to obtain the rates for $\mu_{n,\diamond}^{*}$, we take $\ell(x,y,\nu):=\nu^{2}(a,w)-2y\nu(a,w)$, $h_{n}:=f_{n}$, $h_{n,0}:=f_{n,0}$, $h_{n}^{\#}:=f_{n}^{\#}$, $\nu_{0}:=\mu_{0}$, $\nu_{n,j}:=\mu_{n,j}$. Under C3 and the Lipschitz continuity of $m$ implied by A3, it can be verified that these loss functions satisfy the setup of Theorem 7. ∎
D.6 Key Lemmas for analysis of calibrated one-step debiased estimator
Let $\alpha_{n,\diamond}^{*}:=\{\alpha_{n,j}^{*}:j\in[J]\}$ and $\mu_{n,\diamond}^{*}:=\{\mu_{n,j}^{*}:j\in[J]\}$ be the isotonic calibrated nuisance estimators obtained via Algorithm 1. Let $\overline{\alpha}_{0}$ and $\overline{\mu}_{0}$ be their respective potentially misspecified limits, as in C2.
The following lemma establishes that the isotonic calibrated nuisance estimators are empirically calibrated for their respective loss functions.
Lemma 8 (Empirical calibration via isotonic calibration).
Then, for all $h_{1},h_{2}:\mathbb{R}\rightarrow\mathbb{R}$, it holds that
$\frac{1}{n}\sum_{i=1}^{n}(h_{1}\circ\mu_{n,j(i)}^{*})(A_{i},W_{i})\left\{Y_{i}-\mu_{n,j(i)}^{*}(A_{i},W_{i})\right\}$ $=0;$
$\frac{1}{n}\sum_{i=1}^{n}\left\{(h_{2}\circ\alpha_{n,j(i)}^{*})(A_{i},W_{i})\alpha_{n,j(i)}^{*}(A_{i},W_{i})-\psi_{n,j(i)}(h_{2}\circ\alpha_{n,j(i)}^{*})\right\}$ $=0.$
Proof.
The proof of this lemma follows from a simplification of the proof of Lemma 4 in van der Laan et al. (2023b). We give the proof for $\alpha_{n,\diamond}^{*}$ as the proof for $\mu_{n,\diamond}^{*}$ follows along similar lines. Recall that $\alpha_{n,\diamond}^{*}=g_{n}\circ\alpha_{n,\diamond}$ for the empirical risk minimizer $g_{n}\in\mathcal{F}_{iso}$ defined in Algorithm 1. Also recall, following Groeneboom and Lopuhaa (1993), that we took $g_{n}$ to be the unique càdlàg piece-wise constant solution of the isotonic regression problem that can only take jumps at observed values $\{\alpha_{n,j(i)}(A_{i},W_{i}):i\in[n]\}$. Hence, $g_{n}$ is a piece-wise constant. We claim, for any function $h:\mathbb{R}\rightarrow\mathbb{R}$ and all $\varepsilon\in\mathbb{R}$ sufficiently close to zero, that the perturbed function $g_{n}+\varepsilon h\circ g_{n}\in\mathcal{F}_{iso}$. To see this, observe that if $g_{n}$ is constant in some interval of its domain if and only if the same holds for $h\circ g_{n}$. Hence, the function $g_{n}+\varepsilon h\circ g_{n}$ is piece-wise constant and jumps at the same points as $g_{n}$. Moreover, by taking $\varepsilon$ sufficiently small, we can ensure that the maximal jump change $\varepsilon\sup_{i\in[n]}|(h\circ g_{n})(\mu_{n,j(i)}(A_{i},W_{i})|$ of $h\circ g_{n}$ is strictly smaller than the maximal jump change of $g_{n}$. Therefore, for this choice of $\varepsilon$, the perturbed function $g_{n}+\varepsilon h\circ g_{n}\in\mathcal{F}_{iso}$ must be nondecreasing and, therefore, an element of $\mathcal{F}_{iso}$. The claim now follows.
Proceeding with the proof of the Lemma, we denote $\xi_{n}(\varepsilon,\alpha_{n,j}):=g_{n}\circ\alpha_{n,j}+\varepsilon(h\circ g_{n})(\alpha_{n,j})$ and, in a slight abuse of notation, let
$R_{n}(\xi_{n}(\varepsilon)):=\sum_{i=1}^{n}\left\{\xi_{n}(\varepsilon,\alpha_{n,j})^{2}(A_{i},W_{i})-2\psi_{n,j(i)}(\xi_{n}(\varepsilon,\alpha_{n,j}))\right\}.$
Now, because $g_{n}$ minimizes $g\mapsto R_{n}(g)$ over $\mathcal{F}_{iso}$ and $g_{n}+\varepsilon h\circ g_{n}\in\mathcal{F}_{iso}$ for $\varepsilon$ sufficiently close to 00, it holds that
$\frac{d}{d\varepsilon}[R_{n}(\xi_{n}(\varepsilon))-R_{n}(g_{n})]\Big{|}_{\varepsilon=0}=0.$
Computing the derivative, using linearity of the functional $\psi_{n,\diamond}$ and that $h\circ g_{n}\circ\alpha_{n,\diamond}=h\circ\alpha_{n,\diamond}^{*}$, this further implies:
$\frac{1}{n}\sum_{i=1}^{n}\left\{(h\circ\alpha_{n,j(i)}^{*})(A_{i},W_{i})\alpha_{n,j(i)}^{*}(A_{i},W_{i})-\psi_{n,j(i)}(h\circ\alpha_{n,j(i)}^{*})\right\}=0.$
The result for $\alpha_{n,\diamond}^{*}$ now follows, noting that $h$ was an arbitrary function.
The next lemma establishes that empirical calibration of the nuisance estimators in the sense of Lemma 8 suffices to linearize the cross-product remainder term. Following the notation of Lemma 4, we denote $s_{n,j}^{*}:=\Pi_{\mu_{n,j}^{*}}\{\alpha_{0}-\alpha_{n,j}^{*}\}$ and $r_{n,j}^{*}:=\Pi_{\alpha_{n,j}^{*}}\{\mu_{0}-\mu_{n,j}^{*}\}$. Denote $A_{n,j}^{*}:z\mapsto s_{n,j}^{*}(a,w)\{y-\mu_{n,j}^{*}(a,w)\}$ and define $B_{n,j}^{*}:=\psi_{0}(r_{0})+\widetilde{\phi}_{r_{n,j}^{*}}-r_{n,j}^{*}\alpha_{n,j}^{*}$. We recall that $s_{0}:=\Pi_{\mu_{0}}\{\alpha_{0}-\overline{\alpha}_{0}\}$ and $r_{0}:=\Pi_{\alpha_{0}}\{\mu_{0}-\overline{\mu}_{0}\}$.
Lemma 9 (Linearization of remainder for calibrated one-step debiased estimator).
We have the outcome-favorable decomposition:
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})A_{n,\diamond}^{*}+O_{p}\left(\|(\Pi_{(\mu_{n,\diamond}^{*},\mu_{0})}-\Pi_{\mu_{n,\diamond}^{*}})(\alpha_{n,\diamond}^{*}-\alpha_{0})\|_{\overline{P}_{0}}\|\mu_{0}-\mu_{n,\diamond}^{*}\|_{\overline{P}_{0}}\right),$
Similarly, we have the Riesz-favorable decomposition:
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})B_{n,\diamond}^{*}+\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(r_{n,j}^{*})-\psi_{0}(r_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}^{*}}\right\}$
$\quad+O_{p}(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}\|(\Pi_{(\alpha_{n,\diamond}^{*},\alpha_{0})}-\Pi_{\alpha_{n,\diamond}^{*}})(\mu_{0}-\mu_{n,\diamond}^{*})\|_{\overline{P}_{0}},$
Proof.
The result follows from combining Lemma 4 with Lemma 8, noting that empirical calibration ensures that $\overline{P}_{n}A_{n,\diamond}^{*}=0$ and $\left\{\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(r_{n,j})-\overline{P}_{0}r_{n,\diamond}\alpha_{n,\diamond}\right\}=0$. In particular, the first score equation in Lemma 8 implies that
$\overline{P}_{n}A_{n,\diamond}^{*}=\frac{1}{n}\sum_{i=1}^{n}s_{n,j(i)}^{*}(A_{i},W_{i})\{\mu_{n,j(i)}^{*}(A_{i},W_{i})-Y_{i}\}=0,$
since each $s_{n,j}^{*}$ is a function of $\mu_{n,j}^{*}$. Similarly, the second score equation in Lemma 8 implies that
$\frac{1}{n}\sum_{i=1}^{n}r_{n,j(i)}^{*}(A_{i},X_{i})\alpha_{n,j(i)}^{*}(A_{i},X_{i})-\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(r_{n,j}^{*})=0,$
since each $r_{n,j}$ is a function of $\alpha_{n,j}^{*}$.
The next lemma establishes that certain empirical processes are negligible.
Lemma 10 (Negligible empirical process remainders).
Assume C2-C4 hold. Then, $(\overline{P}_{n}-P_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}-\widetilde{\phi}_{0,\overline{\mu}_{0}}\right\}+(\overline{P}_{n}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=o_{p}(n^{-\frac{1}{2}})$. Moreover, under the setup of Theorem 5 and if D1 and C4 also hold, $(\overline{P}_{n}^{\#}-P_{0})\left\{m(\cdot,\mu_{n,\diamond}^{\#*})-m(\cdot,\overline{\mu}_{0})\right\}+(\overline{P}_{n}^{\#}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=o_{p}(n^{-\frac{1}{2}})$.
Proof.
Condition C2 gives that $\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$ and $\|\alpha_{n,\diamond}^{*}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$. Note $(\overline{P}_{n}-P_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}-\widetilde{\phi}_{0,\overline{\mu}_{0}}\right\}=(\overline{P}_{n}-P_{0})\left\{{\phi}_{0,\mu_{n,\diamond}^{*}}-{\phi}_{0,\overline{\mu}_{0}}\right\}$, where $\mu\mapsto\phi_{0,\mu}$ is defined in A3. By the Lipschitz continuity $\mu\mapsto\phi_{0,\mu}$ in A3, we also have that $\|{\phi}_{0,\mu_{n,\diamond}^{*}}-{\phi}_{0,\overline{\mu}_{0}}\|_{\overline{P}_{0}}=\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$. Furthermore, by C3,
$\|D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\|_{\overline{P}_{0}}=\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}+\|\alpha_{n,\diamond}^{*}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}=o_{p}(1).$
The first part of the lemma will follow from a direct application of Lemma 7. By C3, $(\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*})\mapsto D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}$ is a bounded Lipschitz transformation of $(\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*})$. With probability one, by definition and C3, $\mu_{n,j}^{*}$ and $\alpha_{n,j}^{*}$ respectively fall in $\mathcal{S}_{n,j}$ and $\mathcal{R}_{n,j}$ for each $j\in[J]$. Hence, with probability one, $D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}$ falls in a bounded Lipschitz transformation of $\mathcal{S}_{n,j}\times\mathcal{R}_{n,j}$. Hence, by Lemma 7, $(\overline{P}_{n}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=o_{p}(n^{-\frac{1}{2}})$. Similarly, by A3, $(\overline{P}_{n}-P_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}-\widetilde{\phi}_{0,\overline{\mu}_{0}}\right\}=(\overline{P}_{n}-P_{0})\left\{{\phi}_{0,\mu_{n,\diamond}^{*}}-{\phi}_{0,\overline{\mu}_{0}}\right\}$ where $P_{0}$-almost surely $|{\phi}_{0,\mu_{n,\diamond}^{*}}(a,w)-{\phi}_{0,\overline{\mu}_{0}}(a,w)|\lesssim|\mu_{n,\diamond}^{*}(a,w)-\overline{\mu}_{0}(a,w)|$. Hence, $\|{\phi}_{0,\mu_{n,\diamond}^{*}}-{\phi}_{0,\overline{\mu}_{0}}\|_{\overline{P}_{0}}\lesssim\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$. Therefore, by C3 and Lemma 7, $(\overline{P}_{n}-P_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}-\widetilde{\phi}_{0,\overline{\mu}_{0}}\right\}=o_{p}(n^{-\frac{1}{2}})$, as desired.
To prove the second part of this lemma, note
$(\overline{P}_{n}^{\#}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=(\overline{P}_{n}^{\#}-\overline{P}_{n})\{D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}+(\overline{P}_{n}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}.$
By Theorem 8 and condition C2, we have $\|\alpha_{n}^{*\#}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}\leq\|\alpha_{n,\diamond}^{*}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3})$ and $\|\mu_{n}^{*\#}-\overline{\mu}_{0}\|_{\overline{P}_{0}}\leq\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3})$. Hence, by C2, $\|\alpha_{n}^{*\#}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$ and $\|\mu_{n}^{*\#}-\overline{\mu}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$. Hence, by C3, we also have $\|D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\|_{\overline{P}_{0}}=o_{p}(1)$. Now, the second term on the right-hand side of the above display is $o_{p}(n^{-\frac{1}{2}})$ by the same argument as before. The first term on the right-hand side is also $o_{p}(n^{-\frac{1}{2}})$ by a similar argument, using the second part of Lemma 7.
For each $j\in[J]$, recall that $A_{n,j}^{*}:z\mapsto s_{n,j}^{*}(a,w)\{y-\mu_{n,j}^{*}(a,w)\}$ and $B_{n,j}^{*}:=\psi_{0}(r_{0})+\widetilde{\phi}_{r_{n,j}^{*}}-r_{n,j}^{*}\alpha_{n,j}^{*}$.
Lemma 11 (Negligible cross-product empirical process remainders).
Under C2-C4, $1(\overline{\mu}_{0}=\mu_{0})(\overline{P}_{n}-P_{0})\{A_{n,\diamond}^{*}-A_{0}\}=o_{p}(n^{-\frac{1}{2}})$ and $1(\overline{\alpha}_{0}=\alpha_{0})(\overline{P}_{n}-P_{0})\{B_{n,\diamond}^{*}-B_{0}\}=o_{p}(n^{-\frac{1}{2}})$. Moreover, under the setup of Theorem 5 and if D1, C4, and C5 also hold, $1(\overline{\mu}_{0}=\mu_{0})(\overline{P}_{n}^{\#}-\overline{P}_{0})\{A_{n,\diamond}^{*\#}-A_{0}\}=o_{p}(n^{-\frac{1}{2}})$, $1(\overline{\alpha}_{0}=\alpha_{0})(\overline{P}_{n}^{\#}-\overline{P}_{0})\{B_{n,\diamond}^{*\#}-B_{0}\}=o_{p}(n^{-\frac{1}{2}})$.
Proof.
Condition C5 combined with Lemma 1 and C2 implies that $1(\overline{\alpha}_{0}=\alpha_{0})\|r_{n,\diamond}^{*}-r_{0}\|_{\overline{P}_{0}}=\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}=o_{p}(1)$ and $1(\overline{\mu}_{0}=\mu_{0})\|s_{n,\diamond}^{*}-s_{0}\|_{\overline{P}_{0}}=\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}=o_{p}(1)$. Hence, by C3, when $1(\overline{\mu}_{0}=\mu_{0})$, we have:
$\|A_{n,\diamond}^{*}-A_{0}\|_{\overline{P}_{0}}=\|s_{n,\diamond}^{*}-s_{0}\|_{\overline{P}_{0}}+\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}=o_{p}(1).$
The outcome $Y$ is assumed bounded almost surely. Hence, by C3, $(\mu_{n,\diamond}^{*},s_{n,\diamond}^{*})\mapsto s_{n,\diamond}^{*}(\mathcal{I}_{Y}-\mu_{n,\diamond}^{*})-s_{0}(\mathcal{I}_{Y}-\mu_{0})$ is a Lipschitz transformation of $\mu_{n,\diamond}^{*}$ and $s_{n,\diamond}^{*}$, where $s_{n,\diamond}^{*}(\mathcal{I}_{Y}-\mu_{n,\diamond}^{*})-s_{0}(\mathcal{I}_{Y}-\mu_{0})$ is bounded with probability tending to one. For each $j\in[J]$, with probability tending to one, by definition, $\mu_{n,j}^{*}$ falls in $\mathcal{S}_{n,j}$ and similarly, by C4, $s_{n,j}^{*}$ falls in $\mathcal{S}_{n,j}$. Hence, with probability tending to one, $A_{n,\diamond}^{*}-A_{0}=s_{n,\diamond}^{*}(\mathcal{I}_{Y}-\mu_{n,\diamond}^{*})-s_{0}(\mathcal{I}_{Y}-\mu_{0})$ falls in a bounded Lipschitz transformation of $\mathcal{S}_{n,j}\times\mathcal{S}_{n,j}$. Hence, by Lemma 7, we conclude
$1(\overline{\mu}_{0}=\mu_{0})(\overline{P}_{n}-\overline{P}_{0})\{A_{n,\diamond}^{*}-A_{0}\}=o_{p}(n^{-\frac{1}{2}}).$
Next, note that
$(\overline{P}_{n}-\overline{P}_{0})\{B_{n,\diamond}^{*}-B_{0}\}$ $=(\overline{P}_{n}-\overline{P}_{0})\{\widetilde{\phi}_{r_{n,j}^{*}}-\widetilde{\phi}_{r_{0}}+r_{0}\alpha_{0}-r_{n,j}^{*}\alpha_{n,j}^{*}\}$
$=(\overline{P}_{n}-\overline{P}_{0})\{\phi_{r_{n,j}^{*}}-\phi_{r_{0}}+r_{0}\alpha_{0}-r_{n,j}^{*}\alpha_{n,j}^{*}\}.$
By C3 and property A3, when $1(\overline{\alpha}_{0}=\alpha_{0})$, we have:
$\|\phi_{r_{n,\diamond}^{*}}-\phi_{r_{0}}+r_{0}\alpha_{0}-r_{n,\diamond}^{*}\alpha_{n,\diamond}^{*}\|_{\overline{P}_{0}}=\|r_{n,\diamond}^{*}-r_{0}\|_{\overline{P}_{0}}+\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}=o_{p}(1).$
By C3 and property A3, $(\alpha_{n,\diamond}^{*},r_{n,\diamond}^{*})\mapsto\phi_{r_{n,\diamond}^{*}}-\phi_{r_{0}}+r_{0}\alpha_{0}-r_{n,\diamond}^{*}\alpha_{n,\diamond}^{*}$ is a Lipschitz transformation of $\mu_{n,\diamond}^{*}$ and $s_{n,\diamond}^{*}$, where $\phi_{r_{n,\diamond}^{*}}-\phi_{r_{0}}+r_{0}\alpha_{0}-r_{n,\diamond}^{*}\alpha_{n,\diamond}^{*}$ is bounded with probability tending to one. For each $j\in[J]$, with probability tending to one, by definition, $\alpha_{n,j}^{*}$ falls in $\mathcal{R}_{n,j}$ and similarly, by C4, $r_{n,j}^{*}$ falls in $\mathcal{R}_{n,j}$. Hence, with probability one, $\phi_{r_{n,j}^{*}}-\phi_{r_{0}}+r_{0}\alpha_{0}-r_{n,j}^{*}\alpha_{n,j}^{*}$ falls in a bounded Lipschitz transformation of $\mathcal{R}_{n,j}\times\mathcal{R}_{n,j}$. Hence, by Lemma 7, we conclude
$1(\overline{\mu}_{0}=\mu_{0})(\overline{P}_{n}-\overline{P}_{0})\{B_{n,\diamond}^{*}-B_{0}\}=o_{p}(n^{-\frac{1}{2}}).$
The bootstrap case follows along similar lines. By C5 and Lemma 1, it holds that
$\|r_{n,\diamond}^{*\#}-r_{0}\|_{\overline{P}_{0}}=O\left(\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}\right);$
$\|s_{n,\diamond}^{*\#}-s_{0}\|_{\overline{P}_{0}}=O\left(\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}\right).$
Under the setup of Theorem 5, we can apply Theorem 8 to conclude
$\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}=O_{p}\left(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}\right)+O_{p}(n^{-1/3});$
$\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}=O_{p}\left(\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}\right)+O_{p}(n^{-1/3}).$
Using D1 and D2, an argument identical to the non bootstrap case, applying the second part of Lemma 7 and arguing as in the proof of Lemma 10 gives the second part of the lemma.
D.7 Proofs of main results
Proof of Theorem 4.
Starting point. By Lemma 3, it holds that
$\tau_{n}^{*}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}+D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}\right\}+\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}+Rem_{\mu_{n,\diamond}^{*}}(\psi_{n,\diamond},\psi_{0}).$
where $Rem_{\mu_{n,\diamond}^{*}}(\psi_{n,\diamond},\psi_{0}):=\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(\mu_{n,j}^{*})-\psi_{0}(\mu_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{0,\mu_{n,j}^{*}}\right\}$. By direction application of C7, we have $Rem_{\mu_{n,\diamond}^{*}}(\psi_{n,\diamond},\psi_{0})=o_{p}(n^{-\frac{1}{2}})$. In addition, by Lemma 10 with C2, C3, and C4, we have that
$(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}+D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}\right\}$ $=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}$
$\quad+(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}+D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}\right\}$
$=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}+o_{p}(n^{-\frac{1}{2}}).$
Thus,
$\tau_{n}^{*}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}+\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}.$
It remains to analyze the cross-product remainder term $\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$. We consider two cases.
Outcome-favorable case. Suppose that C6 holds with $\|\mu_{n,j}-\mu_{0}\|=o_{p}(n^{-1/4})$ for each $j\in[J]$, so that $\overline{\mu}_{0}=\mu_{0}$ in C2. By Lemma 9, we have the outcome-favorable decomposition:
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})A_{n,\diamond}^{*}+O_{p}\left(\|(\Pi_{(\mu_{n,\diamond}^{*},\mu_{0})}-\Pi_{\mu_{n,\diamond}^{*}})(\alpha_{n,\diamond}^{*}-\alpha_{0})\|_{\overline{P}_{0}}\|\mu_{0}-\mu_{n,\diamond}^{*}\|_{\overline{P}_{0}}\right),$
By C2 with $\overline{\mu}_{0}=\mu_{0}$, C3, and a direct application of Lemma 11, we find
$(\overline{P}_{n}-P_{0})A_{n,\diamond}^{*}$ $=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+(\overline{P}_{n}-P_{0})\left\{A_{n,\diamond}^{*}-A_{0}\right\}$
$=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+o_{p}(n^{-\frac{1}{2}}).$
Next, by C5 and Lemma 1, it holds that
$\|(\Pi_{(\mu_{n,\diamond}^{*},\mu_{0})}-\Pi_{\mu_{n,\diamond}^{*}})(\alpha_{n,\diamond}^{*}-\alpha_{0})\|_{\overline{P}_{0}}=O_{p}\left(\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{n}}\right).$
Thus, by C6, we find
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+O_{p}\left(\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{n}}^{2}\right)$
$=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+o_{p}(n^{-\frac{1}{2}}).$
Riesz-favorable case. Suppose that C6 holds with $\|\alpha_{n,j}-\alpha_{0}\|=o_{p}(n^{-1/4})$ for each $j\in[J]$, so that $\overline{\alpha}_{0}=\alpha_{0}$ in C2. By Lemma 9, we have the Riesz-favorable decomposition:
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})B_{n,\diamond}^{*}+\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(r_{n,j}^{*})-\psi_{0}(r_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}^{*}}\right\}$
$\quad+O_{p}(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}\|(\Pi_{(\alpha_{n,\diamond}^{*},\alpha_{0})}-\Pi_{\alpha_{n,\diamond}^{*}})(\mu_{0}-\mu_{n,\diamond}^{*})\|_{\overline{P}_{0}},$
By C7, it holds that $\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(r_{n,j}^{*})-\psi_{0}(r_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}^{*}}\right\}=o_{p}(n^{-\frac{1}{2}})$. Furthermore, by C2 with $\overline{\alpha}_{0}=\alpha_{0}$, C3, and a direct application of Lemma 11, we find
$(\overline{P}_{n}-P_{0})B_{n,\diamond}^{*}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+(\overline{P}_{n}-P_{0})\left\{B_{n,\diamond}^{*}-B_{0}\right\}$
$=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Finally, by C5 and Lemma 1, it holds that
$\|(\Pi_{(\alpha_{n,\diamond}^{*},\alpha_{0})}-\Pi_{\alpha_{n,\diamond}^{*}})(\mu_{0}-\mu_{n,\diamond}^{*})\|_{\overline{P}_{0}}=O_{p}\left(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{n}}\right).$
Thus, by C6,
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+O_{p}(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}^{2}$
$=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Completing the argument. Combining the two cases treated above, we find
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+O_{p}(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}^{2})$
$=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Thus, it holds that
$\tau_{n}^{*}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}+1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Note that $\overline{P}_{0}\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=\overline{P}_{0}\{{\phi}_{0,\overline{\mu}_{0}}-\psi_{0}(\mu_{0})+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}$ if either $\overline{\mu}_{0}=\mu_{0}$ or $\overline{\alpha}_{0}=\alpha_{0}$. Recalling that $\chi_{0}={\phi}_{0,\overline{\mu}_{0}}-\psi_{0}(\mu_{0})+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}+1(\overline{\alpha}_{0}\neq\alpha_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})B_{0}$ and that $\chi_{0}$ is mean zero, we obtain the desired asymptotic linearity result:
$\tau_{n}^{*}-\tau_{0}=P_{n}\chi_{0}+o_{p}(n^{-\frac{1}{2}}).$
If both nuisance functions are consistently estimated ($\overline{\alpha}_{0}=\alpha_{0}$ and $\overline{\mu}_{0}=\mu_{0}$), then the above display reduces to
$\tau_{n}^{*}-\tau_{0}=P_{n}D_{0}+o_{p}(n^{-\frac{1}{2}}).$
By Theorem 1, the above means that $\tau_{n}^{*}$ is asymptotically linear with influence function being the efficient influence function of $\Psi$ under the nonparametric statistical model $\mathcal{M}$. As a consequence, by van der Vaart (2000), it holds that $\tau_{n}^{*}$ is a regular and efficient estimator of $\psi_{0}$ under the nonparametric statistical model $\mathcal{M}$. ∎
Proof of Theorem 5.
Starting point. By Lemma 3, it holds that
$\tau_{n}^{*\#}-\tau_{0}$ $=(\overline{P}_{n}^{\#}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*\#}}+D_{\mu_{n,\diamond}^{*\#},\alpha_{n,\diamond}^{*\#}}\right\}+\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}+Rem_{\mu_{n,\diamond}^{*\#}}(\psi_{n,\diamond},\psi_{0})$
$=(\overline{P}_{n}^{\#}-\overline{P}_{0})\left\{m(\cdot,\mu_{n,\diamond}^{*\#})+D_{\mu_{n,\diamond}^{*\#},\alpha_{n,\diamond}^{*\#}}\right\}+\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}},$
where we used that $\widetilde{\phi}_{0,\mu_{n,\diamond}^{*\#}}=m(\cdot,\mu_{n,\diamond}^{*\#})$ and $Rem_{\mu_{n,\diamond}^{*\#}}(\psi_{n,\diamond},\psi_{0})=0$ for the setup of Section 5.2.
Arguing as in the Proof of Theorem 4, we can apply the second part of Lemma 10 with C2, C3, D1, C4, and C4, to find that
$(\overline{P}_{n}^{\#}-\overline{P}_{0})\left\{m(\cdot,\mu_{n,\diamond}^{*\#})+D_{\mu_{n,\diamond}^{*\#},\alpha_{n,\diamond}^{*\#}}\right\}=(\overline{P}_{n}^{\#}-\overline{P}_{0})\left\{m(\cdot,\overline{\mu}_{0})+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}+o_{p}(n^{-\frac{1}{2}}).$
Under the setup of Theorem 5, we can apply Theorem 8 to conclude
$\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}=O_{p}\left(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}\right)+O_{p}(n^{-1/3});$
$\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}=O_{p}\left(\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}\right)+O_{p}(n^{-1/3}).$
Outcome-favorable case. Suppose that C6 holds with $\|\mu_{n,j}-\mu_{0}\|=o_{p}(n^{-1/4})$ for each $j\in[J]$, so that $\overline{\mu}_{0}=\mu_{0}$ in C2. An argument identical to that of the Proof of Theorem 4 using C5 and the second part of Lemma 11 establishes that
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}$ $=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}^{\#}-P_{0})A_{0}+O_{p}\left(\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}^{2}\right).$
By Theorem 8, it holds that $\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}^{2}=O_{P}(n^{-2/3})+\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}^{2}$, and thus, by C6, we have that $\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}^{2}=o_{p}(n^{-\frac{1}{2}})$. We conclude that
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}$ $=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}^{\#}-P_{0})A_{0}+o_{p}(n^{-\frac{1}{2}}).$
Riesz-favorable case. Suppose that C6 holds with $\|\alpha_{n,j}-\alpha_{0}\|=o_{p}(n^{-1/4})$ for each $j\in[J]$, so that $\overline{\alpha}_{0}=\alpha_{0}$ in C2. An argument identical to that of the Proof of Theorem 4 using C5 and the second part of Lemma 11 establishes that
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}^{\#}-P_{0})B_{0}+O_{p}\left(\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}^{2}\right).$
By Theorem 8, it holds that
$\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}^{2}=O_{P}(n^{-2/3})+\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{n}}^{2},$
and, thus, by C6, $\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{n}}^{2}=o_{p}(n^{-\frac{1}{2}})$. We conclude that
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}^{\#}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Completing the argument. Combining the two cases treated above, we find
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}^{\#}-P_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}^{\#}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Thus, it holds that
$\tau_{n}^{*\#}-\tau_{0}=(\overline{P}_{n}^{\#}-\overline{P}_{0})\chi_{0}+o_{p}(n^{-\frac{1}{2}}).$
By Theorem 4, we also have that $\tau_{n}^{*}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\chi_{0}+o_{p}(n^{-\frac{1}{2}})$, from which we can then write that $\tau_{n}^{*\#}-\tau_{n}^{*}=\tau_{n}^{*\#}-\tau_{n}^{*}+\tau_{n}^{*}-\tau_{n}^{*}=(\overline{P}_{n}^{\#}-\overline{P}_{n})\chi_{0}+o_{p}(n^{-\frac{1}{2}})$.
For each $j\in[J]$, by the bootstrap CLT (van der Vaart and Wellner, 1996), we have $(P_{n,j}^{\#}-P_{n,j})\chi_{0}|P_{n,j}$ converges in distribution to $N(0,P_{0}\chi_{0}^{2})$. Hence, by the mutual independence of $\{P_{n,j}:j\in[J]\}$, we also have $(P_{n,j}^{\#}-P_{n,j})\chi_{0}|\{P_{n,j}:j\in[J]\}$ converges in distribution to $N(0,P_{0}\chi_{0}^{2})$. Again by mutual independence of $\{P_{n,j}:j\in[J]\}$, we have $(\overline{P}_{n}^{\#}-\overline{P}_{n})\chi_{0}|\{P_{n,j}:j\in[J]\}$ converges in distribution to $N(0,P_{0}\chi_{0}^{2})$. Thus, the law of $\tau_{n}^{*\#}-\tau_{n}^{*}|\{P_{n,j}:j\in[J]\}$ converges in probability to the law of $N(0,P_{0}\chi_{0}^{2})$.
Proof of Theorem 6.
If $\overline{\alpha}_{0}=\alpha_{0}$ and $\overline{\mu}\neq\mu_{0}$, it follows that
$\chi_{0}$ $=m(\cdot,\overline{\mu}_{0})-\psi_{0}(\mu_{0})+D_{\overline{\mu}_{0},\alpha_{0}}+B_{0}$
$=m(\cdot,\overline{\mu}_{0})-\psi_{0}(\mu_{0})+\alpha_{0}\{\mathcal{I}_{y}-\overline{\mu}_{0}\}+m(\cdot,r_{0})-r_{0}\alpha_{0}$
$=m(\cdot,\overline{\mu}_{0}+r_{0})+\alpha_{0}\{\mathcal{I}_{Y}-\overline{\mu}_{0}-r_{0}\}-\psi_{0}(\mu_{0})$
Now, we have that $\psi_{0}(\mu_{0})=\psi_{0}(\overline{\mu}_{0}+r_{0})$ by the orthogonality conditions defining $r_{0}$. Thus, using that $\overline{\mu}_{0}+r_{0}=\overline{\Pi}_{0,P_{0}}\mu_{0}$, we have
$\chi_{0}$ $=m(\cdot,\overline{\mu}_{0}+r_{0})-\psi_{0}(\overline{\mu}_{0}+r_{0})+\alpha_{0}\{\mathcal{I}_{Y}-\overline{\mu}_{0}-r_{0}\}$
$=m(\cdot,\overline{\Pi}_{0,P_{0}}\mu_{0})-\psi_{0}(\overline{\Pi}_{0,P_{0}}\mu_{0})+\alpha_{0}\{\mathcal{I}_{Y}-\overline{\Pi}_{0,P_{0}}\mu_{0}\},$
where the right-hand side of the above display is the $P_{0}$-efficient influence function of the parameter $P\mapsto\psi_{P}(\overline{\Pi}_{0,P}\mu_{P})$. The latter claim follows from the proof of Theorem 1 after replacing the outcome $Y$ with the residual $Y-\overline{\mu_{0}}$.
If $\overline{\alpha}_{0}\neq\alpha_{0}$ and $\overline{\mu}=\mu_{0}$, it follows that
$\chi_{0}$ $=m(\cdot,\mu_{0})-\psi_{0}(\mu_{0})+D_{\mu_{0},\overline{\alpha}_{0}}+A_{0}$
$=m(\cdot,\mu_{0})-\psi_{0}(\mu_{0})+D_{\mu_{0},\overline{\alpha}_{0}}+s_{0}\{\mathcal{{I}_{Y}}-\mu_{0}\}$
$=m(\cdot,\mu_{0})-\psi_{0}(\mu_{0})+\{\overline{\alpha}_{0}+s_{0}\}\{\mathcal{{I}_{Y}}-\mu_{0}\}$
Thus, using that $\overline{\alpha}_{0}+s_{0}=\overline{\Pi}_{0,P_{0}}^{*}\alpha_{0}$, we have $\chi_{0}=m(\cdot,\mu_{0})-\psi_{0}(\mu_{0})+\overline{\Pi}_{0,P_{0}}^{*}\alpha_{0}(\mathcal{I}_{Y}-\mu_{0})$. We claim that the right-hand side of the above display is the $P_{0}$–efficient influence function of the parameter $P\mapsto\overline{\psi}_{0,P}(\mu_{P})=\langle\overline{\Pi}_{0,P}^{*}\alpha_{P},\mu_{P}\rangle_{P}$. To see this, we note that $\overline{\psi}_{0,P}(\mu_{P})=\langle\overline{\Pi}_{0,P}^{*}\alpha_{P},\mu_{P}\rangle_{P}=\langle\overline{\alpha}_{0}+\Pi_{\mu_{0}}\{\alpha_{P}-\overline{\alpha}_{0}\},\mu_{P}\rangle_{P}=\langle\overline{\alpha}_{0}+s_{0},\mu_{P}\rangle_{P}$ is a linear functional of $\mu_{P}$ with Riesz representer $\overline{\alpha}_{0}+s_{0}$. The claim then follows from the proof of Theorem 1.
Since $\psi_{n}^{*}$ is $P_{0}$-asymptotically linear with the influence function being the $P_{0}$-EIF of the oracle parameter $\Psi_{0}$, it follows that $\psi_{n}^{*}$ is a locally regular and efficient estimator for $\Psi_{0}$ at $P_{0}$. The stated limiting distribution result follows directly from the definition of regularity.