### v2

# Doubly robust inference via calibration

Abstract
Doubly robust estimators are widely used for estimating average treatment effects and other linear summaries of regression functions. While consistency requires only one of two nuisance functions to be estimated consistently, asymptotic normality typically require sufficiently fast convergence of both. In this work, we correct this mismatch: we show that calibrating the nuisance estimators within a doubly robust procedure yields doubly robust asymptotic normality for linear functionals. We introduce a general framework, calibrated debiased machine learning (calibrated DML), and propose a specific estimator that augments standard DML with a simple isotonic regression adjustment. Our theoretical analysis shows that the calibrated DML estimator remains asymptotically normal if either the regression or the Riesz representer of the functional is estimated sufficiently well, allowing the other to converge arbitrarily slowly or even inconsistently. We further propose a simple bootstrap method for constructing confidence intervals, enabling doubly robust inference without additional nuisance estimation. In a range of semi-synthetic benchmark datasets, calibrated DML reduces bias and improves coverage relative to standard DML. Our method can be integrated into existing DML pipelines by adding just a few lines of code to calibrate cross-fitted estimates via isotonic regression.
Keywords: debiasing, isotonic calibration, double robustness, targeted learning, causal inference
1 Introduction
1.1 Motivation
Nonparametric inference on the causal effects of dynamic and stochastic treatment interventions using flexible statistical learning techniques is a fundamental problem in causal inference. These inferential targets fall within a broader class of target parameters that can be expressed as linear functionals of the outcome regression, that is, the conditional mean of the outcome given treatment assignment and baseline covariates (Chernozhukov et al., 2018b; Rotnitzky et al., 2021; Hirshberg and Wager, 2021). To achieve inference for such parameters, various debiased machine learning (DML) frameworks are available, including one-step estimation (Pfanzagl and Wefelmeyer, 1985; Bickel et al., 1993), estimating equations and double machine learning (Robins et al., 1995, 1994; van der Laan and Robins, 2003; Chernozhukov et al., 2018a), and targeted minimum loss estimation (TMLE) (van der Laan and Rubin, 2006; van der Laan and Rose, 2011). When performing inference on linear functionals, debiased machine learning methods frequently produce estimators that demonstrate robustness against inconsistent or slow estimation of certain nuisance functions, a property known as double robustness (Bang and Robins, 2005; Smucler et al., 2019). For example, when estimating the average treatment effect (ATE), doubly robust estimators remain consistent as long as either the outcome regression or the propensity score is consistently estimated, even if the other is inconsistently estimated. This robustness property extends more generally to linear functionals of the outcome regression, necessitating consistent estimation of only the outcome regression itself or the Riesz representer of the linear functional (Chernozhukov et al., 2018b; Rotnitzky et al., 2021).
Extending the double robustness property of debiased machine learning estimators to statistical inference is a challenging problem. Achieving asymptotic normality for such estimators typically requires consistent estimation of both the outcome regression and the Riesz representer at sufficiently fast rates. When one of the nuisance estimators is inconsistent or estimated too slowly, desirable properties such as asymptotic linearity, asymptotic normality, and $n^{\frac{1}{2}}$–consistency can be lost, compromising the validity of inferences (Benkeser et al., 2017). In particular, in such cases, typical confidence intervals and p-values can be expected to be incorrect. One approach to achieving doubly robust inference is through the use of parametric methods, such as maximum likelihood estimation, for estimating nuisance parameters. These methods exhibit model double robustness, maintaining asymptotic linearity when at least one of the nuisance parametric models is correctly specified (Vermeulen and Vansteelandt, 2015, 2016). However, the applicability of these methods for nonparametric inference is limited due to their reliance on strong assumptions, as both models are then likely to be misspecified in practice (Kang and Schafer, 2007).
To overcome the limitations of parametric methods, several estimators exhibiting sparsity double robustness, a weaker form of model double robustness, have been proposed (Avagyan and Vansteelandt, 2017; Athey et al., 2018; Bradic et al., 2019b, a; Smucler et al., 2019; Tan, 2020; Zhang et al., 2021; Bradic et al., 2024). For instance, in the context of high-dimensional generalized linear regression models, Avagyan and Vansteelandt (2017) and Bradic et al. (2019b, a) introduced doubly robust $\ell_{1}$–regularized regression estimators of the ATE. These estimators achieve asymptotic linearity when either the outcome regression or propensity score exhibits sufficient sparsity in the covariates, allowing for misspecification of one of the nuisance models. Additionally, Athey et al. (2018) and Tan (2020) proposed singly robust estimators that are robust to misspecification of the outcome regression model, assuming sufficient sparsity in the propensity score. A significant drawback of these estimators is their reliance on strong sparsity assumptions for the nuisance functions, which may fail to hold in practice. Smucler et al. (2019) developed a unifying framework for constructing sparsity doubly robust estimators using $\ell_{1}$–regularization techniques. They relax the strong sparsity assumptions by requiring that nuisance estimators based on possibly misspecified models converge to approximately sparse functions sufficiently quickly. However, as the authors note, there is limited justification to expect an inconsistent nuisance estimator derived from a misspecified model to tend to a sparse or approximately sparse limiting function, even if the true nuisance function depends on only a small subset of covariates. Furthermore, these methods are constrained by their reliance on $\ell_{1}$–regularized generalized linear models for the nuisance functions, limiting their ability to fully leverage the advantages offered by modern learning techniques.
The seminal work of van der Laan (2014) introduced a debiasing technique—based on the TMLE framework (van der Laan and Rose, 2011)—for constructing estimators that yield doubly robust inference even when nuisances are estimated using generic machine learning techniques. The resulting estimators are doubly robust asymptotically linear (DRAL), in the sense that they remain asymptotically linear even when one of the nuisance functions is estimated too slowly or even inconsistently. Their approach has since been applied to nonparametric inference of the ATE (Benkeser et al., 2017), mean outcomes under informative missingness (Díaz and van der Laan, 2017), counterfactual survival curves (Díaz, 2019), and semiparametric inference on the ATE in partially linear regression models (Dukes et al., 2021). For inference on the ATE, these estimators involve iteratively debiasing initial estimators of the propensity score and outcome regression to linearize the doubly robust remainder term of the resulting ATE estimator.
The TMLE-based framework for doubly robust inference has some limitations. First, the iterative debiasing procedure it uses is difficult to study theoretically and can be computationally demanding to implement, posing challenges for its application to large-scale data or complex data structures. While Bonvini et al. (2024) proposed an alternative, non-iterative nonparametric DRAL estimator of the ATE based on higher-order influence functions, their procedure requires careful tuning of bivariate kernel smoothing parameters, for which practical guidance remains to be developed. Second, when considering new parameters of interest, a new debiasing algorithm must be derived on a case-by-case basis, requiring careful analysis of the relevant remainder terms and novel targeting algorithms. To date, there is no unified approach for obtaining doubly robust inference for a generic linear functional of the outcome regression.
1.2 Contributions of this work
We develop new methods for doubly robust inference on linear functionals of the outcome regression. Our goal is to construct procedures that allow valid inference even if one of the two nuisance functions is estimated arbitrarily slowly or even inconsistently. To do so, we first establish a link between calibration, a technique typically used in prediction and classification tasks (Zadrozny and Elkan, 2001; Niculescu-Mizil and Caruana, 2005b), and doubly robust asymptotic linearity. We then introduce a general framework called calibrated debiased machine learning (calibrated DML). As outlined in Figure 1, calibrated DML starts with standard cross-fitted estimates of the outcome regression and Riesz representer, obtained using any machine learning algorithm, and then applies a simple additional step: calibration of these nuisance estimates using, for example, isotonic regression (Barlow and Brunk, 1972). This calibration step corrects the bias in the nuisance estimates in a way that enables a debiased one-step estimator to achieve doubly robust asymptotic linearity. Confidence intervals can be constructed using the closed-form influence function of the calibrated DML estimator, without requiring knowledge of which, if any, nuisance estimators converge sufficiently quickly. We show that no correction is needed to the standard influence-function-based variance estimator when both nuisance estimators are consistent for the true nuisances, even if one converges arbitrarily slowly. In the general case where one nuisance is inconsistently estimated, to avoid the additional nuisance estimation required in van der Laan (2014) and related approaches, we propose a computationally efficient, bootstrap-assisted procedure that remains valid as long as at least one nuisance estimator is consistent at an appropriate rate.
Estimate NuisancesCalibrateDebias
Figure 1: We show calibrating the nuisances before debiasing ensures doubly robust asymptotic normality.
Table 1: Properties of G-computation plug-in, IPW, AIPW and calibrated DML (proposed approach) estimators of the ATE when nuisances are estimated using machine learning, in settings where only the propensity score vs. only the outcome regression vs. both nuisances are estimated well (i.e., consistently and at a sufficient rate). Report whether each estimator is consistent, asymptotically normal at root-sample size rate, and nonparametric efficient.
($\ast$) Our estimator is efficient if one nuisance is estimated well, provided the other is estimated consistently (at any rate).
only propensity score estimated well only outcome regression estimated well both nuisance functions estimated well
estimator      consistent normal efficient consistent normal efficient consistent normal efficient
G-computation      × ×
IPW      × ×
AIPW      × × × ×
Calibrated DML (ours)      $\ast$ $\ast$
Table 1 summarizes the properties of a specific implementation of calibrated DML in the well-studied setting of ATE estimation. This implementation employs isotonic regression for calibration and a cross-fitted augmented inverse probability weighted (AIPW) estimator for debiasing. Compared to the standard AIPW estimator, calibrated DML remains asymptotically normal—and potentially efficient—even when only one nuisance function is estimated consistently. It yields even larger improvements over non-doubly robust estimators based on the G-computation formula or inverse probability weighting (IPW).
Our approach is motivated by the DRAL estimators of the ATE proposed in van der Laan (2014) and Benkeser et al. (2017). While we share the objective of constructing nuisance estimators to achieve doubly robust asymptotic linearity by solving certain score equations, our strategy and implementation for doing so differs significantly, even when focusing on estimating the ATE. Specifically, our approach is a noniterative and computationally efficient two-step procedure that can be easily implemented using commonly available isotonic regression software. Moreover, our method is not dependent on the specific form of the linear functional but rather on the availability of initial nuisance estimators and the ability to evaluate an empirical estimator of the linear functional as a black box. Furthermore, even in the special case of the ATE, our theoretical results refine those presented by van der Laan (2014) and Benkeser et al. (2017), establishing that our estimator achieves asymptotic linearity not only under consistent estimation of one nuisance function alone but also when both nuisance functions are estimated consistently, with one estimated too slowly.
Prior work has empirically investigated the finite-sample benefits of calibrating propensity score estimates in IPW and AIPW estimators of the ATE (Gutman et al., 2022; Deshpande and Kuleshov, 2023; Ballinari and Bearth, 2024; van der Laan et al., 2024b; Rabenseifner et al., 2025). However, these studies have focused mainly on the calibration of a single nuisance function—the propensity score—and a single parameter, the ATE. Additionally, none of these works has provided theoretical justification for the improved performance achieved through propensity score calibration beyond the observation that calibration can reduce the estimation error of nuisance functions. Our contribution extends this line of research by demonstrating that, for the general class of linear functionals, calibration of both the outcome regression and the Riesz representer is critical for further debiasing, both in finite samples and asymptotically, and for constructing DRAL estimators. In the special case of the ATE, we find that the outcome regression should also be calibrated.
This paper is organized as follows. In Section 2, we introduce the class of regression functionals we consider and provide background on a standard rate-doubly-robust debiased estimator. In Section 3, we formulate the problem of doubly robust inference, introduce a model-agnostic linear expansion for the bias of one-step debiased estimators, and demonstrate how further debiasing of nuisances can be employed to construct DRAL estimators. In Section 4, we present a novel link between nuisance estimator calibration and the doubly robust inference properties of debiased machine learning estimators, use this link to build our general calibrated DML framework, and propose an calibrated DML estimator based on isotonic regression with accompanying bootstrap-based confidence intervals. In Section 5, we establish the theoretical properties of our proposed calibrated DML estimators and confidence intervals. Finally, we empirically investigate the performance of calibrated DML in Section 6 and provide concluding remarks in Section 7.
2 Problem setup
2.1 Data structure and target parameter
[…]
Our objective is to obtain doubly robust nonparametric inference for a feature $\tau_{0}:=\psi_{0}(\mu_{0})$ of the outcome regression $\mu_{0}$ arising from a continuous linear functional $\psi_{0}:\mathcal{H}\rightarrow\mathbb{R}$. Here, $\psi_{0}$ represents the evaluation at $P_{0}$ of a smooth mapping $P\mapsto\psi_{P}(\cdot)$ defined on the statistical model $\mathcal{M}$. In other words, we explicitly allow the functional of interest to itself depend on the unknown data-generating distribution $P_{0}$. For each $P\in\mathcal{M}$, we view $\psi_{P}$ as an element of the space of bounded linear functionals on the normed linear space $(\mathcal{H},\|\cdot\|_{L^{2}(P_{0,A,W})})$, equipped with the operator norm. We require $\psi_{0}$ to satisfy the following conditions:
  1. A1)
    (linearity and boundedness) it holds that $\psi_{0}(c\mu_{1}+\mu_{2})=c\psi_{0}(\mu_{1})+\psi_{0}(\mu_{2})$ for each $\mu_{1},\mu_{2}\in\mathcal{H}$ and $c\in\mathbb{R}$ and that $\sup_{\mu\in\mathcal{H}\backslash\{0\}}|\psi_{0}(\mu)|/\|\mu\|<\infty$;
  2. A2)
    (Hellinger continuity) $P\mapsto\psi_{P}$ is a continuous map at $P_{0}$ with domain $\mathcal{M}$ equipped with the Hellinger distance and codomain $\{\psi_{P}:P\in\mathcal{M}\}$ equipped with the operator norm;
  3. A3)
    (pathwise differentiability) for each $\mu\in\mathcal{H}$, $P\mapsto\psi_{P}(\mu)$ is pathwise differentiable at $P_{0}$ relative to $\mathcal{M}$ with efficient influence function $\widetilde{\phi}_{0,\mu}:=\phi_{0,\mu}-P_{0}\phi_{0,\mu}$, where $\phi_{0,\mu}\in L^{2}(P_{0})$ is such that, for some $L\in(0,\infty)$, the inequality $\|\phi_{0,\mu_{1}}-\phi_{0,\mu_{2}}\|\leq L\|\mu_{1}-\mu_{2}\|$ holds $P_{0}$–almost surely for each $\mu_{1},\mu_{2}\in\mathcal{H}$;
We also assume that we can construct a well-behaved estimator $\psi_{n}$ of $\psi_{0}$ based on the available data, such that $\psi_{n}(\mu)=\psi_{0}(\mu)+(P_{n}-P_{0})\widetilde{\phi}_{0,\mu}+o_{p}(n^{-\frac{1}{2}})$ locally uniformly around each $\mu\in\mathcal{H}$—this requirement is stated precisely as Condition C7 later.
The target parameter $\Psi(P):=\psi_{P}(\mu_{P})$ we consider encompasses various interesting parameters that exhibit linear dependence on the outcome regression. Indeed, many estimands of interest in causal inference, such as counterfactual means and the ATE, correspond to functionals of the form
$\mu\mapsto\psi_{0}(\mu)=E_{0}\left\{m(Z,\mu)\right\}$ (1)
for a known map $m:\mathcal{Z}\times\mathcal{H}\rightarrow\mathbb{R}$. For example, the counterfactual mean outcome under an intervention that sets treatment to level $a_{0}\in\mathcal{A}$ is obtained by taking $m:(z,\mu)\mapsto\mu(a_{0},w)$, whereas the ATE is obtained by taking $m:(z,\mu)\mapsto\mu(1,w)-\mu(0,w)$. For such functionals, $\psi_{0}$ satisfies A1 if $m$ is linear in its last argument and there exists $c<\infty$ such that, for each $\mu\in\mathcal{H}$, $\|m(\cdot,\mu)\|\leq c\|\mu\|$. Furthermore, A3 is satisfied with $\phi_{0,\mu}:=m(\cdot,\mu)$, and an asymptotically linear estimator of $\psi_{0}(\mu)$ is given by $\psi_{n}(\mu):=\frac{1}{n}\sum_{i=1}^{n}m(Z_{i},\mu)$. Specifically, for the ATE functional, this yields $\psi_{n}(\mu)=\frac{1}{n}\sum_{i=1}^{n}\{\mu(1,W_{i})-\mu(0,W_{i})\}$. Additional examples of such functionals can be found in Chernozhukov et al. (2018b) and Rotnitzky et al. (2021).
By the Riesz representation theorem, there exists a unique Riesz representer $\alpha_{0}\in\mathcal{H}$ such that $\psi_{0}(\mu)=\langle\alpha_{0},\mu\rangle$ for all $\mu$. The Riesz representer can be written as a minimizer of the objective function $\alpha\mapsto\|\alpha\|^{2}-2\psi_{0}(\alpha)$ over $\mathcal{H}$ (Chernozhukov et al., 2018b, 2021). It plays an important role in the construction of asymptotically linear and nonparametric efficient estimators of $\tau_{0}$. In particular, $\alpha_{0}$ appears in the efficient influence function of the parameter $P\mapsto\psi_{P}(\mu_{P})$ at $P=P_{0}$, viewed as a functional on the nonparametric statistical model $\mathcal{M}$, and this plays a critical role in the debiasing of plug-in estimators based on the functional $\psi_{n}$. Below, for any $\mu,\alpha\in\mathcal{H}$, we denote by $D_{\mu,\alpha}$ the mapping $z\mapsto\alpha(a,w)\left\{y-\mu(a,w)\right\}$.
Theorem 1 (pathwise differentiability).
Under A1A3, $P\mapsto\psi_{P}(\mu_{P})$ is pathwise differentiable at $P_{0}$ under the model $\mathcal{M}$ and has nonparametric efficient influence function $D_{0}:=\widetilde{\phi}_{0,\mu_{0}}+D_{\mu_{0},\alpha_{0}}$ with $\widetilde{\phi}_{0,\mu_{0}}$ defined in A3. Furthermore, for any $\mu,\alpha\in\mathcal{H}$, $\psi_{0}$ admits the representation
$\psi_{0}(\mu)-\psi_{0}(\mu_{0})=-P_{0}D_{\mu,\alpha}+\langle\mu-\mu_{0},\alpha_{0}-\alpha\rangle\,.$
Below we provide examples of pathwise differentiable parameters covered by our framework.
Example 1 (counterfactual means under general interventions).
For simplicity, suppose $\mathcal{A}$ is discrete, and denote the propensity score by $\pi_{0}(a\mid w):=P_{0}(A=a\mid W=w)$. Let $Y^{a_{0}}$ denote the potential outcome that would be observed if treatment $a_{0}\in\mathcal{A}$ were administered. Under causal conditions (Rubin, 1974), the mean of $Y^{a_{0}}$ is identified by $\tau_{0}:=E_{0}\{\mu_{0}(a_{0},W)\}$, which corresponds to the linear functional $\psi_{0}:\mu\mapsto E_{0}\{\mu(a_{0},W)\}$. Doubly robust inference for this parameter was considered in van der Laan (2014) and Benkeser et al. (2017). More generally, the mean of the counterfactual outcome under a stochastic intervention that, given $W=w$, draws a treatment value $A^{*}$ at random from the distribution with probability mass function $a\mapsto\omega_{0}(a|w)$ depending only on $P_{0,A,W}$ can be identified by $\tau_{0}:=\iint\mu_{0}(a,w)\omega_{0}(a|w)P_{0,A,W}(da,dw)$. If $(a,w)\mapsto\alpha_{0}(a,w):=\omega_{0}(a|w)/\pi_{0}(a|w)$ has finite $\mathcal{H}$–norm, then conditions A1A3 hold and $\alpha_{0}$ is the nonparametric Riesz representer of $\tau_{0}$.
Example 2 (Partial covariance and semiparametric ATE estimation).
Consider the reduced data structure $Z=(W,Y)$ with $\mathcal{A}:=\emptyset$, and let $X\in\mathbb{R}$ be another random variable. Since the treatment $A$ does not appear in this setting, we write the outcome regression as $\mu_{0}(w):=E_{0}[Y\mid W=w]$. The partial covariance between $X$ and $Y$, adjusted for $W$, is defined as $\tau_{0}=E_{0}\left[\{X-e_{0}(W)\}\{Y-\mu_{0}(W)\}\right]$, where $e_{0}(w):=E_{0}[X\mid W=w]$. This parameter arises in semiparametric ATE estimation in partially linear models, where $X$ is a real-valued treatment. To apply our framework, we observe that this quantity simplifies to $\tau_{0}=E_{0}[XY]-E_{0}[X\mu_{0}(W)]$. The first term can be efficiently estimated by the empirical plug-in $\frac{1}{n}\sum_{i=1}^{n}X_{i}Y_{i}$. The second is a linear functional of the regression function $\mu_{0}$, given by $\psi_{0}:\mu\mapsto E_{0}[X\mu(W)]$. The linear functional $\psi_{0}$ has Riesz representer $e_{0}$, and can be efficiently estimated by the empirical plug-in $\psi_{n}:\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}X_{i}\mu(W_{i})$. Consequently, the term $E_{0}[X\mu_{0}(W)]$ corresponds to an estimand within our framework, and the methods we develop apply to doubly robust inference for the partial covariance parameter.
Technically, the functional $\psi_{0}$ depends on the joint distribution of $X$ and $Z$, and not solely on the distribution $P_{0}$ of $Z$; however, this does not affect the applicability of our approach. We study inference for a specific instance of this parameter in Appendix E. Related work on doubly robust inference for this parameter includes Dukes et al. (2021) and Bonvini et al. (2024).
In Appendix E, we show that the average treatment effect under covariate-dependent outcome missingness also falls within our framework. Doubly robust inference for this estimand was previously considered in Díaz and van der Laan (2017).
2.2 One-step debiased estimation and rate double robustness
Suppose we are given arbitrary estimators $\mu_{n}$ and $\alpha_{n}$ of the outcome regression $\mu_{0}$ and Riesz representer $\alpha_{0}$. For example, as in Chernozhukov et al. (2018b, 2021), the estimators $\mu_{n}$ and $\alpha_{n}$ can be learned directly using any supervised machine learning algorithm based on the least-squares loss $(z,\mu)\mapsto\{y-\mu(a,w)\}^{2}$ and the Riesz loss $(z,\alpha)\mapsto\alpha^{2}(a,w)-2\psi_{n}(\alpha)$, respectively. Given these nuisance estimators, the plug-in estimator $\psi_{n}(\mu_{n})$ could be used to estimate $\tau_{0}$. However, when $\mu_{n}$ is obtained using flexible statistical learning tools, this estimator typically lacks both $n^{\frac{1}{2}}$–consistency and asymptotic normality due to excessive bias that arises from the first-order dependence of $\psi_{n}(\mu_{n})-\psi_{0}(\mu_{0})$ on the nuisance estimation error $\mu_{n}-\mu_{0}$ (van der Laan and Rose, 2011; Chernozhukov et al., 2018b). To overcome this sensitivity, debiasing methods must typically be used to eliminate the first-order bias of the plug-in estimator $\psi_{n}(\mu_{n})$. For example, the one-step debiased estimator (Chernozhukov et al., 2018b, 2021)
$\tau_{n}:=\psi_{n}(\mu_{n})+\frac{1}{n}\sum_{i=1}^{n}\alpha_{n}(A_{i},W_{i})\{Y_{i}-\mu_{n}(A_{i},W_{i})\}$
could be used. The incorporation of this debiasing step renders the resulting one-step debiased estimator $\tau_{n}$ insensitive not only to the nuisance estimation error $\mu_{n}-\mu_{0}$ but also to $\alpha_{n}-\alpha_{0}$, thereby allowing for fast estimation rates and inference for $\tau_{0}$, even when $\alpha_{0}$ and $\mu_{0}$ are estimated at comparatively slow rates. The above estimator is a generalization of the well-known augmented inverse probability weighted estimator (Robins et al., 1994, 1995; Bang and Robins, 2005) and has also been referred to as an augmented balancing weighted (ABW) estimator (Bruns-Smith et al., 2023).
Suppose that both $\alpha_{n}$ and $\mu_{n}$ tend to possibly incorrect limits $\overline{\alpha}_{0}$ and $\overline{\mu}_{0}$, in the sense that $\|\alpha_{n}-\overline{\alpha}_{0}\|=o_{p}(1)$ and $\|\mu_{n}-\overline{\mu}_{0}\|=o_{p}(1)$, with either $\overline{\mu}_{0}=\mu_{0}$ or $\overline{\alpha}_{0}=\alpha_{0}$. To understand the asymptotic properties of the one-step debiased estimator $\tau_{n}$, it is useful to scrutinize the expansion in Theorem 1, from which the doubly robust decomposition of the estimation error
$\tau_{n}-\tau_{0}=(P_{n}-P_{0})(\widetilde{\phi}_{0,\mu_{n}}+D_{\mu_{n},\alpha_{n}})+\text{Rem}_{\mu_{n}}(\psi_{n},\psi_{0})+\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ (2)
can be derived, where $D_{\mu_{n},\alpha_{n}}$ is an estimator of the efficient influence function component $D_{\mu_{0},\alpha_{0}}$ in Theorem 1, $\text{Rem}_{\mu_{n}}(\psi_{n},\psi_{0}):=\psi_{n}(\mu_{n})-\psi_{0}(\mu_{n})-(P_{n}-P_{0})\widetilde{\phi}_{0,\mu_{n}}$ is a linearization remainder, and $\widetilde{\phi}_{0,\mu_{n}}$ is as defined in A3. Under weak conditions, the first term in the decomposition is asymptotically linear, satisfying $(P_{n}-P_{0})(\widetilde{\phi}_{0,\mu_{n}}+D_{\mu_{n},\alpha_{n}})=(P_{n}-P_{0})\overline{D}_{0}+o_{p}(n^{-\frac{1}{2}})$, where the influence function $\overline{D}_{0}:=\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}$ reduces to the efficient influence function in Theorem 1 when both $\overline{\mu}_{0}=\mu_{0}$ and $\overline{\alpha}_{0}=\alpha_{0}$. The asymptotic linearity of $\psi_{n}$ in a locally uniform sense implies that the second term is asymptotically negligible, in that $\text{Rem}_{\mu_{n}}(\psi_{n},\psi_{0})=o_{p}(n^{-\frac{1}{2}})$, and is, in fact, identically zero for functionals of the form given in (1). Consequently, even in the case of inconsistent estimation of one nuisance function, it can typically be established, under regularity conditions, that
$\tau_{n}-\tau_{0}=(P_{n}-P_{0})\overline{D}_{0}+o_{p}(n^{-\frac{1}{2}})+\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle\,.$ (3)
As a result, the bias of the one-step debiased estimator is primarily driven by the cross-product (or mixed) bias term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$, whose absolute value is bounded above by $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|$ in view of the Cauchy-Schwarz inequality. Most of the existing literature aims to show that $\tau_{n}$ is asymptotically linear with influence function $\overline{D}_{0}$. When $\tau_{n}$ admits the above representation, this is only possible if $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle=o_{p}(n^{-\frac{1}{2}})$. To ensure that this rate condition holds, the product rate condition $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{2}})$ is often assumed, which typically requires consistent estimation of both nuisance functions at sufficiently fast rates (Robins et al., 1994, 1995; van der Laan and Robins, 2003; Bang and Robins, 2005; Chernozhukov et al., 2018a). In particular, this condition holds if both $\|\alpha_{n}-\alpha_{0}\|=o_{p}(n^{-\frac{1}{4}})$ and $\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{4}})$.
As a consequence of (3), the one-step debiased estimator $\tau_{n}$ is said to be rate doubly robust (Rotnitzky et al., 2021) in the nuisance estimation errors in that asymptotic linearity and efficiency can be achieved if $\|\alpha_{n}-\alpha_{0}\|=o_{p}(1)$, $\|\mu_{n}-\mu_{0}\|=o_{p}(1)$ and $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{2}})$. Moreover, $\tau_{n}$ is doubly robust in terms of consistency: it converges in probability to $\tau_{0}$ as long as $\|\alpha_{n}-\alpha_{0}\|=o_{p}(1)$ or $\|\mu_{n}-\mu_{0}\|=o_{p}(1)$. While it maintains consistency under inconsistent estimation of either the outcome regression $\mu_{0}$ or Riesz representer $\alpha_{0}$, the one-step debiased estimator typically lacks desirable properties necessary for inference, such as $n^{\frac{1}{2}}$–consistency and asymptotic normality, in such scenarios. This is also the case when both nuisance estimators are estimated consistently but at insufficiently fast rates so that the condition $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-\frac{1}{2}})$ does not hold. Consequently, despite rate double robustness, $\tau_{n}$ is typically only asymptotically linear when $\mu_{0}$ and $\alpha_{0}$ are sufficiently structured (e.g., smooth or sparse) to be estimated at fast enough rates, and this represents a stringent condition that may fail to hold in practice.
3 Preliminaries on doubly robust inference
3.1 Statistical objective: doubly robust asymptotic linearity
In this work, we aim to construct estimators of $\tau_{0}$ that are doubly robust asymptotically linear (DRAL), meaning they remain asymptotically linear whenever either
  1. (i)
    $\|\alpha_{n}-\alpha_{0}\|=o_{p}(n^{-1/4})$ and $\mu_{n}$ converges in probability to some $\overline{\mu}_{0}$, or
  2. (ii)
    $\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-1/4})$ and $\alpha_{n}$ converges in probability to some $\overline{\alpha}_{0}$.
The DRAL property improves upon rate double robustness, as it allows one nuisance estimator to converge arbitrarily slowly or even be inconsistent. It does this by relaxing the usual product rate condition $\|\alpha_{n}-\alpha_{0}\|\cdot\|\mu_{n}-\mu_{0}\|=o_{p}(n^{-1/2})$ to $\|\alpha_{n}-\alpha_{0}\|^{2}\wedge\|\mu_{n}-\mu_{0}\|^{2}=o_{p}(n^{-1/2})$.
To illustrate the improvement over rate double robustness, suppose that $\mu_{0}$ and $\alpha_{0}$ are known only to be Hölder smooth with respective exponents $\beta>0$ and $\gamma>0$. In such cases, achieving asymptotic linearity under rate double robustness generally requires that $\beta+\gamma>d$, where $d$ is the dimension of the covariate vector $W$ (Robins et al., 2008). Thus, under rate double robustness, if one of the nuisance functions is insufficiently smooth so that the minimum nuisance smoothness satisfies $\beta\wedge\gamma\ll d/2$, achieving asymptotic linearity typically necessitates the other nuisance function to exhibit a high degree of smoothness, which may not be satisfied by the data-generating mechanism in many applications. In contrast, achieving asymptotic linearity under the DRAL property only places a requirement on the maximum nuisance smoothness, $\beta\vee\gamma>d/2$. As a result, DRAL estimators can maintain asymptotic linearity even when one of the nuisance functions is arbitrarily non-smooth, as long as the other nuisance function has smoothness exponent larger than $d/2$.
As illustrated in the previous section, the lack of normality of the one-step debiased estimator under inconsistent or insufficiently fast estimation of one of the nuisance functions occurs because in such cases the cross-product remainder term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ becomes sensitive to the estimation rate of the consistent nuisance estimator. This sensitivity leads to the cross-product remainder term contributing to the asymptotic bias and variance of $n^{\frac{1}{2}}(\tau_{n}-\tau_{0})$. Therefore, to construct a DRAL estimator, it is critical to ensure that the bias from the cross-product term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ in (2) is corrected for, making it either asymptotically linear or negligible, even if one of the nuisance estimators converges to its target too slowly or inconsistently, in the sense that $\|\alpha_{n}-\alpha_{0}\|\|\mu_{n}-\mu_{0}\|\neq o_{p}(n^{-\frac{1}{2}})$. In the remainder of this section, we examine this bias term in greater detail, illustrating how additional debiasing efforts can render it asymptotically linear in such cases. Readers primarily interested in the details of the proposed method rather than the core requirement underlying the DRAL property may wish to proceed directly to Section 4.
3.2 General strategy: debiasing to linearize the cross-product remainder term
We now illustrate how additional debiasing of the nuisance estimators $\alpha_{n}$ and $\mu_{n}$ linearizes the cross-product remainder term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$. Our key idea, rooted in van der Laan (2014), is to express the cross-product remainder term $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ solely in terms of the most favorable nuisance estimator and the corresponding nuisance function. Here, ‘most favorable’ refers to the nuisance estimator that estimates its intended target at the fastest rate. We can then show that if the most favorable nuisance estimator satisfies certain empirical orthogonality (or score) equations, the cross-product remainder term is debiased and asymptotically linear under relatively weak conditions. This strategy was originally used in van der Laan (2014) and Benkeser et al. (2017) to construct DRAL estimators of the ATE for the special case where one nuisance is estimated inconsistently. However, the regimes in which both nuisances are consistently estimated but one is estimated too slowly were not studied. In this subsection, in addition to studying a broad class of statistical parameters, we establish the asymptotic linearity of the remainder in general, and provide a refined analysis to establish the asymptotic negligibility of the remainder in such regimes.
To provide intuition and elucidate key ideas, we first present an informal analysis of the cross-product remainder term for the case where $\mu_{n}$ is most favorable, which will be useful when the outcome regression is consistently estimated but the Riesz representer is estimated inconsistently or at a slower rate. We first introduce notation. For any vector-valued function $v:\mathcal{A}\times\mathcal{W}\rightarrow\mathbb{R}^{k}$ with $k\in\mathbb{N}$, we define the conditional expectation projection operator $\Pi_{v}:\mathcal{H}\rightarrow\mathcal{H}$ as $\Pi_{v}f:=\operatorname*{argmin}_{\theta\in\Theta_{v}}\|f-\theta\|$, where $\Theta_{v}$ is the set of functions of the form $h\circ v$ for any $h:\mathbb{R}\rightarrow\mathbb{R}$. Whenever $v$ and $f$ are nonrandom functions, we have that $\Pi_{v}f:(a,w)\mapsto E_{0}\{f(A,W)\,|\,v(A,W)=v(a,w)\}$. For brevity, we denote the orthogonal projections of the error terms $\alpha_{0}-\alpha_{n}$ and $\mu_{0}-\mu_{n}$ onto $\mu_{n}$ and $\alpha_{n}$, respectively, as $s_{n,0}:=\Pi_{\mu_{n}}(\alpha_{0}-\alpha_{n})$ and $r_{n,0}:=\Pi_{\alpha_{n}}(\mu_{0}-\mu_{n})$. Following the strategy of van der Laan (2014), when $\mu_{n}$ is consistent for $\mu_{0}$, and $\alpha_{n}$ is inconsistent or converges to $\alpha_{0}$ slower than $\mu_{n}$ tends to $\mu_{0}$, we aim to express the cross-product remainder $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ solely in terms of $\mu_{n}$ and $\mu_{0}$ by applying the law of total expectation to condition on the random vector $(\mu_{n}(A,W),\mu_{0}(A,W))$ within the $P_{0}$–expectation. Doing so, we find that the cross-product remainder term can be written as
$\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ $=\ \langle\Pi_{(\mu_{n},\mu_{0})}(\alpha_{0}-\alpha_{n}),\mu_{n}-\mu_{0}\rangle$
[…]
where, informally, the higher-order terms arise because $\Pi_{(\mu_{n},\mu_{0})}\approx\Pi_{\mu_{n}}$ whenever $\mu_{n}\approx\mu_{0}$ under regularity conditions. The key bias term precluding asymptotic linearity of the one-step debiased estimator under inconsistent or insufficiently fast estimation of $\alpha_{0}$ is driven by the population moment expression $\langle s_{n,0},\mu_{n}-\mu_{0}\rangle$ for the outcome nuisance estimator $\mu_{n}$. To analyze this term further, we note that
[…]
$=\ (P_{n}-P_{0})A_{n,0}-P_{n}A_{n,0}\,,$
where $A_{n,0}:z\mapsto s_{n,0}(a,w)\{y-\mu_{n}(a,w)\}$. The first term is a drift term that is asymptotically linear under conditions, satisfying $(P_{n}-P_{0})A_{n,0}=P_{n}A_{0}+o_{p}(n^{-\frac{1}{2}})$ with $A_{0}:z\mapsto s_{0}(a,w)\{y-\mu_{0}(a,w)\}$ and $s_{0}:=\Pi_{\mu_{0}}(\alpha_{0}-\overline{\alpha}_{0})$. The second term is typically a first-order bias term that precludes the asymptotic linearity of the cross-product remainder in this case.
As a consequence, to debias the cross-product remainder with respect to the estimation error of $\alpha_{n}$, it suffices that the residuals $Y_{1}-\mu_{n}(A_{1},W_{1}),\ldots,Y_{n}-\mu_{n}(A_{n},W_{n})$ for $\mu_{0}$ be orthogonalized in $L^{2}(P_{n})$ to the projected residuals $s_{n,0}(A_{1},W_{1}),\ldots,s_{n,0}(A_{n},W_{n})$ for $\alpha_{0}$, in the sense that
$P_{n}A_{n,0}=\frac{1}{n}\sum_{i=1}^{n}s_{n,0}(A_{i},W_{i})\{Y_{i}-\mu_{n}(A_{i},W_{i})\}=0\,.$ (4)
A similar analysis can be applied to the cross-product remainder term for the case where the Riesz representer estimator $\alpha_{n}$ is most favorable by instead conditioning on $(\alpha_{n},\alpha_{0})$. In particular, when $\alpha_{n}$ is consistent for $\alpha_{0}$ but $\mu_{n}$ is either inconsistent or converges to $\mu_{0}$ slower than $\alpha_{n}$ tends to $\alpha_{0}$, the estimator $\alpha_{n}$ must satisfy the empirical orthogonality condition
$\psi_{n}(r_{n,0})-\frac{1}{n}\sum_{i=1}^{n}r_{n,0}(A_{i},W_{i})\alpha_{n}(A_{i},W_{i})=0$ (5)
for the cross-product remainder term to be debiased with respect to the estimation error of $\mu_{n}$.
We now formalize the heuristic justification for why (4) and (5) suffice to debias the cross-product remainder. In the next section, we introduce a simple post-hoc method for constructing nuisance estimators that exactly satisfy these properties. For now, we assume such estimators $\mu_{n}$ and $\alpha_{n}$ are available. We define $B_{0}:=\psi_{0}(r_{0})+\widetilde{\phi}_{r_{0}}-r_{0}\alpha_{0}$ with $r_{0}:=\Pi_{\alpha_{0}}(\mu_{0}-\overline{\mu}_{0})$, and $B_{n,0}:=\psi_{0}(r_{n})+\widetilde{\phi}_{r_{n,0}}-r_{n,0}\alpha_{n}$, and recall the definitions $A_{0}:z\mapsto s_{0}(a,w)\{y-\mu_{0}(a,w)\}$ and $A_{n,0}:z\mapsto s_{n,0}(a,w)\{y-\mu_{n}(a,w)\}$. We also denote by $\Delta_{\alpha,n}:=\alpha_{n}-\alpha_{0}$ and $\Delta_{\mu,n}:=\mu_{n}-\mu_{0}$ the nuisance estimation errors. The following theorem relies on a condition that relates the estimation error in the projection operators to the nuisance estimation errors $\|\mu_{n}-\mu_{0}\|$ and $\|\alpha_{n}-\alpha_{0}\|$. We note that the theorem itself does not require the nuisance estimators $\mu_{n}$ and $\alpha_{n}$ to converge to the functions $\overline{\mu}_{0}$ and $\overline{\alpha}_{0}$, but all subsequent results and the following discussion do.
  1. B1)
    Coupling between errors:
    1. i)
      $\|(\Pi_{\mu_{n}}-\Pi_{\mu_{0}})\Delta_{\alpha,n}\|+\|(\Pi_{\mu_{n},\mu_{0}}-\Pi_{\mu_{n}})\Delta_{\alpha,n}\|=O_{p}(\|\mu_{n}-\mu_{0}\|)$ ;
    2. ii)
      $\|(\Pi_{\alpha_{n}}-\Pi_{\alpha_{0}})\Delta_{\mu,n}\|+\|(\Pi_{\alpha_{n},\alpha_{0}}-\Pi_{\alpha_{n}})\Delta_{\mu,n}\|=O_{p}(\|\alpha_{n}-\alpha_{0}\|)$
Theorem 2 (Expansion of cross-product remainder under orthogonality).
If $\mu_{n}$ satisfies (4), the outcome regression-favorable decomposition
$\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ $=P_{n}A_{0}+(P_{n}-P_{0})(A_{n,0}-A_{0})-\langle(\Pi_{(\mu_{n},\mu_{0})}-\Pi_{\mu_{n}})\Delta_{\alpha,n},\,\Delta_{\mu,n}\rangle$
holds. If condition B1B1i holds, then $\langle(\Pi_{(\mu_{n},\mu_{0})}-\Pi_{\mu_{n}})\Delta_{\alpha,n},\,\Delta_{\mu,n}\rangle=O_{p}\left(\|\mu_{n}-\mu_{0}\|^{2}\wedge\|\mu_{n}-\mu_{0}\|\|\alpha_{n}-\alpha_{0}\|\right)$.
If $\alpha_{n}$ satisfies (5), the Riesz representer-favorable decomposition
$\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ $=P_{n}B_{0}+(P_{n}-P_{0})(B_{n,0}-B_{0})-\langle(\Pi_{(\alpha_{n},\alpha_{0})}-\Pi_{\alpha_{n}})\Delta_{\mu,n},\,\Delta_{\alpha,n}\rangle+Rem_{\psi_{n}}(r_{n,0})$
holds with $Rem_{\psi_{n}}(r_{n,0}):=\psi_{n}(r_{n,0})-\psi_{0}(r_{n,0})-(P_{n}-P_{0})\widetilde{\phi}_{r_{n,0}}$. If condition B1B1ii holds, then $\langle(\Pi_{(\alpha_{n},\alpha_{0})}-\Pi_{\alpha_{n}})\Delta_{\mu,n},\,\Delta_{\alpha,n}\rangle=O_{p}\left(\|\alpha_{n}-\alpha_{0}\|^{2}\wedge\|\mu_{n}-\mu_{0}\|\|\alpha_{n}-\alpha_{0}\|\right)$.
The second-order terms in both decompositions are second-order in the estimation error of their respective favorable nuisance function. As a consequence, when all conditions of the theorem hold, the bias term is second-order in the estimation rate of whichever nuisance function estimator converges fastest to its intended target. If both nuisance functions are consistently estimated, so that $\overline{\alpha}_{0}=\alpha_{0}$ and $\overline{\mu}_{0}=\mu_{0}$, then $A_{0}=B_{0}=0$ and the cross-product remainder can be expressed purely in terms of second-order remainders that are $o_{p}(n^{-\frac{1}{2}})$ under usual rate conditions. As a consequence, provided $\mu_{n}$ and $\alpha_{n}$ simultaneously satisfy (4) and (5) and the remainders are $o_{p}(n^{-\frac{1}{2}})$, Theorem 2 establishes that the cross-product remainder $\langle\alpha_{0}-\alpha_{n},\mu_{n}-\mu_{0}\rangle$ is a DRAL estimator of 0 with a possibly degenerate influence function $1(\overline{\alpha}_{0}\neq\alpha_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})B_{0}$. Furthermore, if the estimation error for $\tau_{n}$ satisfies the asymptotic expansion in (3), then the one-step debiased estimator $\tau_{n}$ of $\tau_{0}$ with suitably orthogonalized nuisance estimators $\mu_{n}$ and $\alpha_{n}$ will be DRAL with influence function $\overline{D}_{0}+1(\overline{\mu}_{0}\neq\mu_{0})A_{0}+1(\overline{\alpha}_{0}\neq\alpha_{0})B_{0}$.
High-level conditions similar to B1 have appeared in prior work on doubly robust inference, including van der Laan (2014), Benkeser et al. (2017), and Dukes et al. (2021). In Appendix A.1, we derive novel sufficient conditions based on bivariate Lipschitz continuity that formalize the heuristic that, under regularity, the joint projection $\Pi_{\mu_{n},\mu_{0}}$ can be well-approximated by the marginal projection $\Pi_{\mu_{n}}$ when $\mu_{n}$ consistently estimates $\mu_{0}$. These conditions are considerably weaker than the high-dimensional Hölder smoothness assumptions commonly imposed in higher-order estimation, requiring only Lipschitz continuity of certain low-dimensional conditional expectations. A related assumption was used in Bonvini et al. (2024).
4 Calibrated debiased machine learning
4.1 Obtaining doubly robust inference via calibration
[…]
To formalize this notion, let $(z,\nu)\mapsto\ell(z,\nu)$ be a loss function for some nuisance function $\nu_{0}\in\mathcal{H}$. An estimator $\nu_{n}$ is said to be $\ell$–empirically calibrated for $\nu_{0}$ if it satisfies the self-consistency property
$\frac{1}{n}\sum_{i=1}^{n}\ell(Z_{i},\nu_{n})=\min_{\theta}\frac{1}{n}\sum_{i=1}^{n}\ell(Z_{i},\theta\circ\nu_{n})\,,$ (6)
where the minimum is taken over all real-valued functions $\theta:\mathbb{R}\to\mathbb{R}$. That is, the empirical risk of $\nu_{n}$ cannot be improved through transformation by any one-dimensional mapping.
Calibration can be performed post hoc by applying a transformation to $\nu_{n}$ that minimizes empirical loss over a class of candidate functions (van der Laan and Alaa, 2025). Perfect empirical calibration can only be achieved through binning, partitioning the initial predictions and assigning calibrated values within each bin to minimize empirical risk (see Gupta et al. (2020); van der Laan and Alaa (2025) for details). Isotonic calibration is a data-adaptive binning method that imposes a monotonicity constraint to regularize the fitted transformation (Zadrozny and Elkan, 2001; Niculescu-Mizil and Caruana, 2005b; van der Laan et al., 2023b). Among available binning approaches, isotonic regression may be preferred because it performs adaptive binning without requiring parameter tuning, while simultaneously learning the optimal monotone transformation of the original predictor (see Appendix D for details).
When the loss $\ell$ is sufficiently smooth, the first-order equations characterizing the empirical risk minimization problem imply that $\ell$–empirical calibration is equivalent to requiring that, for each map $\theta:\mathbb{R}\rightarrow\mathbb{R}$,
[…]
The following theorem establishes that the empirical calibration of nuisance estimators is sufficient for the orthogonality conditions (4) and (5) to be exactly met, and is a direct consequence of (7) for suitable loss functions. In the theorem, we make use of the following condition:
  1. C1)
    (bounded linearity of estimated functional) it holds $P_{0}$–almost surely that $\psi_{n}(c\mu_{1}+\mu_{2})=c\psi_{n}(\mu_{1})+\psi_{n}(\mu_{2})$ for each $\mu_{1},\mu_{2}\in\mathcal{H}$ and $c\in\mathbb{R}$ and that $\sup_{h\in\mathcal{H}}\psi_{n}(h)|/\|h\|_{P_{n}}<\infty$;
Lemma 1 (Empirical calibration debiases cross-product remainder).
Suppose that Condition C1 holds. Suppose also that $\mu_{n}$ is empirically calibrated with respect to the least-squares loss $(z,\mu)\mapsto\{y-\mu(a,w)\}^{2}$ and that $\alpha_{n}$ is empirically calibrated with respect to the Riesz loss $(z,\alpha)\mapsto\alpha^{2}(a,w)-2\psi_{n}(\alpha)$, so that the calibration conditions
[…]
hold for any $\theta:\mathbb{R}\rightarrow\mathbb{R}$. Then, $\mu_{n}$ and $\alpha_{n}$ satisfy the empirical orthogonality conditions (4) and (5).
Lemma 1 demonstrates that empirical calibration suffices to simultaneously debias the cross-product remainder of the debiased machine learning estimator $\tau_{n}$ with respect to the nuisance estimation errors for $\mu_{0}$ and $\alpha_{0}$, so that the error decompositions in Theorem 2 can then be applied to study the behavior of $\tau_{n}$. To establish doubly robust asymptotic normality, it suffices that the orthogonality conditions in (4) and (5) are solved at rate $o_{p}(n^{-1/2})$, in which case approximate $\ell$-calibration suffices.
The DRAL properties established by Lemma 1 reflect that adjustment for either $\mu_{0}(A,W)$ or $\alpha_{0}(A,W)$ suffices to correct for confounding bias. Notably, when $\psi_{0}$ is a counterfactual mean, adjustment for $\alpha_{0}(A,W)$ is equivalent to adjustment for the propensity score $\pi_{0}(A,W)$, a well-known balancing score (Imai and Ratkovic, 2014). When $\mu_{n}$ is consistent at a sufficiently fast rate, the first calibration condition implies that the plug-in estimator $\psi_{n}(\mu_{n})$ is debiased for $\psi_{0}$ under a model that uses $\mu_{0}(A,W)$ as a dimension reduction of $(A,W)$ (Benkeser et al., 2020). Similarly, when $\alpha_{n}$ is consistent, the second calibration condition implies that the balancing weighted outcome estimator $\frac{1}{n}\sum_{i=1}^{n}\alpha_{n}(A_{i},W_{i})Y_{i}$ is debiased for $\psi_{0}$ under a model that uses $\alpha_{0}(A,W)$ as a dimension reduction. The calibrated DML estimator inherits the debiasedness of both the calibrated plug-in and weighted outcome estimators, leveraging either $\alpha_{0}(A,W)$ or $\mu_{0}(A,W)$ for bias correction provided that at least one nuisance estimator is consistent. We formalize this intuition in Appendix C.
Beyond enabling doubly robust inference, empirical calibration improves estimator stability and predictive accuracy. For squared error loss, it ensures the outcome regression $\mu_{n}(a,w)$ equals the average outcome among individuals with the same predicted value, while for the Riesz loss, it enforces covariate balance across estimated inverse propensity score strata (Deshpande and Kuleshov, 2023; van der Laan et al., 2024b).
4.2 Proposed estimator using isotonic calibration
We refer to the implementation of debiased machine learning with calibrated nuisance functions as calibrated DML. Lemma 1 establishes that calibration of nuisance functions bestows DML estimators with the DRAL property. We now describe a particular implementation of our general framework for calibrated DML, which outlines how to construct DRAL estimators using isotonic calibration and one-step debiased estimation.
Our post-hoc procedure employs isotonic regression (Barlow and Brunk, 1972) to empirically calibrate any given pair of nuisance estimators, $\mu_{n}$ and $\alpha_{n}$, such that they satisfy the empirical orthogonality equations (4) and (5). In view of Lemma 1, this calibration ensures that the cross-product remainder term is asymptotically linear even when one nuisance function is poorly estimated, leading the resulting debiased machine learning estimator to be DRAL. Our procedure is automatic, requiring no knowledge of the specific form of the linear functional $\psi_{0}$; it merely requires the ability to evaluate as a black box an asymptotically linear estimator $\psi_{n}(\mu)$ of $\psi_{0}(\mu)$ for any given fixed $\mu$. Furthermore, our method is non-iterative and can be efficiently implemented as a two-step procedure using widely available software for isotonic regression. As a consequence, calibrated DML approach can be easily integrated into any DML pipeline by simply adding a few lines of code to preprocess cross-fitted estimates and ensure calibration before debiasing.
Given initial outcome regression and Riesz representer estimators $\mu_{n}$ and $\alpha_{n}$, we use isotonic calibration to obtain
$\mu_{n}^{*}:=f_{n}\circ\mu_{n},\textnormal{ with }\ f_{n}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}Y_{i}-f(\mu_{n}(A_{i},W_{i}))\big{\}}^{2},$ (8)
$\alpha_{n}^{*}:=g_{n}\circ\alpha_{n},\textnormal{ with }\ g_{n}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}g(\alpha_{n}(A_{i},W_{i}))\big{\}}^{2}-2\psi_{n}(g\circ\alpha_{n})\,,$
where $\mathcal{F}_{\text{iso}}$ denotes the space of real-valued monotone non-decreasing functions defined over $\mathbb{R}$. Our proposed isotonic-calibrated (isocalibrated) DML estimator is then simply
$\tau_{n}^{*}:=\psi_{n}(\mu_{n}^{*})+\frac{1}{n}\sum_{i=1}^{n}\alpha_{n}^{*}(A_{i},W_{i})\{Y_{i}-\mu_{n}^{*}(A_{i},W_{i})\}\,,$ (9)
the debiased machine learning estimator based on these calibrated nuisance function estimators. Isotonic regression ensures that the calibrated nuisance estimators $\mu_{n}^{*}$ and $\alpha_{n}^{*}$ are $\ell$-empirically calibrated with respect to their respective loss functions, and thus satisfy the conditions of Lemma 1. This follows from a key property of isotonic regression: the resulting solutions $f_{n}$ and $g_{n}$ behave as histogram estimators over a data-dependent partition of the input space. In light of the infinite-dimensional score equations they solve, the calibrated estimators $\mu_{n}^{*}$ and $\alpha_{n}^{*}$ can be viewed as infinite-dimensional targeted minimum loss-based estimators of $\mu_{0}$ and $\alpha_{0}$; see, e.g., van der Laan and Gruber (2016), Luedtke et al. (2017), Qiu et al. (2021), and van der Laan et al. (2024a).
To obtain strong theoretical guarantees under weak conditions, we recommend cross-fitting the initial nuisance estimators (Algorithm 1). Cross-fitting is a widely used technique for mitigating overfitting and relaxing complexity constraints in nuisance estimation (van der Laan et al., 2011; Chernozhukov et al., 2018a). We note, however, that the isotonic calibration step in (8) should not be cross-fitted. Since the class of monotone functions is Donsker, the relevant empirical process conditions can be explicitly verified (Rabenseifner et al., 2025). As a result, applying isotonic calibration to the full sample does not compromise the theoretical guarantees provided by cross-fitting the initial nuisance estimators. Algorithm 1 outlines the procedure for the cross-fitted variant of the calibrated nuisance constructions. Specifically, it involves exhaustively partitioning the available data into $J$ disjoint folds, where for each fold $j$ initial estimators $\alpha_{n,j}$ and $\mu_{n,j}$ are computed using only the data in the complementary folds. The resulting estimators are then combined to obtain out-of-fold nuisance function estimators for each observation. The pooled out-of-fold estimators are then used to learn the calibrators $f_{n},g_{n}\in\mathcal{F}_{\text{iso}}$ in a similar manner as in (8). Finally, these calibrators are composed with the initial fold-specific nuisance estimators to construct a calibrated one-step debiased estimator of $\tau_{0}$. One benefit of calibrating the cross-fitted estimates using the full dataset is that it helps restore full-sample efficiency by leveraging observations excluded from each cross-fit fold.
0:  dataset $\mathcal{D}_{n}=\{O_{i}:i=1,\ldots,n\}$, number $J$ of cross-fitting splits
1:  partition $\mathcal{D}_{n}$ into datasets $\mathcal{C}^{(1)},\mathcal{C}^{(2)},\ldots,\mathcal{C}^{(J)}$;
2:  for $s=1,\ldots,J$ do
3:      get initial estimators $\alpha_{n,s}$ of $\alpha_{0}$ and $\mu_{n,s}$ of $\mu_{0}$ from $\mathcal{E}^{(s)}:=\mathcal{D}_{n}\backslash\mathcal{C}^{(s)}$;
4:      get initial estimators $\mu\mapsto\psi_{n,s}(\mu)$ of $\mu\mapsto\psi_{0}(\mu)$ from $\mathcal{D}_{n}$, where sample-splitting is performed as needed to satisfy Condition C7;
5:  end for
6:  for $s=1,\ldots,J$, set $j(i):=s$ for each $i\in\mathcal{C}^{(s)}$;
7:  compute calibrators $f_{n}$ and $g_{n}$ using pooled out-of-fold estimates as
$f_{n}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}Y_{i}-f(\mu_{n,j(i)}(A_{i},W_{i}))\big{\}}^{2}$
$g_{n}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}g(\alpha_{n,j(i)}(A_{i},W_{i}))^{2}-2\psi_{n,j(i)}(g\circ\alpha_{n,j(i)})\big{\}};$
8:  for $s=1,\ldots,J$, set $\mu_{n,s}^{*}:=f_{n}\circ\mu_{n,s}$ and $\alpha_{n,s}^{*}:=g_{n}\circ\alpha_{n,s}$;
9:  compute calibrated one-step debiased estimator
$\tau_{n}^{*}:=\frac{1}{J}\sum_{j=1}^{J}\left[\psi_{n,j}(\mu_{n,j}^{*})+\frac{1}{n}\sum_{i=1}^{n}\alpha_{n,j(i)}^{*}(A_{i},W_{i})\left\{Y_{i}-\mu_{n,j(i)}^{*}(A_{i},W_{i})\right\}\right].$ (10)
10:  return  $\tau_{n}^{*}$
Algorithm 1 Calibrated DML using isotonic calibration
The solution to an isotonic regression problem is typically nonunique. As in Groeneboom and Lopuhaa (1993), we take $f_{n}$ and $g_{n}$ to be the unique càdlàg piecewise constant solution of the isotonic regression problem with jumps occurring only at the values that $\mu_{n}$ and $\alpha_{n}$ take when evaluated on the observed data. The computation of $g_{n}$ in (8), used to solve the isotonic regression problem, may involve a nonstandard loss function. However, the calibrator $g_{n}$ can equivalently be represented, for the same loss function, as the empirical risk minimizer over all univariate regression trees of unbounded depth subject to the constraint that the trees are monotone nondecreasing. Hence, the isotonic solutions of (8) can be computed using any software that supports generalized regression trees with monotonicity constraints and custom losses, such as the R and Python implementations of xgboost (Chen and Guestrin, 2016). For details on implementing gradient-boosted trees to estimate Riesz representers, we refer to Lee and Schuler (2025). Even so, the following example illustrates that, for a wide range of causal parameters of interest, the estimator can be computed using standard implementations of isotonic regression.
Example 1 (continued).
Consider the $G$-computation identification $\tau_{0}(a_{0}):=E_{0}\{\mu_{0}(a_{0},W)\}$ of the mean of the counterfactual outcome $Y^{a_{0}}$ for $a_{0}\in\mathcal{A}$. In this case, the Riesz representer $\alpha_{0}$, given by $(a,w)\mapsto 1(a=a_{0})/\pi_{0}(a_{0}|w)$, is determined by the propensity score $\pi_{0}$. Given estimators $\mu_{n}$ of $\mu_{0}$ and $\pi_{n}$ of $\pi_{0}$, an isocalibrated DML estimator of $\tau_{0}(a_{0})$ (without cross-fitting) is given by
$\tau_{n}^{*}(a_{0}):=\frac{1}{n}\sum_{i=1}^{n}\left[\mu_{n}^{*}(a_{0},W_{i})+\frac{1(A_{i}=a_{0})}{\pi_{n}^{*}(a_{0}|W_{i})}\{Y_{i}-\mu_{n}^{*}(a_{0},W_{i})\}\right],$
where, for each $w\in\mathcal{W}$, we set $\mu_{n}^{*}(a_{0},w):=(f_{n,a_{0}}\circ\mu_{n})(a_{0},w)$ and $\pi_{n}^{*}(a_{0}|w):=(g_{n,a_{0}}\circ\pi_{n})(a_{0}|w)$ with
$f_{n,a_{0}}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\textnormal{iso}}}\sum_{i=1}^{n}1(A_{i}=a_{0})\left\{Y_{i}-f(\mu_{n}(a_{0},W_{i}))\right\}^{2};$
$g_{n,a_{0}}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\textnormal{iso}}}\sum_{i=1}^{n}\left\{1(A_{i}=a_{0})-g(\pi_{n}(a_{0}\mid W_{i}))\right\}^{2}.$
Here, we stratify the outcome regression calibration step in (8) by treatment, ensuring empirical calibration within each treatment level and potentially improving finite-sample performance. We also calibrate the propensity scores rather than the inverse propensity weights. The first-order conditions characterizing the isotonic regression solutions show that the inverse propensity weight calibration step in (8) is equivalent to isotonic propensity score calibration (van der Laan et al., 2024b). In particular, $g_{n,a_{0}}\in\operatorname*{argmin}_{g\in\mathcal{F}_{\textnormal{iso}}}\sum_{i=1}^{n}\left[\left\{\frac{1(A_{i}=a_{0})}{g(\pi_{n}(A_{i}|W_{i}))}\right\}^{2}-\frac{2}{g(\pi_{n}(a_{0}|W_{i}))}\right].$ The calibrated propensities $\{g_{n,a_{0}}(\pi_{n}(A_{i}\mid W_{i}))\}_{i=1}^{n}$ may be zero for observations with $A_{i}\neq a_{0}$; however, this is not problematic, as the AIPW estimator only requires evaluating these weights for observations with $A_{i}=a_{0}$.
For parameters that can be expressed as smooth transformations of counterfactual means, we can use the above to construct DRAL estimators of $\{\tau_{0}(a_{0}):a_{0}\in\mathcal{A}\}$, say $\{\tau_{n}^{*}(a_{0}):a_{0}\in\mathcal{A}\}$, and then compute the corresponding plug-in estimator of the estimand of interest, which will itself be DRAL. Inference can be obtained using either the delta method or the bootstrap, as detailed in the following sections. For example, DRAL estimators of the ATE $\tau_{0}(1)-\tau_{0}(0)$ and relative ATE $\tau_{0}(1)/\tau_{0}(0)$ are given by $\tau_{n}^{*}(1)-\tau_{n}^{*}(0)$ and $\tau_{n}^{*}(1)/\tau_{n}^{*}(0)$, respectively. Python and R implementations of the calibration procedure are provided in Appendix F.
Our approach shares some similarities with the automatic debiased machine learning estimators of Chernozhukov et al. (2018b, 2021, 2022), where similar tailored loss functions were used to construct cross-fitted estimators of the Riesz representer and outcome regression. However, our framework differs in its objective and use of these loss functions. Specifically, we make use of these loss functions to calibrate user-supplied cross-fitted nuisance estimators to ensure that the resulting one-step debiased estimator is DRAL. In contrast, Chernozhukov et al. (2018b) uses these loss functions to obtain nuisance estimators for the standard one-step debiased estimator. Notably, calibrated DML can be incorporated within the automatic DML framework by applying isotonic regression to calibrate the cross-fitted nuisance estimators.
4.3 Bootstrap-assisted confidence intervals
In previous works, doubly robust confidence intervals were obtained through a standard Wald construction based on the asymptotic normality of the corrected estimator, using adjusted estimators of the asymptotic variance (Benkeser et al., 2017; Dukes et al., 2021; Bonvini et al., 2024). Such confidence interval are valid if either nuisance estimator is consistent at a sufficiently fast rate. A drawback of previous approaches is that they require consistently estimating additional nuisance functions not required by the estimator itself. Here, we propose a new method that avoids this requirement. It uses a variant of the bootstrap to construct doubly robust confidence intervals associated with the estimator described in Algorithm 1 (Efron and Tibshirani, 1994).
Besides avoiding estimating new nuisances, our proposed bootstrap-assisted inference procedure is easy to describe and computationally efficient. Holding the initial cross-fitted nuisance estimators fixed, our procedure recomputes the calibrated DML estimator on bootstrap samples, only refitting the isotonic regressions and recomputing the empirical mean used to evaluate the DML estimator. In this way, each bootstrap replicate can be solved in near-linear time using standard software (Barlow and Brunk, 1972). Consequently, computing our bootstrap procedure is typically much less expensive than fitting the original DML estimator.
We focus our description of the automated algorithm on the case where the linear functional $\psi_{0}$ is of the form $\mu\mapsto E_{0}\{m(Z,\mu)\}$ for a known map $m:\mathcal{Z}\times\mathcal{H}\rightarrow\mathbb{R}$. In such cases, we consider the empirical plug-in estimator $\psi_{n}$ of $\psi_{0}$, defined as $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(Z_{i},\mu)$. A similar algorithm can be devised for the general case by appropriately bootstrapping the estimation of the functional $\psi_{n}$. As in the previous subsection, we first describe our approach in the simpler case where the initial estimators are not cross-fitted. The general procedure allowing for cross-fitting is provided in Algorithm 2. To this end, let $\alpha_{n}$ and $\mu_{n}$ be initial estimators $\alpha_{0}$ and $\mu_{0}$ computed once using the original sample. Our approach involves bootstrapping the meta isotonic fitting step of (8) used originally to obtain the calibrators $f_{n}$ and $g_{n}$. Specifically, we sample $\{Z_{i}^{\#}:i=1,\ldots,n\}$ with replacement from the original sample $\{Z_{i}:i=1,\ldots,n\}$ and compute the bootstrapped calibrators
$f_{n}^{\#}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}Y_{i}^{\#}-f(\mu_{n}(A_{i}^{\#},W_{i}^{\#}))\big{\}}^{2};$
$g_{n}^{\#}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}g(\alpha_{n}(A_{i}^{\#},W_{i}^{\#}))^{2}-2m(Z_{i}^{\#},g\circ\alpha_{n})\big{\}}\,.$
Next, we obtain the bootstrap-calibrated nuisance estimators $\alpha_{n}^{*\#}:=f_{n}^{\#}\circ\alpha_{n}$ and $\mu_{n}^{*\#}:=f_{n}^{\#}\circ\mu_{n}$, and compute the bootstrapped estimator $\tau_{n}^{*\#}$ of $\tau_{0}$ from $\mu_{n}^{*\#}$ and $\alpha_{n}^{*\#}$ following the same steps as in (9). This procedure is repeated $K$ times to obtain $K$ bootstrap estimators $\tau_{n,1}^{*\#},\ldots,\tau_{n,K}^{*\#}$ of $\tau_{0}$. The choice $K\approx 10,000$ has been recommended for use in practice (Efron and Tibshirani, 1994).
To construct a bootstrap-assisted confidence interval, we first compute the bootstrap average $\overline{\tau}_{n}^{*\#}:=\frac{1}{K}\sum_{k=1}^{K}\tau_{n,k}^{*\#}$. For given confidence level $\rho\in(0,1)$, we may consider the interval
$\text{CI}_{n}(\rho):=\left(\tau_{n}^{*}-q(\rho)\sigma_{n}^{\#},\,\tau_{n}^{*}+q(\rho)\sigma_{n}^{\#}\right),$
where $\sigma_{n}^{\#}$ is the empirical standard deviation of $\tau_{n,1}^{*\#},\ldots,\tau_{n,K}^{*\#}$ and, for each $\gamma\in(0,1)$, $q(\gamma)$ denotes the $\gamma/2$–quantile of the standard normal distribution. Alternatively, we could consider a percentile bootstrap interval $\left(\tau_{n}^{*}-q_{n}(\rho),\,\tau_{n}^{*}-q_{n}(2-\rho)\right)$, where for each $\gamma\in(0,1)$ $q_{n}(\gamma)$ denotes the empirical $\gamma/2$–quantile of the centered bootstrapped estimates $\tau_{n,1}^{*\#}-\overline{\tau}_{n}^{*\#},\ldots,\tau_{n,K}^{*\#}-\overline{\tau}_{n}^{*\#}$; this interval may be more robust to finite sample deviations from normality. We refer to Algorithm 2 for the pseudo-code implementing this approach with cross-fitting.
Algorithm 2 Bootstrap-assisted doubly robust CIs
0:  cross-fitted estimators $(\alpha_{n,j},\mu_{n,j})$ excluding fold $\mathcal{C}^{(j)}$ for $j=1,\ldots,J$, isocalibrated DML estimator $\tau_{n}^{*}$; confidence level $1-\rho$; bootstrap replication number $K$;
1:  for $k=1,\ldots,K$ do
2:      for $s=1,\ldots,J$, draw bootstrap sample $\mathcal{C}_{k}^{(s)\#}$ from $\mathcal{C}_{k}^{(s)}$;
3:      set $\mathcal{D}_{n,k}^{\#}:=\{Z_{1}^{\#},\dots,Z_{n}^{\#}\}=\{\mathcal{C}_{k}^{(s)\#}:s=1,\ldots,J\}$;
4:      for $s=1,\ldots,J$, set $j(i)=s$ for each $i\in\{1,\ldots,n\}$ such that $Z_{i}^{\#}\in\mathcal{C}_{k}^{(s)\#}$;
5:      compute calibrators $f^{\#(k)}_{n}$ and $g^{\#(k)}_{n}$ using pooled out-of-fold estimates as
$f_{n}^{\#(k)}$ $\in\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}Y_{i}^{\#}-f(\mu_{n,j(i)}(A_{i}^{\#},W_{i}^{\#}))\big{\}}^{2};$
$g_{n}^{\#(k)}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\big{\{}g(\alpha_{n,j(i)}(A_{i}^{\#},W_{i}^{\#}))^{2}-2m(Z_{i}^{\#},g\circ\alpha_{n,j(i)})\big{\}}\,;$
6:      for $s=1,\ldots,J$, set $\mu_{n,s}^{(k)\#}:=f_{n}^{(k)\#}\circ\mu_{n,s}$ and $\alpha_{n,s}^{(k)\#}:=g_{n}^{(k)\#}\circ\alpha_{n,s}$;
7:      compute calibrated one-step debiased estimator
$\tau_{n,k}^{*\#}:=\frac{1}{n}\sum_{i=1}^{n}\left[m(Z_{i}^{\#},\mu_{n,j(i)}^{(k)\#})+\alpha_{n,j(i)}^{(k)\#}(A_{i}^{\#},W_{i}^{\#})\left\{Y_{i}^{\#}-\mu_{n,j(i)}^{(k)\#}(A_{i}^{\#},W_{i}^{\#})\right\}\right];$
8:  end for
9:  set $\sigma_{n}^{\#2}:=\frac{1}{K}\sum_{k=1}^{K}(\tau_{n,k}^{*\#}-\overline{\tau}_{n}^{*\#})^{2}$ with $\overline{\tau}_{n}^{*\#}:=\frac{1}{K}\sum_{k=1}^{K}\tau_{n,k}^{*\#};$
10:  construct interval $\text{CI}_{n}(\rho):=(\tau_{n}^{*}-q(\rho)\sigma_{n}^{\#},\tau_{n}^{*}+q(\rho)\sigma_{n}^{\#});$
11:  return  bootstrap-assisted confidence interval $\text{CI}_{n}(\rho)$ for $\tau_{0}$.
5 Theoretical properties
5.1 Doubly robust asymptotic linearity and inference
We now present our main theorem on the doubly robust asymptotic linearity of the isocalibrated DML estimator proposed in Alg. 1. The limiting distribution of the scaled and centered version of the cross-fitted isocalibrated DML estimator involves the additional nuisance functions $s_{0}$ and $r_{0}$, defined pointwise as
$s_{0}(a,w)$ $:=E_{0}\left\{\alpha_{0}(A,W)-\overline{\alpha}_{0}(A,W)\,|\,\mu_{0}(A,W)=\mu_{0}(a,w)\right\};$
$r_{0}(a,w)$ $:=E_{0}\left\{\mu_{0}(A,W)-\overline{\mu}_{0}(A,W)\,|\,\alpha_{0}(A,W)=\alpha_{0}(a,w)\right\}.$
Let $P_{n,j}$ denote the empirical measure corresponding to the $j$-th fold $\mathcal{C}^{(j)}$ in the cross-fitting procedure. For each $j=1,\ldots,J$, define the nuisance estimation errors by $\Delta_{\alpha,n,j}:=\alpha_{n,j}^{*}-\alpha_{0}$ and $\Delta_{\mu,n,j}:=\mu_{n,j}^{*}-\mu_{0}$. Define the projection residual as $r_{n,j}^{*}:=\Pi_{\alpha_{n,j}^{*}}(\mu_{0}-\mu_{n,j}^{*}).$ We rely on the following conditions:
  1. C2)
    At least one nuisance is well-estimated: For some $(\overline{\alpha}_{0},\overline{\mu}_{0})\in\mathcal{H}^{2}$, either of the following holds:
    1. i)
      $\|\alpha_{n,j}^{*}-\alpha_{0}\|=o_{p}(n^{-1/4})$ and $\|\mu_{n,j}^{*}-\overline{\mu}_{0}\|=o_{P}(1)$ for $j=1,\ldots,J$;
    2. ii)
      $\|\mu_{n,j}^{*}-\mu_{0}\|=o_{p}(n^{-1/4})$ and $\|\alpha_{n,j}^{*}-\overline{\alpha}_{0}\|=o_{P}(1)$ for $j=1,\ldots,J$;
Under C2C2i, we have $\overline{\alpha}_{0}=\alpha_{0}$, and under C2C2ii, we have $\overline{\mu}_{0}=\mu_{0}$.
  1. C3)
    Projection error bounds: For each $j\in[J]$, both of the following hold:
    1. i)
      $\|(\Pi_{\mu_{n,j}^{*}}-\Pi_{\mu_{0}})\Delta_{\alpha,n,j}\|+\|(\Pi_{\mu_{n,j}^{*},\mu_{0}}-\Pi_{\mu_{n,j}^{*}})\Delta_{\alpha,n,j}\|=O_{p}(\|\mu_{n,j}^{*}-\mu_{0}\|)$;
    2. ii)
      $\|(\Pi_{\alpha_{n,j}^{*}}-\Pi_{\alpha_{0}})\Delta_{\mu,n,j}\|+\|(\Pi_{\alpha_{n,j}^{*},\alpha_{0}}-\Pi_{\alpha_{n,j}^{*}})\Delta_{\mu,n,j}\|=O_{p}(\|\alpha_{n,j}^{*}-\alpha_{0}\|)$.
  2. C4)
    Cross-fitted version of C1: For each $j\in[J]$, it holds $P_{0}$–almost surely that $\psi_{n,j}(c\mu_{1}+\mu_{2})=c\psi_{n,j}(\mu_{1})+\psi_{n,j}(\mu_{2})$ for each $\mu_{1},\mu_{2}\in\mathcal{H}$ and $c\in\mathbb{R}$ and that $\sup_{h\in\mathcal{H}}\psi_{n,j}(h)|/\|h\|_{P_{n,j}}<\infty$;
  3. C5)
    Boundedness of nuisance functions: for some $M\in(0,\infty)$, the functions $\mu_{n,j}^{*}$, $\alpha_{n,j}^{*}$, $\mu_{0}$, $\alpha_{0}$, $\overline{\mu}_{0}$ and $\overline{\alpha}_{0}$ are uniformly bounded by $M$ with $P_{0}$–probability tending to one for each $j=1,\ldots,J$;
  4. C6)
    Complexity of calibrators: with $P_{0}$–probability tending to one, the isotonic calibrators $f_{n}$ and $g_{n}$ take at most $k(n)>0$ distinct values for some deterministic sequence $k(n)=o(n^{\frac{1}{2}}/\log n)$;
  5. C7)
    Negligibility of remainders: For all $j\in[J]$ and $(f_{n,j},f_{0})\in\{(\mu_{n,j}^{*},\overline{\mu}_{0}),(r_{n,j}^{*},r_{0})\}$,
    1. (a)
      $\psi_{n,j}(f_{n,j})-\psi_{0}(f_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{0,f_{n,j}}=o_{p}(n^{-\frac{1}{2}})$.
    2. (b)
      $(P_{n,j}-P_{0})\{\widetilde{\phi}_{0,f_{n,j}}-\widetilde{\phi}_{0,f_{0}}\}=o_{p}(n^{-\frac{1}{2}})$.
In the following theorem, we let
$\chi_{0}={\phi}_{0,\overline{\mu}_{0}}-\psi_{0}(\overline{\mu}_{0})+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}+1(\overline{\alpha}_{0}\neq\alpha_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})B_{0}$
with $A_{0}:=(w,a,y)\mapsto s_{0}(a,w)\{y-\mu_{0}(a,w)\}$ and $B_{0}:=\widetilde{\phi}_{0,r_{0}}-r_{0}\alpha_{0}+\psi_{0}(r_{0})$. The estimator $\tau_{n}^{*}$ of $\tau_{0}$ is called asymptotically linear with influence function $\chi_{0}$ if $\tau_{n}^{*}-\tau_{0}=(P_{n}-P_{0})\chi_{0}+o_{p}(n^{-1/2})$.
Theorem 3 (Main result: DRAL of $\tau_{n}^{*}$).
Suppose A1A3C3C7, and a nuisance is estimated well (C2).
  1. i)
    The isocalibrated DML estimator $\tau_{n}^{*}$ of $\tau_{0}$ is asymptotically linear with influence function $\chi_{0}$.
  2. ii)
    If also $(\overline{\alpha}_{0},\overline{\mu}_{0})=(\alpha_{0},\mu_{0})$, then $\chi_{0}=D_{0}$, so that $\tau_{n}^{*}$ is regular and nonparametric efficient.
Any asymptotically linear estimator is asymptotically normal. Hence, the above shows $n^{\frac{1}{2}}(\tau_{n}^{*}-\tau_{0})$ tends in distribution to a mean-zero normal random variable with variance $\sigma_{0}^{2}:=var_{0}\{\chi_{0}(Z)\}$, provided $\|\mu_{n,j}^{*}-\mu_{0}\|\wedge\|\alpha_{n,j}^{*}-\alpha_{0}\|=o_{p}(n^{-1/4})$ and other conditions hold. Given a consistent estimator of $\sigma_{0}^{2}$, inference can be conducted using Wald-type confidence intervals and tests. The form of the influence function $\chi_{0}$ differs by either $A_{0}$ or $B_{0}$ from the efficient influence function $\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}$ under sampling from any distribution $\overline{P}_{0}$ compatible with the limits of the (possibly inconsistent) nuisance estimators. In the special case where $\psi_{0}$ equals $\mu\mapsto E_{0}\{m(Z,\mu)\}$ for a known map $m:\mathcal{Z}\times\mathcal{H}\rightarrow\mathbb{R}$, the influence function component $B_{0}$ simplifies to $m(\cdot,r_{0})-r_{0}\alpha_{0}$.
When both nuisance estimators are consistent, Theorem 3 indicates that the isocalibrated DML estimator is a regular, asymptotically linear, and efficient estimator of $\tau_{0}$ under weaker conditions than standard doubly robust estimators, allowing one nuisance function to be estimated arbitrarily slowly. This relaxation can also be shown to hold for the DRAL estimators proposed in van der Laan (2014), Benkeser et al. (2017), Dukes et al. (2021), and Bonvini et al. (2024). Even in the special case of the ATE parameter, Theorem 3 assumes substantially weaker conditions than prior works on doubly robust inference (van der Laan, 2014; Benkeser et al., 2017; Dukes et al., 2021). Specifically, these previous works imposed that the nuisance estimators converge to their (possibly incorrect) limits quickly, requiring that $\|\mu_{n,j}^{*}-\overline{\mu}_{0}\|\|\alpha_{n,j}^{*}-\overline{\alpha}_{0}\|=o_{P}(n^{-\frac{1}{2}})$, which offers no improvement over rate double robustness when both nuisance estimators are consistent. Higher-order influence function-based estimators, such as those proposed in Robins et al. (2008), have also been shown to achieve asymptotic efficiency under weaker conditions than standard doubly robust estimators. However, these estimators are usually not model-agnostic, relying, for example, on Hölder smoothness assumptions on the nuisance functions, and require the selection of tuning parameters for which there is limited practical guidance.
One potential concern with DRAL estimators is that they may fail to be regular under certain conditions (Dukes et al., 2021). Irregularity is undesirable because it can lead to poor finite-sample performance, with asymptotic approximations requiring arbitrarily large samples to be accurate (Leeb and Pötscher, 2005). However, Theorem 3 shows that irregularity arises only when one nuisance is inconsistently estimated. Otherwise, $n^{\frac{1}{4}}$–rate consistency of one nuisance estimator suffices for regularity, even if the other is consistent at an arbitrarily slow rate. This rate-free consistency condition is mild. Indeed, many machine learning methods—including neural networks, random forests, and gradient boosting—are universal approximators that can consistently estimate any smooth function (Hornik et al., 1989; Anthony and Bartlett, 2009; Zhang and Yu, 2005; Biau et al., 2008). Nevertheless, when one of the nuisance functions is inconsistent, we show in Appendix C that the calibrated DML estimator remains regular and efficient for an oracle projection parameter—defined with respect to the true distribution $P_{0}$—which formally quantifies the degree of irregularity (van der Laan et al., 2023a).
It can be shown that Theorem 3 remains true even when only the consistent nuisance estimator is calibrated. In particular, if $\overline{\mu}_{0}=\mu_{0}$ and we take $\alpha_{n,j}^{*}=\overline{\alpha}_{0}:=0$ for each $j=1,\ldots,J$, the isotonic calibrated plug-in estimator $\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(\mu_{n,j}^{*})$ is asymptotically linear under similar conditions (van der Laan et al., 2025). In the case of the ATE, the proposed calibrated DML estimator is asymptotically equivalent to the superefficient TML estimator that leverages the outcome regression for confounding bias adjustment (Benkeser et al., 2020). Similarly, if $\overline{\alpha}_{0}=\alpha_{0}$ and we take $\mu_{n,j}^{*}=\overline{\mu}_{0}:=0$, the balancing weighted outcome estimator $\frac{1}{n}\sum_{i=1}^{n}\alpha_{n,j(i)}^{*}(A_{i},W_{i})Y_{i}$ is asymptotically linear, leveraging the Riesz representer for confounding bias adjustment. As a result, isotonic calibration offers a simple and automated debiasing approach for constructing asymptotically linear (although irregular) inverse-probability-of-treatment weighting and G-computation plug-in estimators, eliminating the need to derive an efficient influence function.
The convergence rate conditions in Theorem 3 are stated in terms of the calibrated nuisance estimators, raising the question of whether calibration might degrade performance. Appendix B.2 addresses this concern, showing that isotonic calibration preserves $L^{2}$ convergence rates up to a typically negligible $O_{p}(n^{-1/3})$ term. Moreover, the calibrated nuisances converge to the best monotone transformation of the original nuisances, thereby relaxing consistency requirements: it suffices that the original estimators be consistent up to a monotone transformation.
We now discuss the key conditions of the theorem above; the remaining conditions are deferred to Appendix A.2. Condition C3 allows us to apply the expansion in Theorem 2 using our calibrated nuisance estimators. In Lemma 2 of Appendix A.1, we show a sufficient condition for this is that the random functions $(\widehat{m},m)\mapsto E_{0}\{\Delta_{\alpha,n,j}\mid\mu_{n,j}^{*}(X)=\widehat{m},\,\mu_{0}(X)=m,\,\mathcal{D}_{n}\}$ and $(\widehat{a},a)\mapsto E_{0}\{\Delta_{\mu,n,j}\mid\alpha_{n,j}^{*}(X)=\widehat{a},\,\alpha_{0}(X)=a,\,\mathcal{D}_{n}\}$ are $P_{0}$–almost surely Lipschitz continuous, with a uniformly bounded Lipschitz constant, for each $j$ and all sufficiently large $n$. This condition is not required when the standard cross-product rate condition $\|\mu_{n,j}^{*}-\mu_{0}\|\,\|\alpha_{n,j}^{*}-\alpha_{0}\|=o_{p}(n^{-1/2})$ holds, in which case asymptotic linearity and efficiency of isocalibrated DML estimators follow under conventional assumptions. Condition C2 is the key property defining DRAL of the calibrated DML estimator. For the partial covariance parameter, Bonvini et al. (2024) recently showed that the estimation rate achieved by isocalibrated DML estimator is minimax optimal under a hybrid model combining smoothness and structure-agnostic assumptions on the nuisance functions—similar to the conditions of Lemma 2. As a consequence, the rate conditions for asymptotic normality imposed by C2 cannot be meaningfully improved without making additional assumptions.
5.2 Doubly robust validity of bootstrap-assisted confidence intervals
We now show the bootstrap-assisted confidence intervals from Algorithm 2 provide doubly robust asymptotically valid inference for $\tau_{0}$. Importantly, that algorithm neither requires having prior knowledge of which nuisance estimators are consistent nor estimating the unknown quantities $r_{0}$ and $s_{0}$ appearing in the influence function of $\tau_{n}^{*}$ (Theorem 3). Hence, this bootstrap procedure is fully automated, only requiring computing the isocalibrated DML estimator within each bootstrap sample.
We recall that Algorithm 2 assumes $\psi_{0}$ has the form $\mu\mapsto E_{0}\{m(Z,\mu)\}$ and $\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}$ is taken to be the empirical plug-in estimator $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(Z_{i},\mu)$. Below, we define $P_{n,j}^{\#}$ as the empirical distribution of an arbitrary bootstrap sample $\mathcal{C}^{\#(j)}$ drawn from the empirical distribution $P_{n,j}$ of $\mathcal{C}^{(j)}$. Moreover, we denote the bootstrapped calibrated nuisance estimators corresponding to $P_{n,j}^{\#}$ as $\mu_{n,j}^{\#}$ and $\alpha_{n,j}^{\#}$, which are obtained as in Algorithm 2. Finally, we define $\tau_{n}^{*\#}$ as the corresponding bootstrapped isocalibrated DML estimator obtained from $\mu_{n,1}^{\#},\ldots,\mu_{n,J}^{\#}$ and $\alpha_{n,1}^{\#},\ldots,\alpha_{n,J}^{\#}$. Denoting by $\Delta_{\alpha,n,j}^{\#}:=\alpha_{n,j}^{\#}-\alpha_{0}$, $\Delta_{\mu,n,j}^{\#}:=\mu_{n,j}^{\#}-\mu_{0}$, and $r_{n,j}^{\#}:=\Pi_{\alpha_{n,j}^{\#}}(\mu_{0}-\mu_{n,j}^{\#})$ for $j=1,\ldots,J$ the nuisance estimation errors, the following conditions are used in the upcoming theorem:
  1. D1)
    Boundedness of bootstrap nuisance estimators: for some $M\in(0,\infty)$, the functions $\mu_{n,j}^{*\#}$ and $\alpha_{n,j}^{*\#}$ are uniformly bounded by $M$ with $P_{0}$–probability tending to one for each $j=1,\ldots,J$;
  2. D2)
    Complexity of bootstrap calibrators: with $P_{0}$–probability tending to one, the bootstrap isotonic calibrators $f_{n}^{\#}$ and $g_{n}^{\#}$ take at most $k(n)>0$ distinct values for some deterministic sequence $k(n)=o(n^{\frac{1}{2}}/\log n)$;
  3. D3)
    Pointwise Lipschitz continuity of the functional: $\mathcal{A}$ is finite, and there exists a constant $L\in(0,\infty)$ such that, for all $\mu\in\mathcal{H}$, $|m(Z,\mu)|\leq L\max_{a\in\mathcal{A}}|\mu(a,W)|$ $P_{0}$-almost surely.
  4. D4)
    Projection bootstrap error bounds: For each $j\in[J]$, both of the following hold:
    1. i)
      $\|(\Pi_{\mu_{n,j}^{\#}}-\Pi_{\mu_{0}})\Delta_{\alpha,n,j}^{\#}\|+\|(\Pi_{\mu_{n,j}^{\#},\mu_{0}}-\Pi_{\mu_{n,j}^{\#}})\Delta_{\alpha,n,j}^{\#}\|=O_{p}(\|\mu_{n,j}^{\#}-\mu_{0}\|)$ ;
    2. ii)
      $\|(\Pi_{\alpha_{n,j}^{\#}}-\Pi_{\alpha_{0}})\Delta_{\mu,n,j}^{\#}\|+\|(\Pi_{\alpha_{n,j}^{\#},\alpha_{0}}-\Pi_{\alpha_{n,j}^{\#}})\Delta_{\mu,n,j}^{\#}\|=O_{p}(\|\alpha_{n,j}^{\#}-\alpha_{0}\|)$
Theorem 4 (Double robustness of confidence intervals).
Under the conditions of Theorem 3 and D1D4, the conditional law of $n^{\frac{1}{2}}(\tau_{n}^{*\#}-\tau_{n}^{*})$ given $\{P_{n,j}:j=1,\ldots,J\}$ converges marginally in $P_{0}$–probability to the law of a mean-zero normal random variable with variance $\sigma_{0}^{2}:=var_{0}\{\chi_{0}(Z)\}$.
Under regularity conditions, as the number of bootstrap samples $K\rightarrow\infty$, the scaled version $n^{\frac{1}{2}}\sigma_{n}^{\#}$ of the bootstrap standard error of Algorithm 2 is a consistent estimator of the limiting standard deviation $\sigma_{0}$ of $n^{\frac{1}{2}}(\tau_{n}^{*}-\tau_{0})$. Hence, by Theorem 4, the bootstrap-assisted confidence interval $\text{CI}_{n}(\rho)$ from Algorithm 2 has doubly robust asymptotically exact coverage, in the sense that $P_{0}\{\tau_{0}\in\text{CI}_{n}(\rho)\}\rightarrow 1-\rho$ as $n\rightarrow\infty$, provided $K\rightarrow\infty$ sufficiently quickly.
As an alternative to the bootstrap, Wald-type confidence intervals can also be constructed. In this case, the unknown quantities $r_{0}$ and $s_{0}$ can be estimated using empirical risk minimization, as discussed in Appendix B.3. If one is willing to assume that both nuisance estimators are consistent, then $r_{0}$ and $s_{0}$ are identically zero, so standard confidence intervals based on the calibrated nuisances are valid.
6 Numerical experiments
6.1 Experiment 1: performance under inconsistent nuisance estimation
We evaluate the numerical performance of the calibrated DML estimator through simulation, focusing on inference for the ATE $\tau_{0}=E_{0}\{\mu_{0}(1,W)-\mu_{0}(0,W)\}$. The estimator is defined as $\tau_{n}^{*}:=\tau_{n}^{*}(1)-\tau_{n}^{*}(0)$, where $\tau_{n}^{*}(a)$ is the calibrated DML estimator of the counterfactual mean $E_{0}\{\mu_{0}(a,W)\}$, constructed as in Example 1 using Algorithm 1. We compare its performance to the standard cross-fitted AIPW estimator from Section 2.1 and the DR-TMLE estimator from Benkeser et al. (2017), described in Appendix B.4 and implemented in the drtmle R package (Benkeser and Hejazi, 2023). To quantify uncertainty, we use bootstrap confidence intervals (Algorithm 2) for calibrated DML and Wald-type intervals based on influence function variance estimates for AIPW and DR-TMLE.
We consider the same data-generating process used in the experiments of Benkeser et al. (2017). Covariates $W_{1}\sim\text{Uniform}(-2,2)$ and $W_{2}\sim\text{Bernoulli}(0.5)$ are drawn independently. The treatment and outcome are binary, with propensity $\pi_{0}(1\mid(w_{1},w_{2})):=\text{expit}(-w_{1}+2w_{1}w_{2})$ and outcome regression $\mu_{0}(a,w_{1},w_{2}):=\text{expit}(0.2a-w_{1}+2w_{1}w_{2})$. To evaluate the doubly robust inference property, we examine three scenarios of inconsistent nuisance estimation for each estimator:
  • (a)
    when both the outcome regression and treatment mechanism are estimated consistently;
  • (b)
    when only the treatment mechanism is estimated consistently;
  • (c)
    when only the outcome regression is estimated consistently.
In settings (b) and (c), to induce inconsistency in the estimation of the outcome regression and treatment mechanisms, we estimate the propensity score $\pi_{0}$ and outcome regressions $\mu_{0}$ using misspecified main-term logistic regression models. In all settings, any consistently estimated nuisance function is obtained using a 5-fold cross-fitted Nadaraya–Watson kernel smoother with bandwidth selected via cross-validation and stratified by both $A$ and $W_{2}$ (Hofmeyr, 2020). For DR-TMLE, we apply the same kernel method to estimate the nuisance functions $r_{0}$ and $s_{0}$. To avoid positivity violations, the cross-fitted propensity score estimates used in the AIPW estimator are truncated to lie within $[c_{n},1-c_{n}]$, where $c_{n}:=25/(n^{1/2}\log n)$, as recommended by Gruber et al. (2022). Further details and code for reproducing these analyses are available in the GitHub repository linked in Appendix F.
For each scenario and sample sizes ranging from 250 to 10,000, we generated and analyzed 5,000 datasets. Empirical estimates of bias, standard error, and 95% confidence interval coverage are presented in Figure 2. Results for the setting in which both nuisance functions are consistently estimated appear in Appendix I.1. In scenarios with inconsistent nuisance estimation, both calibrated DML and DR-TMLE exhibited DRAL behavior, showing lower bias and better coverage than AIPW. Relative to DR-TMLE, calibrated DML exhibited less bias, achieved better coverage, and had smaller or comparable variance. In settings where both nuisance functions were consistently estimated, all estimators performed similarly, as expected.
(a) Only outcome regression is estimated consistently.
(b) Only propensity score is estimated consistently.
Figure 2: Empirical bias, standard error, and 95% confidence interval coverage for calibrated DML (IC-DML), DR-TMLE and AIPW estimators under singly consistent estimation of the outcome regression and treatment mechanism.
6.2 Experiment 2: benchmarking on semi-synthetic datasets
In this experiment, we evaluate the performance of our method on a collection of semi-synthetic datasets commonly used to benchmark causal inference methods. These datasets feature moderately high-dimensional covariates, limited overlap, confounding, and treatment imbalance (Table 2). We include the Twins, Lalonde PSID, and Lalonde CPS datasets from the realCause generative modeling framework (Neal et al., 2020), which are derived from original data sources (Louizos et al., 2017; LaLonde, 1986). We also use several datasets from the 2017 Atlantic Causal Inference Challenge (ACIC-2017, settings 17–24) (Hahn et al., 2019), which are based on covariates from the Infant Health and Development Program (IHDP) (Brooks-Gunn et al., 1992), and 680 datasets from the scaling component of ACIC-2018 (Shimoni et al., 2018), with sample sizes of 1,000, 2,500, 5,000, and 10,000. The final dataset (IHDP) is a semi-synthetic version of the IHDP data introduced in Hill (2011) (Setting B) and further analyzed in Shalit et al. (2017). In these benchmarks, the covariates are fixed and derived from real data, and the inferential target is the sample average treatment effect, defined as $\tau_{n,0}:=\frac{1}{n}\sum_{i=1}^{n}(Y_{1,i}-Y_{0,i}),$ the empirical mean of the individual treatment effects in the observed sample.
Table 2: Sample size of total study ($n_{total}$), expected size of treatment arm $(n_{treated})$, and covariate dimension $(d)$ for the semi-synthetic datasets of second experiment. For the ACIC-2017 synthetic dataset, we employed simulation setting 24, characterized by additive errors, a large signal, large noise, and strong confounding.
Dataset $n_{total}$ $n_{treated}$ $d$ source
ACIC-2017 4302 $2233$ 58 Hahn et al. (2019)
ACIC-2018 varied varied varied Shimoni et al. (2018)
IHDP 672 123 25 Shalit et al. (2017)
Lalonde CPS 16177 $167$ 9 Neal et al. (2020)
Lalonde PSID 2675 $175$ 9 Neal et al. (2020)
Twins 11984 8935 76 Neal et al. (2020)
We compare the calibrated DML estimator to the AIPW estimator, with both defined as in the previous section. DR-TMLE was excluded from this experiment due to the computational expense of the iterative debiasing step, as discussed in the Introduction. Cross-fitted nuisance functions (propensity score and outcome regression) are estimated using gradient-boosted regression trees via xgboost (Chen et al., 2015), with the maximum tree depth selected by 10-fold cross-validation. Under smoothness conditions, these nuisance estimators will be consistent, though the rate of convergence may be slow in moderate dimensions (Zhang and Yu, 2005).
For each collection of $k$ dataset realizations, Table 4 reports empirical absolute bias, root mean square error (RMSE), and nominal 95% confidence interval coverage. The calibrated DML estimator consistently achieved superior coverage performance. For instance, in the ACIC-2017 settings, it improved AIPW’s coverage rates from 27% to 64%, 32% to 90%, 56% to 81%, and 32% to 90%. In terms of point estimation, calibrated DML also generally exhibited lower or comparable absolute bias across datasets. Notably, in the ACIC-2017 settings 20 and 24, the calibrated DML estimator outperformed the AIPW estimator, achieving bias reductions by factors of approximately 8. In ACIC-2017 settings 24 and 20, the calibrated DML estimator achieved lower RMSEs than AIPW (0.25 vs. 0.35 and 1.4 vs. 2.0, respectively). While AIPW performs better in terms of RMSE in the Lalonde CPS and Lalonde PSID datasets, this came at the cost of substantially poorer coverage.
Overall, the calibrated DML estimator outperformed the AIPW estimator in terms of confidence interval coverage, RMSE, and absolute bias across a range of semi-synthetic datasets.
Table 3: Evaluation of Absolute Bias, Root Mean Square Error (RMSE), and 95% Confidence Interval coverage on semi-synthetic benchmark datasets. Both outcome regression and propensity scores are estimated using an ensemble of gradient-boosted trees. For comparability across rows, the absolute bias and RMSE is scaled by a constant equal to the average effect size across dataset realizations for that benchmark. For clarity, we report only sample size 10000 results for ACIC-2018 and the settings with strong confounding for ACIC-2017. The full results can be found in Appendix I.2.
Dataset Coverage RMSE Absolute Bias
AIPW C-DML AIPW C-DML AIPW C-DML
ACIC-2017 (18) 0.27 0.64 0.70 0.58 0.21 0.28
ACIC-2017 (20) 0.32 0.90 2.00 1.40 1.60 0.20
ACIC-2017 (22) 0.56 0.81 0.11 0.10 0.004 0.035
ACIC-2017 (24) 0.32 0.90 0.34 0.25 0.30 0.04
ACIC-2018 (10000) 0.55 0.68 690 680 17 13
IHDP 0.57 0.57 0.46 0.46 0.13 0.13
Lalonde CPS 0.16 0.75 0.22 0.34 0.14 0.084
Lalonde PSID 0.46 0.84 0.19 0.44 0.04 0.04
Twins 0.51 0.54 0.24 0.23 0.22 0.21
7 Conclusion
In this work, we introduced a unified machine learning framework for doubly robust inference on linear functionals of the outcome regression. These functionals belong to a subset of parameters characterized by the mixed bias property (Rotnitzky et al., 2021). For the case of linear functionals, our framework provides an affirmative answer to the open question raised by Smucler et al. (2019), regarding the feasibility of achieving doubly robust inference for mixed-bias parameters using black-box nuisance function estimators. Another significant contribution of this work lies in establishing a close relationship between calibration and the doubly robust inference properties of debiased machine learning estimators. Importantly, our proof techniques rely solely on the nuisance estimators satisfying the calibration score equations stated in Section 5.1. While we applied our calibrated DML framework with isotonic calibration, these score equations can also be satisfied when nuisance functions are calibrated using other histogram-type estimators, such as honest regression trees (Wager and Athey, 2018), the highly adaptive lasso (van der Laan, 2017; Benkeser and van der Laan, 2016; Fang et al., 2021), and Venn-Abers calibration (Vovk and Petej, 2012; van der Laan and Alaa, 2024, 2025). In general, both our theoretical and empirical results support the use of calibrated nuisance estimators in causal inference to mitigate bias arising from inconsistent nuisance estimation and to ensure calibrated inferences.
There remain several open questions along this line of research. First, while we outline an extension of our framework to general parameters characterized by the mixed bias property (Rotnitzky et al., 2021) in Appendix E, we leave the full development of this approach to future work. Using this extension, we propose calibrated DML estimators for the partial covariance parameter, which arises in partially linear regression models, analogous to those proposed by Dukes et al. (2021) and Bonvini et al. (2024). Second, it may be of interest to develop an isotonic calibrated targeted minimum loss estimator that inherits both the plug-in property and doubly robust asymptotic linearity, as considered in van der Laan (2014) and Benkeser et al. (2017). Such an estimator could outperform non-plug-in alternatives such as one-step debiased estimators (Porter et al., 2011). A potential approach for this could be to iteratively apply the isotonic calibration and plug-in debiasing steps until the relevant score equations for both DRAL and the plug-in property are solved. Alternatively, the score-preserving TMLE of Pimentel et al. (2025) could be applied to the calibrated nuisance estimates to ensure that the targeting step preserves the empirical calibration scores. Lastly, extending our framework to accommodate multiple and sequentially robust estimators in longitudinal data structures presents further exciting research opportunities (Rotnitzky et al., 2017; Luedtke et al., 2017).
Acknowledgements. Research reported in this publication was supported by NIH grants DP2-LM013340 and R01-HL137808, and NSF grant DMS-2210216. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding agencies.
Appendix A Discussion of conditions
A.1 Sufficient conditions for B1
High-level conditions similar to B1 have appeared in prior work on doubly robust inference, including in van der Laan (2014), Benkeser et al. (2017), and Dukes et al. (2021). Heuristically, we may expect the projection $\Pi_{\mu_{n},\mu_{0}}\Delta_{\alpha,n}$ to be approximated well by $\Pi_{\mu_{n}}\Delta_{\alpha,n}$ as long as $\mu_{n}$ converges to $\mu_{0}$ in an appropriate sense, as the ‘information’ contained in $\{\mu_{n}(A,W),\mu_{0}(A,W),Z_{1},\ldots,Z_{n}\}$ should then be similar to that in $\{\mu_{n}(A,W),Z_{1},\ldots,Z_{n}\}$. Below, we establish sufficient conditions for B1 that formalize this intuition.
The following lemma establishes that Lipschitz continuity of two bivariate functions is sufficient for B1 to hold, a novel result that is relevant also to prior works requiring a similar condition. We make use of the following condition:
  1. B1enumi
    Lipschitz continuity: Denoting $\mathcal{D}_{n}:=\{Z_{1},\ldots,Z_{n}\}$, it holds that, for all $n$ large enough:
    1. i)
      $(\widehat{m},m)\mapsto E_{0}\{\Delta_{\alpha,n}(A,W)\,|\,\mu_{n}(A,W)=\widehat{m},\mu_{0}(A,W)=m,\mathcal{D}_{n}\}$ is $P_{0}$–almost surely Lipschitz continuous with uniformly bounded constant;
    2. ii)
      $(\widehat{a},a)\mapsto E_{0}\{\Delta_{\mu,n}(A,W)\,|\,\alpha_{n}(A,W)=\widehat{a},\alpha_{0}(A,W)=a,\mathcal{D}_{n}\}$ is $P_{0}$–almost surely Lipschitz continuous with uniformly bounded constant.
Lemma 2 (Bivariate Lipschitz continuity).
Conditions B1B1i and B1B1ii are implied, respectively, by conditions B1enumiB1enumii and B1enumiB1enumiii.
Conditions analogous to B1enumiB1enumii and B1enumiB1enumiii were also imposed in Bonvini et al. (2024) to study their proposed DRAL estimator of the ATE, and in Wang et al. (2023) to analyze regressions involving estimated features. Conditions B1enumiB1enumii and B1enumiB1enumiii can be relaxed to requiring Hölder continuity, in which case a weaker form of the bias expansion in Theorem 2, based on the Hölder exponent, can be established (Bonvini et al., 2024). Condition B1enumi implies that a small change in the value of one nuisance function at a point should, on average, lead to a small change in the estimation error for the other nuisance function at that point.
Lipschitz (and more generally, Hölder smoothness) assumptions on the nuisance functions and estimators are commonly made in the literature for higher-order influence function-based (HOIF) estimators, such as those proposed in Robins et al. (2008)—see Bonvini et al. (2024) for a discussion of the relationship between DRAL and HOIF estimators. These estimators have been shown to achieve higher-order debiasing, similarly as established in Theorem 2, and attain efficiency under weaker conditions compared to standard doubly robust estimators. However, these estimators are not model-agnostic, as they rely on high-dimensional kernel smoothing and require the selection of tuning parameters, for which there is limited practical guidance. In contrast to HOIF estimators, which typically impose Hölder conditions on the potentially high-dimensional nuisance functions, Condition B1enumi requires Lipschitz continuity of a bivariate function, which is a relatively mild smoothness condition. Therefore, we hypothesize that this condition, or the higher-level conditions in B1, are plausibly satisfied by complex estimators and real-world data-generating processes without necessitating higher-dimensional Lipschitz continuity. Our experimental results provide empirical support for this hypothesis.
A.2 Discussion of conditions of Theorem 3
Condition C2 ensures that the (possibly inconsistent) nuisance estimators have in-probability limits in $\mathcal{H}$ and is commonly required for doubly robust inference (Benkeser et al., 2017; Smucler et al., 2019; Dukes et al., 2021). Condition C5 requires bounded nuisance functions and estimators. The boundedness of the Riesz representer $\alpha_{0}$ in C5 is a general form of the strong positivity assumption (Chernozhukov et al., 2022). If the $P_{0}$–support of the outcome $Y$ is bounded, then it is typically not difficult to ensure boundedness of $\mu_{0}$ and its estimator. Condition C6 imposes a weak upper bound on the number of plateaus in the (piecewise constant) isotonic regression solution; this is expected to hold in any setting with noisy outcomes where isotonic regression cannot perfectly fit the data. In view of Deng et al. (2021), continuous differentiability of $t\mapsto E_{0}\{Y\,|\,\mu_{n,j}(A,W)=t,\mathcal{D}_{n}\}$ implies that the number of observations in a given plateau of an isotonic regression solution grows at a rate $n^{\frac{2}{3}}$ with probability tending to one; heuristically, this suggests that the number of plateaus grows at rate $n^{\frac{1}{3}}=o(n^{\frac{1}{2}})$. Regardless, this condition can be ensured by constraining the maximum number of plateaus of the isotonic regression solution via the maximum tree depth argument in xgboost.
Condition C3 allows us to apply the expansion in Theorem 2 using our calibrated nuisance estimators. In Lemma 2 of Appendix A.1, we show that a sufficient condition for this is that the random functions $(\widehat{m},m)\mapsto E_{0}\{\Delta_{\alpha,n,j}\mid\mu_{n,j}^{*}(X)=\widehat{m},\,\mu_{0}(X)=m,\,\mathcal{D}_{n}\}$ and $(\widehat{a},a)\mapsto E_{0}\{\Delta_{\mu,n,j}\mid\alpha_{n,j}^{*}(X)=\widehat{a},\,\alpha_{0}(X)=a,\,\mathcal{D}_{n}\}$ are $P_{0}$–almost surely Lipschitz continuous, with a uniformly bounded Lipschitz constant, for each $j=1,\ldots,J$ and all sufficiently large $n$. This condition is not required when the standard cross-product rate condition $\|\mu_{n,j}^{*}-\mu_{0}\|\,\|\alpha_{n,j}^{*}-\alpha_{0}\|=o_{p}(n^{-1/2})$ holds, in which case asymptotic linearity and efficiency of isocalibrated DML estimators follow under conventional assumptions. Condition C2 ensures that the isocalibrated DML estimator is DRAL, allowing for valid inference even with inconsistent or insufficiently fast estimation of one of the nuisance functions, as long as the other nuisance function is consistently estimated at a sufficiently fast rate. Notably, for the partial covariance parameter, Bonvini et al. (2024) recently showed that the estimation rate achieved by isocalibrated DML estimator is minimax optimal under a hybrid model combining smoothness and structure-agnostic assumptions on the nuisance functions—similar to those assumed in Lemma 2. As a consequence, the rate conditions for asymptotic normality imposed by C2 cannot be meaningfully improved without making additional structural assumptions. Condition C7 arises from the need to estimate the linear functional. Condition C7a requires that the linearization remainder term of $\psi_{n,j}(r_{n,j}^{*})$ as an estimator of $\psi_{0}(r_{n,j}^{*})$ be asymptotically negligible. This condition is trivially satisfied for functionals of the form given in (1), that is, with $\psi_{0}:=\mu\mapsto E_{0}\{m(Z,\mu)\}$ for some map $(z\mu)\mapsto m(z,\mu)$, when $\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}$ is taken to be the empirical plug-in estimator $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(Z_{i},\mu)$. Condition C7b is a weak empirical process condition that can be verified using our proof techniques. For example, when $\mathcal{A}$ is a finite set, the first expression holds under the Lipschitz continuity condition $|\phi_{0,\mu_{n,j}^{*}}(z)-\phi_{0,\overline{\mu}_{0}}(z)|\leq L\sup_{a_{0}\in\mathcal{A}}|\mu_{n,j}^{*}(a_{0},w)-\overline{\mu}_{0}(a_{0},w)|$ $P_{0}$-almost everywhere for some $L<\infty$, along with the convergence condition $\sup_{a_{0}\in\mathcal{A}}|\mu_{n,j}^{*}(a_{0},w)-\overline{\mu}_{0}(a_{0},w)|=o_{p}(1)$ (see Lemmas 6 and 12 in the Appendix).
Appendix B Additional details on doubly robust inference
B.1 Benefits and trade-offs in empirical calibration
In the context of debiased machine learning, empirical calibration is valuable not only for enabling doubly robust inference but also for improving estimator stability and predictive accuracy. In our framework, empirical calibration is defined relative to a loss function $\ell(z,\nu)$ for a nuisance function $\nu\in\mathcal{H}$, and requires that $\nu_{n}$ satisfies the self-consistency condition
$\frac{1}{n}\sum_{i=1}^{n}\ell(Z_{i},\nu_{n})=\min_{\theta}\frac{1}{n}\sum_{i=1}^{n}\ell(Z_{i},\theta\circ\nu_{n}),$
where the minimum is taken over all univariate transformations $\theta:\mathbb{R}\to\mathbb{R}$. This condition ensures that the predictions $\nu_{n}$ are optimal with respect to empirical risk and cannot be improved by recalibration.
In the case of squared error loss $\ell(z,\mu)=\{y-\mu(a,w)\}^{2}$, calibration requires that $\mu_{n}(a,w)$ equals the average outcome among individuals with the same predicted value. That is,
$\mu_{n}(a,w)=\frac{\sum_{i=1}^{n}1\{\mu_{n}(A_{i},W_{i})=\mu_{n}(a,w)\}Y_{i}}{\sum_{i=1}^{n}1\{\mu_{n}(A_{i},W_{i})=\mu_{n}(a,w)\}},$
ensuring that $\mu_{n}$ does not systematically over- or under-predict outcomes. For the Riesz loss $\ell(z,\alpha)=\alpha^{2}(a,w)-2\psi_{n}(\alpha)$, associated with estimating a counterfactual mean $\psi_{0}=E_{0}\{\mu_{0}(1,W)\}$, calibration of the Riesz representer $\alpha_{n}$ implies approximate covariate balance across estimated inverse propensity strata. Specifically, the calibration condition
$\frac{1}{n}\sum_{i=1}^{n}A_{i}\alpha_{n}(1,W_{i})f(\alpha_{n}(1,W_{i}))=\frac{1}{n}\sum_{i=1}^{n}f(\alpha_{n}(1,W_{i}))$
must hold for all functions $f$, ensuring that higher weights contribute meaningfully to balancing, rather than inflating variance (Deshpande and Kuleshov, 2023; van der Laan et al., 2024b).
To ensure good performance of the empirically calibrated estimator, it is important that calibration holds not only on the observed data but also at the population level. Notably, perfect empirical calibration can be achieved trivially by overfitting the labels—for example, by interpolating the outcomes—so it is crucial to use calibration procedures that generalize well. Formally, a function $\nu$ is said to be perfectly $\ell$-calibrated with respect to a loss function $\ell(z,\nu)$ if it minimizes the expected loss over all one-dimensional transformations of itself:
$E_{P}[\ell(Z,\nu)]=\inf_{\theta}E_{P}[\ell(Z,\theta\circ\nu)],$
where the infimum is taken over all functions $\theta:\mathbb{R}\to\mathbb{R}$. The notion of $\ell$-calibration generalizes common forms of calibration, such as regression and quantile calibration (Roth, 2022), which correspond to the least-squares and pinball (hinge) losses, respectively (Deng et al., 2023). Related ideas have also been explored in the context of predictive modeling (Whitehouse et al., 2024).
B.2 Convergence rates of isotonic calibrated nuisances
The convergence rate conditions in Theorem 3 are stated with respect to the calibrated nuisance estimators, rather than the original ones. This raises a natural concern: could calibration worsen the convergence rate? The following informal lemma alleviates this concern by showing that the isotonic calibrated estimators perform nearly as well—in mean squared error—as the best monotone non-decreasing transformation of the uncalibrated, cross-fitted nuisance estimators, up to a typically negligible error of order $O_{p}(n^{-2/3})$. Formal statements and regularity conditions are given in Section G.4.
Lemma 3.
Suppose that $\psi_{0}$ has the form (1) with mapping $m:\mathcal{A}\times\mathcal{W}\times\mathcal{H}\rightarrow\mathbb{R}$, and that $\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}$ is taken to be the plug-in estimator $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(Z_{i},\mu)$. Under regularity conditions, it holds that
$\sum_{j=1}^{J}\|\mu_{n,j}^{*}-\mu_{0}\|^{2}\,$ $=\,\min_{f\in\mathcal{F}_{\textnormal{iso}}}\sum_{j=1}^{J}\|f\circ\mu_{n,j}-\mu_{0}\|^{2}+O_{p}(n^{-\frac{2}{3}})\,;$
$\sum_{j=1}^{J}\|\alpha_{n,j}^{*}-\alpha_{0}\|^{2}\,$ $=\,\min_{g\in\mathcal{F}_{\textnormal{iso}}}\sum_{j=1}^{J}\|g\circ\alpha_{n,j}-\alpha_{0}\|^{2}+O_{p}(n^{-\frac{2}{3}})\,.$
Since the identity transformation is itself monotone non-decreasing, the above lemma implies that the isotonic calibrated nuisance estimators typically perform no worse than the uncalibrated ones, and may perform better when the uncalibrated estimators are poorly calibrated or only consistent up to a monotone transformation. Thus, beyond providing doubly robust asymptotic linearity, isotonic calibration relaxes the consistency requirements on the nuisance estimators: it suffices that the initial cross-fitted nuisance estimators be consistent up to a monotone transformation.
B.3 Estimation of nuisance functions in limiting distribution
Using calibrated DML with the bootstrap-assisted inference method of Algorithm 2, it is not necessary to obtain estimates of the univariate nuisance functions $f_{0,\overline{\mu}_{0}}$ and $g_{0,\overline{\alpha}_{0}}$, which appear in the limited distribution of Theorem 3. Nonetheless, if desired, we note that these nuisance functions can be estimated using univariate empirical risk minimization. For example, one may wish to implement the TMLE approach to doubly robust inference discussed in the previous section.
To estimate these nuisances, observe that $f_{0,\overline{\mu}_{0}}$ is the Riesz-representer of the linear functional $\alpha\mapsto\langle\alpha_{0}-\overline{\alpha}_{0},\alpha\rangle_{P_{0}}=\psi_{0}(\alpha)-\langle\overline{\alpha}_{0},\alpha\rangle_{P_{0}}$ with respect to the linear space $\left\{f\circ\overline{\mu}_{0}:f:\mathbb{R}\rightarrow\mathbb{R},\|f\circ\mu\|<\infty\right\}$. Hence, as a consequence of the Riesz-defined above Theorem 1, the function $f_{0,\overline{\mu}_{0}}$ can be obtained via univariate empirical risk minimization, after substituting $\overline{\mu}_{0}$ and $\overline{\alpha}_{0}$ with their respective estimators. Specifically, the function $f_{0,\overline{\mu}_{0}}$ can be directly estimated via offset univariate empirical risk minimization over $\{f\circ\mu_{n,\diamond}^{*}\in L^{2}(\overline{P}_{0,A,W})\}$ based on an empirical estimator of the Riesz-loss risk:
$f\mapsto\frac{1}{n}\sum_{i=1}^{n}\left[\left\{\alpha_{n,j(i)}^{*}-(f\circ\mu_{n,j(i)}^{*})\right\}^{2}(A_{i},W_{i})-2\left\{(f\circ\mu_{n,j(i)}^{*}\cdot\alpha_{n,j(i)}^{*})(A_{i},W_{i})-\psi_{n,j(i)}\left(f\circ\mu_{n,j(i)}^{*}\right)\right\}\right].$
Similarly, noting that $\overline{\mu}_{0,\alpha}(a,w)=E_{0}\left[Y-\overline{\mu}_{0}(A,W)|\alpha(A,W)=\alpha(a,w)\right]$, the function $g_{0,\overline{\alpha}_{0}}$ can be estimated by regressing the residual $\{Y_{i}-\mu_{n,j(i)}^{*}(A_{i},W_{i})\}_{i\in[n]}$ on $\{\alpha_{n,j(i)}^{*}(A_{i},W_{i})\}_{i\in[n]}$ using, for instance, locally adaptive regression splines (Mammen and Van De Geer, 1997) or random forests (Breiman, 2001).
B.4 Doubly robust inference via TMLE
For the special case of doubly robust inference on the average treatment effect, van der Laan (2014) and Benkeser et al. (2017) propose a debiasing procedure based on targeted minimum loss estimation (TMLE) to refine given nuisance estimators to asymptotically satisfy (4) and (5). The idea behind the TMLE-based approach is to iteratively debias the initial nuisance estimators $\mu_{n}$ and $\alpha_{n}$ using a reweighted generalized linear regression algorithm with offset to obtain targeted estimators $\mu_{n}^{*}$ and $\alpha_{n}^{*}$ that satisfy the empirical score equations
$\frac{1}{n}\sum_{i=1}^{n}s_{n}^{*}(A_{i},W_{i})\{Y_{i}-\mu_{n}^{*}(A_{i},W_{i})\}\,=\,o(n^{-\frac{1}{2}})\,;$ (11)
$\frac{1}{n}\sum_{i=1}^{n}r_{n}^{*}(A_{i},W_{i})\alpha_{n}^{*}(A_{i},W_{i})-\psi_{n}(r_{n}^{*})\,=\,o(n^{-\frac{1}{2}})\,.$ (12)
Here, $s_{n}^{*}$ and $r_{n}^{*}$ are estimators of the (data-dependent) nuisance functions $s_{n,0}^{*}:=\Pi_{\mu_{n}^{*}}(\alpha_{0}-\alpha_{n}^{*})$ and $r_{n,0}^{*}:=\Pi_{\alpha_{n}^{*}}(\mu_{0}-\mu_{n}^{*})$, which can be obtained, for example, using univariate empirical risk minimization with features $\mu_{n}^{*}(A_{1},W_{1}),\ldots,\mu_{n}^{*}(A_{n},W_{n})$ and $\alpha_{n}^{*}(A_{1},W_{1}),\ldots,\alpha_{n}^{*}(A_{n},W_{n})$, respectively—we provide details in Appendix B.3. Under certain regularity conditions, it can be shown that (11) and (12) imply that (4) and (5) are satisfied asymptotically, so that the one-step debiased estimator based on $\mu_{n}^{*}$ and $\alpha_{n}^{*}$ is indeed DRAL.
As discussed in the Introduction, the TMLE-based algorithm for doubly robust inference involves estimating two univariate nuisance functions at each iteration, with the final iteration involving estimation of nuisances $r_{n,0}^{*}$ and $s_{n,0}^{*}$; the process can be computationally quite prohibitive. In addition, the final estimation of $r_{n,0}^{*}$ and $s_{n,0}^{*}$ possibly induces additional (unnecessary) finite-sample biases on the one-step debiased estimator $\tau_{n}$ by only approximately solving the empirical moment equations (4) and (5). We also note that neither van der Laan (2014) nor Benkeser et al. (2017) formally demonstrate that their iterative refinement procedures maintain the convergence properties of the original nuisance estimators in pursuit of satisfying the empirical orthogonality equations. Such a formal analysis is important to ensure that, under consistent estimation of both nuisance functions, their refinement procedure does not harm the quality of estimators and inferences based on the one-step debiased estimator. In the following section, we propose a novel procedure for constructing one-step debiased estimators that is free from these issues and, unlike prior works, simultaneously tackles a whole class of target parameters.
Appendix C Estimator irregularity under inconsistent nuisance estimation
As emphasized earlier, Theorem 3 establishes that the isotonic calibrated one-step debiased estimator is DRAL, and that it is regular and efficient when both nuisance functions are consistently estimated, even if one is estimated at an insufficient rate. We recall that an asymptotically linear estimator $\widehat{\psi}_{n}$ of $\psi_{0}$ with influence function $f_{0}$ is said to be regular at $P_{0}$ for the functional $\Psi:\mathcal{M}\to\mathbb{R}$ if, for every one-dimensional regular parametric submodel $\{P_{0,t}:t\in(-\epsilon,\epsilon)\}\subset\mathcal{M}$ through $P_{0}$ at $t=0$, it holds that
$n^{\frac{1}{2}}\,\{\widehat{\psi}_{n}-\Psi(P_{0,t_{n}})\}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,P_{0}f_{0}^{2})$
under sampling from $P_{0,t_{n}}$ with $t_{n}:=n^{-\frac{1}{2}}$. Regularity of $\widehat{\psi}_{n}$ implies that inference for $\psi_{0}$ is uniformly valid under local perturbations of the data-generating distribution. Under inconsistent estimation of one of the nuisance functions, the resulting estimator is typically irregular under a nonparametric model, as it is asymptotically linear with an influence function differing from the nonparametric efficient influence function of the target parameter (van der Vaart, 2000; Dukes et al., 2021). Irregularity is typically viewed as undesirable due to its potential to cause poor performance in finite samples, since it is possible to find a distribution such that an arbitrarily large sample size may be required for the asymptotic limiting distribution to accurately approximate finite-sample behavior (Leeb and Pötscher, 2005). This criticism is arguably less relevant in our case since standard one-step debiased estimators are, under inconsistent estimation of one nuisance function, typically neither regular nor $n^{-\frac{1}{2}}$–consistent, let alone asymptotically linear (Dukes et al., 2021). We stress here that the set of laws under which the isocalibrated DML estimator is regular and asymptotically efficient is a strict superset of the corresponding set of laws for standard one-step debiased estimators, and furthermore, that the isocalibrated DML estimator is asymptotically linear and permits valid pointwise (as opposed to uniform) inference at $n^{\frac{1}{2}}$–rate at laws at which the standard estimators neither converge at $n^{\frac{1}{2}}$–rate nor readily allow even pointwise inference.
The following theorem characterizes the local asymptotic behavior of the isocalibrated DML estimator in the irregular case where one of the nuisance functions is inconsistently estimated. It establishes that the estimator is a regular and efficient estimator of a projection-based oracle parameter $\Psi_{0}$, which agrees with the original target parameter $P\mapsto\Psi(P):=\psi_{P}(\mu_{P})$ at $P_{0}$. As a result, the isocalibrated DML procedure offers regular and locally uniformly valid inference for the corresponding projection estimand, even when sampling from least favorable local perturbations of $P_{0}$.
To define this oracle parameter, we first introduce additional notation. We define the affine spaces $\overline{\Theta}_{0,P}:=\{\overline{\mu}_{0}+f\circ\alpha_{0}:f\}\cap L^{2}(P)$ and $\overline{\Theta}_{0,P}^{*}:=\{\overline{\alpha}_{0}+g\circ\mu_{0}:g\}\cap L^{2}(P)$, where the indexing functions $f$ and $g$ vary over all one-dimensional, real-valued transformations. We denote the orthogonal projection of $\mu_{P}$ onto $\overline{\Theta}_{0,P}$ by $\overline{\Pi}_{0,P}\mu_{P}:=\operatorname*{argmin}_{\mu\in\overline{\Theta}_{0,P}}\|\mu_{P}-\mu\|_{P}$, and the orthogonal projection of $\alpha_{P}$ onto $\overline{\Theta}_{0,P}^{*}$ by $\overline{\Pi}_{0,P}^{*}\alpha_{P}:=\operatorname*{argmin}_{\alpha\in\overline{\Theta}_{0,P}^{*}}\|\alpha_{P}-\alpha\|_{P}$. We may then define the projection-based oracle parameter $\Psi_{0}$, which varies depending on which nuisance function is estimated inconsistently, as
$P\mapsto\Psi_{0}(P)$ :=$:=\begin{cases}\overline{\psi}_{0,P}(\mu_{P})&\text{if }\overline{\mu}_{0}=\mu_{0}\\ \psi_{P}(\overline{\Pi}_{0,P}\mu_{P})&\text{if }\overline{\alpha}_{0}=\alpha_{0}\end{cases}$
with $\overline{\psi}_{0,P}:h\mapsto\langle\overline{\Pi}_{0,P}^{*}\alpha_{P},h\rangle_{P}$ representing an oracle approximation of the linear functional $\psi_{P}:h\mapsto\langle\alpha_{P},h\rangle_{P}$. The orthogonality conditions characterizing the projections $\overline{\Pi}_{0,P}^{*}\alpha_{0}$ and $\overline{\Pi}_{0,P}\mu_{0}$ imply that $\Psi_{0}(P_{0})=\Psi(P_{0})$, so that the oracle parameter agrees with the target parameter at $P_{0}$. In the case where $\mu_{0}$ is consistently estimated, $\Psi_{0}$ reduces to the parameter $P\mapsto\overline{\psi}_{0,P}(\mu_{P})$, which leverages the outcome regression $\mu_{0}(A,W)$ as a dimension reduction of $(A,W)$. Alternatively, in the case where $\alpha_{0}$ is consistently estimated, $\Psi_{0}$ reduces to the parameter $P\mapsto\psi_{P}(\overline{\Pi}_{0,P}\mu_{P})$, which leverages the Riesz representer $\alpha_{0}(A,W)$ as a dimension reduction of $(A,W)$.
Theorem 5 (Behavior under local perturbations).
If the conditions of Theorem 3 hold and that either $\overline{\mu}_{0}=\mu_{0}$ or $\overline{\alpha}_{0}=\alpha_{0}$, $\tau_{n}^{*}$ is a regular and asymptotically linear at $P_{0}$ for the oracle projection parameter $\Psi_{0}$ with influence function $\chi_{0}$. As such, under sampling from any local perturbation $P_{0,t_{n}}$ of $P_{0}$, $n^{\frac{1}{2}}\{\tau_{n}^{*}-\Psi_{0}(P_{0,t_{n}})\}$ tends to a mean-zero normal random variable with variance $\sigma_{0}^{2}$ as defined in Theorem 3.
Theorem 5 demonstrates that even under sampling from least favorable local perturbations, the isocalibrated DML estimator still provides regular and locally uniformly valid inference for the projection-based oracle parameter $\Psi_{0}$. Notably, the magnitude of the oracle bias $\Psi_{0}(P)-\Psi(P)$ for $\Psi$ is driven by the degree of nuisance estimator inconsistency via projection distances $\|\overline{\Pi}_{\mu_{0}}\mu_{P}-\mu_{P}\|_{P}$ and $\|\overline{\Pi}_{\alpha_{0}}\alpha_{P}-\alpha_{P}\|_{P}$. Consequently, under sampling from a local perturbation, when the nuisance functions are only mildly inconsistently estimated, the local asymptotic bias of the isocalibrated DML estimator for the target parameter $\Psi$ may still be relatively small. A similar result was derived in van der Laan et al. (2023a) for superefficient and irregular estimators obtained using data-driven model selection techniques. As a special case, Theorem 5 implies that $\tau_{n}^{*}$ is regular and asymptotically linear for the actual target parameter $\Psi$ over all local perturbations that do not perturb the consistently estimated nuisance function, which agrees with Theorem 2 of Dukes et al. (2021).
Appendix D Background on calibration
In this section, we provide background on calibration using binning and isotonic regression. For additional details, we refer to van der Laan and Alaa (2025) for a general treatment of histogram binning and isotonic calibration; see also Gupta et al. (2020); Gupta and Ramdas (2021) for histogram binning, van der Laan et al. (2023b) and van der Laan et al. (2024b) for applications of isotonic calibration in causal inference, and Whitehouse et al. (2024) for calibration with respect to general loss functions.
D.1 Empirical calibration via histogram binning
A simple, distribution-free method for calibrating nuisance estimators is histogram binning, such as uniform-mass (quantile) binning (Gupta et al., 2020; Gupta and Ramdas, 2021). Given an estimated function $\nu_{n}$, the calibration dataset $\mathcal{C}_{n}=\{Z_{i}\}_{i=1}^{n}$, and a loss function $\ell(z,\nu)$, the range of predictions $\nu_{n}(Z_{i})$ is partitioned into $K$ disjoint bins $\{B_{k}\}_{k=1}^{K}$ in an outcome-independent manner—for example, using quantiles of $\{\nu_{n}(Z_{i})\}_{i=1}^{n}$. The calibrated estimator $\nu_{n}^{*}:=\theta_{n}\circ\nu_{n}$ is then obtained by minimizing the empirical loss within each bin:
$\theta_{n}(t):=\operatorname*{argmin}_{c\in\mathbb{R}}\sum_{i=1}^{n}\mathbbm{1}\{\nu_{n}(Z_{i})\in B_{k(t)}\}\,\ell(Z_{i},c),$
where $k(t)$ denotes the index of the bin containing $t\in\nu_{n}(\mathcal{Z})$. For squared error loss, this reduces to computing the average of the observed outcomes within each bin. Since any transformation of a predictor preserves its piecewise constant structure, the empirical risk cannot be reduced by further transforming its predictions, and hence it is empirically $\ell$-calibrated.
While histogram binning guarantees empirical calibration, it does not ensure good out-of-sample performance. With a fine partition $(K=n)$, the estimator may interpolate the data and overfit, leading to poor generalization. Conversely, with a coarse partition $(K=1)$, the estimator is well-calibrated but lacks predictive power, collapsing to a constant predictor. This trade-off highlights the importance of choosing $K$ carefully: too few bins limit expressiveness, while too many increase variance and degrade calibration. Under appropriate choices of $K$, histogram binning yields asymptotically calibrated estimators, with the conditional $\ell^{2}$-calibration error satisfying $\operatorname{Cal}_{\ell^{2}}(\nu_{n}^{*})=O_{p}\left(\frac{K\log(n/K)}{n}\right)$ (Whitehouse et al., 2024; van der Laan and Alaa, 2025). Thus, tuning $K$, for example via cross-validation, is essential to balance calibration and accuracy. In the next section, we adopt a data-adaptive alternative—isotonic calibration—which flexibly enforces monotonicity while maintaining distribution-free guarantees (Zadrozny and Elkan, 2001; Niculescu-Mizil and Caruana, 2005b; van der Laan et al., 2023b).
D.2 Empirical calibration via isotonic regression
Isotonic calibration (Zadrozny and Elkan, 2002; Niculescu-Mizil and Caruana, 2005a) is a data-adaptive histogram binning method that learns the bins using isotonic regression, a nonparametric technique traditionally used for estimating monotone functions (Barlow and Brunk, 1972; Groeneboom and Lopuhaa, 1993). Specifically, the bins are determined by minimizing an empirical mean squared error criterion under the constraint that the calibrated predictor is a non-decreasing monotone transformation of the original predictor. Isotonic calibration is motivated by the heuristic that, for a good predictor $z\mapsto\nu_{n}(z)$, the optimal transformation of the predictor should be approximately monotone as a function of $\nu_{n}$. For example, for a perfect predictor, the optimal transformation should be the identity function. Even when this map is not approximately monotone, isotonic regression ensures that the quality of the initial predictor is not harmed, since the identity transformation—being monotone—can always be learned. Isotonic calibration is distribution-free—it does not rely on monotonicity assumptions—and, unlike traditional histogram binning, it is tuning-parameter-free and naturally preserves the mean squared error of the original predictor (as the identity transformation is monotone) (van der Laan et al., 2023b).
For clarity, we focus on the regression case where $\ell$ denotes the squared error loss. Formally, isotonic calibration takes a fitted nuisance estimator $\nu_{n}$ and a calibration dataset $\mathcal{C}_{n}$ to produce a calibrated model $\nu_{n}^{*}:=\theta_{n}\circ\nu_{n}$, where $\theta_{n}:\mathbb{R}\rightarrow\mathbb{R}$ is an isotonic step function obtained by solving the optimization problem:
$\theta_{n}\in\operatorname*{argmin}_{\theta\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\left\{Y_{i}-\theta(\nu_{n}(W_{i},A_{i}))\right\}^{2},$ (13)
where $\mathcal{F}_{\text{iso}}$ denotes the set of all univariate, piecewise constant functions that are monotonically nondecreasing. Following Groeneboom and Lopuhaa (1993), we consider the unique càdlàg piecewise constant solution to this problem, which has jumps only at observed values $\{\nu_{n}(W_{i},A_{i}):i\in[n]\}$. The first-order optimality conditions of this convex problem imply that the isotonic solution $\theta_{n}$ acts as a binning calibrator over a data-adaptive set of bins determined by the jump points of the step function. Thus, isotonic calibration yields perfect empirical calibration. Specifically, for any transformation $g:\mathbb{R}\rightarrow\mathbb{R}$, the perturbed step function $\varepsilon\mapsto\theta_{n}+\varepsilon(g\circ\theta_{n})$ remains isotonic for all sufficiently small $\varepsilon$ such that $|\varepsilon|\sup_{t\in\nu_{n}(\mathcal{Z})}|(g\circ\theta_{n})(t)|$ is less than the maximum jump size of $\theta_{n}$, given by $\sup_{t\in\nu_{n}(\mathcal{Z})}|\theta_{n}(t)-\theta_{n}(t-)|$. Since $\theta_{n}$ minimizes the empirical mean squared error over all isotonic functions, it follows that for each $g:\mathbb{R}\rightarrow\mathbb{R}$, the following condition holds:
$\frac{d}{d\varepsilon}\frac{1}{2}\sum_{i=1}^{n}\left\{Y_{i}-\theta_{n}(\nu_{n}(W_{i},A_{i}))-\varepsilon g(\theta_{n}(\nu_{n}(W_{i},A_{i})))\right\}^{2}\Big{|}_{\varepsilon=0}=\sum_{i=1}^{n}g(\nu_{n}^{*}(Z_{i}))\{Y_{i}-\nu_{n}^{*}(Z_{i})\}=0.$
These orthogonality conditions are equivalent to perfect empirical calibration. In particular, by taking $g(t)=1\{t=\theta_{n}(\nu_{n}(z))\}$, we conclude that the isotonic calibrated predictor $\nu_{n}^{*}$ is empirical calibrated.
Appendix E Examples and extensions of calibrated DML
E.1 Extensions
In this section, we provide a brief overview of how our doubly robust inference framework can be extended to a general class of parameters with the mixed bias property, as introduced in Rotnitzky et al. (2021). While our primary focus in this paper remains on the class of linear functionals discussed in Section 2.1, it is important to emphasize that our technique and results can be readily generalized to parameters exhibiting the mixed bias property.
We focus on debiasing mixed bias remainders of the following form using isotonic calibration:
$\int w(z)\{\alpha_{n}(w,a)-\alpha_{0}(w,a)\}\{\gamma_{n}(w,a)-\gamma_{0}(w,a)\}P_{0}(dz)$
where $z\mapsto w(z)$ is a known weighting function, and $\alpha_{0}:=\operatorname*{argmin}_{\alpha}E_{0}[w(Z)\{\alpha(A,W)\}^{2}-2m_{1}(A,W,\alpha)]$ and $\gamma_{0}:=\operatorname*{argmin}_{\gamma}E_{0}[w(Z)\{\gamma(A,W)\}^{2}-2m_{2}(A,W,\gamma)]$ are weighted Riesz representers for the linear functionals $\alpha\mapsto E_{0}[m_{1}(A,W,\alpha)]$ and $\gamma\mapsto E_{0}[m_{2}(A,W,\gamma)]$. Mixed bias parameters, as defined in Rotnitzky et al. (2021), have remainders of this form. More generally, parameters may involve second-order remainders consisting of multiple such mixed bias terms, each of which can be linearized separately using the results of this section.
Given estimators $\alpha_{n}$ and $\gamma_{n}$ of $\alpha_{0}$ and $\gamma_{0}$, we propose using the isocalibrated nuisance estimators $\alpha_{n}^{*}=g_{n,1}\circ\alpha_{n}$ and $\gamma_{n}^{*}=g_{n,2}\circ\gamma_{n}$, where
$g_{n,1}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}w(Z_{i})\big{\{}g(\alpha_{n}(A_{i},W_{i}))\big{\}}^{2}-2m_{1}(Z_{i},g\circ\alpha_{n}),$ (14)
$g_{n,2}$ $\in\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}w(Z_{i})\big{\{}g(\gamma_{n}(A_{i},W_{i}))\big{\}}^{2}-2m_{2}(Z_{i},g\circ\gamma_{n}).$
Notably, the first-order conditions of the minimizing solutions guarantee that
$\frac{1}{n}\sum_{i=1}^{n}\left[w(Z_{i})(g\circ\alpha_{n}^{*})(W_{i},A_{i})\alpha_{n}^{*}(W_{i},A_{i})-m_{1}(W_{i},A_{i},g\circ\alpha_{n}^{*})\right]$ $=0,$
$\frac{1}{n}\sum_{i=1}^{n}\left[w(Z_{i})(g\circ\gamma_{n}^{*})(W_{i},A_{i})\gamma_{n}^{*}(W_{i},A_{i})-m_{2}(W_{i},A_{i},g\circ\gamma_{n}^{*})\right]$ $=0$
for all transformations $g:\mathbb{R}\rightarrow\mathbb{R}$. As in Algorithms 1 and 2, the initial nuisance estimators can be cross-fitted, and confidence intervals can be constructed using the bootstrap. Our proof techniques can be readily extended to establish that isotonic calibration guarantees doubly robust asymptotic linearity of mixed bias remainders.
E.2 Additional examples
Example 4 (Partial covariance).
Let $\pi_{n}$ and $m_{n}$ be estimators of $w\mapsto E_{0}[A\mid W=w]$ and $w\mapsto E_{0}[Y\mid W=w]$, respectively. The partial covariance parameter
$E_{0}\left[\{A-\pi_{0}(W)\}\{Y-m_{0}(W)\}\right]=E_{0}\left[\{A-\pi_{0}(W)\}^{2}\{\mu_{0}(1,W)-\mu_{0}(0,W)\}\right]$
arises in the context of estimating the ATE under treatment effect homogeneity. Doubly robust inference for this parameter was studied in Dukes et al. (2021) and Bonvini et al. (2024). This parameter is a linear functional of the outcome regression, but the functional depends on the nuisance function $\pi_{0}$. In this case, a modification of our procedure yields improved properties. A doubly robust estimator of the expected covariate parameter is given by
$\frac{1}{n}\sum_{i=1}^{n}\{A_{i}-\pi_{n}(W_{i})\}\{Y_{i}-m_{n}(W_{i})\}.$
The second-order remainder of this estimator involves the cross-product term (Robinson, 1988):
$\langle\pi_{n}(W)-\pi_{0}(W),m_{n}-m_{0}(W)\rangle.$
This mixed bias term is exactly of the form addressed in Section 2.2, involving an inner product between a regression function and a Riesz representer. The regression function $m_{0}$ corresponds to the squared error loss $(z,m)\mapsto\{y-m(w)\}^{2}$, while the function $\pi_{0}$ is a Riesz representer of the linear functional $\alpha\mapsto E_{0}[A\alpha(W)]$, with Riesz loss corresponding to the usual squared error loss $(z,\alpha)\mapsto\{a-\alpha(w)\}^{2}$.
Thus, a calibrated DML estimator of the expected covariate parameter is given by
$\frac{1}{n}\sum_{i=1}^{n}\{A_{i}-\pi_{n}^{*}(W_{i})\}\{Y_{i}-m_{n}^{*}(W_{i})\},$
where $\pi_{n}^{*}=g_{n}\circ\pi_{n}$ and $m_{n}^{*}=f_{n}\circ m_{n}$, with
$g_{n}$ $=\operatorname*{argmin}_{g\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\left\{A_{i}-g(\pi_{n}(W_{i}))\right\}^{2},$
$f_{n}$ $=\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\left\{Y_{i}-f(m_{n}(W_{i}))\right\}^{2}.$
Doubly robust asymptotic linearity of this estimator follows from conditions analogous to those in Section 5.
Example 5 (average treatment effect under covariate-dependent outcome missingness).
Suppose that the observed data unit consists of $(W,S,\Delta,\Delta Y)$, where $S\in\{0,1\}$ is a binary treatment, $\Delta\in\{0,1\}$ is a missingness indicator taking value $1$ if $Y$ is observed, and $\Delta Y$ is the observed outcome, equal to $Y$ if $\Delta=1$ and 00 otherwise. Under causal conditions, the ATE under covariate-dependent outcome missingness is identified by the estimand $\tau_{0}:=E_{0}\{E_{0}(\Delta Y\,|\,S=1,\Delta=1,W)-E_{0}(\Delta Y\,|\,S=0,\Delta=1,W)\}$. Doubly robust inference for this estimand was considered in Díaz and van der Laan (2017). Using the notation of this section, we can write $Z:=(W,A,Y^{\prime})$, where $A:=(S,\Delta)$ is a multi-valued interventional variable and $Y^{\prime}:=\Delta Y$ is an outcome with missingness. The linear functional corresponding with $\tau_{0}$ is $\psi_{0}:\mu\mapsto E_{0}\{\mu((1,1),W)-\mu((0,1),W)\}$. The nonparametric Riesz representer of $\psi_{0}$ is given by $\alpha_{0}:(s,\delta,w)\mapsto(2s-1)\delta/P_{0}(\Delta=1,S=s\,|\,W=w)$.
Appendix F Implementation of calibrated DML
An R package calibratedDML as well as Python code implementing CDML can be found on GitHub at the following link: https://github.com/Larsvanderlaan/calibratedDML.
The following subsections provide R and Python code for calibrating inverse propensity weights and outcome regressions. These calibrated models can be directly used within DML frameworks to construct isocalibrated DML estimators of the average treatment effect. Here, we implement isotonic regression using xgboost, allowing control over the maximum tree depth and the minimum number of observations in each constant segment of the isotonic regression fit.
F.1 R Code
# Function: isoreg_with_xgboost
# Purpose: Fits isotonic regression using XGBoost.
# Inputs:
# - x: A vector or matrix of predictor variables.
# - y: A vector of response variables.
# - max_depth: Maximum depth of the trees in XGBoost (default = 15).
# - min_child_weight: Minimum sum of instance weights (Hessian)
# needed in a child node (default = 20).
# Returns:
# - A function that takes a new predictor variable x
# and returns the model’s predicted values.
isoreg_with_xgboost <- function(x, y, max_depth = 15, min_child_weight = 20) {
# Create an XGBoost DMatrix object from the data
data <- xgboost::xgb.DMatrix(data = as.matrix(x), label = as.vector(y))
# Set parameters for the monotonic XGBoost model
params <- list(max_depth = max_depth,
min_child_weight = min_child_weight,
monotone_constraints = 1, # Enforce monotonic increase
eta = 1, gamma = 0,
lambda = 0)
# Train the model with one boosting round
iso_fit <- xgboost::xgb.train(params = params,
data = data,
nrounds = 1)
# Prediction function for new data
fun <- function(x) {
data_pred <- xgboost::xgb.DMatrix(data = as.matrix(x))
pred <- predict(iso_fit, data_pred)
return(pred)
return(fun)
# Function: calibrate_inverse_weights
# Purpose: Calibrates inverse weights using isotonic regression
# with XGBoost for two propensity scores.
# Inputs:
# - A: Binary indicator variable.
# - pi1: Estimated propensity score for treatment group A = 1.
# - pi0: Estimated propensity score for control group A = 0.
# Returns:
# - A list containing calibrated inverse weights for each group:
# - alpha1_star: Calibrated inverse weights for A = 1.
# - alpha0_star: Calibrated inverse weights for A = 0.
calibrate_inverse_weights <- function(A, pi1, pi0) {
# Calibrate pi1 using monotonic XGBoost
calibrator_pi1 <- isoreg_with_xgboost(pi1, A)
pi1_star <- calibrator_pi1(pi1)
# Set minimum truncation level for treated group
c1 <- min(pi1_star[A == 1])
pi1_star = pmax(pi1_star, c0)
alpha1_star <- 1 / pi1_star
# Calibrate pi0 using monotonic XGBoost
calibrator_pi0 <- isoreg_with_xgboost(pi0, 1 - A)
pi0_star <- calibrator_pi0(pi0)
# Set minimum truncation level for control group
c0 <- min(pi0_star[A == 0])
pi0_star = pmax(pi0_star, c0)
alpha0_star <- 1 / pi0_star
# Return calibrated inverse weights for both groups
return(list(alpha1_star = alpha1_star, alpha0_star = alpha0_star))
# Function: calibrate_outcome_regression
# Purpose: Calibrates outcome regression predictions using isotonic
# regression with XGBoost.
# Inputs:
# - Y: Observed outcomes.
# - mu1: Predicted outcome for treated group.
# - mu0: Predicted outcome for control group.
# - A: Binary indicator variable.
# Returns:
# - A list containing calibrated predictions for each group:
# - mu1_star: Calibrated predictions for A = 1.
# - mu0_star: Calibrated predictions for A = 0.
calibrate_outcome_regression <- function(Y, mu1, mu0, A) {
# Calibrate mu1 using monotonic XGBoost for treated group
calibrator_mu1 <- isoreg_with_xgboost(mu1[A==1], Y[A==1])
mu1_star <- calibrator_mu1(mu1)
# Calibrate mu0 using monotonic XGBoost for control group
calibrator_mu0 <- isoreg_with_xgboost(mu0[A==0], Y[A==0])
mu0_star <- calibrator_mu0(mu0)
# Return calibrated values for both groups
return(list(mu1_star = mu1_star, mu0_star = mu0_star))
F.2 Python code
import xgboost as xgb
import numpy as np
def isoreg_with_xgboost(x, y, max_depth=15, min_child_weight=20):
␣␣␣␣FitsisotonicregressionusingXGBoostwithmonotonicconstraintstoensure
␣␣␣␣non-decreasingpredictionsasthepredictorvariableincreases.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣x(np.array):Avectorormatrixofpredictorvariables.
␣␣␣␣␣␣␣␣y(np.array):Avectorofresponsevariables.
␣␣␣␣␣␣␣␣max_depth(int,optional):MaximumdepthofthetreesinXGBoost.
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣Defaultis15.
␣␣␣␣␣␣␣␣min_child_weight(float,optional):Minimumsumofinstanceweights
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣neededinachildnode.Defaultis20.
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣function:Apredictionfunctionthattakesanewpredictorvariablex
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣andreturnsthemodel’spredictedvalues.
␣␣␣␣Example:
␣␣␣␣␣␣␣␣$>$x=np.array([[1],[2],[3]])
␣␣␣␣␣␣␣␣$>$y=np.array([1,2,3])
␣␣␣␣␣␣␣␣$>$model=isoreg_with_xgboost(x,y)
␣␣␣␣␣␣␣␣$>$model(np.array([[1.5],[2.5]]))
# Create an XGBoost DMatrix object from the data
data = xgb.DMatrix(data=np.asarray(x), label=np.asarray(y))
# Set parameters for the monotonic XGBoost model
params = {
’max_depth’: max_depth,
’min_child_weight’: min_child_weight,
’monotone_constraints’: "(1)", # Enforce monotonic increase
’eta’: 1,
’gamma’: 0,
’lambda’: 0
# Train the model with one boosting round
iso_fit = xgb.train(params=params, dtrain=data, num_boost_round=1)
# Prediction function for new data
def predict_fn(x):
␣␣␣␣␣␣␣␣Predictsoutputfornewinputdatausingthetrainedisotonicregressionmodel.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣x(np.array):Newpredictorvariablesasavectorormatrix.
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣np.array:Predictedvalues.
data_pred = xgb.DMatrix(data=np.asarray(x))
pred = iso_fit.predict(data_pred)
return pred
return predict_fn
def calibrate_inverse_weights(A, pi1, pi0):
␣␣␣␣CalibratesinverseweightsusingisotonicregressionwithXGBoostfortwopropensityscores.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣A(np.array):Binaryindicatorvariable.
␣␣␣␣␣␣␣␣pi1(np.array):Propensityscorefortreatmentgroup(A=1).
␣␣␣␣␣␣␣␣pi0(np.array):Propensityscoreforcontrolgroup(A=0).
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣dict:Containscalibratedinverseweights:
␣␣␣␣␣␣␣␣␣␣␣␣␣␣-alpha1_star:InverseweightsforA=1.
␣␣␣␣␣␣␣␣␣␣␣␣␣␣-alpha0_star:InverseweightsforA=0.
calibrator_pi1 = isoreg_with_xgboost(pi1, A)
pi1_star = calibrator_pi1(pi1)
c1 = np.min(pi1_star[A == 1])
pi1_star = np.maximum(pi1_star, c1)
alpha1_star = 1 / pi1_star
calibrator_pi0 = isoreg_with_xgboost(pi0, 1 - A)
pi0_star = calibrator_pi0(pi0)
c0 = np.min(pi0_star[A == 0])
pi0_star = np.maximum(pi0_star, c0)
alpha0_star = 1 / pi0_star
return {’alpha1_star’: alpha1_star, ’alpha0_star’: alpha0_star}
def calibrate_outcome_regression(Y, mu1, mu0, A):
␣␣␣␣CalibratesoutcomeregressionpredictionsusingisotonicregressionwithXGBoost.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣Y(np.array):Observedoutcomes.
␣␣␣␣␣␣␣␣mu1(np.array):Predictedoutcomeforthetreatedgroup(A=1).
␣␣␣␣␣␣␣␣mu0(np.array):Predictedoutcomeforthecontrolgroup(A=0).
␣␣␣␣␣␣␣␣A(np.array):Binarytreatmentindicator(1fortreatment,0forcontrol).
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣dict:Calibratedpredictionsforeachgroup:
␣␣␣␣␣␣␣␣␣␣␣␣␣␣-’mu1_star’:CalibratedpredictionsforA=1.
␣␣␣␣␣␣␣␣␣␣␣␣␣␣-’mu0_star’:CalibratedpredictionsforA=0.
# Calibrate mu1 using treated group data (A = 1)
calibrator_mu1 = isoreg_with_xgboost(mu1[A == 1], Y[A == 1])
mu1_star = calibrator_mu1(mu1) # Apply calibrator to all mu1 values
# Calibrate mu0 using control group data (A = 0)
calibrator_mu0 = isoreg_with_xgboost(mu0[A == 0], Y[A == 0])
mu0_star = calibrator_mu0(mu0) # Apply calibrator to all mu0 values
return {’mu1_star’: mu1_star, ’mu0_star’: mu0_star}
Appendix G Proofs of supporting theory and notation
G.1 Notation
Let $J\in\mathbb{N}$ denote a fixed number of cross-fitting splits. Let $\mathcal{D}_{n}^{1},\mathcal{D}_{n}^{2},\ldots,\mathcal{D}_{n}^{J}$ be a partition of the available data $\mathcal{D}_{n}$ into $J$ datasets of approximately equal size, corresponding to index sets $\mathcal{I}^{1}_{n},\mathcal{I}^{2}_{n},\ldots,\mathcal{I}^{J}_{n}$. For each $i\in[n]$, let $j(i)\in[J]$ be the index of the data fold containing observation $i$, so that $i\in\mathcal{I}_{n}^{j(i)}$. For ease of presentation, we will use the following cross-fitting notation. For a collection of fold-specific functions $\nu_{n,\diamond}:=\{\nu_{n,j}:j\in[J]\}$ that may depend on $n$, we use the fold-averaged empirical process notation $\overline{P}_{0}\nu_{n,\diamond}:=\frac{1}{J}\sum_{j=1}^{J}P_{0}\nu_{n,j}$ and $\overline{P}_{n}\nu_{n,\diamond}:=\frac{1}{J}\sum_{j=1}^{J}P_{n,j}\nu_{n,j}$. This notation allows us to view the collections of calibrated cross-fitted estimators $\alpha_{n,\diamond}^{*}:=\{\alpha_{n,j}^{*}:j\in[J]\}$ and $\mu_{n,\diamond}^{*}:=\{\mu_{n,j}^{*}:j\in[J]\}$ as extended functions defined on $\mathcal{A}\times\mathcal{W}\times[J]$.
For a uniformly bounded function class $\mathcal{F}$, let $N(\epsilon,\mathcal{F},L_{2}(P))$ denote the $\epsilon-$covering number (van der Vaart and Wellner, 1996) of $\mathcal{F}$ relative to the $L^{2}(P)$ metric and define the uniform entropy integral of $\mathcal{F}$ by
$\mathcal{J}(\delta,\mathcal{F}):=\int_{0}^{\delta}\sup_{Q}\sqrt{\log N(\epsilon,\mathcal{F},L_{2}(Q))}\,d\epsilon\ ,$
where the supremum is taken over all discrete probability distributions $Q$. In contrast to the definition provided in van der Vaart and Wellner (1996), we do not define the uniform entropy integral relative to an envelope function for the function class $\mathcal{F}$. We can do this since all function classes we consider are uniformly bounded. Thus, any uniformly bounded envelope function will only change the uniform entropy integral as defined in van der Vaart and Wellner (1996) by a constant.
Let $k(n)=o(\sqrt{n})$ be the upper bound on the number of jump points of $f_{n}$ and $g_{n}$ in C6. Since C6 states that $k(n)$ is an upper bound with probability tending to one, we can, without loss of generality, assume that $k(n)$ is an upper bound. Let $M/2<\infty$ be an upper bound on the nuisance functions and estimators in C5. Let $\mathcal{F}_{\text{iso}}\subset\{f:[-M,M]\rightarrow\mathbb{R}|\text{$f$ is monotone nondecreasing}\}$ consist of all univariate, isotonic functions with supremum norm smaller than $M<\infty$. Let $\mathcal{F}_{n}\subset\{f:[-M,M]\rightarrow\mathbb{R}|\text{$f$ is piecewise constant with at most $k(n)$ jumps}\}$ consist of all univariate functions with supremum norm smaller than $M<\infty$, that are left-continuous with right-hand limits and have at most $k(n)<\infty$ jump points (i.e., discontinuities). For $j\in[J]$, we define $\mathcal{S}_{n,j}:=\left\{f\circ\mu_{n,j}|f\in\mathcal{F}_{n}\cup\mathcal{F}_{\text{iso}}\right\}$ and $\mathcal{R}_{n,j}:=\left\{g\circ\alpha_{n,j}|g\in\mathcal{F}_{n}\cup\mathcal{F}_{\text{iso}}\right\}$. By C6 and C5, we know $\mu_{n,j}^{*}=f_{n}\circ\mu_{n,j}$ and $s_{n,j}^{*}$ fall in $\mathcal{S}_{n,j}$ (with probability tending to one). Similarly, we know $\alpha_{n,j}^{*}=g_{n}\circ\alpha_{n,j}$ and $r_{n,j}^{*}$ fall in $\mathcal{R}_{n,j}$. Note that both $\mathcal{S}_{n,j}$ and $\mathcal{R}_{n,j}$ are fixed function classes conditional on the $j$-th training set $\mathcal{D}_{n}\backslash\mathcal{C}_{j}$.
For two quantities $x$ and $y$, we use the expression $x\lesssim y$ to mean that $x$ is upper bounded by $y$ times a universal constant that may only depend on global constants that appear in C5. We denote by $\mathcal{I}_{Y}$ the coordinate projection map $z=(w,a,y)\mapsto y$; for example, $P_{0}\{\mathcal{I}_{Y}-\mu_{0}\}=E_{0}[Y-\mu_{0}(A,W)]$.
G.2 Expansions for doubly robust estimators
In this section, we let $\alpha_{n,\diamond}=\{\alpha_{n,j}:j\in[J]\}$ and $\mu_{n,\diamond}=\{\mu_{n,j}:j\in[J]\}$ be arbitrary fold-dependent estimators of $\alpha_{0}$ and $\mu_{0}$ that are potentially uncalibrated. We denote the associated one-step debiased estimator by
$\tau_{n}:=\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(\mu_{n,j})+\frac{1}{n}\sum_{i=1}^{n}\alpha_{n,j(i)}(A_{i},W_{i})\{Y_{i}-\mu_{n,j(i)}(A_{i},W_{i})\}.$
The following expansion of the estimation error $\tau_{n}-\tau_{0}$ is a standard starting point in the analysis of doubly robust estimators.
Lemma 4 (Standard doubly robust bias expansion).
It holds that
$\tau_{n}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}}+D_{\mu_{n,\diamond},\alpha_{n,\diamond}}\right\}+\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}+Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}).$
where $Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}):=\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(\mu_{n,j})-\psi_{0}(\mu_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{0,\mu_{n,j}}\right\}$.
Proof.
We can write
$\frac{1}{J}\sum_{j=1}^{J}\{\psi_{n,j}(\mu_{n,j})-\psi_{0}(\mu_{n,j})\}=(\overline{P}_{n}-\overline{P}_{0})\widetilde{\phi}_{0,\mu_{n,\diamond}}+Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}),$
where $Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}):=\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(\mu_{n,j})-\psi_{0}(\mu_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{0,\mu_{n,j}}\right\}$. The estimation error of $\tau_{n}$ can be decomposed as:
$\tau_{n}-\tau_{0}$ $=\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(\mu_{n,j})+\overline{P}_{n}D_{\mu_{n,\diamond},\alpha_{n,\diamond}}-\psi_{0}(\mu_{0})$
$=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}}+D_{\mu_{n,\diamond},\alpha_{n,\diamond}}\right\}+\overline{P}_{0}D_{\mu_{n,\diamond},\alpha_{n,\diamond}}+\frac{1}{J}\sum_{j=1}^{J}\{\psi_{0}(\mu_{n,j})-\psi_{0}(\mu_{0})\}+Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}).$
By the Reisz representation $\psi_{0}(\cdot)=\langle\alpha_{0},\cdot\rangle_{P_{0}}$, it holds that $\psi_{0}(\mu_{n,j})-\psi_{0}(\mu_{0})=\langle\alpha_{0},\mu_{n,j}-\mu_{0}\rangle_{P_{0}}$. Moreover, by the law of iterated expectations, we have
[…]
$=\langle\alpha_{n,\diamond},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}.$
Hence, $\overline{P}_{0}D_{\mu_{n,\diamond},\alpha_{n,\diamond}}+\frac{1}{J}\sum_{j=1}^{J}\{\psi_{0}(\mu_{n,j})-\psi_{0}(\mu_{0})\}=\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}.$ Therefore,
$\tau_{n}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}}+D_{\mu_{n,\diamond},\alpha_{n,\diamond}}\right\}+\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}+Rem_{\mu_{n,\diamond}}(\psi_{n,\diamond},\psi_{0}).$
The next lemma establishes the leading order behavior of the cross-product remainder term in both the outcome- and Riesz-favorable settings. We recall some of the notation from Section 2.2. Let $s_{n,j}:=\Pi_{(\mu_{n,j})}\{\alpha_{0}-\alpha_{n,j}\}$ and $r_{n,j}:=\Pi_{(\alpha_{n,j})}\{\mu_{0}-\mu_{n,j}\}$ denote the orthogonal projections of the error terms $\alpha_{0}-\alpha_{n,j}$ and $\mu_{0}-\mu_{n,j}$ onto $\mu_{n,j}$ and $\alpha_{n,j}$, respectively. Define $A_{n,j}:z\mapsto s_{n,j}(a,w)\{y-\mu_{n,j}(a,w)\}$, and let $B_{n,j}:=\psi_{0}(r_{0})+\widetilde{\phi}_{r_{n,j}}-r_{n,j}\alpha_{n,j}$.
Lemma 5 (Decompositions of cross-product remainder).
We have the outcome-favorable decomposition:
$\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})A_{n,\diamond}-\overline{P}_{n}A_{n,\diamond}+O_{p}\left(\|(\Pi_{(\mu_{n,\diamond},\mu_{0})}-\Pi_{\mu_{n,\diamond}})(\alpha_{n,\diamond}-\alpha_{0})\|_{\overline{P}_{0}}\|\mu_{0}-\mu_{n,\diamond}\|_{\overline{P}_{0}}\right),$
Similarly, we have the Riesz-favorable decomposition:
$\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})B_{n,\diamond}-\left\{\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(r_{n,j})-\overline{P}_{n}r_{n,\diamond}\alpha_{n,\diamond}\right\}$
$\quad+\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}}\right\}$
$\quad+O_{p}(\|\alpha_{n,\diamond}-\alpha_{0}\|_{\overline{P}_{0}}\|(\Pi_{(\alpha_{n,\diamond},\alpha_{0})}-\Pi_{\alpha_{n,\diamond}})(\mu_{0}-\mu_{n,\diamond})\|_{\overline{P}_{0}},$
Proof.
By definition, $\langle\alpha_{n,\diamond}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}\rangle_{\overline{P}_{0}}=\frac{1}{J}\sum_{j=1}^{J}\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$. Next, by the law of total expectation, we have, for each $j\in[J]$, that
$\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$ $=\langle\Pi_{(\mu_{n,j},\mu_{0})}\{\alpha_{n,j}-\alpha_{0}\},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$
$=\langle\Pi_{\mu_{n,j}}\{\alpha_{n,j}-\alpha_{0}\},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}+\langle(\Pi_{(\mu_{n,j},\mu_{0})}-\Pi_{\mu_{n,j}})(\alpha_{n,j}-\alpha_{0}),\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$
$=\langle s_{n,j},\mu_{n,j}-\mu_{0}\rangle_{P_{0}}+\langle(\Pi_{(\mu_{n,j},\mu_{0})}-\Pi_{\mu_{n,j}})(\alpha_{n,j}-\alpha_{0}),\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$
$=\langle s_{n,j},\mu_{n,j}-\mu_{0}\rangle_{P_{0}}+O_{p}\left(\|(\Pi_{(\mu_{n,j},\mu_{0})}-\Pi_{\mu_{n,j}})(\alpha_{n,j}-\alpha_{0})\|\|\mu_{0}-\mu_{n,j}\|_{P_{0}}\right),$
where the final equality follows from the Cauchy-Schwarz inequality. The leading term on the right-hand side of the above display can be expressed as
[…]
$=(P_{n,j}-P_{0})A_{n,j}-P_{n,j}A_{n,j}.$
Hence,
$\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$ $=(P_{n,j}-P_{0})A_{n,j}-P_{n,j}A_{n,j}+O_{p}\left(\|(\Pi_{(\mu_{n,j},\mu_{0})}-\Pi_{\mu_{n,j}})(\alpha_{n,j}-\alpha_{0})\|\|\mu_{0}-\mu_{n,j}\|_{P_{0}}\right),$
Averaging the above display over $j\in[J]$, the first equality then follows.
We now derive the second equality. Similarly, by the law of total expectation and the Cauchy-Schwarz inequality, we have, for each $j\in[J]$, that
$\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$ $=\langle\alpha_{n,j}-\alpha_{0},\Pi_{(\alpha_{n,j},\alpha_{0})}(\mu_{0}-\mu_{n,j})\rangle_{P_{0}}$
$=\langle\alpha_{n,j}-\alpha_{0},\Pi_{\alpha_{n,j}}\{\mu_{0}-\mu_{n,j}\}\rangle_{P_{0}}+\langle\alpha_{n,j}-\alpha_{0},(\Pi_{(\alpha_{n,j},\alpha_{0})}-\Pi_{\alpha_{n,j})})(\mu_{0}-\mu_{n,j})\rangle_{P_{0}}$
$=\langle\alpha_{n,j}-\alpha_{0},r_{n,j}\rangle_{P_{0}}+O_{p}(\|\alpha_{n,j}-\alpha_{0}\|\|(\Pi_{(\alpha_{n,j},\alpha_{0})}-\Pi_{\alpha_{n,j})})(\mu_{0}-\mu_{n,j})\|,$
where $\langle\alpha_{n,j}-\alpha_{0},r_{n,j}\rangle_{P_{0}}=\langle r_{n,j},\alpha_{n,j}\rangle_{P_{0}}-\psi_{0}(r_{n,j})$. The leading term on the right-hand side of the above display can be expressed as
$\langle r_{n,j},\alpha_{n,j}\rangle_{P_{0}}-\psi_{0}(r_{n,j})$ $=P_{0}r_{n,j}\alpha_{n,j}-\psi_{0}(r_{n,j})$
$=P_{n,j}r_{n,j}\alpha_{n,j}-\psi_{n,j}(r_{n,j})+\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})+(P_{0}-P_{n,j})\left\{r_{n,j}\alpha_{n,j}\right\}$
$=P_{n,j}r_{n,j}\alpha_{n,j}-\psi_{n,j}(r_{n,j})+(P_{0}-P_{n,j})\left\{r_{n,j}\alpha_{n,j}-\widetilde{\phi}_{r_{n,j}}\right\}$
$\quad+\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}}$
$=P_{n,j}r_{n,j}\alpha_{n,j}-\psi_{n,j}(r_{n,j})+(P_{n,j}-P_{0})\left\{\widetilde{\phi}_{r_{n,j}}-r_{n,j}\alpha_{n,j}\right\}$
$\quad+\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}}.$
We note that $(P_{n,j}-P_{0})B_{n,j}=(P_{n,j}-P_{0})\left\{\widetilde{\phi}_{r_{n,j}}-r_{n,j}\alpha_{n,j}\right\}$ since $\psi_{0}(r_{0})$ is a constant and, therefore, satisfies $(P_{n,j}-P_{0})(\psi_{0}(r_{0}))=0$. Hence,
$\langle\alpha_{n,j}-\alpha_{0},\mu_{0}-\mu_{n,j}\rangle_{P_{0}}$ $=P_{n,j}r_{n,j}\alpha_{n,j}-\psi_{n,j}(r_{n,j})+(P_{n,j}-P_{0})\left\{\widetilde{\phi}_{r_{n,j}}-r_{n,j}\alpha_{n,j}\right\}$
$\quad+\psi_{n,j}(r_{n,j})-\psi_{0}(r_{n,j})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}}$
$\quad+O_{p}(\|\alpha_{n,j}-\alpha_{0}\|\|(\Pi_{(\alpha_{n,j},\alpha_{0})}-\Pi_{\alpha_{n,j})})(\mu_{0}-\mu_{n,j})\|,$
Averaging the above displays over $j\in[J]$, the second equality then follows. ∎
G.3 Maximal inequalities for empirical processes and the bootstrap
Lemma 6.
Suppose $\mathcal{A}$ is finite, and let $m:\mathcal{Z}\times\mathcal{H}\rightarrow\mathbb{R}$ be a functional that is linear in its second argument. Assume there exists $L\in(0,\infty)$ such that $|m(Z,h)|\leq L\max_{a\in\mathcal{A}}|h(a,W)|$ almost surely. Let $\nu\in\mathcal{H}$ be a fixed, uniformly bounded function, and let $\mathcal{F}$ be a class of functions from $\mathbb{R}$ to $\mathbb{R}$ with $\sup_{f\in\mathcal{F}}\|f\|_{\infty}\leq B$ and $\mathcal{J}(\infty,\mathcal{F})<\infty$. Define the function classes $\mathcal{F}_{\nu}:=\{f\circ\nu:f\in\mathcal{F}\}$ and $\mathcal{F}_{m,\nu}:=\{m(\cdot,f):f\in\mathcal{F}_{\nu}\}$. Then, there exists an $M<\infty$ depending only on $L$ such that, for all $\delta>0$, $\mathcal{J}(\delta,\mathcal{F}_{\nu})\leq M\mathcal{J}(\delta,\mathcal{F})$ and $\mathcal{J}(\delta,\mathcal{F}_{m,\nu})\leq M\mathcal{J}(\delta,\mathcal{F})$.
Proof.
Since $|m(Z,h)|\leq L\max_{a\in\mathcal{A}}|h(a,W)|$ almost surely, we have
$N(\varepsilon,\mathcal{F}_{m,\nu},L^{2}(Q))\lesssim\sum_{a\in\mathcal{A}}N(\varepsilon,\mathcal{F}_{\nu}^{(a)},L^{2}(Q)),$
where we define the function class $\mathcal{F}_{\nu}^{(a)}:=\{w\mapsto f(a,w):f\in\mathcal{F}_{\nu}\}$ and use the finiteness of $\mathcal{A}$. Hence,
$\mathcal{J}(\delta,\mathcal{F}_{m,\nu})\lesssim\sum_{a\in\mathcal{A}}\mathcal{J}(\delta,\mathcal{F}_{\nu}^{(a)}).$
By the definition of $\mathcal{F}_{\nu}$, we have $\mathcal{F}_{\nu}^{(a)}=\{w\mapsto f(\nu(a,w)):f\in\mathcal{F}\}$. Now, for all $a\in\mathcal{A}$,
$\mathcal{J}(\delta,\mathcal{F}_{\nu}^{(a)})$ $=\int_{0}^{\delta}\sup_{Q}\sqrt{N(\varepsilon,\mathcal{F}_{\nu}^{(a)},L^{2}(Q))}\,d\varepsilon$
$=\int_{0}^{\delta}\sup_{Q_{a,\nu}}\sqrt{N(\varepsilon,\mathcal{F},L^{2}(Q_{a,\nu}))}\,d\varepsilon$
$=\mathcal{J}(\delta,\mathcal{F}),$
where $Q_{a,\nu}$ denotes the pushforward measure of the random variable $\nu(a,W)$. It follows that $\mathcal{J}(\delta,\mathcal{F}_{m,\nu})\lesssim\mathcal{J}(\delta,\mathcal{F})$. By an identical argument, we also have $\mathcal{J}(\delta,\mathcal{F}_{\nu})\lesssim\mathcal{J}(\delta,\mathcal{F})$. ∎
Lemma 7.
For $j\in[J]$, it holds that $\mathcal{J}(\delta,\mathcal{R}_{n,j})+\mathcal{J}(\delta,\mathcal{S}_{n,j})\lesssim\sqrt{\delta}\left(1+\sqrt{k(n)\delta\log(1/\delta)}\right)$.
Proof.
We establish this explicitly for $\mathcal{S}_{n,j}$ below, as the result for $\mathcal{R}_{n,j}$ follows from an identical argument. To this end, we have
$\mathcal{J}(\delta,\mathcal{S}_{n,j})$ $=\int_{0}^{\delta}\sup_{Q}\sqrt{N(\varepsilon,\mathcal{S}_{n,j},\norm{\,\cdot\,}_{Q})}\,d\varepsilon=\int_{0}^{\delta}\sup_{Q}\sqrt{N(\varepsilon,\mathcal{F}_{n}\cup\mathcal{F}_{\text{iso}},\norm{\,\cdot\,}_{Q\circ\mu_{n,j}^{-1}})}\,d\varepsilon$
$\leq\int_{0}^{\delta}\sup_{Q}\sqrt{N(\varepsilon,\mathcal{F}_{n},\norm{\,\cdot\,}_{Q\circ\mu_{n,j}^{-1}})}\,d\varepsilon+\int_{0}^{\delta}\sup_{Q}\sqrt{N(\varepsilon,\mathcal{F}_{\text{iso}},\norm{\,\cdot\,}_{Q\circ\mu_{n,j}^{-1}})}\,d\varepsilon$
$\leq\mathcal{J}(\delta,\mathcal{F}_{n})+\mathcal{J}(\delta,\mathcal{F}_{iso}),$
where $Q\circ\mu_{n,j}^{-1}$ is the push-forward probability measure for the random variable $\mu_{n,j}(W)$ conditional on $\mathcal{E}_{n}^{(j)}=\mathcal{D}_{n}\backslash\mathcal{C}^{(j)}$. Since $\mathcal{J}(\delta,\mathcal{F}_{n})$ consists of univariate functions with at most $k(n)$ constant segments, it has VC-subgraph dimension $O(k(n))$. As a consequence, $\mathcal{J}(\delta,\mathcal{F}_{n})\lesssim\delta\sqrt{k(n)\log(1/\delta)}$, as desired. Moreover, by standard metric entropy results for monotone functions, we have $\mathcal{J}(\delta,\mathcal{F}_{iso})\lesssim\sqrt{\delta}$. ∎
Lemma 8.
For each $j\in[J]$, let $\mathcal{H}_{n,j}=\mathcal{F}_{n,j}\times\mathcal{F}_{n,j}\times\mathcal{F}_{n,j}\times\mathcal{F}_{n,j}$ for a function class $\mathcal{F}_{n,j}$ that is deterministic conditional on the $j$-th training set $\mathcal{D}_{n}\backslash\mathcal{C}^{(j)}$. Assume that $\mathcal{J}(\delta,\mathcal{F}_{n,j})\leq t_{n}(\delta)$ for some map $\delta\mapsto t_{n}(\delta)$. For each $j\in[J]$, let $T_{n,j}:\mathbb{R}^{4}\rightarrow\mathbb{R}$ be a bounded Lipchitz function that is deterministic conditional on $\mathcal{D}_{n}\backslash\mathcal{C}^{(j)}$. Then, for all $\delta>0$,
$E_{0}^{n}\left[\sup_{h_{\diamond}\in\mathcal{H}_{n,\diamond}:\|T_{n,\diamond}(h_{\diamond}(\cdot))\|_{\overline{P}_{0}}\leq\delta}\left(\overline{P}_{n}-\overline{P}_{0}\right)T_{n,\diamond}(h_{\diamond}(\cdot))\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right).$
Furthermore, for any $\sqrt{n}\delta^{2}>t_{n}(\delta)$,
$E_{0}^{n}\left[\sup_{h_{\diamond}\in\mathcal{H}_{n,\diamond}:\|T_{n,\diamond}(h_{\diamond}(\cdot))\|_{\overline{P}_{0}}\leq\delta}\left(\overline{P}_{n}^{\#}-\overline{P}_{n}\right)T_{n,\diamond}(h_{\diamond}(\cdot))\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right).$
Proof.
To establish the first bound, by the triangle inequality, it suffices to show, for each $j\in[J]$, that
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|T_{n,j}(h)(\cdot)\|\leq J\delta}\left(P_{n,j}-P_{0}\right)T_{n,j}(h)\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right).$
By Theorem 2.10.20 of van der Vaart and Wellner (1996) on preservation of entropy integrals under Lipschitz transformation and Lemma 7, we have that $\mathcal{H}_{n,j}$ is a uniformly bounded function class with uniform entropy integral satisfying $\mathcal{J}(\delta,\mathcal{H}_{n,j})\lesssim 4\mathcal{J}(\delta,\mathcal{F}_{n,j})\lesssim t_{n}(\delta)$. Hence, by Theorem 2.1 of van der Vaart and Wellner (2011),
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|T_{n,j}(h(\cdot))\|\leq J\delta}\left(P_{n,j}-P_{0}\right)T_{n,j}(h)\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right).$
Next, noting that $P_{n,j}^{\#}$ is the empirical distribution of $O(n/J)$ i.i.d. samples from $P_{n,j}$, we can apply the same argument conditional on $P_{n,j}$ with $\delta_{n}:=\sup_{h\in\mathcal{H}_{n,j}:\|T_{n,j}(h(\cdot))\|_{P_{0}}\leq\delta}\|T_{n,j}(h(\cdot))\|_{P_{n,j}}$ to find:
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|T_{n,j}(h(\cdot))\|_{P_{0}}\leq J\delta}\left(P_{n,j}^{\#}-P_{n,j}\right)T_{n,j}(h)|P_{n,j}\right]$ $\lesssim n^{-\frac{1}{2}}\{\mathcal{J}(\delta_{n},\mathcal{S}_{n,j})+\mathcal{J}(\delta_{n},\mathcal{R}_{n,j})\}\left(1+\frac{\mathcal{J}(\delta_{n},\mathcal{S}_{n,j})+\mathcal{J}(\delta_{n},\mathcal{R}_{n,j})}{\sqrt{n}\delta_{n}^{2}}\right)$
$\lesssim n^{-\frac{1}{2}}t_{n}(\delta_{n})\left(1+\frac{t_{n}(\delta_{n})}{\sqrt{n}\delta_{n}^{2}}\right)$
We can show that $\varepsilon\mapsto t_{n}(\sqrt{\varepsilon})\left(1+\frac{t_{n}(\sqrt{\varepsilon})}{\sqrt{n}\varepsilon}\right)$ is a concave function. Thus, taking the expectation over both sides of the previous display and applying Jenson’s inequality, we find for $\widetilde{\delta}_{n}:=\sqrt{E\delta_{n}^{2}}$ that
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|T_{n,j}(h(\cdot))\|_{P_{0}}\leq J\delta}\left(P_{n,j}^{\#}-P_{n,j}\right)T_{n,j}(h)\right]\lesssim n^{-\frac{1}{2}}t_{n}(\widetilde{\delta}_{n})\left(1+\frac{t_{n}(\widetilde{\delta}_{n})}{\sqrt{n}\widetilde{\delta}_{n}^{2}}\right)$
Moreover, by Theorem 2.2 of van de Geer (2014), we have $\widetilde{\delta}_{n}^{2}=E\delta_{n}^{2}\lesssim\delta^{2}$ for any $\delta>0$ with $\sqrt{n}\delta^{2}>t_{n}(\delta)$. Combining this with the previous display, we obtain
$E_{0}^{n}\left[\sup_{h\in\mathcal{H}_{n,j}:\|T_{n,j}(h(\cdot))\|_{P_{0}}\leq\delta}\left(P_{n,j}^{\#}-P_{n,j}\right)T_{n,j}(h)\right]\lesssim n^{-\frac{1}{2}}t_{n}(\delta)\left(1+\frac{t_{n}(\delta)}{\sqrt{n}\delta^{2}}\right),$
for any $\delta^{2}>n^{-\frac{1}{2}}t_{n}(\delta)$. ∎
Lemma 9.
For each $j\in[J]$, let $T_{n,j}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ be a Lipschitz continuous function that is deterministic conditional on $\mathcal{D}_{n}\backslash\mathcal{C}^{(j)}$. Moreover, let $\rho_{n,j}:=T_{n,j}(\mu_{n,j}^{*},s_{n,j}^{*},\alpha_{n,j}^{*},r_{n,j}^{*})$ be a random function by applying the Lipschitz transformation $T_{n,j}$ to the estimated nuisance functions $(\mu_{n,j}^{*},s_{n,j}^{*},\alpha_{n,j}^{*},r_{n,j}^{*})$. Suppose $\|\rho_{n,\diamond}\|_{\overline{P}_{0}}=O_{P}(\varepsilon_{n})$ for a deterministic sequence $\varepsilon_{n}\rightarrow 0$. Then, $\left(\overline{P}_{n}-\overline{P}_{0}\right)\rho_{n,\diamond}=o_{p}(n^{-\frac{1}{2}})$. Moreover, when $\overline{P}_{n},\alpha_{n,j}^{*}$, $\mu_{n,j}^{*}$, $\rho_{n,j}$ are replaced with the bootstrap estimators $\overline{P}_{n}^{\#},\alpha_{n,j}^{*\#}$, $\mu_{n,j}^{*\#}$, and $\rho_{n,j}^{\#}$, we also have $\left(\overline{P}_{n}^{\#}-\overline{P}_{n}\right)\rho_{n,\diamond}^{\#}=o_{p}(n^{-\frac{1}{2}})$ whenever $\|\rho_{n,\diamond}^{\#}\|_{\overline{P}_{0}}=O_{P}(\varepsilon_{n})$.
Proof.
Since $\|\rho_{n,\diamond}\|_{\overline{P}_{0}}=O_{P}(\varepsilon_{n})$ for some $\varepsilon_{n}\rightarrow 0$, there exists a deterministic sequence $\delta_{n}\rightarrow 0$ such that $\|\rho_{n,\diamond}\|_{\overline{P}_{0}}\lesssim\delta_{n}$ with probability tending to one. Let $A_{n}$ denote the event that $\|\rho_{n,\diamond}\|_{\overline{P}_{0}}\lesssim\delta_{n}$ holds. On this event, we have that
$\left(\overline{P}_{n}-\overline{P}_{0}\right)\rho_{n,\diamond}\leq\sup_{h_{\diamond}\in\mathcal{H}_{n,\diamond}:\|T_{n,\diamond}(h_{\diamond}(\cdot))\|_{\overline{P}_{0}}\leq\delta_{n}}\left(\overline{P}_{n}-\overline{P}_{0}\right)T_{n,\diamond}(h_{\diamond}(\cdot)).$
Moreover, Lemma 8 applied on $A_{n}$ implies
$E_{0}^{n}\left[1(A_{n})\sup_{h_{\diamond}\in\mathcal{H}_{n,\diamond}:\|T_{n,\diamond}(h_{\diamond}(\cdot))\|_{\overline{P}_{0}}\leq\delta_{n}}\left(\overline{P}_{n}-\overline{P}_{0}\right)T_{n,\diamond}(h_{\diamond}(\cdot))\right]=o(n^{-\frac{1}{2}}).$
Hence, by Markov’s inequality, $1(A_{n})\left(\overline{P}_{n}-\overline{P}_{0}\right)\rho_{n,\diamond}=o(n^{-\frac{1}{2}})$. Since $A_{n}$ occurs with probability tending to one, we also have $\left(\overline{P}_{n}-\overline{P}_{0}\right)\rho_{n,\diamond}=o_{p}(n^{-\frac{1}{2}})$, as desired. The statement for the bootstrap empirical process follows from an identical argument, using the second part of Lemma 8, noting that $P_{n,j}^{\#}$ is the empirical distribution of $\approx n/J$ i.i.d. samples from $P_{n,j}$. ∎
G.4 Convergence rates for isotonic calibrated nuisance estimators
Let $\alpha_{n,\diamond}^{*}:=\{\alpha_{n,j}^{*}:j\in[J]\}$ and $\mu_{n,\diamond}^{*}:=\{\mu_{n,j}^{*}:j\in[J]\}$ be the isotonic calibrated nuisance estimators obtained via Algorithm 1. Let $\overline{\alpha}_{0}$ and $\overline{\mu}_{0}$ be their respective potentially misspecified limits, as in C2. In this section, we determine the mean square error rates for the calibrated nuisance estimators $\alpha_{n,j}^{*}$ and $\mu_{n,j}^{*}$ with respect to $\alpha_{0}$ and $\mu_{0}$. We focus our convergence rate results on the case where the linear functional $\psi_{0}$ takes the form $\mu\mapsto E_{0}\{m(Z,\mu)\}$ for a known map $m:\mathcal{Z}\times\mathcal{H}\rightarrow\mathbb{R}$. Additionally, we assume that the linear functional estimator $\psi_{n}$ is taken as the empirical plug-in estimator $\mu\mapsto\frac{1}{n}\sum_{i=1}^{n}m(Z_{i},\mu)$. Our convergence rate results can be extended to general estimators $\psi_{n}$, provided that certain regularity conditions hold for the second-order remainder.
We begin by providing a general result on the convergence rates of isotonic calibrated nuisance estimators for Riesz representers of bounded linear functionals (Chernozhukov et al., 2021), of which both the outcome regression $\mu_{0}$ and the Riesz representer $\alpha_{0}$ are special cases. Defined over $z\in\mathcal{Z}$ and $\nu\in\mathcal{H}$, consider the Riesz loss function $\ell(z,\nu):=\nu^{2}(a,w)-2m(z,\nu)$ for some mapping $m$ that is linear in its second argument. Let $\{\nu_{n,j}:j\in[J]\}$ be cross-fitted estimators of the population risk minimizer $\nu_{0}:=\operatorname*{argmin}_{\nu\in\mathcal{H}}P_{0}\ell(\cdot,\nu)$. Define $\nu_{n,\diamond}^{*}:=h_{n}\circ\nu_{n,\diamond}$ and $\nu_{n,\diamond}^{*\#}:=h_{n}^{\#}\circ\nu_{n,\diamond}$ where we define the isotonic regression solutions:
$h_{n}$ $:=\operatorname*{argmin}_{h\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\ell(Z_{i},h\circ\nu_{n,j(i)});$
$h_{n}^{\#}$ $:=\operatorname*{argmin}_{h\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\ell(Z_{i}^{\#},h\circ\nu_{n,j(i)});$
$h_{n,0}$ $:=\operatorname*{argmin}_{h\in\mathcal{F}_{iso}}\sum_{j=1}^{J}E_{0}[\ell(Z,h\circ v_{n,j(i)})]=\operatorname*{argmin}_{h\in\mathcal{F}_{iso}}\|h\circ\nu_{n,\diamond}-\nu_{0}\|_{\overline{P}_{0}}.$
We make the following conditions on the loss $\ell$ and the estimators. In the following, for each $\nu\in\mathcal{H}$, define the function classes $\mathcal{F}_{\ell,\nu,\mathrm{iso}}:=\{\ell(z,f\circ\nu):f\in\mathcal{F}_{\mathrm{iso}}\}$ and $\mathcal{F}_{\nu,\mathrm{iso}}:=\{\ell(z,f\circ\nu):f\in\mathcal{F}_{\mathrm{iso}}\}$.
  1. (F1)
    $h_{n}$, $h_{n}^{\#}$, and $h_{n,0}$ lie in a fixed, uniformly bounded subset of $\mathcal{F}_{\mathrm{iso}}$ with probability tending to one.
  2. (F2)
    There exists $L<\infty$ such that $|E_{0}[m(\cdot,h)]|\leq L\|h\|_{P_{0}}$ for all $h\in\mathcal{H}$.
  3. (F3)
    $\mathcal{A}$ is finite, and there exists a constant $L\in(0,\infty)$ such that, for all $\mu\in\mathcal{H}$, $|m(Z,\mu)|\leq L\max_{a\in\mathcal{A}}|\mu(a,W)|$ almost surely.
Theorem 6.
Then, $\|h_{n}\circ\nu_{n,\diamond}-h_{n,0}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}=O_{p}(n^{-1/3})$. Hence,
$\|h_{n}\circ\nu_{n,\diamond}-\nu_{0}\|_{\overline{P}_{0}}$ $\leq\inf_{h\in\mathcal{F}_{iso}}\|h\circ\nu_{n,\diamond}-\nu_{0}\|+O_{p}(n^{-1/3})$
Moreover, the bootstrap calibrated nuisance estimators satisfy, for any function $\overline{\nu}_{0}\in L^{2}(P_{0})$ and as $n\rightarrow\infty$, the following root mean square error bounds:
$\|h_{n}^{\#}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}$ $\leq\|h_{n}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3}).$
Proof.
Let ~$\widetilde{\mathcal{F}}_{iso}$ be a uniformly bounded subset of $\mathcal{F}_{iso}$ that contains $h_{n}$, $h_{n}^{\#}$, and $h_{n,0}$ with probability tending to one (which exists by F1). Define $\mathcal{F}_{n,j}:=\{f\circ\nu_{n,j}:f\in\widetilde{\mathcal{F}}_{iso}\}$. Since $h_{n}$ and $h_{n,0}$ fall in $\mathcal{F}_{iso}$ with probability tending to one by F1, it holds that $\nu_{n,j}^{*}$ and $\nu_{n,0}$ fall in $\mathcal{F}_{n,j}$ with probability tending to one. By Condition F3, we may choose a constant $L^{\prime}<\infty$ such that $L_{n}\lesssim L$ with probability at least $1-\varepsilon$. Without loss of generality, we take $L_{n}=L^{\prime}$ and condition on this event, as it does not affect the $O_{p}$ statements in the theorem.
We first show that the excess risk implied by the loss is quadratic. By assumption F2, the functional $\nu\mapsto P_{0}m(\cdot,\nu)$ is a bounded linear functional. Moreover, $\ell$ is exactly the Riesz loss associated with this functional (Chernozhukov et al., 2022), and therefore $\nu_{0}:=\operatorname*{argmin}_{\nu\in\mathcal{H}}P_{0}\ell(\cdot,\nu)$ is the nonparametric Riesz representer for it. Using the Riesz representation property as in Chernozhukov et al. (2022), we can show that $\nu\mapsto P_{0}\ell(\cdot,\nu)=P_{0}\{\nu^{2}-2\nu\nu_{0}\}$. This risk is strongly convex and satisfies, for any uniformly bounded, convex function class $\mathcal{F}$, the inequality
$\|\nu-\nu_{0,\mathcal{F}}\|^{2}_{P_{0}}\lesssim P_{0}\{\ell(\cdot,\nu)-\ell(\cdot,\nu_{0,\mathcal{F}})\},$
where $\nu_{0,\mathcal{F}}:=\arg\min_{\nu\in\mathcal{F}}E_{0}[\ell(X,Y,\nu)]$.
We prove the convergence rates for the bootstrap estimators, as the rates for the non-bootstrap estimators follow along similar lines. To this end, for $j\in[J]$, we define the random rate $\delta_{n,j}^{\#}:=\|h_{n}^{\#}\circ\nu_{n,j}-h_{n,0}\circ\nu_{n,j}\|_{P_{0}}$ and let $(\delta_{n}^{\#})^{2}:=\frac{1}{J}\sum_{j=1}^{J}(\delta_{n,j}^{\#})^{2}$. By convexity of $\mathcal{F}_{iso}$, the quadratic excess risk bound of the previous paragraph, and $h_{n,0}$ being a population risk minimizer, we have that
$(\delta_{n}^{\#})^{2}\lesssim\overline{P}_{0}\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\overline{P}_{0}\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond}).$
Furthermore, since $h_{n}^{\#}\circ\nu_{n}$ is the bootstrap empirical risk minimizer, it holds that
$\overline{P}_{n}^{\#}\left\{\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})\right\}\leq 0.$
Combining the above two displays, we obtain that:
$\overline{P}_{0}\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\overline{P}_{0}\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})$ $=(\overline{P}_{0}-\overline{P}_{n}^{\#})\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})+\overline{P}_{n}^{\#}\left\{\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})\right\}$
$\quad-(\overline{P}_{0}-\overline{P}_{n}^{\#})\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})$
$\leq(\overline{P}_{0}-\overline{P}_{n}^{\#})\left\{\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})\right\},$
and, therefore,
$(\delta_{n}^{\#})^{2}$ $\lesssim(\overline{P}_{0}-\overline{P}_{n}^{\#})\left\{\ell(\cdot,h_{n}^{\#}\circ\nu_{n,\diamond})-\ell(\cdot,h_{n,0}\circ\nu_{n,\diamond})\right\}$
$\leq\frac{1}{J}\sum_{j=1}^{J}\left[\gamma_{n,j}(\delta_{n,j}^{\#})+\gamma_{n,j}^{\#}(\delta_{n,j}^{\#})\right].$
where we define:
$\gamma_{n,j}(\delta)$ $:=\sup_{\nu_{1},\nu_{2}\in\mathcal{F}_{n,j}:\|\nu_{1}-\nu_{2}\|\leq\delta}\left|({P}_{0}-{P}_{n,j})\left\{\ell(\cdot,\nu_{1})-\ell(\cdot,\nu_{2})\right\}\right|;$
$\gamma_{n,j}^{\#}(\delta)$ $:=\sup_{\nu_{1},\nu_{2}\in\mathcal{F}_{n,j}:\|\nu_{1}-\nu_{2}\|\leq\delta}\left|(P_{n,j}-P_{n,j}^{\#})\left\{\ell(\cdot,\nu_{1})-\ell(\cdot,\nu_{2})\right\}\right|.$
By Assumption F3 and the boundedness of $\mathcal{F}_{n,j}$, the function class $\{\ell(\cdot,\nu_{1})-\ell(\cdot,\nu_{2}):\nu_{1},\nu_{2}\in\mathcal{F}_{n,j}\}$ has controlled uniform entropy integral, where this class is fixed conditional on the training data of the nuisances. Specifically, by the preservation properties of the uniform entropy integral (van der Vaart and Wellner, 1996), the function class $\mathcal{F}_{n,j}\times\mathcal{F}_{n,j}$ has a uniform entropy integral bounded, up to a constant, by $\mathcal{J}(\delta,\mathcal{F}_{\nu_{n,j},\mathrm{iso}})+\mathcal{J}(\delta,\mathcal{F}_{\nu_{n,j},\mathrm{iso}})$ for each $j\in[J]$. This quantity is itself, up to a constant, bounded above by $\mathcal{J}(\delta,\mathcal{F}_{\mathrm{iso}})$ (see, e.g., the proof of Lemma 7). Thus, by applying Lemma 6 and following an argument identical to the proof of Lemma 8 with $t_{n}(\delta):=\sqrt{\delta}$, we obtain, for any fixed $\delta>0$, conditional on $\mathcal{E}_{n,j}:=\mathcal{D}_{n}\backslash\mathcal{C}_{n,j}$:
$E_{0}^{n}[\gamma_{n,j}(\delta)\mid\mathcal{E}_{n,j}]$ $\lesssim n^{-\frac{1}{2}}\sqrt{\delta}\left(1+\frac{\sqrt{\delta}}{\sqrt{n}\delta^{2}}\right);$
$E_{0}^{n}[\gamma_{n,j}^{\#}(\delta)\mid\mathcal{E}_{n,j}]$ $\lesssim n^{-\frac{1}{2}}\sqrt{\delta}\left(1+\frac{\sqrt{\delta}}{\sqrt{n}\delta^{2}}\right),$
where the bound on the right-hand side is deterministic.
We are now ready to obtain a convergence rate bound for $\delta_{n}^{\#}=\|h_{n}^{\#}\circ\nu_{n,\diamond}-h_{n,0}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}$. To do so, we use a peeling argument similar to the proof of Theorem 3.2.5 in van der Vaart and Wellner (1996). We will also use the following fact: if, for given vectors $x,y,u,v\in\mathbb{R}^{J}$, both $\sum_{j=1}^{J}x_{j}\leq\sum_{j=1}^{J}y_{j}$ and $\sum_{j=1}^{J}u_{j}\leq\sum_{j=1}^{J}v_{j}$ hold then there exists some $j^{*}\in[J]$ such that both $x_{j^{*}}\leq y_{j^{*}}$ and $u_{j^{*}}\leq v_{j^{*}}$. To verify this claim, suppose the contrapositive that $x_{j}>y_{j}$ and $u_{j}>v_{j}$ for all $j\in[J]$. Then, we must have $\sum_{j=1}^{J}x_{j}>\sum_{j=1}^{J}y_{j}$ and $\sum_{j=1}^{J}u_{j}<\sum_{j=1}^{J}v_{j}$, which is a contradiction. The claim follows. Proceeding with the proof and letting $\varepsilon_{n}\in\mathbb{R}$ be a deterministic rate to be set later, note
$P(2^{k+1}\varepsilon_{n}\geq\delta_{n}^{\#}\geq 2^{k}\varepsilon_{n})$ $=P\left(2^{k+1}\varepsilon_{n}\geq\delta_{n}^{\#}\geq 2^{k}\varepsilon_{n},\sum_{j\in[J]}(\delta_{n,j}^{\#})^{2}\leq\sum_{j\in[J]}\left[\gamma_{n,j}(\delta_{n,j}^{\#})+\gamma_{n,j}^{\#}(\delta_{n,j}^{\#})\right]\right)$
$\leq P\left(2^{k+1}\varepsilon_{n}\geq\delta_{n,j_{n}}\geq 2^{k}\varepsilon_{n},\delta_{n,j_{n}}^{2}\leq\left[\theta_{n,j_{n}}(\delta_{n,j_{n}})+\theta_{n,j_{n}}^{\#}(\delta_{n,j_{n}})\right]\text{ for some }j_{n}\in[J]\right)$
$\leq\sum_{j\in[J]}P\left(2^{k+1}\varepsilon_{n}\geq\delta_{n,j}^{\#}\geq 2^{k}\varepsilon_{n},(\delta_{n,j}^{\#})^{2}\leq\left[\gamma_{n,j}(\delta_{n,j}^{\#})+\gamma_{n,j}^{\#}(\delta_{n,j}^{\#})\right]\right)$
$\leq\sum_{j\in[J]}P\left(2^{k+1}\varepsilon_{n}\geq\delta_{n,j}^{\#}\geq 2^{k}\varepsilon_{n},2^{2k}\varepsilon_{n}^{2}\leq\left[\gamma_{n,j}(2^{k+1}\varepsilon_{n})+\gamma_{n,j}^{\#}(2^{k+1}\varepsilon_{n})\right]\right)$
$\leq\sum_{j\in[J]}E_{0}^{n}\left[P\left(2^{2k}\varepsilon_{n}^{2}\leq\left[\gamma_{n,j}(2^{k+1}\varepsilon_{n})+\gamma_{n,j}^{\#}(2^{k+1}\varepsilon_{n})\right]\right)|\mathcal{E}_{n,j}\right]$
$\lesssim E_{0}^{n}\left[\frac{E_{0}^{n}[\widetilde{\gamma}_{n,j}^{\#}(2^{k+1}\varepsilon_{n})|\mathcal{E}_{n,j}]+E_{0}^{n}[\widetilde{\gamma}_{n,j}^{\#}(2^{k+1}\varepsilon_{n})|\mathcal{E}_{n,j}]}{2^{2k}\varepsilon_{n}^{2}}\right]$
$\leq C\frac{n^{-\frac{1}{2}}\sqrt{2^{k+1}\varepsilon_{n}}(1+\frac{\sqrt{2^{k+1}\varepsilon_{n}}}{\sqrt{n}\varepsilon_{n}^{2}})}{2^{2k}\varepsilon_{n}^{2}},$
for some fixed constant $C>0$. Taking $\varepsilon_{n}:=n^{-1/3}$ in the above, it is straightforward to show, for some other fixed constant $C^{\prime}>0$, that
$P(2^{k+1}n^{-1/3}\geq\delta_{n}^{\#}\geq 2^{k}n^{-1/3})$ $\lesssim\frac{n^{-\frac{1}{2}}\sqrt{2^{k+1}\varepsilon_{n}}(1+\frac{\sqrt{2^{k+1}\varepsilon_{n}}}{\sqrt{n}\varepsilon_{n}^{2}})}{2^{2k}\varepsilon_{n}^{2}}\leq C^{\prime}\frac{1}{2^{k}}.$
Therefore, for any constant $K>0$, we have
$P(\delta_{n}^{\#}\geq 2^{K}n^{-1/3})$ $=\sum_{k=K}^{\infty}P(2^{k+1}n^{-1/3}\geq\delta_{n}^{\#}\geq 2^{k}n^{-1/3})\leq C^{\prime}\sum_{k=K}^{\infty}\frac{1}{2^{k}},$
where $\sum_{k=K}^{\infty}\frac{1}{2^{k}}$ tends to zero as $K\rightarrow\infty$. Hence, for all $\varepsilon>0$, we can find a constant $K>0$ such that $P(\delta_{n}^{\#}\geq 2^{K}n^{-1/3})\leq\varepsilon$. By definition, it follows that $\delta_{n}^{\#}=O_{p}(n^{-1/3})$ and, therefore, $\|h_{n}^{\#}\circ\nu_{n,\diamond}-h_{n,0}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}^{2}=o_{p}(n^{-2/3})$. An analogous argument shows $\|h_{n}\circ\nu_{n,\diamond}-h_{n,0}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}^{2}=O_{p}(n^{-2/3})$. Next, the triangle inequality implies that $\|h_{n}^{\#}\circ\nu_{n,\diamond}-h_{n}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}^{2}=O_{p}(n^{-2/3})$. Therefore, for any function $\overline{\nu}_{0}$, it holds that
$\|h_{n}^{\#}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}$ $\leq\|h_{n}^{\#}\circ\nu_{n,\diamond}-h_{n}\circ\nu_{n,\diamond}\|_{\overline{P}_{0}}+\|h_{n}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}$
$\leq O_{p}(n^{-1/3})+\|h_{n}\circ\nu_{n,\diamond}-\overline{\nu}_{0}\|_{\overline{P}_{0}}.$
The following theorem is a direct application of Theorem 6. In the following, we define the population minimizers corresponding to the calibrators $f_{n}$ and $g_{n}$ as:
$f_{n,0}$ $:=\operatorname*{argmin}_{f\in\mathcal{F}_{iso}}\sum_{j=1}^{J}\|\mu_{0}-f\circ\mu_{n,j}\|_{P_{0}}^{2};$
$g_{n,0}$ $:=\operatorname*{argmin}_{g\in\mathcal{F}_{iso}}\sum_{j=1}^{J}\left\{\|g\circ\alpha_{n,j}\|_{P_{0}}^{2}-2\psi_{0}(g\circ\alpha_{n,j})\right\}.$
Theorem 7.
Suppose that Conditions C5 holds, that Condition C1 holds with $\psi_{n,j}$ and $P_{n,j}$ for each $j\in[J]$, and that Condition D3 also holds. Then, $\|\alpha_{n,\diamond}^{*}-g_{n,0}\circ\alpha_{n}\|_{\overline{P}_{0}}=O_{p}(n^{-1/3})$ and $\|\mu_{n,\diamond}^{*}-f_{n,0}\circ\mu_{n}\|_{\overline{P}_{0}}=O_{p}(n^{-1/3})$. Hence,
$\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}$ $\leq\min_{g\in\mathcal{F}_{iso}}\|g\circ\alpha_{n,\diamond}-\alpha_{0}\|+O_{p}(n^{-1/3});$
$\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}$ $\leq\min_{f\in\mathcal{F}_{iso}}\|f\circ\mu_{n,\diamond}-\mu_{0}\|+O_{p}(n^{-1/3}).$
Additionally, assume D1 holds. Then, the bootstrap calibrated nuisance estimators satisfy, for any functions $\overline{\alpha}_{0},\overline{\mu}_{0}\in L^{2}(P_{0})$ and as $n\rightarrow\infty$, the following root mean square error bounds:
$\|\alpha_{n}^{*\#}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}$ $\leq\|\alpha_{n,\diamond}^{*}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3});$
$\|\mu_{n}^{*\#}-\overline{\mu}_{0}\|_{\overline{P}_{0}}$ $\leq\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3}).$
Proof.
The proof follows from two direct applications of Theorem 6. To obtain the claimed rates for $\alpha_{n,\diamond}^{*}$, we take
$\ell(x,y,\nu):=\nu^{2}(a,w)-2m(z,\nu),\quad h_{n}:=g_{n},\quad h_{n,0}:=g_{n,0},\quad h_{n}^{\#}:=g_{n}^{\#},\quad\nu_{0}:=\alpha_{0},\quad\nu_{n,j}:=\alpha_{n,j}.$
Similarly, to obtain the rates for $\mu_{n,\diamond}^{*}$, we take
$\ell(x,y,\nu):=\nu^{2}(a,w)-2y\nu(a,w),\quad h_{n}:=f_{n},\quad h_{n,0}:=f_{n,0},\quad h_{n}^{\#}:=f_{n}^{\#},\quad\nu_{0}:=\mu_{0},\quad\nu_{n,j}:=\mu_{n,j}.$
In both cases, Condition F1 holds by Conditions C5 and D1. For the squared error loss, Condition F2 holds trivially, while for the Riesz loss, it holds by Condition A1. For the squared error loss, Condition F3 holds by the pointwise Lipschitz continuity of the map $m\mapsto m^{2}-2ym$ and the uniform boundedness of the outcome $Y$. For the Riesz loss, Condition F3 holds by D3.
G.5 Key Lemmas for analysis of calibrated one-step debiased estimator
Let $\alpha_{n,\diamond}^{*}:=\{\alpha_{n,j}^{*}:j\in[J]\}$ and $\mu_{n,\diamond}^{*}:=\{\mu_{n,j}^{*}:j\in[J]\}$ be the isotonic calibrated nuisance estimators obtained via Algorithm 1. Let $\overline{\alpha}_{0}$ and $\overline{\mu}_{0}$ be their respective potentially misspecified limits, as in C2.
The following lemma establishes that the isotonic calibrated nuisance estimators are empirically calibrated for their respective loss functions.
Lemma 10 (Empirical calibration via isotonic calibration).
Assume C1. Then, for all $h_{1},h_{2}:\mathbb{R}\rightarrow\mathbb{R}$, it holds that
$\frac{1}{n}\sum_{i=1}^{n}(h_{1}\circ\mu_{n,j(i)}^{*})(A_{i},W_{i})\left\{Y_{i}-\mu_{n,j(i)}^{*}(A_{i},W_{i})\right\}$ $=0;$
$\frac{1}{n}\sum_{i=1}^{n}\left\{(h_{2}\circ\alpha_{n,j(i)}^{*})(A_{i},W_{i})\alpha_{n,j(i)}^{*}(A_{i},W_{i})-\psi_{n,j(i)}(h_{2}\circ\alpha_{n,j(i)}^{*})\right\}$ $=0.$
Proof.
The proof of this lemma follows from a simplification of the proof of Lemma 4 in van der Laan et al. (2023b). We give the proof for $\alpha_{n,\diamond}^{*}$ as the proof for $\mu_{n,\diamond}^{*}$ follows along similar lines. Recall that $\alpha_{n,\diamond}^{*}=g_{n}\circ\alpha_{n,\diamond}$ for the empirical risk minimizer $g_{n}\in\mathcal{F}_{iso}$ defined in Algorithm 1. Also recall, following Groeneboom and Lopuhaa (1993), that we took $g_{n}$ to be the unique càdlàg piece-wise constant solution of the isotonic regression problem that can only take jumps at observed values $\{\alpha_{n,j(i)}(A_{i},W_{i}):i\in[n]\}$.
By Condition C1, the empirical risk used to define the isotonic solution $g_{n}$ is strongly convex and continuous on $\mathcal{F}_{\text{iso}}$ with respect to the empirical $L^{2}$ norm $\left\{\sum_{i=1}^{n}\left[(g\circ\alpha_{n,j(i)})(A_{i},W_{i})\right]^{2}\right\}^{1/2}$. By Theorem 5.5 of Alexanderian (2019), a strongly convex and continuous functional defined on a closed and bounded convex subset of a Hilbert space admits a unique minimizing solution. Consequently, the solution $g_{n}$ is uniquely defined on the set $\{\alpha_{n,j(i)}(A_{i},W_{i}):1\leq i\leq n\}$, and, by assumption, it is the uniquely defined piecewise constant solution with jumps occurring only at these values.
It holds that $g_{n}$ is a piece-wise constant. We claim, for any function $h:\mathbb{R}\rightarrow\mathbb{R}$ and all $\varepsilon\in\mathbb{R}$ sufficiently close to zero, that the perturbed function $g_{n}+\varepsilon h\circ g_{n}\in\mathcal{F}_{iso}$. To see this, observe that $g_{n}$ is constant in some interval of its domain if and only if the same holds for $h\circ g_{n}$. Hence, the function $g_{n}+\varepsilon h\circ g_{n}$ is piece-wise constant and jumps at the same points as $g_{n}$. Moreover, by taking $\varepsilon$ sufficiently small, we can ensure that the maximal jump change $\varepsilon\sup_{i\in[n]}|(h\circ g_{n})(\mu_{n,j(i)}(A_{i},W_{i})|$ of $h\circ g_{n}$ is strictly smaller than the maximal jump change of $g_{n}$. Therefore, for this choice of $\varepsilon$, the perturbed function $g_{n}+\varepsilon h\circ g_{n}\in\mathcal{F}_{iso}$ must be nondecreasing and, therefore, an element of $\mathcal{F}_{iso}$. The claim now follows.
Proceeding with the proof of the Lemma, we denote $\xi_{n}(\varepsilon):=g_{n}+\varepsilon(h\circ g_{n})$ and $\xi_{n}(\varepsilon,\alpha_{n,j}):=g_{n}\circ\alpha_{n,j}+\varepsilon(h\circ g_{n})(\alpha_{n,j})$ and, in a slight abuse of notation, let
$R_{n}(\xi_{n}(\varepsilon)):=\sum_{i=1}^{n}\left\{\xi_{n}(\varepsilon,\alpha_{n,j(i)})^{2}(A_{i},W_{i})-2\psi_{n,j(i)}(\xi_{n}(\varepsilon,\alpha_{n,j}))\right\}.$
Now, because $g_{n}$ minimises $g\mapsto R_{n}(g)$ over $\mathcal{F}_{\mathrm{iso}}$ and $g_{n}+\varepsilon h\circ g_{n}\in\mathcal{F}_{\mathrm{iso}}$ for $\varepsilon$ sufficiently close to 00, $g_{n}$ also minimizes $g\mapsto R_{n}(g)$ over the one-dimensional submodel $\varepsilon\mapsto g_{n}+\varepsilon h\circ g_{n}$ in a neighborhood of $\varepsilon=0$. Thus, the first-order optimality condition along this submodel implies that
$\frac{d}{d\varepsilon}[R_{n}(\xi_{n}(\varepsilon,\alpha_{n,j(i)}))-R_{n}(g_{n})]\Big{|}_{\varepsilon=0}=0.$
Computing the derivative, using linearity of the functional $\psi_{n,\diamond}$ and that $h\circ g_{n}\circ\alpha_{n,\diamond}=h\circ\alpha_{n,\diamond}^{*}$, this further implies:
$\frac{1}{n}\sum_{i=1}^{n}\left\{(h\circ\alpha_{n,j(i)}^{*})(A_{i},W_{i})\alpha_{n,j(i)}^{*}(A_{i},W_{i})-\psi_{n,j(i)}(h\circ\alpha_{n,j(i)}^{*})\right\}=0.$
The result for $\alpha_{n,\diamond}^{*}$ now follows, noting that $h$ was an arbitrary function.
The next lemma establishes that empirical calibration of the nuisance estimators in the sense of Lemma 10 suffices to linearize the cross-product remainder term. Following the notation of Lemma 5, we denote $s_{n,j}^{*}:=\Pi_{\mu_{n,j}^{*}}\{\alpha_{0}-\alpha_{n,j}^{*}\}$ and $r_{n,j}^{*}:=\Pi_{\alpha_{n,j}^{*}}\{\mu_{0}-\mu_{n,j}^{*}\}$. Denote $A_{n,j}^{*}:z\mapsto s_{n,j}^{*}(a,w)\{y-\mu_{n,j}^{*}(a,w)\}$ and define $B_{n,j}^{*}:=\psi_{0}(r_{0})+\widetilde{\phi}_{r_{n,j}^{*}}-r_{n,j}^{*}\alpha_{n,j}^{*}$. We recall that $s_{0}:=\Pi_{\mu_{0}}\{\alpha_{0}-\overline{\alpha}_{0}\}$ and $r_{0}:=\Pi_{\alpha_{0}}\{\mu_{0}-\overline{\mu}_{0}\}$.
Lemma 11 (Linearization of remainder for calibrated one-step debiased estimator).
We have the outcome-favorable decomposition:
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})A_{n,\diamond}^{*}+O_{p}\left(\|(\Pi_{(\mu_{n,\diamond}^{*},\mu_{0})}-\Pi_{\mu_{n,\diamond}^{*}})(\alpha_{n,\diamond}^{*}-\alpha_{0})\|_{\overline{P}_{0}}\|\mu_{0}-\mu_{n,\diamond}^{*}\|_{\overline{P}_{0}}\right),$
Similarly, we have the Riesz-favorable decomposition:
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})B_{n,\diamond}^{*}+\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(r_{n,j}^{*})-\psi_{0}(r_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}^{*}}\right\}$
$\quad+O_{p}(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}\|(\Pi_{(\alpha_{n,\diamond}^{*},\alpha_{0})}-\Pi_{\alpha_{n,\diamond}^{*}})(\mu_{0}-\mu_{n,\diamond}^{*})\|_{\overline{P}_{0}},$
Proof.
The result follows from combining Lemma 5 with Lemma 10, noting that empirical calibration ensures that $\overline{P}_{n}A_{n,\diamond}^{*}=0$ and $\left\{\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(r_{n,j}^{*})-\overline{P}_{0}r_{n,\diamond}^{*}\alpha_{n,\diamond}^{*}\right\}=0$. In particular, the first score equation in Lemma 10 implies that
$\overline{P}_{n}A_{n,\diamond}^{*}=\frac{1}{n}\sum_{i=1}^{n}s_{n,j(i)}^{*}(A_{i},W_{i})\{\mu_{n,j(i)}^{*}(A_{i},W_{i})-Y_{i}\}=0,$
since each $s_{n,j}^{*}$ is a function of $\mu_{n,j}^{*}$. Similarly, the second score equation in Lemma 10 implies that
$\frac{1}{n}\sum_{i=1}^{n}r_{n,j(i)}^{*}(A_{i},X_{i})\alpha_{n,j(i)}^{*}(A_{i},X_{i})-\frac{1}{J}\sum_{j=1}^{J}\psi_{n,j}(r_{n,j}^{*})=0,$
since each $r_{n,j}^{*}$ is a function of $\alpha_{n,j}^{*}$.
The next lemmas establishes the negligibility of certain empirical processes.
Lemma 12.
Consider the setup of section 5.2, where $\psi_{0}(h)=E_{0}[m(Z,h)]$ for a functional $m$. Assume A3 , C2 -C3, D1-D4. Then, it holds that $\frac{1}{J}\sum_{j=1}^{J}(P_{n,j}-P_{0})\{m(\cdot,\mu_{n,j}^{*})-m(\cdot,\overline{\mu}_{0})\}=o_{p}(n^{-1/2})$, $1\{\overline{\mu}_{0}=\mu_{0}\}\,\frac{1}{J}\sum_{j=1}^{J}(P_{n,j}-P_{0})\{m(\cdot,r_{n,j}^{*})-m(\cdot,r_{0})\}=o_{p}(n^{-1/2})$, $\frac{1}{J}\sum_{j=1}^{J}(P_{n,j}^{\#}-P_{n,j})\{m(\cdot,\mu_{n,j}^{\#})-m(\cdot,\overline{\mu}_{0})\}=o_{p}(n^{-1/2})$, and $1\{\overline{\mu}_{0}=\mu_{0}\}\,\frac{1}{J}\sum_{j=1}^{J}(P_{n,j}^{\#}-P_{n,j})\{m(\cdot,r_{n,j}^{\#})-m(\cdot,r_{0})\}=o_{p}(n^{-1/2})$.
Proof.
By Conditions C5 and D1, we have that $\mu_{n,j}^{*},\mu_{n,j}^{\#}\in\mathcal{R}_{n,j}=\{f\circ\mu_{n,j}:f\in\mathcal{F}_{n}\cup\mathcal{F}_{\mathrm{iso}}\}$ (with probability tending to one) for each $j\in[J]$. For each $j\in[J]$, we aim to apply Lemma 6 conditional on the $j$th training fold with $\mathcal{F}:=\mathcal{F}_{n}\cup\mathcal{F}_{\mathrm{iso}}$, $\nu:=\mu_{n,j}$, and $m:=m$. By the proof of Lemma 7, we have
$\mathcal{J}(\delta,\mathcal{F}_{n}\cup\mathcal{F}_{\mathrm{iso}})\lesssim\sqrt{\delta}\left(1+\sqrt{k(n)\delta\log(1/\delta)}\right).$
Hence, by Lemma 6, it follows that
$\mathcal{J}(\delta,\mathcal{F}_{m,\mu_{n,j}})\lesssim\mathcal{J}(\delta,\mathcal{F}_{n}\cup\mathcal{F}_{\mathrm{iso}})\lesssim\sqrt{\delta}\left(1+\sqrt{k(n)\delta\log(1/\delta)}\right),$
where $\mathcal{F}_{m,\mu_{n,j}}:=\{m(\cdot,\theta\circ\mu_{n,j}):\theta\in\mathcal{F}_{n}\cup\mathcal{F}_{\mathrm{iso}}\}$.
The functions $m(\cdot,\mu_{n,j}^{*})-m(\cdot,\overline{\mu}_{0})$, $m(\cdot,\mu_{n,j}^{\#})-m(\cdot,\overline{\mu}_{0})$, $m(\cdot,r_{n,j}^{*})-m(\cdot,\overline{\mu}_{0})$, and $m(\cdot,r_{n,j}^{\#})-m(\cdot,\overline{\mu}_{0})$ fall in $\mathcal{F}_{m,\mu_{n,j}}-\mathcal{F}_{m,\mu_{n,j}}$ with probability tending to one, where the latter function class has uniform entropy integral also upper bounded by $\sqrt{\delta}\left(1+\sqrt{k(n)\delta\log(1/\delta)}\right)$.
For this class of linear functionals, the function $\phi_{0,\mu}$ in Condition A3 is given by $m(\cdot,\mu)$. Hence, by the Lipschitz continuity condition in A3, together with Conditions C2, C3, and D4, we have
$\|m(\cdot,\mu_{n,j}^{*})-m(\cdot,\overline{\mu}_{0})\|=o_{p}(1)\quad\text{and}\quad 1\{\overline{\mu}_{0}=\mu_{0}\}\|m(\cdot,r_{n,j}^{*})-m(\cdot,r_{0})\|=o_{p}(1).$
$\|m(\cdot,\mu_{n,j}^{\#})-m(\cdot,\overline{\mu}_{0})\|=o_{p}(1)\quad\text{and}\quad 1\{\overline{\mu}_{0}=\mu_{0}\}\|m(\cdot,r_{n,j}^{\#})-m(\cdot,r_{0})\|=o_{p}(1).$
Moreover, by C6 and D2, $k(n)=o(n^{-1/2}/\log n)$. Hence, applying the local maximal inequalities in 8 and arguing as in the proof of Lemma 9, we have, for each $j\in[J]$, $(P_{n,j}-P_{0})\{m(\cdot,\mu_{n,j}^{*})-m(\cdot,\overline{\mu}_{0})\}=o_{p}(n^{-1/2})$, $1\{\overline{\mu}_{0}=\mu_{0}\}(P_{n,j}-P_{0})\{m(\cdot,r_{n,j}^{*})-m(\cdot,r_{0})\}=o_{p}(n^{-1/2})$, $(P_{n,j}^{\#}-P_{n,j})\{m(\cdot,\mu_{n,j}^{\#})-m(\cdot,\overline{\mu}_{0})\}=o_{p}(n^{-1/2})$ , and $1\{\overline{\mu}_{0}=\mu_{0}\}(P_{n,j}^{\#}-P_{n,j})\{m(\cdot,r_{n,j}^{\#})-m(\cdot,r_{0})\}=o_{p}(n^{-1/2})$. The result then follows.
Lemma 13 (Negligible empirical process remainders).
Assume A3-C6 hold. Then, $(\overline{P}_{n}-P_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}-\widetilde{\phi}_{0,\overline{\mu}_{0}}\right\}+(\overline{P}_{n}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=o_{p}(n^{-\frac{1}{2}})$. Moreover, under the setup of Theorem 4 and if D1-D4 also hold, $(\overline{P}_{n}^{\#}-P_{0})\left\{m(\cdot,\mu_{n,\diamond}^{\#*})-m(\cdot,\overline{\mu}_{0})\right\}+(\overline{P}_{n}^{\#}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=o_{p}(n^{-\frac{1}{2}})$.
Proof.
Condition C2 gives that $\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$ and $\|\alpha_{n,\diamond}^{*}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$. Note $(\overline{P}_{n}-P_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}-\widetilde{\phi}_{0,\overline{\mu}_{0}}\right\}=(\overline{P}_{n}-P_{0})\left\{{\phi}_{0,\mu_{n,\diamond}^{*}}-{\phi}_{0,\overline{\mu}_{0}}\right\}$, where $\mu\mapsto\phi_{0,\mu}$ is defined in A3. Hence, $(\overline{P}_{n}-P_{0})\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}-\widetilde{\phi}_{0,\overline{\mu}_{0}}\}=o_{p}(n^{-1/2})$ is a direct consequence of C7a. Furthermore, by C5,
$\|D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\|_{\overline{P}_{0}}=\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}+\|\alpha_{n,\diamond}^{*}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}=o_{p}(1).$
The first part of the lemma will follow from a direct application of Lemma 9. By C5, $(\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*})\mapsto D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}$ is a bounded Lipschitz transformation of $(\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*})$. With probability one, by definition and C5, $\mu_{n,j}^{*}$ and $\alpha_{n,j}^{*}$ respectively fall in $\mathcal{S}_{n,j}$ and $\mathcal{R}_{n,j}$ for each $j\in[J]$. Hence, with probability one, $D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}$ falls in a bounded Lipschitz transformation of $\mathcal{S}_{n,j}\times\mathcal{R}_{n,j}$. Hence, by Lemma 9, $(\overline{P}_{n}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=o_{p}(n^{-\frac{1}{2}})$.
To prove the second part of this lemma, note
$(\overline{P}_{n}^{\#}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=(\overline{P}_{n}^{\#}-\overline{P}_{n})\{D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}+(\overline{P}_{n}-\overline{P}_{0})\{D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}.$
By Theorem 7 and condition C2, we have $\|\alpha_{n}^{*\#}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}\leq\|\alpha_{n,\diamond}^{*}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3})$ and $\|\mu_{n}^{*\#}-\overline{\mu}_{0}\|_{\overline{P}_{0}}\leq\|\mu_{n,\diamond}^{*}-\overline{\mu}_{0}\|_{\overline{P}_{0}}+O_{p}(n^{-1/3})$. Hence, by C2, $\|\alpha_{n}^{*\#}-\overline{\alpha}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$ and $\|\mu_{n}^{*\#}-\overline{\mu}_{0}\|_{\overline{P}_{0}}=o_{p}(1)$. Hence, by C5, we also have $\|D_{\mu_{n,\diamond}^{\#*},\alpha_{n,\diamond}^{\#*}}-D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\|_{\overline{P}_{0}}=o_{p}(1)$. Now, the second term on the right-hand side of the above display is $o_{p}(n^{-\frac{1}{2}})$ by the same argument as before. The first term on the right-hand side is also $o_{p}(n^{-\frac{1}{2}})$ by a similar argument, using the second part of Lemma 9. Finally, the expression $(\overline{P}_{n}^{\#}-P_{0})\{m(\cdot,\mu_{n,\diamond}^{\#*})-m(\cdot,\overline{\mu}_{0})\}=o_{p}(n^{-1/2})$ follows directly from Lemma 12
For each $j\in[J]$, recall that $A_{n,j}^{*}:z\mapsto s_{n,j}^{*}(a,w)\{y-\mu_{n,j}^{*}(a,w)\}$ and $B_{n,j}^{*}:=\psi_{0}(r_{0})+\widetilde{\phi}_{r_{n,j}^{*}}-r_{n,j}^{*}\alpha_{n,j}^{*}$.
Lemma 14 (Negligible cross-product empirical process remainders).
Under A3-C6, $1(\overline{\mu}_{0}=\mu_{0})(\overline{P}_{n}-P_{0})\{A_{n,\diamond}^{*}-A_{0}\}=o_{p}(n^{-\frac{1}{2}})$ and $1(\overline{\alpha}_{0}=\alpha_{0})(\overline{P}_{n}-P_{0})\{B_{n,\diamond}^{*}-B_{0}\}=o_{p}(n^{-\frac{1}{2}})$. Moreover, under the setup of Theorem 4 and if D1-D4 also hold, $1(\overline{\mu}_{0}=\mu_{0})(\overline{P}_{n}^{\#}-\overline{P}_{0})\{A_{n,\diamond}^{*\#}-A_{0}\}=o_{p}(n^{-\frac{1}{2}})$, $1(\overline{\alpha}_{0}=\alpha_{0})(\overline{P}_{n}^{\#}-\overline{P}_{0})\{B_{n,\diamond}^{*\#}-B_{0}\}=o_{p}(n^{-\frac{1}{2}})$.
Proof.
Condition C3 and C2 implies that $1(\overline{\alpha}_{0}=\alpha_{0})\|r_{n,\diamond}^{*}-r_{0}\|_{\overline{P}_{0}}=\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}=o_{p}(1)$ and $1(\overline{\mu}_{0}=\mu_{0})\|s_{n,\diamond}^{*}-s_{0}\|_{\overline{P}_{0}}=\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}=o_{p}(1)$. Hence, by C5, when $1(\overline{\mu}_{0}=\mu_{0})$, we have:
$\|A_{n,\diamond}^{*}-A_{0}\|_{\overline{P}_{0}}=\|s_{n,\diamond}^{*}-s_{0}\|_{\overline{P}_{0}}+\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}=o_{p}(1).$
The outcome $Y$ is assumed bounded almost surely. Hence, by C5, $(\mu_{n,\diamond}^{*},s_{n,\diamond}^{*})\mapsto s_{n,\diamond}^{*}(\mathcal{I}_{Y}-\mu_{n,\diamond}^{*})-s_{0}(\mathcal{I}_{Y}-\mu_{0})$ is a Lipschitz transformation of $\mu_{n,\diamond}^{*}$ and $s_{n,\diamond}^{*}$, where $s_{n,\diamond}^{*}(\mathcal{I}_{Y}-\mu_{n,\diamond}^{*})-s_{0}(\mathcal{I}_{Y}-\mu_{0})$ is bounded with probability tending to one. For each $j\in[J]$, with probability tending to one, by definition, $\mu_{n,j}^{*}$ falls in $\mathcal{S}_{n,j}$ and similarly, by C6, $s_{n,j}^{*}$ falls in $\mathcal{S}_{n,j}$. Hence, with probability tending to one, $A_{n,\diamond}^{*}-A_{0}=s_{n,\diamond}^{*}(\mathcal{I}_{Y}-\mu_{n,\diamond}^{*})-s_{0}(\mathcal{I}_{Y}-\mu_{0})$ falls in a bounded Lipschitz transformation of $\mathcal{S}_{n,j}\times\mathcal{S}_{n,j}$. Hence, by Lemma 9, we conclude
$1(\overline{\mu}_{0}=\mu_{0})(\overline{P}_{n}-\overline{P}_{0})\{A_{n,\diamond}^{*}-A_{0}\}=o_{p}(n^{-\frac{1}{2}}).$
We recall that $(\overline{P}_{n}-\overline{P}_{0})\{{\phi}_{r_{n,j}^{*}}-{\phi}_{r_{0}}\}=o_{p}(n^{-1/2})$ by C7b. Next, note that
$(\overline{P}_{n}-\overline{P}_{0})\{B_{n,\diamond}^{*}-B_{0}\}$ $=(\overline{P}_{n}-\overline{P}_{0})\{\widetilde{\phi}_{r_{n,j}^{*}}-\widetilde{\phi}_{r_{0}}+r_{0}\alpha_{0}-r_{n,j}^{*}\alpha_{n,j}^{*}\}$
$=(\overline{P}_{n}-\overline{P}_{0})\{\phi_{r_{n,j}^{*}}-\phi_{r_{0}}+r_{0}\alpha_{0}-r_{n,j}^{*}\alpha_{n,j}^{*}\}$
$=(\overline{P}_{n}-\overline{P}_{0})\{r_{0}\alpha_{0}-r_{n,j}^{*}\alpha_{n,j}^{*}\}+o_{p}(n^{-1/2}).$
By C5, when $1(\overline{\alpha}_{0}=\alpha_{0})$, we have:
$\|r_{0}\alpha_{0}-r_{n,\diamond}^{*}\alpha_{n,\diamond}^{*}\|_{\overline{P}_{0}}=\|r_{n,\diamond}^{*}-r_{0}\|_{\overline{P}_{0}}+\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}=o_{p}(1).$
By C5, $(\alpha_{n,\diamond}^{*},r_{n,\diamond}^{*})\mapsto r_{0}\alpha_{0}-r_{n,\diamond}^{*}\alpha_{n,\diamond}^{*}$ is a Lipschitz transformation of $\mu_{n,\diamond}^{*}$ and $s_{n,\diamond}^{*}$, where $r_{0}\alpha_{0}-r_{n,\diamond}^{*}\alpha_{n,\diamond}^{*}$ is uniformly bounded with probability tending to one. For each $j\in[J]$, with probability tending to one, by definition, $\alpha_{n,j}^{*}$ falls in $\mathcal{R}_{n,j}$ and similarly, by C6, $r_{n,j}^{*}$ falls in $\mathcal{R}_{n,j}$. Hence, with probability one, $r_{0}\alpha_{0}-r_{n,j}^{*}\alpha_{n,j}^{*}$ falls in a bounded Lipschitz transformation of $\mathcal{R}_{n,j}\times\mathcal{R}_{n,j}$. Hence, by Lemma 9, we conclude
$1(\overline{\mu}_{0}=\mu_{0})(\overline{P}_{n}-\overline{P}_{0})\{B_{n,\diamond}^{*}-B_{0}\}=o_{p}(n^{-\frac{1}{2}}).$
The bootstrap case follows along similar lines, where we apply Lemma 12 to conclude that
$\frac{1}{J}\sum_{j=1}^{J}(P_{n,j}^{\#}-P_{n,j})\{m(\cdot,r_{n,j}^{\#})-m(\cdot,r_{0})\}=o_{p}(n^{-1/2}),$
and the bootstrap variants of the remainder conditions in Condition C7b also hold. In particular, by C3, it holds that
$\|r_{n,\diamond}^{*\#}-r_{0}\|_{\overline{P}_{0}}=O\left(\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}\right);$
$\|s_{n,\diamond}^{*\#}-s_{0}\|_{\overline{P}_{0}}=O\left(\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}\right).$
Under the setup of Theorem 4, we can apply Theorem 7 to conclude
$\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}=O_{p}\left(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}\right)+O_{p}(n^{-1/3});$
$\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}=O_{p}\left(\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}\right)+O_{p}(n^{-1/3}).$
Using D1-D4, an argument identical to the non bootstrap case, applying the second part of Lemma 9 and arguing as in the proof of Lemma 13 gives the second part of the lemma.
Appendix H Proofs of main results
H.1 Proof of Theorem 1
Proof of Theorem 1.
We will first show that $(P,\mu)\mapsto\psi_{P}(\mu)$ is total pathwise differentiable in the sense of Appendix A.4. in Luedtke (2024). To do so, we apply Lemma S6 in Luedtke (2024). To do so, we need to show that the total operator $(P,\mu)\mapsto\psi_{P}(\mu)$ is partially differentiable in its first argument and continuously partially differentiable in its second argument at $(P_{0},\mu_{0})$, where these notions of differentiability are defined in Appendix A.4 of Luedtke (2024). The former differentiability holds since, by assumption A3, $P\mapsto\psi_{P}(\mu_{0})$ is pathwise differentiable at $P_{0}$. To establish continuous partial differentiability in the second argument, note that $\mu\mapsto\psi_{P}(\mu)$ is a linear functional and, thus, the partial derivative of $(P,\mu)\mapsto\psi_{P}(\mu)$ in its second argument is simply $\mu\mapsto\psi_{P}(\mu)$. Hence, to establish continuous partial differentiability, it remains to show that $P\mapsto\psi_{P}$ is a continuous functional at $P_{0}$ with the domain equipped with the Hellinger distance and the range equipped with the operator norm. This is true by assumption A2. By Lemma S6 in Luedtke (2024), the total pathwise derivative at $(P_{0},\mu_{0})$ in the direction of $(s,h)\in T_{\mathcal{M}}(P_{0})\times\mathcal{H}$ is then given by $\langle\widetilde{\phi}_{0,\mu_{0}},s\rangle+\langle\alpha_{0},h\rangle$.
Under our assumptions, $P\mapsto\mu_{P}$ is pathwise differentiable at $P_{0}$ in a Hilbert-valued sense by Example 5 of Luedtke and Chung (2024), with pathwise derivative $d\mu_{0}(s)$ in the direction $s\in T_{\mathcal{M}}(P_{0})$ given by $h\mapsto\langle h\{\mathcal{I}_{Y}-\mu_{0}\},s\rangle$, where $\mathcal{I}_{Y}:o\mapsto y$. As a consequence, by the chain rule, $P\mapsto\psi_{P}(\mu_{P})$ is pathwise differentiable at $P_{0}$ with pathwise derivative given by $\langle\widetilde{\phi}_{0,\mu_{0}},s\rangle+\langle\alpha_{0}\{\mathcal{I}_{Y}-\mu_{0}\},s\rangle.$ It follows that the efficient influence function is $\widetilde{\phi}_{0,\mu_{0}}+D_{\mu_{0},\alpha_{0}}$. ∎
H.2 Proofs of Theorem 2, Lemma 2, and Lemma 1
Proof of Theorem 2.
Theorem 2 follows from Lemma 5 by taking the fold-specific nuisance estimators and fold-specific functional estimators to be identical, and by applying the orthogonality conditions in (4) and (5), under which the second term in both the outcome-favorable and Riesz-favorable decompositions vanishes. The Cauchy–Schwarz inequality, combined with Condition B1, establishes that the second-order remainders behave as $O_{p}\left(\|\mu_{n}-\mu_{0}\|^{2}\wedge\|\mu_{n}-\mu_{0}\|\|\alpha_{n}-\alpha_{0}\|\right)\quad\text{and}\quad O_{p}\left(\|\alpha_{n}-\alpha_{0}\|^{2}\wedge\|\mu_{n}-\mu_{0}\|\|\alpha_{n}-\alpha_{0}\|\right).$ See also the proofs in Appendix G.5, which derive an analogous expansion for the isocalibrated DML estimator. ∎
Proof of Lemma 2.
Denote $X:=(W,A)$. Let $g:\mathbb{R}^{2}\rightarrow\mathbb{R}$ be a Lipschitz continuous function with constant $L>0$. By Lipschitz continuity, we have that
$\left|g(\mu_{n}(x),\mu_{0}(x))-E[g(\mu_{n}(X),\mu_{0}(X))|\mu_{n}(X)=\mu_{n}(x)]\right|$ $=\left|E[g(\mu_{n}(x),\mu_{0}(x))-g(\mu_{n}(x),\mu_{0}(X))|\mu_{n}(X)=\mu_{n}(x)]\right|$
$\leq E[\left|g(\mu_{n}(x),\mu_{0}(x))-g(\mu_{n}(x),\mu_{0}(X))\right||\mu_{n}(X)=\mu_{n}(x)]$
$\leq LE[\left|\mu_{0}(x)-\mu_{0}(X)\right||\mu_{n}(X)=\mu_{n}(x)].$
On the event $\{\mu_{n}(X)=\mu_{n}(x)\}$, we know
$|\mu_{0}(x)-\mu_{0}(X)|$ $\leq|\mu_{0}(x)-\mu_{n}(x)|+|\mu_{n}(x)-\mu_{n}(X)|+|\mu_{0}(X)-\mu_{n}(X)|$
$\leq|\mu_{0}(x)-\mu_{n}(x)|+|\mu_{0}(X)-\mu_{n}(X)|.$
Therefore,
$\left|g(\mu_{n}(x),\mu_{0}(x))-E[g(\mu_{n}(X),\mu_{0}(X))|\mu_{n}(X)=\mu_{n}(x)]\right|$ $\lesssim E[\left|\mu_{0}(x)-\mu_{0}(X)\right|\mid\mu_{n}(X)=\mu_{n}(x)]$
$\lesssim E|\mu_{0}(x)-\mu_{n}(x)|+E[|\mu_{0}(X)-\mu_{n}(X)|\mid\mu_{n}(X)=\mu_{n}(x)].$
Now, for some function $f\in L^{2}(P_{X})$, suppose that $(\mu_{n}(x),\mu_{0}(x))\mapsto(\Pi_{\mu_{n},\mu_{0}}f)(x)$ is Lipschitz continuous. Then, defining $g:(\widehat{m},m)\mapsto E_{0}[f(X)|\mu_{n}(X)=\widehat{m},\mu_{0}(X)=m,\mathcal{D}_{n}]$ and noting by the law of iterated expectation that $\Pi_{\mu_{n}}\Pi_{\mu_{n},\mu_{0}}f=\Pi_{\mu_{n}}f$, we obtain the following pointwise error bound:
$\left|\Pi_{\mu_{n},\mu_{0}}f-\Pi_{\mu_{n}}f\right|\lesssim\left|\mu_{n}-\mu_{0}\right|+\Pi_{\mu_{n}}(|\mu_{n}-\mu_{0}|).$
Since $\|\Pi_{\mu_{n}}(|\mu_{n}-\mu_{0}|)\|_{L^{2}(P_{0})}\leq\|\mu_{n}-\mu_{0}\|_{L^{2}(P_{0})}$ by the properties of projections, it follows that
$\|\Pi_{\mu_{n},\mu_{0}}f-\Pi_{\mu_{n}}f\|_{L^{2}(P)}\lesssim\|\mu_{n}-\mu_{0}\|_{L^{2}(P)}.$
Taking $f$ to be $\Delta_{\alpha_{n}}$, we conclude that Condition B1B1i implies Condition B1enumiB1enumii. By an identical argument, Condition B1B1ii implies Condition B1enumiB1enumiii. The result then follows.
Proof of Lemma 1.
[…]
$\frac{d}{d\varepsilon}\psi_{n}(Z_{i},\alpha_{n}+\varepsilon\theta\circ\alpha_{n})\big{|}_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}\psi_{n}(Z_{i},\varepsilon\theta\circ\alpha_{n})\big{|}_{\varepsilon=0}$
$=\frac{d}{d\varepsilon}\varepsilon\psi_{n}(Z_{i},\theta\circ\alpha_{n})\big{|}_{\varepsilon=0}$
$=\psi_{n}(Z_{i},\theta\circ\alpha_{n}).$
See also Lemma 10, which establishes empirical calibration for the isocalibrated nuisance estimators. ∎
H.3 Proof of Theorem 3
Proof of Theorem 3.
Starting point. By Lemma 4, it holds that
$\tau_{n}^{*}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}+D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}\right\}+\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}+Rem_{\mu_{n,\diamond}^{*}}(\psi_{n,\diamond},\psi_{0}).$
where $Rem_{\mu_{n,\diamond}^{*}}(\psi_{n,\diamond},\psi_{0}):=\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(\mu_{n,j}^{*})-\psi_{0}(\mu_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{0,\mu_{n,j}^{*}}\right\}$. By direction application of C7, we have $Rem_{\mu_{n,\diamond}^{*}}(\psi_{n,\diamond},\psi_{0})=o_{p}(n^{-\frac{1}{2}})$. In addition, by Lemma 13 with C2, C5, and C6, we have that
$(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}+D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}\right\}$ $=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}$
$\quad+(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*}}+D_{\mu_{n,\diamond}^{*},\alpha_{n,\diamond}^{*}}-\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}\right\}$
$=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}+o_{p}(n^{-\frac{1}{2}}).$
Thus,
$\tau_{n}^{*}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}+\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}+o_{p}(n^{-\frac{1}{2}}).$
It remains to analyze the cross-product remainder term $\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$. We consider two cases.
Outcome-favorable case. Suppose that C2 holds with $\|\mu_{n,j}-\mu_{0}\|=o_{p}(n^{-1/4})$ for each $j\in[J]$, so that $\overline{\mu}_{0}=\mu_{0}$ in C2. By Lemma 11, we have the outcome-favorable decomposition:
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})A_{n,\diamond}^{*}+O_{p}\left(\|(\Pi_{(\mu_{n,\diamond}^{*},\mu_{0})}-\Pi_{\mu_{n,\diamond}^{*}})(\alpha_{n,\diamond}^{*}-\alpha_{0})\|_{\overline{P}_{0}}\|\mu_{0}-\mu_{n,\diamond}^{*}\|_{\overline{P}_{0}}\right),$
By C2 with $\overline{\mu}_{0}=\mu_{0}$, C5, and a direct application of Lemma 14, we find
$(\overline{P}_{n}-P_{0})A_{n,\diamond}^{*}$ $=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+(\overline{P}_{n}-P_{0})\left\{A_{n,\diamond}^{*}-A_{0}\right\}$
$=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+o_{p}(n^{-\frac{1}{2}}).$
Next, by C3, it holds that
$\|(\Pi_{(\mu_{n,\diamond}^{*},\mu_{0})}-\Pi_{\mu_{n,\diamond}^{*}})(\alpha_{n,\diamond}^{*}-\alpha_{0})\|_{\overline{P}_{0}}=O_{p}\left(\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{n}}\right).$
Thus, by C2, we find
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+O_{p}\left(\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{n}}^{2}\right)$
$=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+o_{p}(n^{-\frac{1}{2}}).$
Riesz-favorable case. Suppose that C2 holds with $\|\alpha_{n,j}-\alpha_{0}\|=o_{p}(n^{-1/4})$ for each $j\in[J]$, so that $\overline{\alpha}_{0}=\alpha_{0}$ in C2. By Lemma 11, we have the Riesz-favorable decomposition:
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=(\overline{P}_{n}-P_{0})B_{n,\diamond}^{*}+\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(r_{n,j}^{*})-\psi_{0}(r_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}^{*}}\right\}$
$\quad+O_{p}(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}\|(\Pi_{(\alpha_{n,\diamond}^{*},\alpha_{0})}-\Pi_{\alpha_{n,\diamond}^{*}})(\mu_{0}-\mu_{n,\diamond}^{*})\|_{\overline{P}_{0}},$
By C7, it holds that $\frac{1}{J}\sum_{j=1}^{J}\left\{\psi_{n,j}(r_{n,j}^{*})-\psi_{0}(r_{n,j}^{*})-(P_{n,j}-P_{0})\widetilde{\phi}_{r_{n,j}^{*}}\right\}=o_{p}(n^{-\frac{1}{2}})$. Furthermore, by C2 with $\overline{\alpha}_{0}=\alpha_{0}$, C5, and a direct application of Lemma 14, we find
$(\overline{P}_{n}-P_{0})B_{n,\diamond}^{*}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+(\overline{P}_{n}-P_{0})\left\{B_{n,\diamond}^{*}-B_{0}\right\}$
$=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Finally, by C3, it holds that
$\|(\Pi_{(\alpha_{n,\diamond}^{*},\alpha_{0})}-\Pi_{\alpha_{n,\diamond}^{*}})(\mu_{0}-\mu_{n,\diamond}^{*})\|_{\overline{P}_{0}}=O_{p}\left(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{n}}\right).$
Thus, by C2,
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+O_{p}(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}^{2}$
$=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Completing the argument. Combining the two cases treated above, we find
$\langle\alpha_{n,\diamond}^{*}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*}\rangle_{\overline{P}_{0}}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+O_{p}(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}^{2})$
$=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Thus, it holds that
$\tau_{n}^{*}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}+1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}-P_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Note that $\overline{P}_{0}\{\widetilde{\phi}_{0,\overline{\mu}_{0}}+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}=\overline{P}_{0}\{{\phi}_{0,\overline{\mu}_{0}}-\psi_{0}(\mu_{0})+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\}$ if either $\overline{\mu}_{0}=\mu_{0}$ or $\overline{\alpha}_{0}=\alpha_{0}$. Recalling that $\chi_{0}={\phi}_{0,\overline{\mu}_{0}}-\psi_{0}(\mu_{0})+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}+1(\overline{\alpha}_{0}\neq\alpha_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})B_{0}$ and that $\chi_{0}$ is mean zero, we obtain the desired asymptotic linearity result:
$\tau_{n}^{*}-\tau_{0}=P_{n}\chi_{0}+o_{p}(n^{-\frac{1}{2}}).$
If both nuisance functions are consistently estimated ($\overline{\alpha}_{0}=\alpha_{0}$ and $\overline{\mu}_{0}=\mu_{0}$), then the above display reduces to
$\tau_{n}^{*}-\tau_{0}=P_{n}D_{0}+o_{p}(n^{-\frac{1}{2}}).$
By Theorem 1, the above means that $\tau_{n}^{*}$ is asymptotically linear with influence function being the efficient influence function of $\Psi$ under the nonparametric statistical model $\mathcal{M}$. As a consequence, by van der Vaart (2000), it holds that $\tau_{n}^{*}$ is a regular and efficient estimator of $\psi_{0}$ under the nonparametric statistical model $\mathcal{M}$. ∎
H.4 Proof of Theorem 4
Proof of Theorem 4.
Starting point. By Lemma 4, it holds that
$\tau_{n}^{*\#}-\tau_{0}$ $=(\overline{P}_{n}^{\#}-\overline{P}_{0})\left\{\widetilde{\phi}_{0,\mu_{n,\diamond}^{*\#}}+D_{\mu_{n,\diamond}^{*\#},\alpha_{n,\diamond}^{*\#}}\right\}+\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}+Rem_{\mu_{n,\diamond}^{*\#}}(\psi_{n,\diamond},\psi_{0})$
$=(\overline{P}_{n}^{\#}-\overline{P}_{0})\left\{m(\cdot,\mu_{n,\diamond}^{*\#})+D_{\mu_{n,\diamond}^{*\#},\alpha_{n,\diamond}^{*\#}}\right\}+\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}},$
where we used that $\widetilde{\phi}_{0,\mu_{n,\diamond}^{*\#}}=m(\cdot,\mu_{n,\diamond}^{*\#})$ and $Rem_{\mu_{n,\diamond}^{*\#}}(\psi_{n,\diamond},\psi_{0})=0$ for the setup of Section 5.2.
Arguing as in the Proof of Theorem 3, we can apply the second part of Lemma 13 with C2, C5, D1, C6, and D2, to find that
$(\overline{P}_{n}^{\#}-\overline{P}_{0})\left\{m(\cdot,\mu_{n,\diamond}^{*\#})+D_{\mu_{n,\diamond}^{*\#},\alpha_{n,\diamond}^{*\#}}\right\}=(\overline{P}_{n}^{\#}-\overline{P}_{0})\left\{m(\cdot,\overline{\mu}_{0})+D_{\overline{\mu}_{0},\overline{\alpha}_{0}}\right\}+o_{p}(n^{-\frac{1}{2}}).$
Under the setup of Theorem 4 and D3 , we can apply Theorem 7 to conclude
$\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}=O_{p}\left(\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{0}}\right)+O_{p}(n^{-1/3});$
$\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}=O_{p}\left(\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}\right)+O_{p}(n^{-1/3}).$
Outcome-favorable case. Suppose that C2 holds with $\|\mu_{n,j}-\mu_{0}\|=o_{p}(n^{-1/4})$ for each $j\in[J]$, so that $\overline{\mu}_{0}=\mu_{0}$ in C2. An argument identical to that of the Proof of Theorem 3 using D4 and the second part of Lemma 14 establishes that
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}$ $=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}^{\#}-P_{0})A_{0}+O_{p}\left(\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}^{2}\right).$
By Theorem 7, it holds that $\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}^{2}=O_{P}(n^{-2/3})+\|\mu_{n,\diamond}^{*}-\mu_{0}\|_{\overline{P}_{0}}^{2}$, and thus, by C2, we have that $\|\mu_{n,\diamond}^{*\#}-\mu_{0}\|_{\overline{P}_{0}}^{2}=o_{p}(n^{-\frac{1}{2}})$. We conclude that
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}$ $=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}^{\#}-P_{0})A_{0}+o_{p}(n^{-\frac{1}{2}}).$
Riesz-favorable case. Suppose that C2 holds with $\|\alpha_{n,j}-\alpha_{0}\|=o_{p}(n^{-1/4})$ for each $j\in[J]$, so that $\overline{\alpha}_{0}=\alpha_{0}$ in C2. An argument identical to that of the Proof of Theorem 3 using D4 and the second part of Lemma 14 establishes that
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}^{\#}-P_{0})B_{0}+O_{p}\left(\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}^{2}\right).$
By Theorem 7, it holds that
$\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{0}}^{2}=O_{P}(n^{-2/3})+\|\alpha_{n,\diamond}^{*}-\alpha_{0}\|_{\overline{P}_{n}}^{2},$
and, thus, by C2, $\|\alpha_{n,\diamond}^{*\#}-\alpha_{0}\|_{\overline{P}_{n}}^{2}=o_{p}(n^{-\frac{1}{2}})$. We conclude that
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}$ $=1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}^{\#}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Completing the argument. Combining the two cases treated above, we find
$\langle\alpha_{n,\diamond}^{*\#}-\alpha_{0},\mu_{0}-\mu_{n,\diamond}^{*\#}\rangle_{\overline{P}_{0}}=1(\overline{\alpha}_{0}\neq\alpha_{0})(\overline{P}_{n}^{\#}-P_{0})A_{0}+1(\overline{\mu}_{0}\neq\mu_{0})(\overline{P}_{n}^{\#}-P_{0})B_{0}+o_{p}(n^{-\frac{1}{2}}).$
Thus, it holds that
$\tau_{n}^{*\#}-\tau_{0}=(\overline{P}_{n}^{\#}-\overline{P}_{0})\chi_{0}+o_{p}(n^{-\frac{1}{2}}).$
By Theorem 3, we also have that $\tau_{n}^{*}-\tau_{0}=(\overline{P}_{n}-\overline{P}_{0})\chi_{0}+o_{p}(n^{-\frac{1}{2}})$, from which we can then write that $\tau_{n}^{*\#}-\tau_{n}^{*}=\tau_{n}^{*\#}-\tau_{n}^{*}+\tau_{n}^{*}-\tau_{n}^{*}=(\overline{P}_{n}^{\#}-\overline{P}_{n})\chi_{0}+o_{p}(n^{-\frac{1}{2}})$.
For each $j\in[J]$, by the bootstrap CLT (van der Vaart and Wellner, 1996), we have $(P_{n,j}^{\#}-P_{n,j})\chi_{0}|P_{n,j}$ converges in distribution to $N(0,P_{0}\chi_{0}^{2})$. Hence, by the mutual independence of $\{P_{n,j}:j\in[J]\}$, we also have $(P_{n,j}^{\#}-P_{n,j})\chi_{0}|\{P_{n,j}:j\in[J]\}$ converges in distribution to $N(0,P_{0}\chi_{0}^{2})$. Again by mutual independence of $\{P_{n,j}:j\in[J]\}$, we have $(\overline{P}_{n}^{\#}-\overline{P}_{n})\chi_{0}|\{P_{n,j}:j\in[J]\}$ converges in distribution to $N(0,P_{0}\chi_{0}^{2})$. Thus, the law of $\tau_{n}^{*\#}-\tau_{n}^{*}|\{P_{n,j}:j\in[J]\}$ converges in probability to the law of $N(0,P_{0}\chi_{0}^{2})$.
H.5 Proof of Theorem 5
Proof of Theorem 5.
If $\overline{\alpha}_{0}=\alpha_{0}$ and $\overline{\mu}\neq\mu_{0}$, it follows that
$\chi_{0}$ $=m(\cdot,\overline{\mu}_{0})-\psi_{0}(\mu_{0})+D_{\overline{\mu}_{0},\alpha_{0}}+B_{0}$
$=m(\cdot,\overline{\mu}_{0})-\psi_{0}(\mu_{0})+\alpha_{0}\{\mathcal{I}_{Y}-\overline{\mu}_{0}\}+m(\cdot,r_{0})-r_{0}\alpha_{0}$
$=m(\cdot,\overline{\mu}_{0}+r_{0})+\alpha_{0}\{\mathcal{I}_{Y}-\overline{\mu}_{0}-r_{0}\}-\psi_{0}(\mu_{0})$
Now, we have that $\psi_{0}(\mu_{0})=\psi_{0}(\overline{\mu}_{0}+r_{0})$ by the orthogonality conditions defining $r_{0}$. Thus, using that $\overline{\mu}_{0}+r_{0}=\overline{\Pi}_{0,P_{0}}\mu_{0}$, we have
$\chi_{0}$ $=m(\cdot,\overline{\mu}_{0}+r_{0})-\psi_{0}(\overline{\mu}_{0}+r_{0})+\alpha_{0}\{\mathcal{I}_{Y}-\overline{\mu}_{0}-r_{0}\}$
$=m(\cdot,\overline{\Pi}_{0,P_{0}}\mu_{0})-\psi_{0}(\overline{\Pi}_{0,P_{0}}\mu_{0})+\alpha_{0}\{\mathcal{I}_{Y}-\overline{\Pi}_{0,P_{0}}\mu_{0}\},$
where the right-hand side of the above display is the $P_{0}$-efficient influence function of the parameter $P\mapsto\psi_{P}(\overline{\Pi}_{0,P}\mu_{P})$. The latter claim follows from the proof of Theorem 1 after replacing the outcome $Y$ with the residual $Y-\overline{\mu_{0}}$.
If $\overline{\alpha}_{0}\neq\alpha_{0}$ and $\overline{\mu}=\mu_{0}$, it follows that
$\chi_{0}$ $=m(\cdot,\mu_{0})-\psi_{0}(\mu_{0})+D_{\mu_{0},\overline{\alpha}_{0}}+A_{0}$
$=m(\cdot,\mu_{0})-\psi_{0}(\mu_{0})+D_{\mu_{0},\overline{\alpha}_{0}}+s_{0}\{\mathcal{{I}_{Y}}-\mu_{0}\}$
$=m(\cdot,\mu_{0})-\psi_{0}(\mu_{0})+\{\overline{\alpha}_{0}+s_{0}\}\{\mathcal{{I}_{Y}}-\mu_{0}\}$
Thus, using that $\overline{\alpha}_{0}+s_{0}=\overline{\Pi}_{0,P_{0}}^{*}\alpha_{0}$, we have $\chi_{0}=m(\cdot,\mu_{0})-\psi_{0}(\mu_{0})+\overline{\Pi}_{0,P_{0}}^{*}\alpha_{0}(\mathcal{I}_{Y}-\mu_{0})$. We claim that the right-hand side of the above display is the $P_{0}$–efficient influence function of the parameter $P\mapsto\overline{\psi}_{0,P}(\mu_{P})=\langle\overline{\Pi}_{0,P}^{*}\alpha_{P},\mu_{P}\rangle_{P}$. To see this, we note that $\overline{\psi}_{0,P}(\mu_{P})=\langle\overline{\Pi}_{0,P}^{*}\alpha_{P},\mu_{P}\rangle_{P}=\langle\overline{\alpha}_{0}+\Pi_{\mu_{0}}\{\alpha_{P}-\overline{\alpha}_{0}\},\mu_{P}\rangle_{P}=\langle\overline{\alpha}_{0}+s_{0},\mu_{P}\rangle_{P}$ is a linear functional of $\mu_{P}$ with Riesz representer $\overline{\alpha}_{0}+s_{0}$. The claim then follows from the proof of Theorem 1.
Since $\psi_{n}^{*}$ is $P_{0}$-asymptotically linear with the influence function being the $P_{0}$-EIF of the oracle parameter $\Psi_{0}$, it follows that $\psi_{n}^{*}$ is a regular and efficient estimator for $\Psi_{0}$ at $P_{0}$. The stated limiting distribution result follows directly from the definition of regularity.
Appendix I Additional details on numerical experiments
I.1 Complete results for experiment 1
Figure 3: Empirical relative bias, standard error, and 95% confidence interval coverage for isocalibrated DML (IC-DML), DR-TMLE and AIPW estimators under both doubly consistent estimation of the outcome regression and treatment mechanism. Both propensity score and outcome regression are estimated consistently.
I.2 Complete results for experiment 2
Table 4: Evaluation of absolute bias, scaled root mean square error (RMSE), and 95% confidence-interval coverage. Both outcome regression and propensity scores are estimated with gradient-boosted trees.
Dataset Bias RMSE Coverage
AIPW C-DML AIPW C-DML AIPW C-DML
ACIC-2017 (17) 0.031 0.026 0.095 0.094 0.95 0.94
ACIC-2017 (18) 0.21 0.28 0.7 0.58 0.27 0.64
ACIC-2017 (19) 0.082 0.057 0.43 0.43 0.96 0.96
ACIC-2017 (20) 1.6 0.2 2 1.4 0.32 0.90
ACIC-2017 (21) 0.006 0.004 0.018 0.017 1 1
ACIC-2017 (22) 0.004 0.035 0.11 0.096 0.56 0.81
ACIC-2017 (23) 0.014 0.01 0.08 0.08 0.97 0.96
ACIC-2017 (24) 0.3 0.04 0.34 0.25 0.32 0.90
ACIC-2018 (1000) 290 260 4300 3900 0.57 0.65
ACIC-2018 (10000) 17 13 690 680 0.55 0.68
ACIC-2018 (2500) 85 99 1300 1300 0.59 0.65
ACIC-2018 (5000) 520 510 4100 4100 0.54 0.64
IHDP 0.13 0.13 0.46 0.46 0.57 0.57
Lalonde CPS 0.14 0.084 0.22 0.34 0.16 0.75
Lalonde PSID 0.04 0.04 0.19 0.44 0.46 0.84
Twins 0.22 0.21 0.24 0.23 0.51 0.54